A Bibliography of Publications in IEEE Computer, 2020–2029

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

04 October 2025 Version 1.62

Title word cross-reference

3 [105]. 5 [77]. Π [504, 967].

-Learner [967]. -RT [504].

1 [2162, 2196, 2424]. 1.0 [2292]. 10-Year [440]. 101 [543]. 12 [945]. 1838 [739]. 19 [208, 223, 313, 315, 385, 485, 1067, 1291, 1370, 2000].

2.0 [2202, 2292]. 2020 [2, 31, 48, 61, 79, 282]. 2021 [254, 388, 395, 439, 889]. 2022 [494, 763, 864]. 2023 [1196, 1243, 1287, 1341, 1425, 2219]. 2024 [1642, 1771, 2084, 2259]. 2025 [2288]. 21st [97, 1505, 1541]. 25 [2104]. 25-Year [78]. 25-Year-Old [2149].

3D [1618, 1742, 2040, 2271, 2537]. 3E [1348].

4.0 [266, 1414, 1657, 2008]. 4600 [1375]. 4C [1934].

5G [2139].

6G [2139].

802 [379].

AAAI [89, 494]. ABET [1899]. Abilities [2471]. Able [189]. Academia [168]. Academic [1481, 1668, 2067, 2486]. Academy [210, 713, 886, 1081, 1911]. Accelerated [2197]. Acceleration [176, 512, 1888, 2530]. Accelerator [2153]. Accelerators [39, 813, 1293, 2337]. Acceptable [2487]. Acceptance [998].

Access [307, 322, 995, 1296, 1442, 2200, 2414]. Accessibility [1867]. Accessible [809, 1850]. Accidents [2104]. Accompaniment [1845]. Accomplishments [724]. Accountability [568, 1332, 1659]. Accountable [1372]. Accuser [501]. ACM [406]. ACM/IEEE [406]. Acronyms [2524]. Across [1660, 2068]. Action [241, 2332, 2335]. Actionable [276]. Actions [2112]. Active [2185]. Activities [46, 2117]. Activity [525, 1745]. Acts [1591, 2508]. Adapting [46, 1977]. Adaptive [67, 73, 227, 366, 2414]. Added [10]. Adding [176, 1268]. Addressing [100, 140, 487, 1819, 1889]. Admission [1081]. Admissions [21]. Ado [2300, 2481, 2519]. Adolescent [1152]. Adopt [478]. Adopted [478]. Adoption [998, 1885, 1924]. Advance [992, 1454, 1924, 2017]. Advance-Warning [2017]. Advanced [372, 1073, 1927]. Advances [525]. Advancing [927, 1035, 2236]. Advantage [908]. Advantages [1742]. Adversarial [540, 1479, 1664, 2388]. Adversity [396]. Advisory [694]. AempOS [1578]. Aerial [2193]. Affairs [209]. Affecting [998, 1332, 1510]. Affordances [1112]. Aftershocks [1146]. Against [135, 946, 1067, 1147, 1522, 1664, 1732, 2051, 2101, 2535]. Age [313, 411, 514, 1630, 1768, 2110, 2429, 2453]. Agenda [227, 1623]. Agent [70]. Agent-Based [70]. Agentic [2270, 2302, 2303, 2333, 2349, 2429, 2507]. Agents [231, 896, 2189, 2202, 2298, 2349, 2507, 2516]. Aggregated [236]. Agile [766, 972]. Agility [1045]. Aging [941]. Ago [8, 36, 65, 95, 128, 158, 191, 221, 259, 286, 327, 356, 401, 435, 469, 499, 528, 557, 584, 619, 654, 687, 723, 756, 792, 828, 863, 898, 929, 959, 990, 1026, 1058, 1096, 1133, 1169, 1209, 1251, 1283, 1318, 1357, 1397, 1430, 1467, 1500, 1534, 1569, 1606, 1645, 1688, 1724, 1758, 1798, 1839, 1879, 1920, 1955, 1989, 2038, 2090, 2133, 2174, 2211, 2248, 2284, 2320, 2360, 2403, 2456, 2502]. Agribusiness [1190]. Agricultural [1234, 2097]. Agriculture [1189, 2091]. Ahead [2386]. Ahrens [1049]. AI [89, 196, 224, 419, 478, 494, 607, 670, 673, 696, 698, 774, 808, 907, 948, 1013, 1228, 1254, 1259, 1263, 1299, 1321, 1324, 1332, 1347, 1360, 1363, 1364, 1370, 1403, 1577, 1589, 1593, 1698, 1699, 1712, 1725, 1730, 1736, 1737, 1779, 1807, 1827, 1893, 1924, 1928, 1936, 1938, 1962, 1967, 2012, 2015, 2043, 2058, 2064, 2065, 2067, 2068, 2096, 2101, 2134, 2160, 2187, 2189, 2197, 2250, 2263, 2265, 2270, 2298, 2302, 2304, 2306, 2349, 2356, 2364, 2368, 2370, 2389, 2391, 2412, 2420, 2427, 2429, 2431, 2440, 2487, 2490, 2505, 2508, 2517, 2520, 2524, 2526, 2527, 2534]. AI [1368, 1733, 1749, 1815, 1822, 1897, 1960, 1974, 2061, 2196, 2234, 2333, 2366, 2372, 2384, 2387, 2414, 2425, 2510]. AI- [1938]. AI-Accelerated [2197]. AI-Assisted [2366]. AI-Based [1733]. AI-Driven [2061]. AI-Generated [2067]. AI-Modular [2510]. AI/ML [1960]. AI/ML-Centric [1960]. AIChat [1385]. Aids [746]. Airborne [1224]. Aircraft [1440, 2271]. Airspace [1440]. Alaska [2417]. Alexa [850]. Algorithm [232, 630, 906, 1339, 1579, 1898]. Algorithmic [1117, 1416, 1486]. Algorithmics [702]. Algorithms [329, 405, 702, 900, 906, 1009, 1261, 1333, 1763, 1881, 1897, 2012, 2356]. Alignment [766]. Allen [347]. Almost [453]. Alone [2526]. Also [1587, 1727, 1802, 2095, 2460]. Alternating [2535]. Altitude [1440]. Always [453, 838, 947, 1702]. Ambassador [179]. America [1014]. American [1308]. Amid [385]. Among [481, 947, 1370, 1447, 2387]. Analogies [1370]. Analogy [805]. Analysis [69, 70, 183, 311, 487, 908, 1015, 1421, 1546, 1770, 1963, 1970, 1971, 2047, 2052, 2144, 2152, 2257, 2490]. Analytics [442, 1566, 2221]. Analyze [34].

Analyzing [2099]. Ancient [971, 1302]. and/or [1230]. Angle [633]. Animators [105]. Annals [163, 257, 298, 358, 399, 454, 470, 548, 574, 594, 623, 676, 700, 734, 754, 832, 1012, 1124, 1167, 1231, 1270, 1322, 1381, 1452, 1557, 1592, 1628, 1667, 1714, 1743, 1784, 1803, 1910, 2021, 2086, 2141, 2192, 2233, 2390, 2439, 2470, 2533]. Anne [1123]. Announced [889]. Ant [140, 1336]. Anthropomorphic [2467]. Anti [82]. Anticipated [360]. Antiquities [1647]. Anybody [725]. Anymore [1388]. Apache [515]. Apart [592]. APIs [1367]. App [444, 2443, 2493]. Appear [735, 2340]. Application [405, 537, 731, 1121, 2503]. Application-Level [537]. Applications [18, 99, 133, 136, 169, 267, 288, 361, 404, 452, 473, 474, 504, 546, 570, 628, 671, 706, 736, 835, 870, 1062, 1187, 1189, 1229, 1271, 1288, 1300, 1369, 1417, 1469, 1575, 1631, 1650, 1661, 1699, 1805, 1885, 1906, 1909, 1946, 1960, 1963, 1982, 2018, 2123, 2190, 2223, 2228, 2266, 2293, 2345, 2373, 2376, 2428, 2472, 2531]. Applying [374, 969]. Approach [369, 372, 506, 565, 630, 1235, 1486, 1577, 1593, 1762, 1965, 2050, 2112, 2114]. Approaches [269, 1332, 1977, 2380, 2532]. Apps [1846]. AR-Based [1581]. Arc [2162, 2224]. Architecting [1076]. Architects [2020]. Architectural [366, 487, 872]. Architecture [73, 74, 93, 136, 197, 373, 397, 477, 588, 707, 771, 939, 1328, 1347, 1497, 1506, 2202, 2285, 2408]. Architectures [236, 504, 595, 703, 1384, 1845, 2416, 2419]. Area [132, 134]. Arithmetic [1093]. Arm [1428]. Arms [701, 2302, 2517]. Army [1014]. Arrays [1888]. Arrest [2159]. Arrive [1273]. Art [817, 1783, 1895]. ARTEMIS [340]. Article [2479]. Articles [19, 66, 107, 112, 171, 173, 190, 205, 207, 306, 308, 328, 339, 370, 414, 416, 445, 448, 482, 484, 509, 535, 538, 586, 608, 610, 635, 638, 659, 661, 767, 769, 798, 802, 841, 845, 876, 1044, 1109, 1115, 1153, 1159, 1193, 1197, 1227, 1295, 1338, 1386, 1484, 1509, 1514, 1547, 1549, 1582, 1597, 1619, 1654, 1677, 1818, 1863, 1900, 2006, 2011, 2069, 2070, 2092, 2437]. Artificial [39, 89, 96, 189, 225, 227, 276, 292, 294, 299, 302, 336, 337, 383, 398, 415, 572, 636, 690, 741, 775, 830, 844, 852, 867, 873, 878, 906, 918, 943, 963, 997, 1083, 1120, 1189, 1243, 1257, 1287, 1326, 1328, 1330, 1341, 1342, 1353, 1378, 1380, 1411, 1474, 1481, 1487, 1513, 1519, 1554, 1585, 1617, 1621, 1633, 1651, 1653, 1668, 1679, 1680, 1693, 1701, 1715, 1729, 1776, 1854, 1856, 1959, 1961, 1966, 1977, 1981, 2003, 2010, 2022, 2046, 2055, 2110, 2200, 2202, 2258, 2286, 2303, 2308, 2317, 2322, 2326, 2330, 2347, 2400, 2436, 2468, 2471]. Artificial [228, 296, 301, 335, 340, 360, 692, 800, 877, 883, 1008, 1009, 1112, 1256, 1293, 1486, 1707, 1747, 1849, 1859, 1925, 1965, 1975, 2051, 2071, 2139, 2477]. Artificial-Intelligence-Based [1961]. Asia [881, 1710]. Ask [1702]. Aspect [2230]. Aspect-Oriented [2230]. Assembly [1581]. Assessing [335, 808, 1405, 1707]. Assessment [189, 732, 1260, 1474]. Assistance [2175]. Assistants [2180]. Assisted [43, 2366, 2469]. Associated [906]. Assurance [363, 364, 626, 878, 1440, 1889, 2368, 2419]. Assurances [873]. Assured [1038, 1729, 1817, 2326]. Assuring [1517]. Asymmetry [846]. Athavale [1196, 2084]. Attack [139, 516, 731, 957, 1815, 2339, 2388]. Attacking [138, 142]. Attacks [71, 135, 230, 372, 540, 1102, 1147, 1664, 1995, 2051, 2101, 2185]. Attention [419, 2149]. Attributes [1326]. Attribution [2438]. Auction [157]. Auditing [197, 627, 906]. Augmented [17, 967, 1620, 1627, 1962, 2134, 2139, 2466]. Augmenting [227]. Australia [1891]. Authenticated [920]. Authenticating [510]. Authentication [1143, 2414]. Author [2127, 2169, 2206, 2243, 2279]. Authorship [1617]. Autism [41]. Autodocumentation [1823]. Automated [50, 111, 626, 875, 1003, 1290, 2152, 2349, 2384].

Automatic [702]. Automation [197, 601, 602, 665, 1624, 2220, 2532]. Automotive [1286, 1963]. Autonomous [334, 362, 514, 625, 627, 629, 632, 633, 704, 727, 764, 804, 967, 1063, 1257, 1258, 1271, 1375, 1454, 2017, 2046, 2050, 2052, 2101, 2157, 2189, 2429]. Autonomy [364, 533, 697, 743, 1440, 1729, 1762, 1817, 2326]. Availability [998, 1660, 1770]. Avatar [41, 2467]. Avatar-Based [41]. Aviation [2223]. Avoid [2161]. Avoiding [1226, 1701]. Awakening [1011]. Award [89, 406, 425, 494, 507, 964, 1024, 1343, 1371, 1404, 1437, 1476, 2085]. Awards [121, 359, 468, 517, 760, 799, 1111, 1172, 1239, 1323, 1412, 1415, 1438, 1473, 1490, 1515, 1561, 1611, 1632, 1682, 1717, 1785, 1866, 1902, 1937, 1978, 2116, 2338, 2383, 2488, 2518]. Aware [1376, 1463, 1886, 2514]. Awareness [305]. Away [2230].

B [1590]. B-Schools [1590]. Back [637, 806, 1302, 2145]. Backyard [1662]. Bad [398]. Balancing [1938]. Banality [1149]. Bandwidth [1864]. Bank [569, 881, 1710]. Banking [656, 709, 803]. Barriers [395, 1187, 1220, 1745, 2184]. Based [13, 41, 42, 70, 132, 135, 230, 266, 302, 446, 627, 630, 639, 702, 730, 771, 782, 1040, 1102, 1230, 1235, 1266, 1372, 1543, 1544, 1579, 1581, 1733, 1814, 1815, 1848, 1961, 2052, 2161, 2290, 2414, 2535]. Battery [1744]. BCI [2149]. Be [104, 236, 451, 789, 1011, 1147, 2007, 2340]. Beacon [1022]. Because [2340]. Becoming [2215]. BeeMe [231]. Been [947]. Before [782]. Begins [1487]. Begun [995]. Behave [413]. Behavior [232, 1512]. Behind [377]. Bell [934]. Belong [599]. Benchmarking [1037, 2440]. Benchmarks [2292]. Beneficial [1740]. Benefit [89, 1546, 1966, 2505]. Benefits [22, 123, 185, 300, 461, 492, 520, 551, 576, 647, 680, 722, 818, 1090, 1128, 1163, 1203, 1245, 1279, 1314, 1361, 1409, 1434, 1464, 1573, 1616, 1644, 1690, 1702, 1707, 1728, 1828, 1843, 1890, 1935, 1973, 2089, 2132, 2210, 2252, 2283, 2319, 2363, 2409, 2457, 2511]. Benoit [1123]. Best [940]. Better [637, 1333, 1898, 2392]. Between [398, 2392, 2526]. Beyond [55, 305, 312, 383, 974, 1867, 2114, 2139, 2304, 2427]. Bias [789, 1367, 1416]. Biases [2003]. Big [25, 52, 116, 146, 178, 200, 223, 238, 293, 378, 420, 490, 513, 530, 776, 814, 849, 885, 1113, 1145, 1148, 1177, 1188, 1219, 1331, 1377, 1409, 1594, 1615, 1621, 1663, 1679, 1779, 1789, 1830, 1860, 1938, 2004, 2056, 2057, 2102, 2186, 2235, 2250, 2327, 2367, 2426, 2480, 2521]. Bike [1186]. Bike-Sharing [1186]. Bill [2269]. Bimonthly [675]. Biological [644]. Biometric [45, 978, 1436, 2255]. Biometrics [670, 1657, 1840, 1847]. Bitcoin [807, 975]. Black [228, 1370, 1788]. Black-Box [1370, 1788]. Blind [108]. Block [1067]. Blockbuster [2263]. Blockchain [13, 18, 104, 166, 195, 197–199, 201, 264–266, 268, 412, 444, 446, 627, 639, 771, 1041, 1294, 1359, 2014, 2187, 2289, 2344, 2380, 2410, 2483]. Blockchain-Based [13, 266, 627, 639, 771]. Blockchain-Enabled [268]. Blockchains [269, 377]. Blue [2339]. Bluetooth [842]. Blunted [1932]. Board [29, 46, 59, 87, 244, 270, 349, 390, 402, 488, 541, 603, 715, 748, 784, 822, 922, 952, 982, 1017, 1051, 1088, 1126, 1161, 1201, 1277, 1945, 2442]. Boards [498, 562, 674, 891, 1309]. Body [566]. Bongshin [1049]. Bono [1702]. Boosting [1117]. Both [263]. Bots [1503]. Bottlenecks [1520]. Boundaries [1867]. Box [305, 1370, 1788]. Boxes [228]. BPCNet [160]. Brain [939]. Brain-Inspired [939]. Brains [176, 774]. Brawn [176]. Brazil [274, 301, 1190]. Break [2356]. Breaking [2184]. Bridges [2526]. Bridging [1105, 2392, 2441]. Brief [278, 455, 2391]. Bring [2296]. Broad [1924]. Broadband [995]. Brokering [1294]. Brooks [1450]. Bruce [2064, 2065]. Bryan [1015]. Buggy [2384]. Build [880, 2474].

Building [197, 243, 809, 1085, 1288, 1483, 1812, 2187, 2335, 2526]. Buildings [49]. Built [774]. Bullet [360, 447]. Bullsh*t [2340]. Bungie [537]. Burst [419]. Business [568, 777, 1188, 1548, 2265]. Byzantine [1302, 1579].

C [344, 1370, 1548]. C-19 [1370]. C-Suites [1548]. CAI [1243, 1287]. Calculations [806]. Calibrated [693]. Call [19, 66, 89, 107, 112, 129, 141, 153, 171, 173, 180, 190, 194, 196, 205, 207, 214, 217, 224, 226, 250, 273, 275, 304, 306, 308, 314, 328, 339, 342, 365, 370, 410, 414, 416, 433, 439, 445, 448, 450, 480, 482, 484, 494, 507, 509, 517, 526, 535, 538, 559, 586, 598, 608, 610, 631, 635, 638, 642, 659, 661, 664, 740, 765, 767, 769, 779, 796, 798, 833, 839, 841, 864, 871, 874, 904, 917, 996, 1028, 1044, 1094, 1115, 1132, 1171]. Call [802, 845, 876, 1109, 1141, 1153, 1159, 1182, 1193, 1197, 1223, 1227, 1292, 1295, 1325, 1329, 1338, 1358, 1371, 1386, 1391, 1415, 1432, 1466, 1484, 1490, 1498, 1509, 1514, 1532, 1547, 1549, 1567, 1582, 1597, 1608, 1654, 1669, 1677, 1694, 1734, 1766, 1797, 1818, 1838, 1863, 1878, 1900, 1933, 1964, 1980, 1988, 2006, 2009, 2011, 2048, 2049, 2069, 2070, 2092, 2098, 2136, 2138, 2176, 2201, 2214, 2237, 2247, 2274, 2291, 2309, 2325, 2334, 2357, 2381, 2402, 2446, 2450, 2455, 2476, 2497, 2501, 2506]. Call [241, 451, 1619, 1897, 2158, 2182, 2339, 2437]. Camera [976]. Campaign [1349]. Campaigns [1922]. Can [104, 263, 291, 398, 485, 701, 704, 789, 815, 877, 908, 1046, 1065, 1147, 1333, 1583, 1897, 1903, 2229, 2384]. Cancel [713]. Candidates [246, 678, 1048, 1489, 1945, 1947, 2442, 2444]. Cannot [2007]. Canvas [1188]. CAP [53]. Capabilities [2052]. Capability [2329]. Capacity [67]. Capitalism [1445]. Capsules [2512]. Capturing [1545]. Car [2047]. Carbon [993, 1041, 1135, 1376]. Carbon-Aware [1376]. Care [42, 201, 636, 1107, 1189, 1508, 1543, 1897, 1899, 2008, 2391, 2510]. Career [1351, 1362, 1423, 1458, 1492, 1526, 1558, 1599, 1629, 1681, 1716, 1750, 1790, 1831, 1871, 1912, 1948, 1983, 2030, 2078, 2124, 2166, 2203, 2240, 2276, 2311, 2350, 2394, 2447, 2494, 2539]. CARES [1934]. Carpal [1512]. Carver [2074]. Cascading [1349]. Case [265, 665, 694, 842, 875, 1370, 1375, 1448, 1666, 1740, 1809, 1810, 1859, 2105, 2222, 2257]. Cases [364, 1114]. Category [1862]. Causability [698]. Cause [2047]. Causes [846, 2046]. Cautionary [165]. Celebrating [1418]. Cellular [2408]. Center [451, 994, 1351, 1362, 1423, 1442, 1458, 1492, 1526, 1558, 1599, 1629, 1681, 1716, 1750, 1790, 1831, 1871, 1912, 1948, 1983, 2030, 2078, 2124, 2166, 2203, 2240, 2276, 2311, 2350, 2394, 2447, 2494, 2539]. Centered [997, 1403, 1486, 1735]. Centering [297]. Centers [1662, 1901, 2025, 2120]. Central [569, 881, 1710]. Centralized [451, 1621]. Centric [135, 565, 1378, 1891, 1905, 1960]. Century [1505, 1541]. Certifiability [1963]. Certification [2489]. Certifications [1144]. Certifying [1960, 1962]. CFP [930, 968, 1059]. CGA [1007, 1074]. Chain [177, 604, 639, 1067, 1102, 1103, 1441]. Chains [15, 1101, 1998]. Chair [820, 979, 1015, 1049, 1123, 1158, 1198]. Challenge [172, 386, 545, 889, 1475]. Challenges [71, 199, 204, 253, 269, 309, 670, 730, 731, 811, 843, 905, 970, 1056, 1222, 1290, 1332, 1373, 1419, 1513, 1517, 1676, 1707, 1737, 1742, 1850, 1887, 1944, 1995, 1996, 2015, 2140, 2223, 2232, 2330, 2333, 2368, 2375, 2416]. Challenging [998, 2120]. Chang [1158]. Change [291, 396, 1206, 1522]. Changes [368, 1974, 2490]. Changing [46, 636, 884, 1139, 1339, 1609, 1862, 2156]. Channel [230]. Chaos [2532]. Charge [2084]. Charging [1814]. Charting [1712]. Chat [1373]. ChatAI [1516]. Chatbots [1962]. ChatGPT [1373, 1385, 1456, 1624, 1740, 2229, 2373, 2481, 2519]. Cheating [701].

Check [305]. Check-the-Box [305]. Checking [663]. Chemistry [398]. Chess [1956]. Chicken [1029]. Chief [1563, 1601, 1675, 1678, 1684]. Children [41]. Chiller [696]. Chilling [1635]. China [1420, 1854]. Chip [1104, 1420, 1428, 1576]. Chiplet [1661]. Chips [415, 1340, 1591, 1786, 2427]. Choice [2127, 2169, 2206, 2243, 2279]. Choreography [2373]. Circular [1178, 1180, 1185, 1188, 1232, 2236]. CISE [940, 1005, 1070]. Citations [261]. Cities [545, 564, 662, 834, 1178, 1180, 1183, 1185, 1188, 1774]. Citizen [2306, 2405]. City [1184, 1187, 1448, 2101, 2290]. City-Scale [2101]. Civil [2223]. Claims [2479]. Clarity [2532]. Class [815, 2220]. Classical [1778]. Classification [230]. Classroom [1859]. Classrooms [1140, 1934]. Climate [1522, 1909]. Clinical [1480, 2008, 2431]. Closer [735]. Closing [1045, 2507]. Cloud [332, 534, 707, 731, 1705, 1867, 1887, 2265, 2423, 2515]. Cloud-Native [731]. Clouds [1506]. Clustering [976]. CMMI [766]. Co [1382, 1577]. Co-Design [1382, 1577]. Coaching [40, 41]. Coast [2193]. Code [408, 974, 1545, 1697, 1853, 2213, 2217, 2219– 2221, 2306, 2384, 2532]. Code/No [2213, 2217, 2219, 2221, 2306]. Coding [50, 702]. Cognition [1347]. Cognitive [1347, 1829, 2308]. Cognizant [2514]. Coin [848]. Collaboration [564, 1039, 1745]. Collaborative [227, 873]. Collapsing [210, 713, 886, 1081, 1911]. Collection [2105]. Collective [840]. College [21]. Colors [741]. Column [675]. Columns [2066]. Combinatorial [1817, 1884, 1970, 2258]. Come [2268]. Comeback [2483]. Commandments [607]. Commerce [510, 1715]. Commercial [848]. Commercialize [1806]. Commitment [1206, 2424]. Committee [601, 1049, 1275]. Common [634, 1905]. Communicating [2484]. Communication [1447, 1814].

Communities [440, 455, 998, 1413].

Community

[515, 820, 979, 1015, 1049, 1123, 1158, 1198]. Compacting [1936]. Company [2227]. Comparative [730]. Compare [2019]. Comparison [474, 1763, 1845]. Compartmentalization [1886]. Compartmentalization-by-Design [1886]. Competence [1771, 2022]. Competency [1040]. Competency-Based [1040]. Compiler [2342]. Complex [334, 409, 694, 1770, 2295]. Complexities [2317]. Complexity [98, 100, 626]. Compliance [305, 311, 341, 381, 2349]. Compliant [1961]. Compliting [620]. Composition [311, 702]. Compositional [768]. Comprehension [694, 2469]. Comprehensive [685, 2181, 2416]. COMPSAC [2, 48, 79, 763, 2288, 2450, 2497]. Computation [1043]. Computational [18, 70, 363, 512, 1140, 1306]. Computations [236, 806, 2434]. Compute [1384, 2232, 2515]. Computer [28, 58, 59, 76, 86, 99, 120, 121, 123, 129, 133, 150, 151, 169, 182, 185, 213, 215, 226, 244, 245, 247, 248, 262, 267, 270, 279, 280, 288, 300, 304, 316, 319, 322, 342, 348, 349, 351, 359, 361, 365, 373, 382, 388– 390, 397, 404, 406, 410, 426, 427, 450, 452, 460, 461, 473, 477, 480, 491, 546, 570, 588, 628, 646, 647, 664, 671, 677, 679, 680, 706, 708, 714, 722, 736, 738, 740, 747, 748, 760, 765, 783]. Computer [392, 429, 463, 492, 519, 520, 522, 526, 550, 551, 553, 559, 575, 576, 578, 598, 611, 612, 614, 624, 631, 667, 682, 717, 750, 786, 796, 799, 810, 818, 821, 824, 835, 839, 856, 859, 869, 871, 890, 904, 921, 951, 981, 1002, 1016, 1019, 1028, 1032, 1047, 1050, 1087, 1090, 1094, 1098, 1111, 1116, 1121, 1125, 1128, 1132, 1138, 1155, 1160, 1163, 1171, 1172, 1195, 1196, 1200, 1203, 1229, 1239, 1417, 1432, 1434, 1438, 1457, 1458, 1464, 1466, 1469, 1575, 1631]. Computer [1212, 1242, 1245, 1249, 1255, 1276, 1279, 1292, 1311, 1312, 1314, 1323, 1325, 1343, 1350, 1358, 1361, 1362, 1390, 1402, 1406, 1412, 1415, 1422, 1423, 1473, 1488, 1490– 1492, 1494, 1498, 1515, 1518, 1521, 1525, 1526, 1528, 1532, 1536, 1558, 1560, 1561, 1567, 1573, 1595, 1598, 1601, 1608, 1611, 1616, 1632, 1634, 1636, 1638, 1644, 1650, 1805, 1858, 1878, 1883, 1890, 1906, 1907, 1931, 1933, 1935, 1940, 1945, 1946, 1958, 1964, 1968, 1973, 1980, 1982, 1988, 1993, 2002, 2018, 2085, 2123]. Computer [1655, 1669, 1690, 1694, 1708, 1719, 1728, 1731, 1734, 1753, 1766, 1769, 1785, 1787, 1790, 1793, 1797, 1821, 1826, 1828, 1838, 1843, 1844, 1861, 1942, 2026, 2030, 2041, 2042, 2044, 2045, 2048, 2049, 2078, 2089, 2098, 2111, 2119, 2132, 2136, 2138, 2146, 2176, 2190, 2199, 2201, 2203, 2210, 2214, 2218, 2228, 2237, 2247, 2252, 2253, 2266, 2274, 2283, 2291, 2293, 2307, 2309, 2345, 2357, 2363, 2376, 2378, 2381, 2402, 2407, 2409, 2428, 2442, 2455, 2457, 2464, 2472, 2476, 2488, 2501, 2511, 2531, 2538]. Computer [5, 11, 22, 29, 35, 46, 64, 92, 93, 103, 153, 156, 188, 214, 217, 220, 250, 258, 285, 313, 325, 355, 395, 400, 434, 467, 497, 527, 556, 583, 622, 649, 652, 721, 755, 791, 806, 827, 833, 862, 897, 902, 928, 958, 989, 1025, 1057, 1093, 1095, 1131, 1168, 1176, 1208, 1250, 1264, 1282, 1297, 1300, 1302, 1317, 1356, 1369, 1396, 1447, 1451, 1475, 1497, 1574, 1576, 1709, 1735, 2020, 2084, 2117, 2130, 2144, 2164, 2195, 2202, 2319, 2325, 2328, 2334, 2338, 2346, 2379, 2385, 2419, 2487]. Computer [1429, 1465, 1499, 1533, 1563, 1568, 1605, 1643, 1687, 1722, 1756, 1796, 1837, 1877, 1919, 1954, 1987, 2036, 2037, 2088, 2131, 2172, 2209, 2246, 2282, 2318, 2324, 2358, 2401, 2446, 2454, 2506]. Computer-Enhanced [1447]. Computerized [908]. Computers [81, 119, 180, 194, 314, 433, 642, 644, 779, 939, 1037, 1141, 1182, 1217, 1329, 1383, 1622, 1652, 1673, 1857, 1869, 1896, 1904, 1999, 2118, 2158, 2182, 2273, 2359, 2421, 2478, 2528]. Computing [38, 96, 102, 141, 159, 192, 222, 287, 357, 436, 500, 511, 544, 558, 567, 600, 669, 688, 710, 757, 759, 790, 829, 837, 868, 899, 950, 960, 1027, 1042, 1097, 1170, 1194, 1199, 1214, 1223, 1231, 1237, 1252, 1305, 1319, 1389, 1398, 1401, 1418, 1431, 1455, 1468, 1501, 1509, 1523, 1535, 1570, 1580, 1607, 1646, 1689, 1725, 1746, 1751, 1759, 1791, 1799, 1832, 1840, 1872, 1880, 1913, 1921, 1949, 1956, 1984, 1990, 2039, 2079, 2091, 2125, 2134, 2175, 2212, 2213, 2249, 2285, 2321, 2361, 2404, 2458, 2503]. Computing [7, 9, 19, 39, 54, 93, 103, 112, 135, 149, 163, 167, 173, 205, 229, 242, 257, 265, 275, 295, 298, 308, 376, 380, 399, 441, 454, 470, 505, 531, 533, 548, 569, 574, 594, 622, 623, 652, 676, 703, 753, 971, 994, 1004, 1056, 1071, 1107, 1145, 1181, 1213, 1253, 1290, 1342, 1346, 1347, 1376, 1504, 1531, 1540, 1544, 1659, 1680, 1691, 1695, 1705, 1761, 1775, 1778, 1888, 1926, 1934, 1996, 2031, 2097, 2120, 2148, 2167, 2204, 2241, 2277, 2292, 2312, 2337, 2343, 2351, 2395, 2430, 2448, 2485, 2495, 2540]. Computing [416, 422, 445, 457, 484, 535, 610, 638, 659, 700, 734, 742, 754, 769, 781, 802, 816, 832, 845, 851, 854, 876, 887, 916, 936, 1109, 1124, 1134, 1153, 1167, 1175, 1193, 1270, 1295, 1322, 1327, 1338, 1374, 1381, 1386, 1452, 1484, 1549, 1555, 1557, 1586, 1592, 1597, 1612, 1619, 1628, 1654, 1667, 1671, 1703, 1714, 1723, 1743, 1757, 1782, 1784, 1803, 1811, 1818, 1825, 1851, 1863, 1894, 1910, 1997, 2006, 2016, 2021, 2053, 2054, 2069, 2070, 2086, 2100, 2113, 2141, 2192, 2233, 2390]. Computing [2154, 2179, 2188, 2239, 2256, 2299, 2365, 2369, 2411, 2415, 2439, 2463, 2470, 2475, 2525, 2533, 2536]. ComputingEdge [320]. Concatenated [1296]. Concept [1898]. Concepts [2509]. Conceptual [1062, 1256]. Concerns [702, 997, 1552]. Concrete [2322]. Cone [1079]. Conference [97, 1243, 1287, 1341, 1353]. Confidence [632]. Configurable [877]. Configuration [877]. Configure [1809]. Conflict [2183]. Confronting [501]. Congratulations [2085]. Congress [864]. Connected [2099]. Connecting [354, 1589]. Connection [27, 57, 83, 118, 148, 181, 212, 317, 346, 424, 459, 573, 609, 1310, 2075, 2310, 2492]. ConnectionRoom [805]. Conquering [396]. Cons [2216]. Consensus [878, 1579]. Consequences [846]. Considerations [881, 1166, 1885]. Consistent [1579]. Constraints [606]. Construction [1898]. Constructs [2435]. Consumers [1370]. Consumption [807, 877]. Contact [842]. Contactless [206]. Container [1544]. Container-Based [1544]. Contemporary [1110, 1779, 1889]. Content [935, 1296, 1516, 1862]. Content-Rich [1296]. Contents [3, 32, 62, 90, 124, 154, 186, 218, 251, 283, 323, 352, 393, 430, 464, 495, 523, 554, 579, 615, 650, 683, 718, 751, 787, 825, 860, 894, 925, 955, 970, 985, 1020, 1054, 1091, 1129, 1164, 1204, 1246, 1280, 1315, 1354, 1394, 1426, 1461, 1495, 1529, 1564, 1602, 1639, 1685, 1720, 1754, 1794, 1835, 1875, 1916, 1952, 1985, 2032, 2033, 2082, 2128, 2170, 2207, 2244, 2280, 2315, 2354, 2398, 2451, 2498]. Contested [2271]. Context [1588, 1697]. Contexts [2490]. Continually [2027]. Continuing [1093]. Continuous [381, 627, 1033, 1143]. Continuum [332, 534, 1107, 1145, 1544, 1981, 2232]. Contracts [18, 771, 2419]. Contrasting [2380]. Contributions [766, 1023, 1853]. Control [231, 236, 819, 884, 1635, 2393, 2414]. Controller [368, 880]. Convenience [2529]. Converged [1680]. Convergence [994, 1294, 1707, 1867, 1939, 2163]. Conversational [383, 636, 905, 1083, 1930]. Cooker [1008]. Cooperation [775]. Cooperative [363, 2507]. Copyleft [2438]. Cord [43]. Core [2509]. Corporate [1812, 2061]. CORRECT [1775]. Correction [56, 318]. Correctness [855, 1898, 2017]. Correctness-by-Construction [1898]. Correlation [475]. Corruptions [2343]. Cost [1546, 2267, 2529, 2534]. Cost-Benefit [1546]. Council [775]. Count [2229]. Counterfeit [198]. Countermeasures [1441]. Counting [724]. Countries [639, 775, 881]. Country [1672]. Course [1775, 2342, 2434]. Court [563]. Courts [471]. Cover [1, 30, 60, 88, 122, 152, 184, 216, 249, 281, 321, 350, 391, 428, 462, 493, 521, 552, 577, 613, 648, 681, 716, 749, 785, 823, 858, 892, 923, 953, 983, 1018, 1052, 1089, 1127, 1162, 1202, 1244, 1278, 1313, 1352, 1392, 1424, 1459, 1493, 1527, 1562, 1600, 1637, 1683, 1718, 1752, 1792, 1833, 1873, 1914, 1950, 2028, 2076, 2080, 2126, 2168, 2205, 2242, 2278, 2313, 2352, 2396, 2449, 2496]. Coverage [16, 633, 1884, 2254]. Covered [1402, 1494, 1536, 1769, 1993, 2041, 2042]. COVID [208, 223, 313, 315, 385, 485, 1067, 1291, 2000]. COVID-19 [208, 223, 313, 315, 385, 485, 1067, 1291, 2000]. COVIDWISE [842]. Coworker [2286]. Crawling [2519]. Create [1672, 2263]. Creative [1959, 2340]. Credit [1364]. Crimes [910]. Criminal [563]. Crisis [2483]. Critical [21, 71, 198, 336, 690, 1365, 1517, 2362, 2392]. Critiques [108]. Cross [474, 2271]. Cross-Entropy [2271]. Cross-Platform [474]. Crossroads [1925, 1959]. Crowd [505]. Crowd-Machine [505]. Crowds [1149]. Crowdsourced [108]. Crowdsourcing [20, 732, 1923]. Crucible [1008]. Crying [1099]. Cryptanalysis [1552]. Crypto [2483]. Cryptocurrencies [13]. Cryptocurrency [412, 2530]. Cryptographic [405, 978, 2153]. Cryptojacking [794]. CS [27, 57, 83, 87, 118, 126, 148, 181, 212, 317, 346, 424, 459, 573, 609, 618, 755, 893, 924, 930, 954, 968, 996, 1005, 1007, 1010, 1013, 1059, 1137, 1310, 2075, 2310, 2492, 2500]. CSAB [1899]. CUF [1033]. CUF-Links [1033]. Cui [1702]. Cultural [1522]. Culture [713]. Currencies [881, 1710]. Currency [569, 2294]. Current [209, 730, 819, 970, 1220, 1710, 2535]. Curve [913]. CVPR [282]. Cyber [456, 643, 819, 1060, 1233, 1334, 1349, 1908, 1932, 2007, 2302, 2303]. Cyber-Physical [1908]. Cyber-Physical-Social [1334]. Cyberarms [1487]. Cyberattack [1183]. Cyberattacks [274, 2535]. Cyberbiosecurity [299]. Cyberconflict [1979]. Cybercrime [1556, 1587]. Cyberdefense [704]. Cyberdidacticism [2308]. Cyberinfrastructure [1073]. Cyberpandemics [161]. Cyberpanics [1922]. Cyberphysical [24, 69, 71, 74, 75, 136, 204, 264, 265, 268, 409, 415, 565, 602, 657, 658, 660, 663, 672, 768, 778, 855, 866, 945, 1117, 1154, 1268, 1288, 1747, 2393]. Cybersecurity [130, 183, 272, 485, 572, 732, 773, 846, 884, 940, 1099, 1151, 1273, 1413, 1510, 1635, 1657, 1711, 1824, 1850, 1891, 2122, 2161, 2386, 2392, 2489]. Cyberspace [1240]. Cyberthreat [1905]. Cyberthreats [164]. Cybertrust [276]. Cyberwarfare [1006].

D [105]. D&I [1528, 2002]. Daily [1512]. Dan [84]. Dapper [67]. Dark [1145, 2343]. Darknet [2257]. Darwin [939]. Darwin-S [939]. Dashboards [1928]. Data [25, 26, 52, 69, 70, 97, 111, 116, 135, 146, 178, 200, 238, 293, 315, 378, 386, 417, 420, 453, 475, 481, 513, 605, 627, 645, 711, 820, 843, 853, 881, 903, 918, 937, 969, 973, 976, 994, 1004, 1037, 1038, 1112, 1145, 1187–1189, 1262, 1269, 1288, 1334, 1339, 1378, 1409, 1416, 1442, 1443, 1480, 1544, 1546, 1556, 1566, 1621, 1630, 1662, 1732, 1779, 1867, 1887, 1901, 1938, 1957, 2025, 2105, 2108, 2120, 2140, 2142, 2181, 2187, 2200, 2215, 2221, 2236, 2264, 2272, 2321, 2329, 2335, 2343, 2400, 2408, 2410, 2431, 2520, 2529]. Data [490, 530, 776, 814, 849, 885, 1113, 1148, 1177, 1219, 1331, 1377, 1594, 1615, 1663, 1789, 1830, 1860, 2004, 2056, 2057, 2102, 2186, 2235, 2327, 2367, 2426, 2480, 2521]. Data-Centric [135, 1378]. Data-Driven [1112, 1334, 1938]. Data-Oriented [937]. Database [481, 1297]. Databases [508, 2453]. Dataport [977, 984, 2491]. Dataset [2254]. DBLP [1297]. dd [1296]. DDS [2181].

Dead [1304]. Deaf [1447]. Deb [1198]. Debug [34]. Debugging [475]. Decade [2255]. Decentralized [367, 1107, 1419, 1885]. Deception [1670, 2356]. Deciphering [1925]. Decision [694, 1845, 1885, 2431, 2477, 2510]. Decision-Making [1845]. Decisions [872]. Decline [1550]. Decoding [2055]. Decryption [1230]. Deep [260, 905, 938, 1260, 1293, 1372, 1478, 1664, 1806]. Deep-Learning-Based [1372]. Deep-Vision [2294]. Deepfake [1433, 1870]. Deepfakes [1367, 1485]. DeepSeek [2300, 2481, 2519]. Defend [135]. Defendants [1881]. Defending [563]. Defense [634]. Defenses [1765, 2051]. Defensive [673]. Defined [590, 2289, 2290, 2422]. Defining [1036]. Definition [709, 1625]. Deforming [720]. Degrees [1513]. Delegate [246, 678, 1048, 1489, 1947, 2444]. Delegate/Director [246, 678, 1048, 1489, 1947, 2444]. Delivering [237]. Delivery [334]. Delusion [1240]. Democratizing [1073, 1585, 2200, 2420]. Dena’ina [2465]. Denial [132, 209, 241]. Denial-of [209]. Denial-of-Service [132, 241]. Deny [1897]. Dependability [672, 1079, 1759, 1760, 1767, 1768]. Dependable [1301, 1761, 1767]. Dependencies [1545]. Dependency [51, 487]. Depends [1767]. Deploying [963]. Deployment [294]. Deployments [2101]. Deprivation [43]. Depth [512]. Derivatives [108]. Description [709]. Deserts [1508]. Deserves [2149]. Design [204, 412, 417, 601, 643, 727, 730, 844, 870, 872, 937, 1334, 1382, 1439, 1576, 1577, 1666, 1698, 1886, 1898, 2202, 2427, 2489]. Designing [228, 1486, 1998]. Desktop [1142]. Detect [398, 976]. Detecting [1262, 1487, 2185]. Detection [71, 139, 372, 386, 695, 696, 1183, 1408, 1545, 1815, 2067, 2099, 2258]. Develop [50, 1420]. Developer [2373]. Developers [34, 1885]. Developing [110, 589, 639, 963, 1996, 2157]. Development [168, 771, 912, 937, 972, 1003, 1063, 1079, 1378, 1585, 1745, 1909, 1972, 2061, 2096, 2197, 2219, 2221, 2298, 2306, 2331, 2361, 2364, 2366, 2368, 2370, 2373, 2471, 2509]. Developments [1706, 2195]. Deviations [1262]. Devices [138, 547, 730, 746, 1022, 1478, 1512, 1519, 1961]. DeViTC [2294]. DevOps [1144]. DevSecOps [1981]. Diabetes [1733]. Diabetic [302]. Diagnosis [660, 2431]. Diagnostics [834]. Dialogue [42]. Dialogue-Based [42]. Dialogues [2373]. Dictator [1240]. Did [675, 1081, 2504]. Diets [20]. Digital [40, 45, 240, 272, 297, 411, 501, 518, 545, 569, 622, 644, 732, 795, 809, 840, 881, 915, 970, 998, 1001, 1062, 1063, 1108, 1216, 1238, 1344, 1414, 1418, 1419, 1439, 1448, 1522, 1618, 1710, 1745, 1773, 1808, 1810, 1812, 1813, 1852, 1891, 1972, 2025, 2215, 2344, 2393, 2474, 2522]. Digitization [1328]. Dilemma [1240, 1508, 2005]. Dilemmas [629]. Dimensions [805]. Direct [1442]. Directing [773, 1712, 2189]. Directions [1737, 1850, 2259]. Director [246, 678, 1048, 1489, 1947, 2444]. Directory [2185]. Disaggregation [1077]. Disappearing [1399]. Disaster [2410]. Disasters [1346]. Discourses [337]. Discovering [836, 2373]. Discovery [253, 1342, 2197, 2427, 2507]. Discriminatory [335]. Discussion [1729, 1923]. Diseases [2417]. Disinfecting [1827]. Disinformation [605, 780, 1613, 1922, 2110, 2225, 2479]. Disorder [41]. Disparate [481]. Displays [354]. Disproportionately [2058]. Disruption [166, 1829, 1959]. Disruptive [1614]. Dissertation [1801]. Distancing [192]. Distillation [2509]. Distraction [1014]. Distributed [69, 70, 73, 209, 241, 326, 442, 581, 903, 927, 1543, 1737, 1778, 1845, 2260]. Distribution [1781, 2181]. Distributor [777]. Diverging [589]. Diverse [12, 2068]. Diversity [738, 810, 869, 911, 918, 933, 1010, 1019, 1155, 1365, 1518, 1634, 1638, 1719, 1753, 1793, 1858, 1907, 1940, 2084]. Divide [998]. Dividend [2534]. Division [246, 678, 1048, 1489, 1947, 2444]. DNA [111, 1443]. Do [375, 413, 453, 529, 699, 949, 1388, 1571, 1604, 2227, 2268, 2474]. Dobberpuhl [84]. Documentation [1823]. Documenting [1363]. Documents [45, 1326, 2469]. Doing [1413]. Dollar [1103]. Domain [1062, 1257]. Done [201]. Donnybrook [703]. Don’t [539]. Doomed [1043]. Doris [2074]. Down [167, 931]. Downturn [1554]. Doxxing [1307]. Dragging [2047]. Drama [202]. Drawing [1895]. Drawings [1328]. Dream [2387]. Drive [1689, 1705]. Driven [74, 873, 1112, 1334, 1421, 1513, 1938, 2061]. Drivers [296, 1745]. Drives [1520]. Driving [106, 395, 545, 626, 918, 967, 1063, 1257, 1813, 2047, 2344]. Drone [2101]. Drones [369, 1248]. DROO [753]. Drugs [198]. DSP [2292]. Dual [1079]. Dual-Cone [1079]. Dumbing [931]. During [315, 663, 1291, 2220]. Dusting [1470]. Dyad [2467]. Dynamic [139, 364, 627, 2414]. Dysfunction [1829].

E-Coaching [40]. E-Commerce [510, 1715]. E-Health-Care [201]. e-Learning [1139]. E-Mails [1975]. e-Personhood [2215]. E-Waste [2177]. Early [1428]. Earth [2515]. Easier [201]. Eats [471]. Eckert [406]. Eclipse [571]. eCloud [534]. Ecological [2467]. Ecology [208]. Economic [1867]. Economics [274, 377, 412, 518, 569, 639, 775, 1118, 1420, 1481, 1485, 1633, 1776, 1943, 2160]. Economies [589, 908]. Economy [1232, 2236]. Ecosystem [374, 1810, 2015]. Ecosystems [1510, 2014]. Edge [93, 135, 332, 415, 534, 753, 1069, 1071, 1181, 1577, 1737, 1751, 1791, 1832, 1872, 1887, 1913, 1949, 1984, 2031, 2079, 2125, 2167, 2204, 2241, 2277, 2312, 2351, 2395, 2448, 2495, 2540]. Edge-to-Cloud [332]. EDHOC [920, 1971]. Editor [1563, 1601, 1675, 1678, 1684, 1918]. Editors [692]. Educate [637]. Educating [1306, 1627, 1806]. Education [103, 206, 313, 449, 468, 514, 568, 843, 908, 945, 1139, 1140, 1176, 1268, 1380, 1411, 1451, 1574, 1576, 1651, 1693, 1745, 1911, 1934, 1977, 2112, 2144, 2221, 2222, 2379]. Educational [517, 815, 1578, 2010, 2157, 2414]. Educator [1451]. Educators [1306]. Effective [344, 1633, 1887]. Effectively [633, 2161]. Effects [1349, 1428]. Efficacy [2424]. Efficiency [2267]. Efficient [504, 733, 902, 1290, 2400]. Efforts [2347, 2441]. EIC [2111]. EIC’S [560]. Eight [1745]. Elderly [998, 1733, 1845, 1846]. Elect [388, 1196]. Election [245, 262, 280, 614, 649, 677, 1002, 1032, 1047, 1488, 2019]. Electric [1814, 2535]. Electrical [1062, 1079]. Electrical/Electronic [1079]. Electricity [138, 478]. Electrocardiogram [1506]. Electronic [622, 1079, 1418, 1543, 2197]. Electronics [1022]. Element [103]. Elementary [1140]. Elements [508]. Elephant [140, 487]. Elliptic [913]. Else [2524]. Elsewhere [126, 618, 755, 2500]. Embedded [73, 340, 409, 838, 902, 1301, 1519, 2292]. Embedding [1545]. Embench [2292]. Embodied [593]. Embracing [256, 1513, 2386, 2434]. Emergence [1093]. Emerging [652, 1056, 1439, 1531, 1614, 1699, 1744, 1885, 2259, 2353, 2397, 2473]. Emotion [1852]. Empire [1302]. Employing [1848]. Employment [2003]. Empowering [2084]. Enable [504, 1296, 1384]. Enabled [268, 1183, 1517, 1666]. Enabler [1660]. Enables [1232, 1862]. Enabling [581, 743, 1104, 1546, 1658, 2427]. Encrypting [1230, 1705]. Energy [134, 268, 269, 504, 807, 842, 877, 1166, 1290, 1405, 1537, 1901, 2120, 2163, 2250, 2260]. Energy-Efficient [504, 1290]. Engagement [2130]. Engages [1368]. Engaging [168, 1609, 2112]. Engineering [237, 375, 479, 511, 514, 542, 544, 549, 567, 571, 600, 669, 710, 759, 770, 790, 820, 837, 867, 868, 870, 950, 1042, 1076, 1122, 1152, 1176, 1236, 1268, 1289, 1305, 1321, 1324, 1389, 1421, 1455, 1523, 1555, 1574, 1576, 1580, 1593, 1622, 1652, 1670, 1746, 1782, 1811, 1868, 1869, 1904, 1981, 1996, 2016, 2118, 2144, 2148, 2262, 2298, 2372, 2375, 2392, 2430, 2434, 2485]. Engineers [549]. Engines [2332]. Enhance [1348]. Enhanced [1447, 1624, 2349]. Enhances [2187]. Enhancing [572, 1566, 1774, 2267]. Enough [1320, 2516, 2526]. Ensembling [1333]. Enterprise [166, 1905, 2198, 2370]. Enterprise-Centric [1905]. Entity [1813]. Entrepreneurship [1804, 1808]. Entropy [1679, 2271]. Environment [568, 663, 2271, 2290]. Environmental [1041]. Environments [334, 1962, 2097, 2414]. Envisioning [2408]. Epidemic [1038, 2417]. Epidemic-Handling [1038]. Epistemic [2308]. Epistemology [1385]. Equifax [145]. Equipment [2535]. Equitable [844]. Equity [592, 918]. Era [415, 1531, 1617, 1812, 2379, 2389, 2471, 2515]. Erasure [2142]. Erich [820]. Erratum [345, 423]. Error [1428]. Errors [693, 2268]. Essence [1713]. Establishing [2236]. Estimation [139, 1763]. Estimations [1428]. Estonia [1810]. Ethereum [1103]. Ethernet [1442]. Ethical [607, 805, 906, 1222, 1410, 1439, 1698, 1699, 1959, 2264, 2272]. Ethics [419, 629, 764, 1439, 1695, 1697, 1938, 2140, 2468]. EU [2015]. Europe [1188, 1808, 2015]. European [337, 1363, 1961, 2380]. Evaluating [1367]. Evaluation [1330, 2157, 2440]. Event [976]. Eventually [453]. Everyone [1966]. Everything [838, 888, 999]. Evidence [2161]. Evidence-Based [2161]. Evolution [534, 915, 1475, 1507, 1967, 2306, 2412, 2483, 2510]. Evolutionary [2356]. Evolving [731, 1705, 2304]. Examining [2003]. Example [1664]. Examples [2388]. Exchange [920]. Executable [108]. Exist [1571]. ExoMars [697]. Expect [1548, 1590]. Expectation [730]. Expectations [337]. Experience [1265, 2434, 2462, 2468]. Experiences [1624]. Experiencing [591]. Experts [1410, 1651, 1693]. Explainability [698, 852, 1897]. Explainable [227, 276, 692, 693, 695, 696, 1330, 1815, 2431]. Explaining [1261]. Explanations [694]. Exploiting [24]. Exploration [743]. Exploring [232, 1713, 1765, 2384, 2509]. Exposed [2177]. Exposures [1905]. Express [1661]. Expressively [896]. Expressiveness [801]. Extend [1744]. Extended [1959]. Extensible [537]. External [1911]. Eye [1893].

Fabrication [761]. Face [1367]. Facial [1222]. Facilitators [1220]. Facing [1911]. Factoring [1023]. Factors [998, 1332]. Factory [1483]. Faded [2230]. Failed [1387]. Failure [1039]. Failures [2039, 2043, 2046, 2295]. FAIRness [808, 1033, 1333]. Fake [1991]. Fall [1249, 1408]. Falsification [761, 1288]. FAQ [949]. Farm [1559]. Farming [1234, 2267]. Farms [2099]. Fast [941]. Fatal [1516]. Father [1236]. Fault [73, 537, 660, 696, 1579, 1889]. Fault-Detection [696]. Fault-Tolerant [73, 1579]. Faults [409]. Fear [1119]. Feature [1660]. Federal [563]. Federated [481, 903, 1183, 1506, 1621, 1765, 2289]. Federation [1697]. Fellow [338]. Female [2184]. Feminist [1666]. Femwork [1666]. Feudalism [2215]. Few [1114, 1264]. Field [1888, 2466]. Field-Programmable [1888]. Fifth [1011]. Fight [1189]. Fighting [1067]. File [1520]. filler [936, 942, 944, 948, 950, 952, 954, 982, 996, 1000, 1005, 1007, 1012, 1017, 1059, 1064, 1066, 1068, 1070, 1074, 1078, 1080, 1082, 1084, 1086, 1088, 1137, 1915, 1933, 1935, 1937, 1940, 1942, 1946, 1948, 1949]. Film [1596, 2462]. Filmmaking [2263]. Finally [2149]. Finance [694, 808, 1909, 2344]. Financial [13, 1856]. Find [697, 1706]. Finding [2388]. Findings [1188]. Fingerprints [45]. Finnish [965]. Fintech [1909]. Firms [113, 2015]. First [806, 1651, 1775, 2286]. Fitness [2520]. Fix [2060, 2384]. Fixing [409]. Flaky [875]. Flavored [1786]. Flavors [712, 778]. Flaws [1516]. FLB [2289]. Flexibility [24]. Flexible [446, 2290]. Flight [2271]. Flipping [506]. Flow [2159]. Fluidified [1405]. Fly [45]. Flying [38]. Focused [451]. Fog [43, 265]. Fog-Assisted [43]. Food [377, 471]. Footprint [993]. Footprinting [1135]. Footprints [1041]. Forecasting [134, 2510]. Forefront [652, 2296]. Foreign [1513]. Forensic [840]. Forgetting [566]. Forgotten [2230]. Formal [144, 405, 657, 658, 662, 665, 733, 2020, 2112, 2419]. Formalizing [2479]. Format [1463]. Forming [1664]. Forrest [395]. Forward [278, 630, 797, 957, 1235, 2145]. Foster [2264]. Fostering [1808]. Found [585]. Foundation [1788]. Foundations [2202]. Four [974, 2109]. FPGA [446]. FPGA-Based [446]. Framework [43, 70, 504, 840, 967, 1289, 1334, 1364, 1403, 1735, 1736, 1934, 1965, 2317, 2372]. Frameworking [1376]. Frameworks [132, 474, 572, 1040]. Frances [347]. Fraud [695]. Frauds [910]. Fred [1450]. Free [311, 455, 1445, 1941]. Freelancer [1289]. Freelancer-Supported [1289]. Frequency [2427]. Friendly [1846]. Front [1, 30, 60, 88, 122, 152, 184, 216, 249, 281, 321, 350, 391, 428, 462, 493, 521, 552, 577, 613, 648, 681, 716, 749, 785, 823, 858, 892, 923, 953, 983, 1018, 1052, 1089, 1127, 1162, 1202, 1244, 1278, 1313, 1352, 1392, 1424, 1459, 1493, 1527, 1562, 1600, 1637, 1683, 1718, 1752, 1792, 1833, 1873, 1914, 1950, 2028, 2076, 2080, 2126, 2168, 2205, 2242, 2278, 2313, 2352, 2396, 2449, 2496]. Frontier [1979, 2142, 2200, 2508]. Frontiers [1926]. Fully [2189]. Function [1069, 1071]. Functional [489, 1275, 1370, 2275, 2441]. Fund [738, 810, 869, 1019, 1155, 1528, 1638, 1719, 1753, 1793, 1858, 1907, 1940, 2002]. Fundamentals [1235]. Fundamentals-Based [1235]. Fusing [2068]. Fusion [969, 1970, 2255, 2258]. Future [15, 449, 513, 549, 590, 591, 644, 673, 730, 746, 801, 998, 1076, 1108, 1181, 1184, 1218, 1436, 1440, 1483, 1552, 1583, 1625, 1630, 1653, 1656, 1713, 1783, 1807, 1812, 1850, 1901, 1925, 1927, 2152, 2259, 2260, 2270, 2273, 2298, 2305, 2344, 2347, 2382, 2412, 2484, 2509]. Fuzzing [672].

G [77]. Gag [1081]. Gain [1081]. Gaining [696]. Gamble [975]. Game [937, 1014, 1060, 1294, 1596, 1862, 1864, 2462, 2465]. Games [782, 815, 934, 949, 1156, 1672]. Gaming [2014, 2149, 2155]. Gap [1045, 1819, 2392]. Gaps [2423, 2469]. Gate [1888]. Gem [10]. Gemini [2519]. Gen [2155]. GenAI [2434]. Gender [846, 1848]. General [26, 944, 1008, 1078, 1084, 1086, 1554, 2333]. Generals [1302]. Generate [1903, 2340]. Generated [2067]. Generation [432, 476, 743, 875, 1293, 1658, 1962, 2408]. Generative [1120, 1380, 1411, 1481, 1486, 1633, 1653, 1679, 1715, 1749, 1822, 1827, 1856, 1859, 1962, 1974, 1975, 1977, 2015, 2058, 2096, 2258, 2263, 2270, 2306, 2308, 2364, 2366, 2368, 2372, 2389, 2391, 2487, 2508, 2527, 2534]. Generator [1516]. Generic [382, 1082, 1678]. Genomic [2272]. Genomics [1944]. GeNose [1370]. George [1207]. Get [866, 949, 2077]. Getting [2149]. Giants [1228, 2272]. Give [2387]. Glance [1407]. Global [889, 1848, 1922, 1979, 2071, 2177, 2530]. Go [699, 2268, 2387]. Goes [1011]. Going [803, 976]. Gold [2272]. Golden [1841].

Gonzo [1445]. Good [236, 398, 1297, 1320]. Governance [18, 515, 1180, 1736, 1776, 1967, 2012, 2061]. Governing [2429]. Government [292, 301]. Governors [1945, 2442]. GPS [139]. Graduate [773]. Granular [481, 2159]. Graph [1589]. Graphics [99, 133, 169, 267, 288, 361, 404, 452, 473, 546, 570, 628, 671, 706, 736, 835, 1049, 1121, 1229, 1300, 1369, 1417, 1469, 1575, 1631, 1650, 1805, 1906, 1946, 1982, 2018, 2123, 2190, 2228, 2266, 2293, 2345, 2376, 2428, 2472, 2531]. Graphs [2330]. Great [1011]. Greater [1587]. Greatest [2424]. Greedy [2388]. Greek [971]. Green [545, 877, 1401, 1403, 1407, 2336]. Green-IN [1407]. Green500 [2145]. Greener [1630]. Grid [130]. Grids [135, 1075, 2535]. Gropp [388]. Group [1336]. Grow [2022]. Growing [1152]. Grownups [413]. Growth [2267]. GTS [2388]. Guarantees [842]. Guard [2193]. Guest [692]. GUI [875]. Guidance [1581]. Guide [166, 2020]. Guided [232]. Guidelines [2222]. Gulf [775]. Guy [2387].

Hack [145]. Hack3D [732]. Hackathons [543, 1477, 2000, 2184]. Hand [1848]. Handle [2453]. Handling [1038]. Handshake [1971]. Happen [368, 782]. Happening [1699]. Hard [1447]. Harder [2400]. Hardware [309, 466, 705, 800, 1056, 1382, 1577, 1864, 2419, 2435, 2537]. Hardware/Software [1382]. Harms [1699, 1736, 2058]. Harnessing [844, 1630]. Hash [1579]. Hazard [2052]. Health [40, 42, 201, 301, 636, 641, 1107, 1114, 1189, 1508, 1543, 1551, 1849, 2008, 2272, 2391, 2505, 2510, 2520]. Health-Care [42, 1107, 1508]. Healthy [844, 2059, 2335]. Hear [746, 1604]. Hearing [746, 1447]. Helm [395]. Helmets [774]. Help [189, 697, 877]. Helping [34]. Here [271]. Heritage [1522, 2465]. Heterogeneity [1680]. Heterogeneous [204, 504, 1713, 1742, 1996]. Heuristics [1711]. Hiding [970, 1571]. Hierarchical [1545]. High [39, 807, 994, 1194, 1342, 1363, 1680, 1764, 1888, 1926, 1962, 2419, 2427]. High-Assurance [2419]. High-Frequency [2427]. High-Integrity [1764]. High-Performance [39, 994, 1194, 1342, 1680, 1888, 1926]. High-Risk [1363]. High-Stakes [1962]. Higher [206, 1745, 1911, 1977]. Highlight [820, 979, 1015, 1049, 1123, 1158, 1198]. Highlights [5, 35, 64, 92, 156, 188, 220, 258, 285, 325, 355, 400, 434, 467, 497, 527, 556, 583, 653, 686, 721, 755, 791, 827, 862, 897, 928, 958, 989, 1025, 1057, 1095, 1131, 1168, 1208, 1250, 1282, 1317, 1356, 1396, 1429, 1465, 1499, 1533, 1568, 1605, 1643, 1687, 1722, 1756, 1796, 1837, 1877, 1919, 1954, 1987, 2036, 2037, 2088, 2131, 2172, 2209, 2246, 2282, 2318, 2358, 2401, 2454]. Highly [73]. Hire [2155]. History [163, 257, 278, 298, 399, 419, 454, 455, 470, 548, 574, 594, 623, 676, 700, 734, 754, 832, 1124, 1167, 1231, 1270, 1322, 1373, 1381, 1452, 1557, 1592, 1784, 1803, 2021, 2086, 2141, 2192, 2233, 2390, 2417, 2439, 2470, 2533]. Hoax [78]. Holistic [626]. Hollywood [2263]. Home [289, 1929]. Homework [560]. Honeypot [2257]. Honoring [102, 141]. Horizons [1194]. HOST [31, 211]. HPC [1867]. Hubris [1456]. Human [26, 225, 227, 231, 525, 564, 696, 698, 744, 915, 997, 1085, 1114, 1291, 1347, 1403, 1486, 1618, 1701, 1735, 1891, 2349, 2427, 2520]. Human-AI [1347]. Human-Artificial [225]. Human-Centered [997, 1403, 1486, 1735]. Human-Centric [1891]. Human-Intelligent [1085]. Human-Machine [564]. Humanism [1238]. Humanity [89, 992, 1435, 1554, 1924, 2505]. Humanoid [232]. Humans [1339, 1368, 1503, 2143]. Humphrey [964, 988, 1024, 1343, 1371, 1404, 1437, 1476]. Hwu [979]. Hybrid [225, 227, 505, 1934]. Hyesoon [979]. Hype [336, 853, 2387].

Hyperloops [591]. Hyperscale [994, 1442]. Hypnotized [408].

ICCQ [254, 290, 331, 1642]. Ideals [233]. Ideas [2509]. Identities [978]. Identity [240, 1143, 1419, 1891, 2429]. IEEE [28, 58, 59, 61, 76, 86, 87, 97, 109, 120, 121, 123, 129, 131, 150, 182, 183, 185, 203, 213, 226, 244– 247, 262, 270, 279, 280, 300, 304, 319, 342, 348, 349, 359, 365, 382, 387–390, 395, 406, 410, 426, 427, 450, 460, 461, 480, 491, 492, 507, 517, 519, 520, 526, 550, 551, 559, 624, 646, 647, 664, 677, 679, 680, 708, 714, 722, 738, 740, 747, 748, 760, 765, 783, 864, 1243, 1341, 2084, 2288]. IEEE [402, 439, 468, 488, 498, 541, 562, 575, 576, 603, 611, 612, 614, 631, 674, 678, 715, 763, 784, 796, 799, 810, 818, 821, 822, 839, 856, 857, 869, 871, 890, 891, 904, 907, 911, 917, 921, 922, 930, 933, 942, 948, 951, 952, 977, 981, 982, 996, 1002, 1005, 1007, 1010, 1013, 1016, 1017, 1019, 1028, 1032, 1047, 1048, 1050, 1051, 1053, 1059, 1087, 1088, 1090, 1094, 1111, 1116, 1125, 1126, 1128, 1132, 1137, 1155, 1160, 1161, 1163, 1171, 1172, 1195, 1200, 1201]. IEEE [966, 980, 1173, 1203, 1223, 1229, 1231, 1237, 1239, 1242, 1245, 1249, 1267, 1276, 1277, 1279, 1292, 1309, 1311, 1314, 1323, 1325, 1335, 1343, 1350, 1351, 1353, 1358, 1361, 1362, 1366, 1390, 1391, 1393, 1402, 1406, 1412, 1415, 1422, 1423, 1425, 1432, 1434, 1438, 1457, 1458, 1460, 1464, 1466, 1472, 1473, 1488–1492, 1494, 1498, 1509, 1515, 1518, 1525, 1526, 1528, 1532, 1536, 1558, 1560, 1561, 1567, 1573, 1584, 1598, 1599, 1601, 1608, 1611, 1616, 1626, 1629, 1632, 1634, 1636]. IEEE [1214, 1217, 1221, 1638, 1644, 1655, 1665, 1669, 1678, 1681, 1684, 1690, 1694, 1696, 1700, 1708, 1716, 1719, 1728, 1731, 1734, 1741, 1748, 1750, 1753, 1766, 1769, 1785, 1787, 1790, 1793, 1797, 1816, 1821, 1828, 1831, 1838, 1843, 1844, 1858, 1865, 1871, 1878, 1883, 1890, 1900, 1907, 1912, 1931, 1933, 1935, 1940, 1945, 1947, 1948, 1958, 1964, 1968, 1973, 1980, 1983, 1988, 1993, 2002, 2026, 2030, 2041, 2042, 2044, 2045, 2048, 2049, 2062, 2063, 2078, 2085, 2089, 2098, 2111, 2115, 2119, 2121, 2124]. IEEE [1682, 1717, 1866, 1902, 1937, 1978, 2116, 2117, 2127, 2132, 2136, 2138, 2146, 2166, 2169, 2176, 2195, 2199, 2201, 2203, 2206, 2210, 2214, 2218, 2231, 2237, 2240, 2243, 2247, 2252, 2253, 2261, 2274–2276, 2279, 2283, 2291, 2297, 2307, 2309, 2311, 2319, 2325, 2328, 2334, 2338, 2348, 2350, 2353, 2357, 2363, 2371, 2378, 2381, 2383, 2394, 2397, 2402, 2407, 2409, 2413, 2436, 2442– 2444, 2447, 2450, 2455, 2457, 2464, 2473, 2476, 2482, 2488, 2491, 2493, 2494, 2497, 2501, 2511, 2518, 2538, 2539]. IEEE [7, 9, 25, 46, 68, 72, 81, 94, 99, 112, 114, 116, 119, 127, 133, 137, 143, 146, 149, 163, 168, 169, 173, 175, 178, 180, 190, 194, 200, 205, 207, 215, 229, 234, 238, 242, 248, 257, 267, 273, 275, 277, 288, 293, 295, 298, 306, 308, 310, 314, 316, 320, 322, 328, 333, 339, 351, 358, 361, 370, 373, 376, 378– 380, 384, 399, 454, 470, 548, 563, 574, 588, 594, 622, 623, 676, 700, 734, 739, 754, 1232, 1899, 1972, 2130, 2259]. IEEE [392, 397, 404, 414, 416, 418, 420, 422, 429, 433, 443, 445, 448, 452, 457, 463, 473, 477, 482, 484, 486, 490, 503, 509, 522, 530, 532, 535, 538, 544, 546, 553, 567, 570, 578, 582, 586, 597, 598, 600, 608, 610, 628, 635, 638, 640, 642, 659, 661, 667, 669, 671, 682, 691, 706, 710, 717, 736, 742, 745, 750, 759, 767, 769, 776, 779, 781, 790, 798, 832, 833, 835, 837, 841, 868, 874, 950, 1012, 1042, 1044, 1115, 1121, 1124, 1167]. IEEE [786, 802, 812, 814, 816, 824, 845, 847, 849, 851, 854, 859, 876, 882, 885, 887, 893, 916, 919, 924, 936, 1034, 1098, 1106, 1109, 1113, 1134, 1138, 1141, 1148, 1150, 1153, 1159, 1175, 1177, 1179, 1182, 1193, 1197, 1199, 1212, 1219, 1255, 1270, 1272, 1295, 1305, 1312, 1322, 1338, 1345, 1379, 1381, 1386, 1389, 1417, 1449, 1452, 1455, 1469, 1482, 1484, 1502, 1511, 1514, 1521, 1523, 1539, 1547, 1549, 1553, 1557, 1563, 1575, 1580, 1582, 1592, 1595, 1597, 1610, 1619, 1622, 1628, 1631]. IEEE [1327, 1329, 1331, 1374, 1377, 1383, 1586, 1594, 1612, 1615, 1648, 1650, 1652, 1654, 1663, 1667, 1673, 1677, 1714, 1743, 1746, 1751, 1772, 1780, 1784, 1789, 1791, 1800, 1803, 1805, 1818, 1820, 1826, 1830, 1832, 1834, 1855, 1857, 1860, 1861, 1863, 1869, 1872, 1874, 1892, 1896, 1904, 1906, 1910, 1913, 1915, 1942, 1946, 1949, 1951, 1976, 1982, 1984, 1999, 2004, 2006, 2013, 2018, 2021, 2023, 2024, 2029, 2056, 2057, 2069, 2070, 2072, 2073, 2077, 2079, 2086, 2092, 2093, 2102, 2106, 2113, 2118, 2123, 2125]. IEEE [1671, 1703, 1723, 1757, 1825, 1851, 1894, 1997, 2053, 2054, 2081, 2100, 2141, 2148, 2151, 2154, 2158, 2164, 2165, 2167, 2179, 2182, 2186, 2188, 2190, 2192, 2194, 2204, 2226, 2228, 2233, 2235, 2239, 2241, 2256, 2266, 2277, 2293, 2299, 2301, 2312, 2314, 2327, 2341, 2345, 2346, 2351, 2359, 2367, 2369, 2374, 2376, 2385, 2390, 2395, 2411, 2418, 2421, 2426, 2428, 2430, 2432, 2433, 2437, 2439, 2446, 2448, 2461, 2463, 2470, 2472, 2478, 2480, 2485, 2495, 2506, 2513, 2521, 2523, 2528, 2531, 2533, 2536, 2540]. IEEE [2365, 2415, 2445, 2475, 2525]. IEEE-CS [996]. IEEEE [968]. IEEE’s [2441]. If [271]. Ignored [1387]. II [713, 1360, 1540, 2224, 2234]. III [886, 1730]. IIoT [446]. Illuminators [1224]. Images [589, 1848]. Immersive [105, 789, 1119, 2468]. Immunology [2425]. Impact [1380, 1416, 1435, 1474, 1581, 1668]. Implementation [2372]. Implications [266, 2046, 2420]. Importance [489, 1306]. Improve [456, 908, 1891]. Improvement [1767]. Improving [417, 537, 632, 1853]. In-Class [815]. In-Memory [1347, 1531]. In-Network [1071]. Incentives [412]. Incidents [819, 2055, 2220]. Include [1375]. Inclusion [738, 810, 911, 918, 933, 1010, 1019, 1155, 1518, 1634, 1638, 1753, 1793, 1858, 1907, 1940]. Inclusive [869, 1850, 2440]. Income [13]. Increase [1370]. Increasing [2184]. Indian [476]. Industrial [34, 1062, 1104, 1154, 1328, 1483, 1676, 1809]. Industry [168, 198, 265, 266, 637, 934, 949, 1022, 1414, 1420, 1481, 1596, 1620, 1657, 1661, 1672, 1771, 1854, 1856, 1961]. Inevitable [166]. Inexpensive [1296]. Infant [301, 1372]. Inference [176, 2068]. Influence [80, 1979]. Influencer [1726]. Information [97, 970, 1154, 1421, 1697, 1821, 1844, 1848, 1883, 1931, 1968, 2026, 2044, 2045, 2119, 2146, 2199, 2218, 2253, 2255, 2307, 2328, 2378, 2407, 2459, 2464, 2538]. Information-Driven [1421]. Information-Operation [1154]. Informed [1824]. Infrastructure [690, 967, 1180, 1181, 1405]. Infrastructures [421, 965, 2392, 2423]. Infusing [1981]. Ingestible [2512]. Initiative [2264]. Initiatives [2071]. Injury [43]. Innovation [395, 421, 1213, 1771, 1808, 1812, 1938, 2084, 2336, 2422]. Innovations [2255]. Input [16, 230, 1463]. Input-Aware [1463]. Inputs [2068]. Insect [2099]. Insights [622, 1738, 1854, 2375]. Insolar [412]. Inspection [685]. Inspired [939, 1215]. Inspiring [1806]. Institutions [421, 1806, 1809, 1977]. Instructional [1266]. Instructors [1899]. Insulin [1733]. Integrated [74, 753, 1965, 2423]. Integrating [1624, 1778, 2014, 2181]. Integration [97, 266, 1154, 1198, 1339, 1713, 1742, 1926, 2193, 2372]. Integrity [1719, 1764, 2019, 2067, 2187]. Intellect [227]. Intelligence [39, 89, 189, 225, 227, 276, 292, 294, 299, 302, 336, 337, 363, 383, 398, 415, 572, 593, 636, 690, 741, 775, 830, 844, 852, 867, 873, 878, 906, 918, 943, 963, 997, 1083, 1120, 1184, 1189, 1243, 1257, 1287, 1326, 1328, 1330, 1342, 1378, 1380, 1411, 1474, 1481, 1487, 1513, 1519, 1554, 1617, 1621, 1633, 1651, 1653, 1668, 1679, 1680, 1693, 1701, 1715, 1729, 1776, 1854, 1856, 1905, 1959, 1961, 1966, 1977, 1981, 2003, 2010, 2022, 2039, 2046, 2055, 2110, 2175, 2200, 2286, 2303, 2308, 2317, 2322, 2326, 2330, 2347, 2400, 2436, 2468, 2471]. Intelligence [228, 296, 301, 335, 340, 360, 692, 800, 877, 883, 1008, 1009, 1015, 1112, 1114, 1256, 1293, 1341, 1353, 1486, 1585, 1707, 1747, 1849, 1859, 1925, 1965, 1975, 2051, 2071, 2139, 2183, 2202, 2258, 2477]. Intelligence-Based [302]. Intelligence-Driven [1513].

Intelligence/Machine [360]. Intelligent [42, 143, 362, 369, 432, 811, 1085, 1107, 1301, 1448, 1502, 2023, 2156, 2180, 2249, 2251, 2340, 2379, 2429]. Intention [2263]. Interaction [1056, 1735]. Interactions [2467]. Interactive [17, 228, 935, 1618, 1862, 2417, 2462]. Interconnect [1661]. Interface [105]. Interfaces [698]. Interference [957]. Interlinked [1659]. Internal [1911]. International [97, 1022, 1697, 1745]. Internationally [1291]. Internet [43, 49, 172, 174, 231, 243, 367, 595, 811, 840, 866, 909, 1104, 1183, 1185, 1237, 1248, 1288, 1483, 1508, 1583, 1621, 1624, 1676, 1707, 1786, 1868, 2008, 2267, 2290, 2423]. Interoperability [411, 489, 2305]. Interoperable [1185]. Interplay [670]. Interpretable [276, 660]. Interpretation [2469]. Intersection [299]. Intersections [1184, 2103]. Intervention [643]. Interview [965, 2064, 2065, 2162, 2224]. Interviews [2003]. Intraoperative [1507]. Introduces [704]. Introduction [692, 2117]. Intrusion [2258]. Intuition [1210, 1670, 2427]. Invention [622, 1418, 2503]. Inversion [165]. Investigation [840]. Invisible [900, 1102, 1271]. IOT [110, 135, 138, 712, 730, 1190, 1286, 1439, 1939, 2122, 2257, 2292]. IoT-Based [135]. IOTA [1543]. IOTA-Based [1543]. IoTility [1110]. IQQC [655, 728]. IRI [97, 439]. ISEMV [2259]. Ising [2337]. Island [2108]. Isn’t [2526]. ISO [2219]. Isomorphic [595]. Issue [101, 153, 162, 214, 217, 250, 330, 438, 472, 502, 561, 726, 762, 831, 901, 932, 962, 1030, 1061, 1100, 1136, 1285, 1400, 1471, 1538, 1572, 1692, 1727, 1802, 1842, 1882, 1980, 1992, 2009, 2095, 2138, 2178, 2201, 2237, 2274, 2287, 2309, 2323, 2334, 2381, 2406, 2460]. Issues [199, 833, 843, 1248, 1441, 1442, 1699, 2067, 2446, 2476, 2506]. ITRI [1809]. IV [1081].

Jim [1049]. Job [87, 541, 562, 674, 1126, 1277, 1309, 1868]. JobBoards [857]. Jobs [29, 59, 244, 270, 349, 390, 402, 488, 498, 603, 715, 748, 784, 822, 891, 922, 952, 982, 1017, 1051, 1088, 1161, 1201]. Journal [151, 215, 248, 316, 322, 351, 392, 429, 463, 522, 553, 578, 624, 667, 682, 717, 750, 786, 824, 859, 893, 924, 1098, 1138, 1195, 1212, 1255, 1312, 1521, 1595, 1826, 1861, 1942, 2164, 2346, 2385, 2486]. Journalism [2405]. Journals [1297]. Journey [253, 1808]. Judson [2162, 2224]. Just [817, 995, 1524, 1822]. Jyotika [1196, 2084].

KDN [2289]. KDN-FLB [2289]. Keep [2104]. Kemosabe [529]. Kerberoasting [2185]. Key [103, 106, 111, 920, 1660, 1885, 2130, 2195, 2305]. Keyboards [256]. KF [630]. Kim [979]. Kinship [6]. Kit [1556]. Know [375, 838, 1893]. Knowledge [566, 1589, 1777, 2289, 2330, 2366, 2410, 2509]. Knowledge-Defined [2289]. Known [772].

L [2074]. Labeling [1337]. Labels [737]. Laboratories [540]. Labs [934]. Lambda [1506]. Landscape [458, 590, 1139, 1288, 2067, 2157, 2490]. Language [476, 1513, 1537, 1589, 1740, 1749, 1788, 1925, 1930, 1936, 2068, 2142, 2150, 2304, 2340, 2384, 2440, 2465, 2469]. Language-Taught [1513]. Large [67, 943, 1198, 1537, 1589, 1740, 1749, 1788, 1925, 1930, 2142, 2150, 2340, 2384, 2469, 2509]. Large-Capacity [67]. Large-Scale [943]. Last [334, 2286]. Last-Mile [334]. Latency [533, 1296, 1864]. Later [2066]. Latin [971]. Law [297, 566]. Laws [1081]. Lay [1370]. Layer [1524]. Lead [1333, 1583]. Leadership [2130]. Leading [1542, 1609]. Leak [453]. Leap [2139]. Learn [157, 413, 485, 896, 1065, 1738, 2143, 2229, 2504]. Learned [855, 1233, 1336, 1454, 2017]. Learner [967]. Learning [142, 170, 260, 360, 540, 660, 692, 695, 697, 753, 870, 872, 883, 903, 906, 938, 965, 967, 969, 993, 1040, 1065, 1107, 1139, 1147, 1151, 1158, 1260, 1333, 1348, 1368, 1372, 1407, 1478, 1517, 1621, 1660, 1664, 1729, 1732, 1765, 1774, 1884, 1963, 1995, 2027, 2052, 2104, 2254, 2273, 2289, 2326, 2379, 2471, 2509]. Learning-Enabled [1517]. Learns [2508]. Led [2084]. Ledger [442, 1543]. Lee [1049]. Legacy [1151]. Legal [906, 1410]. Legends [1477]. Lens [2219]. Lessons [1233, 1236, 1336, 1477]. Let [888, 2387]. Letters [255, 261, 373, 397, 477, 588, 617, 1918, 2087, 2229]. Level [537, 1334, 1577]. Levels [1263]. Leveraging [1884, 1891, 1898, 2375]. Libraries [1296]. License [311, 341, 381, 2438]. Lie [1037]. Lies [2304, 2386]. Life [41, 96, 385, 412, 697, 1302, 1744, 2423]. Lifecycle [1038, 1079, 1884, 1961, 2096, 2364, 2366, 2373]. Lifelong [967]. Lights [1344]. Lightweight [855, 920]. Like [413, 478, 934, 2143]. Limitations [1151, 1581, 2003]. Limits [743]. Lineage [375]. Lines [408]. Linguistic [1513]. Linkages [1033]. Linked [1187, 1188]. Linking [508]. Links [1033]. Listening [725]. Lithography [1658]. Litigation [1156]. Live [508, 1433]. Lives [2317]. LLM [2221]. LLM-Powered [2221]. LLMs [1537, 2375]. Local [1848]. Localization [483]. Location [1230]. Location-Based [1230]. Logic [2112]. Logical [475]. Logins [510]. London [1186]. Long [419, 1093, 2058]. Look [21, 806, 2145]. Looking [2145]. Loop [2507]. Loss [20]. Lost [585]. Love [2524]. Low [13, 842, 1440, 1475, 1963, 2213, 2217, 2219– 2221, 2267, 2306, 2515]. Low-Altitude [1440]. Low-Code [2219–2221, 2306]. Low-Code/No-Code [2219, 2221, 2306]. Low-Cost [2267]. Low-Income [13]. Low-Power [1475]. Low-Risk [1963]. Lower [1537]. LSTM [630]. LSTM-TD [630].

M.D. [850]. Machine [142, 170, 360, 505, 540, 564, 692, 695, 697, 744, 817, 855, 870, 872, 883, 903, 906, 965, 969, 993, 1015, 1065, 1107, 1114, 1147, 1151, 1333, 1407, 1454, 1517, 1660, 1729, 1732, 1738, 1774, 1884, 1963, 1995, 2017, 2052, 2142, 2254, 2273, 2326, 2379]. Machine-Learned [855, 1454, 2017]. Machine-Learning-Based [2052]. Machines [514, 804, 1895]. Machining [1210]. Magazine [94, 127, 133, 137, 143, 169, 175, 190, 277, 328, 333, 443, 532, 586, 597, 745, 812, 882, 1563, 2092, 2093]. Magazines [5, 35, 64, 92, 156, 188, 220, 258, 285, 325, 355, 400, 434, 467, 497, 527, 556, 583, 653, 686, 721, 755, 791, 827, 862, 897, 928, 958, 989, 1025, 1057, 1095, 1131, 1168, 1208, 1250, 1282, 1317, 1356, 1396, 1429, 1465, 1499, 1533, 1568, 1605, 1643, 1687, 1722, 1756, 1796, 1837, 1877, 1919, 1954, 1987, 2036, 2037, 2088, 2131, 2172, 2209, 2246, 2282, 2318, 2358, 2401, 2454]. Magic [2377]. Maiga [1158]. Mails [1975]. Mainstream [2530]. Maintenance [1259, 2441, 2471]. Major [507, 531, 1415, 1490]. Make [539, 1641]. Makers [1885]. Making [641, 1216, 1698, 1740, 1845, 2458, 2477, 2479]. Malls [458]. Malpractice [549]. Mamba [2416]. Manage [704, 1233]. Management [379, 1301, 1364, 1387, 1885, 2007, 2061, 2161, 2303]. Manager [67]. Managing [51, 177, 294, 906]. Mandy [1198]. Mania [886]. Manipulation [1959]. Manual [23]. Manufacturing [106, 272, 590, 665, 732, 1943]. Map [272, 2275, 2417]. Mapping [483, 2144]. Mark [1641]. Market [910, 1868]. Marketing [1779, 1938, 2270]. Marketplaces [195]. Markets [138, 197]. Married [866]. Mars [697, 2387]. Mass [1014, 1972]. Mastery [2424]. Masthead [4, 33, 63, 91, 125, 155, 187, 219, 252, 284, 324, 353, 394, 431, 465, 496, 524, 555, 580, 616, 651, 684, 719, 752, 788, 826, 861, 895, 926, 956, 986, 1021, 1055, 1092, 1130, 1165, 1205, 1247, 1281, 1316, 1355, 1395, 1427, 1462, 1496, 1530, 1565, 1603, 1640, 1686, 1721, 1755, 1795, 1836, 1876, 1917, 1953, 1986, 2034, 2035, 2083, 2129, 2171, 2208, 2245, 2281, 2316, 2355, 2399, 2452, 2499]. Materials [2197, 2269]. Matter [1266, 1711]. Matters [16, 1266]. Maturity [2329]. Mauchly [406]. Maximum [157]. May [1273]. Maybe [1974]. Me [539, 689, 2387]. Meaningful [2474]. Measurability [2238]. Measure [456, 1711]. Measured [789]. Measurement [139, 1767, 1768, 2109]. Measurements [1428]. Measuring [879, 2254]. Mechanical [147]. MedEd [2222]. Media [82, 805, 836, 1149, 1232, 1551, 1829, 1852, 1862, 2159]. Mediated [2467]. Medical [42, 547, 607, 698, 1233, 1439, 1961, 2222]. Medicine [1439, 1707]. Medicines [1253]. Meet [1508]. Meets [223, 770, 1140, 1849, 2217]. Megacities [1181]. Megatrends [1446, 2505]. Mei [979]. Member [123, 185, 300, 461, 492, 551, 576, 647, 680, 722, 818, 1245, 1279, 1434, 1573, 1616]. Membership [427, 520, 612, 708]. Memex [23]. Memoriam [84, 2486]. Memories [1744]. Memory [67, 795, 1077, 1081, 1347, 1384, 1442, 1531, 1661, 1783]. Mental [1551]. Mentality [193]. Merely [1932]. Merging [1589]. Merwin [425]. Message [560]. Meta [2412, 2509]. Meta-Learning [2509]. Metadata [1480, 1668]. Metamorphic [1889]. Metaverse [888, 1085, 1119, 1241, 1266, 1274, 1456, 1524, 1550, 2110, 2212, 2217, 2223, 2259, 2404, 2412]. Metaverse-Based [1266]. Metaverses [1192, 2222]. Method [1330, 2290]. Methodologies [1260]. Methodology [2435]. Methods [144, 405, 658, 662, 1817, 2112]. Metric [886]. Metrics [912, 2435]. Metrology [961, 1658]. Metropolises [1184]. Microarchitecture [466, 979]. Microprocessors [1428]. Microprogramming [979]. Microservice [34]. Mile [334]. Milestone [758]. Mind [1893, 2156]. Mindset [374, 479]. Minefield [2272]. Minimalism [2308]. Mining [807, 1186]. Mirror [735]. Misinformation [2110]. Missing [969]. Mission [697]. Missions [743]. Mistakes [1701]. Mitigating [1510, 1827, 1965, 2185]. Mitigation [819, 2108]. ML-Centric [1960]. Mobile [444, 474, 516, 1478, 1512, 1846]. Mobility [1291]. Mockery [1991]. Modality [1847]. Model [367, 564, 630, 663, 733, 777, 1079, 1188, 1256, 1348, 1443, 1621, 1788, 1887, 1934, 1936, 1995, 2143, 2329, 2469]. Model-Assisted [2469]. Modeling [663, 665, 972, 1479, 1994, 2159]. Models [943, 1262, 1537, 1589, 1740, 1749, 1770, 1788, 1925, 1930, 2068, 2142, 2150, 2160, 2265, 2304, 2340, 2384, 2420, 2440, 2509, 2526]. Modern [103, 1774, 1888, 2202, 2292, 2331]. Modernization [2260]. Modular [2510]. Modularity [483, 1065]. Mondira [1198]. Monetize [1556]. Monetizing [437]. Money [666]. Monitoring [40, 132, 1372, 1507, 1510, 1543, 1544]. Monkey [140]. Monopoly [2366]. Moon [2387]. Moonshot [117]. Moore [566]. Morse [1015]. Mortality [301]. Mostly [815]. Motion [1618]. Motivation [1710]. Motivations [2000]. Move [739]. moveSMART [1140]. Movie [2458]. Movies [2263]. Moving [305, 976, 1701]. MRI [1480]. MTTKRP [1463]. Much [2300, 2481, 2519]. Multiagent [606, 2271]. Multibillion [1103]. Multidrone [536]. Multilayer [230]. Multilayer-Perception-Based [230]. Multilayered [834]. Multilevel [840, 1770]. Multimaster [1579]. Multimedia [94, 127, 597, 2024, 2093, 2433]. Multimodal [483, 2068, 2255]. Multiple [230, 1072, 1293]. Multiple-Deep [1293]. Multiple-Input [230]. Multiprocessor [1104, 1578]. Multipurpose [746]. Multirobot [368]. Multistep [630]. Multistep-Forward [630]. Multitenant [1705]. Multiyear [1474].

Must [1893, 2474]. My [510, 1749, 1929].

MySQL [2453]. Myth [337].

Name [2286]. Named [2408]. Narrow [100]. National [337, 1274, 1891]. Nations [1340]. Native [474, 731, 2417]. Natural [1346]. Navigating [1222, 1513, 2015, 2067]. Nearer [1004]. Necessities [1647]. Need [934, 1191, 1388, 2263, 2529]. Needs [873, 1771]. Neighborhood [134]. Neighbors [2290]. Neither [451]. Network [134, 379, 412, 1071, 1293, 1294, 1520, 1545, 1815]. Network-Based [1815]. Networking [11, 387, 1116, 1215, 1958, 2289, 2290, 2408]. Networking-Based [2290]. Networks [179, 627, 987, 994, 1365, 1664, 1737]. Neuhold [820]. Neumann [1531]. Neural [179, 987, 1293, 1365, 1478, 1664, 1815, 2537]. Neurodiversity [2386]. Neuron [633]. Neurophysiological [1507]. Neurosymbolic [2532]. Never [585]. News [2405]. Next [432, 707, 743, 915, 1293, 1658, 2408, 2434, 2508]. Next-Generation [432, 1293, 2408]. NFTs [1623]. Niceties [1647]. Niche [2530]. Niches [1809]. Night [1408]. NIST [2122]. No [447, 1635, 2213, 2217, 2262]. No-Code [2219, 2221, 2306]. Node [1579]. Nodes [1184]. Nomination [1024]. Nominations [89, 141, 196, 224, 468, 494, 507, 517, 1371, 1404, 1415, 1437, 1476, 1490]. Non [2490]. Non-AI [2490]. Nonfungible [910, 1118]. Nonlinearity [602]. Nor [451]. Normative [1326]. NoSQL [508]. Notable [1189]. Noteworthy [1114]. Notice [1454]. Novel [42, 369, 444, 1556]. NP [1777]. NSF [1035]. Nudge [1486].

Obituary [347, 2074]. Object [369, 1224]. Objects [735, 976]. Obligations [2438]. Observing [1291]. Obtaining [108]. Oceans [938]. OERs [1348]. Off [1470]. Offense [512]. Office [289, 1336]. Officers [1945, 2442]. Offices [2107]. Offloading [753]. Oh [1749]. OJ [954]. Old [1451, 2149]. Olympics [1060]. Ombudsman [2196, 2234]. On-the-Fly [45]. One [1339]. Ongoing [172, 1657]. Online [26, 471, 568, 630, 780, 1011, 1149, 1445, 2014]. Onslaught [1932]. Open [51, 113, 144, 151, 168, 177, 199, 215, 239, 248, 291, 309, 311, 316, 322, 341, 351, 381, 392, 429, 455, 463, 515, 522, 553, 571, 578, 624, 641, 667, 682, 705, 709, 717, 750, 777, 786, 803, 824, 859, 893, 924, 974, 1036, 1098, 1138, 1187, 1188, 1191, 1195, 1212, 1255, 1265, 1312, 1336, 1413, 1521, 1595, 1625, 1661, 1706, 1773, 1826, 1853, 1861, 1941, 1942, 2005, 2107, 2160, 2164, 2191, 2335, 2346, 2385, 2420, 2438, 2517]. Open-Source [2005]. OpenAI [1903]. OpenChain [341]. Opening [656]. Opens [974]. OpenStack [848]. Opera [780]. Operating [1578]. Operation [69, 1154]. Operational [15, 811, 1151]. Operations [1271, 1440, 2193, 2270]. Opioid [836]. Opportunistic [1039]. Opportunities [71, 199, 269, 670, 731, 905, 1119, 1609, 1712, 1807, 1870, 1944, 2222, 2330, 2368, 2416]. Opportunity [253, 1224]. Optane [1304]. Optical [1142]. Optimal [134]. Optimization [606, 753]. Optimizing [1463]. Oracle [1889]. Oracles [1454, 2017]. Orbit [2515]. Orbital [93]. Organization [1387]. Organizations [766, 965]. Oriented [937, 2230]. Originality [2294]. Origins [622, 2268]. OT [1939]. Other [800]. Our [1609, 2156, 2317]. Outlines [1972]. Outlook [407, 797, 1211, 1649, 2137]. Outlooks [12]. Outrun [1269]. Outward [115]. Overcome [1513]. Overcoming [1187, 1520]. Oversight [2429]. Overview [1552, 1971, 2487]. Own [2227, 2263]. Ownership [1659]. Oxymoron [47].

P [384, 1777]. P2851 [2275]. Pacific [881]. Pain [189]. Pains [1152]. Pairings [913]. Pandemic [485, 2000, 2105].

Pandemic-Related [2000]. Pandemics [449]. Panel [1693, 1908]. Pant [1198]. Paper [871]. Papers [129, 180, 194, 226, 273, 275, 304, 314, 342, 365, 410, 433, 439, 450, 480, 526, 559, 598, 631, 642, 664, 740, 765, 779, 796, 839, 864, 874, 904, 917, 996, 1028, 1094, 1132, 1141, 1171, 1182, 1223, 1249, 1292, 1325, 1329, 1358, 1432, 1466, 1498, 1532, 1567, 1608, 1669, 1694, 1734, 1766, 1797, 1838, 1878, 1933, 1964, 1988, 2048, 2049, 2098, 2136, 2158, 2176, 2182, 2214, 2247, 2291, 2325, 2357, 2402, 2450, 2455, 2497, 2501]. Paradigm [884, 1063, 1104, 1583, 1862, 2308]. Paradigms [730, 2368]. Paradox [1739, 2183]. Parallel [326, 581, 927, 1037, 1123]. Parallels [485]. Parasocial [2467]. Parlor [1822]. Parma [1448]. Part [210, 713, 886, 1254, 1360, 1540, 1730, 2162, 2196, 2224, 2234, 2424]. Partially [1387]. Participation [848, 1391, 2000, 2184]. Partition [74]. Partition-Driven [74]. Partners [1368]. Parts [536]. Passé [240]. Passive [1234, 2105]. Passphrase [599]. Passwordless [240]. Passwords [240]. Past [531, 1552, 1701, 1812, 2152]. Patching [193]. Patent [1156]. Path [878, 1439, 1587, 1713, 1846]. Pathfinding [606]. Pathway [364]. Patient [1510, 1543, 1544]. Patients [43, 302]. Pattern [1015]. Patterning [1658]. Patterns [836, 870]. Paving [2489]. Pay [865]. Paying [1235]. Payments [2530]. Pedagogy [1139]. Pedestrian [2047]. Peer [268, 1617]. Peer-to-Peer [268]. Pencil [1823]. People [13, 398, 1346, 1447, 1733]. Perceived [1819]. Perception [230, 2050, 2052]. Perceptions [694, 2003]. Perfect [1284]. Performance [39, 994, 1037, 1194, 1342, 1477, 1520, 1680, 1763, 1781, 1888, 1926]. Perils [1380]. Periocular [1847]. Persistent [67, 372]. Person [969]. Personal [978]. Personalized [20, 1733]. Personhood [2180, 2215]. Perspective [637, 660, 1108, 1259, 1363, 1479, 1698, 1735]. Perspectives [82, 626, 773, 1275, 1364, 1961, 2001]. Pervasive [19, 112, 167, 173, 205, 229, 233, 275, 308, 380, 416, 445, 484, 535, 610, 638, 659, 769, 802, 845, 876, 936, 1109, 1153, 1193, 1223, 1295, 1338, 1386, 1484, 1509, 1549, 1597, 1619, 1654, 1818, 1863, 1926, 2006, 2069, 2070, 2113, 2188, 2256, 2299, 2369, 2411, 2463, 2536]. Pharmaceutical [198]. PharmaCrypt [198]. Phasor [139]. Phenomenon [947]. Photonics [914]. Physical [103, 256, 1140, 1334, 1908]. Physics [1824, 2535]. Physics-Based [2535]. Physics-Informed [1824]. Physiological [2140]. Pi [1268]. Pill [2339]. Pioneering [903]. Pipelines [1145, 1544, 1660]. Pitfalls [2161]. Pixar [1903]. Pixels [2356]. Places [844]. Placing [1003]. Plagiarism [386]. Planning [564, 2271]. Plateau [993]. Platform [42, 474]. Platforms [506, 2219, 2306]. Plausible [875]. Play [157, 865, 1340, 2149]. Playbook [2389]. Player [2424]. Playing [1956, 2404]. Plug [2149]. Podast [1472]. Podcast [1157, 1173, 1225, 1267, 1303, 1335, 1366, 1406, 1444, 1505, 1541, 1584, 1626, 1665, 1696, 1741]. Poisoning [1147, 1732]. Policies [2487]. Policy [337]. Political [2135]. Population [1849]. Portal [160]. Position [949]. Positions [1945, 2442]. Post [1479, 1531, 2483]. Post-2022 [2483]. Post-Quantum [1479]. Post-von [1531]. Postcoronavirus [206]. Postsilicon [466]. Pot [371]. Potential [13, 1405, 1624, 2173, 2230, 2373]. Poultry [2267]. Poverty [1189]. Power [135, 136, 533, 1075, 1475, 1630, 2227, 2535]. Powered [2221]. Practical [13, 236, 1063, 1139, 1859, 2112]. Practice [645, 1105, 1328, 1421, 1701, 1770, 2380]. Practices [940, 965, 1364, 1387, 1745, 2272, 2349]. Practitioner [2001, 2389]. Precision [1234, 2512]. Predicting [301, 992, 1435, 1680, 2295, 2382]. Prediction [43, 202, 303, 630]. Predictions [587, 1921, 1924, 2147]. Predictive [2270, 2441]. Prefetching [1864]. Preliminary [1326]. Preparedness [208]. Preprocessing [1333, 1378]. Present [513, 1552, 1812, 2152]. Preserved [134]. Preserving [903, 1043, 1765, 2290]. President [388, 395, 1196, 2084]. President-Elect [388, 1196]. Pressure [1008]. Preterm [1372]. Prevent [701]. Prevention [819, 836, 2046]. Principles [1666, 1680, 1701]. Printing [2040]. Prioritization [1905]. Prisoner [2005]. Privacy [14, 44, 72, 114, 134, 137, 167, 175, 203, 234, 263, 277, 310, 315, 333, 418, 443, 486, 503, 532, 582, 640, 645, 691, 745, 812, 842, 847, 881, 882, 903, 919, 973, 1034, 1038, 1043, 1075, 1106, 1122, 1150, 1179, 1221, 1272, 1298, 1345, 1379, 1449, 1482, 1505, 1510, 1511, 1541, 1553, 1610, 1648, 1700, 1739, 1765, 1772, 1780, 1800, 1820, 1834, 1855, 1874, 1886, 1892, 1915, 1951, 1976, 2013, 2029, 2072, 2073, 2077, 2081, 2106, 2140, 2151, 2187, 2194, 2226, 2290, 2301, 2469, 2477, 2490, 2529]. Privacy [2165, 2314, 2341, 2374, 2418, 2461, 2513, 2523]. Privacy-Assured [1038]. Privacy-Preserved [134]. Privacy-Preserving [903, 1043, 1765, 2290]. Private [1621]. Prize [202, 303]. Probabilistic [1777]. Probabilities [2007]. Problem [228, 685, 800, 969, 1889, 2108, 2527]. Problematic [1975]. Problems [1410]. Process [937, 1151]. Processes [1003, 2372, 2471]. Processing [369, 1123, 1478, 1506, 1697]. Processors [2020]. Product [1079, 1884, 2061, 2298]. Production [24]. Productivity [1587, 2349, 2370]. Products [374]. Profession [542]. Professional [85, 107, 171, 207, 273, 306, 339, 370, 414, 448, 482, 509, 608, 635, 661, 767, 841, 1044, 1115, 1159, 1197, 1227, 1514, 1547, 1582, 1677, 1816, 1865, 2011, 2062, 2063, 2198, 2231, 2348, 2371]. Professionalism [278]. Professionals [538, 798, 874, 1144, 1702, 1748, 1806, 1900, 2115, 2261, 2297, 2413, 2482]. Professor [385, 1308]. Program [773, 1336, 1588, 2107, 2122, 2384]. Programmable [1888]. Programmers [1627]. Programming [367, 1895, 2230]. Programs [50, 940, 2487]. Progress [730, 1213]. Project [641, 1140, 1387]. Projects [2335]. Promise [17, 49, 508, 1517]. Promises [309, 1380, 1419]. Prompting [1633]. Proof [1777]. Proofs [1117, 2410]. Propagate [1809]. Propagating [1613]. Properties [855, 2017]. Proposal [301, 1980, 2012, 2138, 2237, 2274, 2309, 2334, 2381]. Proposals [153, 214, 217, 250, 2009, 2201]. Pros [2216]. Prosecutions [1588]. Prospects [1120, 2001]. Prosperity [1011, 1554]. Protect [2392]. Protecting [774, 1346, 1522]. Protection [26]. Protocols [367, 405, 537, 730, 1546]. Provisioning [1737]. Public [197, 294, 296, 449, 1621, 1774, 1809, 1810]. Publication [567]. Published [2077]. PUF [730]. PUF-Based [730]. Puke [539]. Puppeteers [900]. Purpose [912]. Pushing [743]. Putting [240]. PVC [1084].

QAnon [947]. Quagmire [1551]. Quality [568, 879, 1326, 1343, 1404, 1437, 1566, 2152, 2219, 2368]. Quantifying [1039]. Quantitative [2114]. Quantum [55, 61, 109, 117, 131, 174, 236, 307, 644, 880, 1023, 1053, 1213, 1273, 1347, 1391, 1393, 1425, 1460, 1479, 1552, 1778, 2139]. Quasi [139]. Quasi-Dynamic [139]. Quest [50, 1378]. Questions [144].

Race [701, 995, 999, 1487, 2302, 2517, 2537]. Racial [1367]. Radio [2427]. Rainbow [942, 966, 1000, 1064, 1157, 1173, 1225, 1267, 1303, 1335, 1366, 1406, 1444, 1472, 1505, 1541, 1584, 1626, 1665, 1696, 1741]. Random [1296]. Ransomware [865, 946, 1556]. Ready [867]. Real [231, 266, 369, 405, 412, 504, 540, 819, 999, 1507, 2055, 2099, 2537]. Real-Time [231, 266, 369, 504, 1507, 2099, 2537]. Real-World [405, 2055]. Realities [2010]. Reality [13, 17, 105, 256, 336, 789, 853, 1620, 1898, 1959, 2466]. Realization [2330]. Realizing [49]. Rearchitecting [458]. Reasoning [2225]. Recipients [2085]. Recognition [525, 632, 1222, 1367, 1848]. Recognize [2294]. Recommendations [693]. Recommender [905, 1733]. Reconfigurable [466]. Reconstruction [2537]. Recovery [545, 1142]. Redefining [1180, 2349]. Reduce [877, 1413]. Reductions [2534]. Redundancy [1365]. Redux [2061]. Reference [939]. Referral [302]. Reflecting [53, 758]. Regenerative [2510]. Regions [1508]. Registration [2220]. RegTech [1174]. Regulating [1960]. Regulation [26, 167, 1180, 1852, 2527]. Regulations [2015]. Regulatory [1220, 1363, 1961]. Regulatory-Compliant [1961]. Reinforcement [1664]. Reinventing [37]. Relapse [836]. Related [1552, 2000]. Reliability [165, 625, 1075, 1763, 1770, 1889, 1994, 2060, 2333, 2356, 2441, 2459]. Reliable [69, 73, 602, 817, 1740, 1996, 2410]. Reminiscences [1451]. Remote [40, 592, 1442, 1507, 1508, 1510, 1543, 1544]. Removing [395]. Rename [888]. Rendering [2537]. Repair [2384]. Replace [815]. Representation [1813]. Reproducibility [761, 927, 1031, 1035, 1036]. Reproducible [581, 1033]. Repurposed [145]. Reputation [444]. Require [1271]. Requirement [852, 1289]. Requirements [771, 1257, 1324, 2152]. Rescue [1668, 2193]. Research [167, 204, 227, 239, 581, 927, 1023, 1031, 1033, 1036, 1226, 1264, 1290, 1376, 1585, 1623, 1701, 1737, 1806, 1809, 1849, 2144, 2255, 2329, 2330]. Researchers [168]. Reshapes [1234]. Reshaping [240, 590, 811, 2317, 2405]. Resign [2012]. Resilience [136, 456, 533, 545, 643, 1260, 1908, 1932]. Resiliency [565]. Resilient [75, 132, 1674, 1704, 2264]. Resolving [2220]. Resource [2349]. Resources [160, 481]. Responsible [227, 1321, 1324, 1328, 1334, 1474, 2071, 2264, 2330]. Results [1275]. Resume [1470]. Resurgence [2483]. Resynthesis [368]. Retail [458]. Rethink [374]. Rethinking [289, 2272, 2331, 2431]. Retinopathy [302]. Retrieval [1962]. Retrieval-Augmented [1962]. Retrospective [78, 440, 2259]. Return [2377]. Reusability [1660]. Reuse [97, 2191]. Revenue [157]. Review [1326, 1617, 2181, 2195]. Reviewers [1991]. Reviews [1991]. Revisited [145]. Revisiting [961, 973, 1224, 1994]. Revolución [235]. Revolution [1657, 1807]. Revolutionizing [2263]. Rewards [572]. Rich [1296]. Richard [425]. Right [1706]. Rigid [606]. RIGOR [301]. Rigorous [417]. Rise [174, 593, 804, 1174, 1380, 1550, 2015, 2420]. Risk [208, 704, 1363, 1364, 1963, 2007, 2191, 2423]. Risks [170, 572, 731, 873, 906, 1233, 1413, 1510, 1707, 1712, 1788, 1827, 1870, 2366, 2393]. Road [272, 1419, 2189, 2275]. Roadmap [1022, 1712, 1972, 1998]. Roadside [967]. Robo [694]. Robo-Advisory [694]. Robotic [504, 2512]. Robotics [743, 2505, 2514]. Robots [2156, 2512]. Robust [2050, 2258]. Robustness [1259, 2101]. Role [483, 918, 1036, 1823, 1888, 1899, 1908, 2330, 2517]. Roles [1339, 1440]. Room [487]. Root [2047]. Rosebush [2162, 2224]. Roundtable [727, 1729, 1908, 1923]. Roundtables [1651, 1693]. Route [2271]. Routing [369]. RT [504]. Rule [1339]. Runtime [504, 1764, 2017]. Rush [2272]. Russian [1932].

S [384, 939, 964, 1343, 1404, 1419, 1437, 1554]. SA [168]. SaaS [2265]. Safe [665, 811]. Safer [2071]. Safety [71, 140, 363, 489, 625, 626, 764, 1275, 1365, 1375, 1517, 1593, 1659, 1747, 1762, 1773, 1774, 2052, 2275, 2441, 2489]. Safety-Critical [71, 1365, 1517]. Saga [2263]. Said [201]. Salvador [975]. Sanctions [2161]. Sanitization [1232]. Santander [1186]. Satellite [1824]. Savings [1405]. Say [675]. Scalable [880, 2537]. Scale [908, 943, 1198, 2101]. Scaled [1474]. Scales [1342]. Scaling [2526]. Scams [910]. Scandal [21]. Scanner [1480]. Scarcity [403]. Scenario [702]. Scenario-Based [702]. Scenarios [1847]. Scene [811]. Schedule [2220]. Scheirer [1015]. Scheme [1905]. Schneier [2064, 2065]. School [289]. Schools [1548, 1590]. Science [75, 97, 103, 291, 544, 567, 600, 669, 697, 710, 759, 837, 843, 853, 868, 950, 1037, 1042, 1073, 1199, 1233, 1269, 1297, 1305, 1389, 1451, 1455, 1523, 1555, 1580, 1622, 1652, 1746, 1782, 1811, 1869, 1904, 2016, 2118, 2148, 2296, 2430, 2485, 2487]. Sciences [511, 790]. Scientific [770, 987, 1342, 1926]. Scientist [1269]. Scientists [1799, 1804, 1806]. Scientometrics [886]. Scope [1706]. Screening [1512]. Script [506]. Search [386, 1601, 1675, 1678, 1684, 1936, 2111, 2193, 2271, 2332, 2388]. Second [1693, 1930]. Secondary [1326]. Sector [294, 1234, 1411, 1810]. Sectors [2163]. Secure [75, 243, 432, 446, 665, 1996, 1998, 2020, 2251, 2260]. Securely [354]. Securing [172, 731]. Security [72, 74, 77, 114, 137, 170, 175, 234, 277, 305, 310, 333, 418, 443, 486, 503, 516, 532, 565, 582, 640, 645, 691, 730, 731, 733, 745, 812, 847, 882, 919, 1034, 1045, 1075, 1106, 1150, 1179, 1221, 1272, 1298, 1345, 1379, 1382, 1441, 1449, 1482, 1505, 1511, 1541, 1553, 1593, 1610, 1648, 1700, 1732, 1742, 1772, 1820, 1855, 1880, 1887, 1892, 1932, 1976, 1990, 1995, 2005, 2013, 2064, 2065, 2072, 2073, 2106, 2151, 2185, 2187, 2194, 2226, 2236, 2238, 2249, 2301, 2341, 2374, 2418, 2429, 2435, 2461, 2513]. Security [14, 44, 47, 203, 263, 343, 604, 625, 652, 672, 1248, 1302, 1421, 1788, 1971, 2001, 2423, 2445, 2489]. See [782]. Seeking [991]. Seeks [1563]. Selection [1463]. Selective [2400]. Self [366, 915, 978, 1705, 1813, 2047, 2189]. Self-Adaptive [366]. Self-Directing [2189]. Self-Driving [1813, 2047]. Self-Encrypting [1705]. Self-Sovereign [978]. Semantic [1545, 1679]. Semi [2050]. Semi- [2050]. Semiconductor [1384, 1420, 1658, 1783, 1809, 1943]. Sense [1210]. Sensing [505, 2520]. Sensitive [139]. Sensor [782, 1382]. Sensor-Based [782]. Sensoring [1559]. Sensors [1072, 1151, 1618, 1928]. Sensory [1959]. Separate [702]. Sequencing [111]. Serfs [2215]. Serious [1775, 2465]. Server [1296, 2453]. Serverless [1506, 1680]. Service [121, 132, 209, 241, 771, 799, 1069, 1071, 1239, 1323, 1412, 1438, 1473, 1515, 1556, 1561, 1583, 1621, 1682, 1717, 1785, 1866, 1902, 1937, 1978, 2116, 2181, 2338, 2383, 2488, 2518]. Services [42, 864, 1445, 1632, 1845, 1856]. Set [1929]. Setting [601]. Settings [1266]. Shadows [1344, 1571]. Shape [1609, 1848]. Shaping [2270, 2347, 2412]. Share [481]. Shared [17, 1296]. Sharing [627, 1186, 1887, 2410]. Shifting [458]. Shoehorning [668]. Shooter [789]. Short [1296]. Short-Latency [1296]. Shortage [1591]. Should [866, 1548, 1590, 1702, 1899, 2012]. Shoulders [1228]. Shrink [993]. Shull [395]. Side [230, 2343]. Side-Channel [230]. Sides [848]. Sign [476, 632]. Signatures [108]. Signing [45]. Silent [2343]. Silicon [621, 914, 2027]. Silver [360, 447]. Simulate [999]. Simulation [1618, 1620]. Simultaneous [483]. Single [113]. Single-Vendor [113]. Site [541]. Situated [231]. Situation [2105, 2290]. Sixth [1210]. Skepticism [1119, 2059]. Skills [1819]. Sleep [43]. Sleep-Deprivation [43]. Slice [1294]. Slow [941]. Slowing [167]. Small [237, 1387, 1621, 1925, 1936]. Smart [49, 106, 130, 140, 265, 374, 547, 564, 593, 662, 771, 834, 1178, 1181, 1183– 1185, 1187, 1188, 1448, 1929, 1998, 2099, 2181, 2223, 2267, 2290, 2336]. Smarter [991, 1271, 2400]. Social [11, 82, 192, 387, 545, 805, 836, 1116, 1149, 1334, 1551, 1829, 1852, 1958, 1979, 2135]. Socially [2514]. Society [5, 11, 22, 28, 29, 35, 46, 58, 59, 64, 76, 86, 92, 120, 121, 123, 129, 150, 156, 182, 185, 188, 213, 220, 226, 244, 245, 247, 258, 262, 270, 279, 280, 285, 300, 304, 319, 325, 342, 348, 349, 355, 359, 365, 382, 388–390, 395, 406, 578, 677, 738, 748, 799, 810, 869, 900, 1019, 1111, 1155, 1196, 1323, 1343, 1362, 1423, 1458, 1492, 1526, 1558, 1753, 1785, 1790, 1793, 1858, 1907, 1940, 1945, 2002, 2030, 2078, 2084, 2195, 2203]. Society [400, 410, 426, 427, 434, 450, 460, 461, 467, 480, 491, 492, 497, 519, 520, 526, 527, 550, 551, 556, 559, 575, 576, 583, 611, 612, 614, 631, 646, 647, 649, 653, 664, 679, 680, 686, 708, 714, 721, 722, 740, 747, 755, 760, 765, 783, 791, 796, 818, 821, 827, 839, 856, 862, 871, 890, 897, 904, 921, 928, 951, 958, 981, 989, 1002, 1016, 1025, 1057, 1087, 1090, 1094, 1095, 1116, 1125, 1128, 1131, 1132, 1160, 1163, 1168, 1171, 1172, 1200]. Society [1028, 1032, 1047, 1050, 1195, 1203, 1208, 1239, 1242, 1245, 1249, 1250, 1276, 1279, 1282, 1292, 1311, 1314, 1317, 1325, 1350, 1356, 1358, 1361, 1390, 1396, 1402, 1406, 1412, 1415, 1422, 1429, 1432, 1434, 1438, 1457, 1464– 1466, 1473, 1488, 1490, 1491, 1494, 1498, 1499, 1515, 1518, 1525, 1528, 1532, 1533, 1536, 1560, 1561, 1567, 1568, 1573, 1598, 1601, 1605, 1608, 1611, 1616, 1632, 1634, 1636, 1638, 1643, 1644, 1687, 1722, 1756, 1796, 1837, 1877, 1919, 1954, 1987, 2036, 2037, 2085, 2088]. Society [1655, 1669, 1690, 1694, 1708, 1719, 1728, 1731, 1734, 1766, 1769, 1787, 1797, 1821, 1828, 1838, 1843, 1844, 1878, 1883, 1890, 1931, 1933, 1935, 1958, 1964, 1968, 1973, 1980, 1988, 1993, 2026, 2041, 2042, 2044, 2045, 2048, 2049, 2089, 2098, 2111, 2119, 2131, 2132, 2136, 2138, 2146, 2172, 2176, 2199, 2201, 2209, 2210, 2214, 2218, 2237, 2246, 2247, 2252, 2253, 2274, 2282, 2283, 2318, 2357, 2358, 2363, 2378, 2381, 2401, 2402, 2407, 2409, 2442, 2454, 2455, 2457, 2464, 2488, 2501, 2511, 2538]. Society [151, 215, 248, 316, 322, 351, 392, 429, 463, 522, 553, 598, 622, 624, 667, 682, 717, 750, 758, 786, 824, 859, 1098, 1138, 1212, 1255, 1312, 1521, 1595, 1826, 1861, 1942, 2117, 2130, 2164, 2291, 2307, 2309, 2319, 2325, 2328, 2334, 2338, 2346, 2385]. Socioeconomic [441, 846]. Sociotechnical [506]. Soft [1428, 1657]. Software [66, 100, 110, 190, 237, 239, 278, 311, 328, 343, 375, 417, 455, 542, 549, 563, 571, 586, 590, 625, 629, 632, 737, 770, 771, 801, 867, 870, 877, 879, 912, 939, 961, 1003, 1036, 1040, 1056, 1065, 1067, 1076, 1101, 1103, 1152, 1236, 1289, 1301, 1321, 1328, 1343, 1382, 1387, 1404, 1437, 1441, 1524, 1539, 1577, 1653, 1670, 1702, 1763, 1768, 1823, 1889, 1941, 1981, 1994, 2005, 2060, 2092, 2096, 2109, 2160, 2191, 2268, 2269, 2290, 2298, 2306, 2331, 2361, 2364, 2366, 2368, 2370, 2372, 2373, 2375, 2389, 2419, 2422, 2432, 2437, 2532, 2534]. Software- [237]. Software-Defined [590, 2290, 2422]. Software-Engineering [870]. Sokol [78]. Solid [1520]. Solution [875, 1587]. Solutionism [337]. Solutions [366, 1475, 1887, 2067, 2198, 2267]. Solvable [1739]. Some [1189, 1451]. Something [934]. Soon [236]. Sooner [1273]. Sora [1903, 2162, 2224]. Sort [1679]. Source [51, 113, 168, 177, 239, 309, 311, 341, 381, 455, 515, 571, 641, 705, 777, 974, 1036, 1102, 1191, 1265, 1336, 1413, 1545, 1625, 1706, 1853, 1941, 2005, 2107, 2160, 2191, 2335, 2420, 2438, 2517]. Sovereign [978]. Sovereignty [1240]. SP [966, 980, 1068]. Space [16, 743, 995, 1156, 1927, 2505]. Spaceborne [1224]. Spaces [140]. Sparse [1463]. Sparsity [720]. Spatial [1813]. Speak [896, 1651, 1693]. Speakers [1142]. Special [153, 214, 217, 250, 440, 833, 1980, 2009, 2138, 2201, 2237, 2274, 2309, 2334, 2381, 2446, 2476, 2506]. Specification [1232, 2152, 2419]. Spectrum [41, 100]. Speculative [957]. Speech [1142]. Speed [1930]. Spinal [43]. Spiral [115]. Spoofing [139]. Spooked [2094]. SQL [2453]. Squares [596]. Squirrel [89, 494]. Stability [1039]. Stablecoins [2344]. Stacked [236]. Stakes [1962]. Standard [739, 1661, 1971]. Standardization [106, 1867, 2347, 2380, 2441]. Standardizing [341]. Standards [46, 237, 379, 489, 549, 563, 601, 1275, 1387, 2117, 2195, 2275, 2529]. Standing [1228]. Stars [2387]. Start [1289]. Start-Ups [1289]. State [139, 209, 297, 315, 645, 817, 1190, 1520, 1783, 2149]. State-of-the-Art [1783]. Stateful [1069]. States [1943, 2019]. Statistics [134, 1037]. Status [1220, 1710]. Staying [1072]. Steady [1093]. STEAM [2466]. Steering [633]. STEM [2524]. Steps [2109]. Still [447, 1571]. Stirring [371]. Stock [2227]. Storage [111, 795, 1232, 1443, 1463, 1661, 1972, 2236, 2321, 2329]. Stores [444, 1660]. Storm [947]. Story [689, 1903, 2162, 2224]. Straggler [685]. Strangers [1284]. Strategic [1349, 2412]. Strategies [296, 819, 1274, 1867, 2130, 2270]. Strategy [946, 1556]. Stream [935, 1506]. Streamed [1862]. Streaming [199, 1862]. Strengthening [1585, 2180, 2185]. Strong [1453]. Stronger [873, 1072]. Structural [879, 2183]. Structure [1809]. Struggle [1420]. Struggling [604]. Student [701, 889, 1975]. Students [1775, 1806, 2112]. Studies [1326]. Study [632, 665, 730, 842, 965, 1421, 1740, 1810, 1859, 2000, 2105, 2257]. Subjects [1701]. Submarines [1957]. Subscribe [14, 44, 114, 137, 175, 203, 267, 277, 288, 310, 333, 384, 418, 443, 503, 582, 640, 691, 812, 847, 882, 980, 1034, 1068, 1106, 1272, 1298, 1648, 1700, 1820, 1855, 1892]. Subscription [404, 486, 745, 1150, 1179, 1221, 1345, 1379, 1449, 1482, 1511, 1610, 1976, 2106, 2151, 2226, 2301, 2341, 2374, 2418, 2432, 2461, 2513]. Succeed [1043]. Succeeding [793]. Success [1039, 2109]. Successors [772]. Suites [1548]. Sum [536]. Supercomputing [2382]. Supervisor [1105]. Supervisory [1220]. Supplies [1901]. Supply [15, 177, 377, 604, 639, 1101–1103, 1441, 1998]. Supply-Chain [604, 1441]. Support [694, 698, 1814, 1845, 2431, 2510]. Supported [1289, 1806]. Supporting [1934]. Supremacy [55]. Surface [731]. Surprise [1215]. Surprise-Inspired [1215]. Surveillance [26, 1436, 1739]. Survey [1260, 1765, 2142, 2391, 2416]. Susatainable [1851]. Sustainability [1235, 1497, 1630, 1867, 2236]. Sustainable [7, 54, 149, 242, 295, 376, 422, 457, 742, 781, 816, 844, 854, 887, 916, 1134, 1175, 1214, 1327, 1359, 1374, 1401, 1403, 1586, 1612, 1671, 1703, 1709, 1723, 1745, 1757, 1825, 1894, 1909, 1927, 1997, 2053, 2054, 2100, 2154, 2179, 2223, 2239, 2260, 2335, 2365, 2415, 2475, 2525]. Swatting [1307]. Swiss [2121]. Symmetry [1117]. Symposium [31, 211]. Sync [1546]. Synchrophasor [132]. Synchrophasor-Based [132]. Syndrome [1512]. Synergy [1072]. Synthesis [111, 1105]. Synthetic [2200]. System [136, 204, 302, 368, 444, 446, 475, 476, 481, 672, 819, 884, 1063, 1186, 1334, 1410, 1448, 1520, 1543, 1578, 1593, 1635, 1814, 2202, 2414, 2510]. System-Level [1334]. Systematic [693, 1262, 1762, 2050]. Systems [24, 34, 69, 73–75, 110, 132, 142, 201, 237, 264– 266, 269, 336, 362, 363, 366, 405, 409, 415, 432, 483, 536, 565, 581, 602, 632, 657, 658, 660, 663, 690, 694, 696, 707, 727, 768, 778, 852, 855, 866, 877, 879, 883, 902, 905, 927, 945, 963, 965, 1022, 1038, 1040, 1104, 1107, 1112, 1117, 1154, 1185, 1258, 1264, 1268, 1288, 1293, 1328, 1365, 1370, 1372, 1378, 1408, 1440, 1454, 1517, 1709, 1747, 1770, 1777, 1885, 1891, 1908, 1944, 1963, 1996, 2017, 2156, 2157, 2181, 2193, 2251, 2295, 2329, 2379, 2393, 2429]. Systems [143, 301, 326, 1334, 1502, 1889, 2023]. Systems-Centric [565]. Systems-Engineering [237]. Systems-on-Chip [1104].

Table

[3, 32, 62, 90, 124, 154, 186, 218, 251, 283, 323, 352, 393, 430, 464, 495, 523, 554, 579, 615, 650, 683, 718, 751, 787, 825, 860, 894, 925, 955, 985, 1020, 1054, 1091, 1129, 1164, 1204, 1246, 1280, 1315, 1354, 1394, 1426, 1461, 1495, 1529, 1564, 1602, 1639, 1685, 1720, 1754, 1794, 1835, 1875, 1916, 1952, 1985, 2032, 2033, 2082, 2128, 2170, 2207, 2244, 2280, 2315, 2354, 2398, 2451, 2498]. Tactics [946]. Tail [2058]. Tailoring [1324]. Taiwan [1809]. Takes [395, 1596]. Taking [908, 1740, 2191]. Tale [165]. Talents [1588]. Talk [1191]. Taming [1077]. Tapping [1145]. Taught [1513]. Taxonomies [1226]. Taxonomy [442, 516]. TBD [1066]. TC [820, 979, 1015, 1082, 1123, 1158, 1198]. TCDE [820]. TCPP [1123]. TCuARCH [979]. TD [630]. Teaching [549, 815, 972, 1348, 1368, 1577, 1934, 2342]. Teaching/Learning [1348]. Team [1261, 1453, 1477]. Teaming [744]. Teams [237]. Tech [223, 1666, 1806, 2250, 2272]. Tech-Enabled [1666]. Technical [13, 440, 820, 979, 1015, 1049, 1123, 1139, 1158, 1198, 1387, 1867]. Techniques [970, 1419, 1421, 1765]. Technological [701, 906]. Technologies [12, 13, 106, 240, 432, 531, 595, 992, 1409, 1543, 1614, 1783, 1809, 1924, 1927, 2223, 2265, 2353, 2397, 2462, 2473]. Technology [15, 195, 587, 592, 817, 903, 998, 1154, 1158, 1180, 1220, 1222, 1435, 1596, 1804, 1885, 1921, 1924, 1972, 2014, 2015, 2147, 2149, 2198, 2412]. Technomoral [1112]. Technoscapes [589]. Telecardiology [1542]. Telehealth [1508, 1510]. Telemedicine [909, 1504, 1540, 1542, 1543]. Telephony [516]. Tell [689]. Temporal [2112]. Ten [607]. Tensor [1463]. Test [875, 2157, 2453]. Testability [883, 2238]. Testbed [136]. Testing [100, 633, 672, 875, 1653, 1994, 2331, 2389]. Tests [875]. Text [1689, 2067]. Textual [970]. Their [508, 1340, 1365, 1806, 1899]. Them [1339, 1413, 2104]. Theology [1011]. Theory [768, 1105, 1294]. Therac [2104]. Therac-25 [2104]. There [447, 1264, 1477, 1635, 1732, 2108, 2150]. Things [43, 49, 172, 243, 367, 593, 595, 840, 866, 1104, 1183, 1185, 1288, 1337, 1483, 1583, 1621, 1624, 1676, 1707, 1786, 1868, 2008, 2267, 2290, 2423, 2474]. Things-Based [2290]. Things-Enabled [1183]. Things-Flavored [1786]. Things/Cyberphysical [1288]. Think [1273, 2227]. Thinking [768, 797, 1140, 2362, 2471]. Thinks [2508]. Thirsty [1691]. Thoughts [26, 98, 1286, 1760, 2459]. Thousand [1588]. Threat [299, 372, 1273, 1288, 1911, 1995, 2271, 2303]. Threat-Contested [2271]. Threatcasting [1656]. Threats [1765, 1807, 1965, 2257]. Three [1263]. TikTok [973, 1456]. Time [38, 96, 159, 192, 222, 231, 266, 287, 357, 369, 436, 500, 504, 558, 620, 666, 688, 757, 829, 899, 960, 972, 1027, 1097, 1170, 1230, 1238, 1252, 1319, 1398, 1431, 1456, 1468, 1501, 1507, 1535, 1570, 1607, 1646, 1689, 1725, 1759, 1799, 1840, 1880, 1921, 1956, 1990, 2039, 2091, 2099, 2134, 2175, 2212, 2213, 2249, 2285, 2321, 2361, 2404, 2424, 2453, 2458, 2503, 2537]. Time- [1230]. Tips [208]. To-Dos [766]. Today [636, 1822, 2304, 2504]. Together [508, 793, 1072]. Token [412, 910]. Tokens [1118, 1930]. Tolerance [537, 1889]. Tolerant [73, 1579]. Tomorrow [451, 601, 636, 1271, 1419]. Too [236]. Tools [311, 1189, 1421, 2040]. Top [80, 170, 1827, 2474, 2487]. Topic [1531, 2144]. Topics [652]. Touch [1737]. Toy [1903]. Trace [475, 838]. Traceability [639, 733].

Traces [230]. Tracing [842]. Track [976]. Tracking [1224]. Trading [134, 268]. Traditional [1139]. Traffic [334, 632, 2181]. Training [305, 993, 1566, 2400]. Trans [378]. Transaction [916, 1113, 1800, 1915, 1951]. Transactions [7, 9, 25, 81, 116, 119, 146, 149, 178, 180, 194, 200, 238, 242, 293, 295, 314, 376, 420, 422, 433, 457, 490, 530, 642, 742, 776, 779, 781, 814, 816, 849, 851, 854, 885, 887, 1134, 1141, 1148, 1175, 1177, 1182, 1199, 1214, 1217, 1219, 1327, 1329, 1331, 1374, 1377, 1383, 1586, 1594, 1612, 1615, 1663, 1671, 1673, 1703, 1723, 1757, 1780, 1789, 1825, 1830, 1834, 1851, 1857, 1860, 1874, 1894, 1896, 1997, 1999, 2004, 2029, 2053, 2054, 2056, 2057, 2077, 2081, 2102]. Transactions [52, 54, 68, 2100, 2154, 2158, 2165, 2179, 2182, 2186, 2235, 2239, 2314, 2327, 2359, 2365, 2367, 2415, 2421, 2426, 2445, 2475, 2478, 2480, 2521, 2523, 2525, 2528]. Transactive [269]. Transformation [166, 1108, 1344, 2393]. Transformations [441]. Transformative [396, 1378, 2103]. Transformer [2294]. Transformers [875]. Transforming [746, 2298, 2370, 2372]. Transit [591]. Transitions [2159]. Translators [1513]. Transparency [299, 335, 1263]. Transparent [641]. Transportation [432, 811, 1185]. Travelers [338]. Tree [2388]. Trends [106, 1120, 1143, 1542, 1676, 1771, 1783, 2120, 2232, 2509]. Triangle [2377]. Tribal [1226]. Tribulation [1456]. Tribute [1450, 1969, 2074]. Trick [1822]. Trips [2466]. Trojan [1102]. TRON [2162, 2224]. Trust [98, 100, 108, 147, 197, 312, 693, 696, 729, 830, 1003, 1046, 1370, 1674, 2187]. Trusted [104, 364, 481]. Trusting [744, 1001]. Trustworthiness [417, 1367]. Trustworthy [1147, 1254, 1256– 1258, 1326, 1330, 1360, 1364, 1368, 1725, 1730, 1733, 1735, 1736, 1965, 2375, 2532]. Truth [98, 1477, 1659]. TSUSC [1080]. Tunnel [1512]. Turns [2336]. Twenty [2066]. Twin [1063, 1813, 2474]. Twinlets [2522]. Twins [411, 518, 1001, 1062, 1216, 1414, 1522, 2025]. Two [228, 693, 848, 1651, 1693]. Type [417, 935, 1733]. Type-2 [1733]. Typology [1192].

U [206]. U.S. [458, 2015, 2380]. Ubiquities [1647]. Ubiquitous [1033, 1271]. Ukraine [1006, 1146]. Ukrainian [1932]. UL [1375]. Ultimate [105, 2191]. Unanticipated [506]. Uncertainty [606, 1301, 2273]. Unconstrained [1847]. Unconventional [997, 2337]. Uncovering [2423]. Uncrewed [1440, 2193]. Undecidable [2527]. Undergraduate [1577]. Understanding [232, 409, 547, 805, 1083, 1497, 1668]. UNESCO [1698]. Unethical [1701]. Unfulfilling [1635]. Unified [1288]. Union [1961, 2380]. Unique [1413]. United [1943, 2019]. Units [139, 1478]. Universal [411, 1661, 2153]. Universities [1388, 2487]. University [701, 1108, 1207, 1241, 2196, 2234]. Unlearn [1738]. Unlearning [1738, 2142, 2150]. Unleashing [1624]. Unlimited [111, 1554]. Unlock [2109, 2173]. Unlocking [111, 1868]. Unravel [2317]. Unraveling [1812, 2000]. Unrest [2135]. Unveiling [2257]. Upholding [2067]. Ups [1289]. Urban [505, 564, 1448, 2097]. Urgent [1346, 2423, 2529]. Usability [694, 1265, 1886]. Usability-Aware [1886]. Usable [47, 2001]. Usage [1186, 1891]. Use [45, 694, 836, 1348, 1365, 1823, 1859, 2161, 2222, 2384, 2487]. Used [2007]. User [693, 694, 852, 998, 1143, 1265, 1479, 1624]. Users [1841, 2058]. Uses [506]. Ushering [2515]. Using [108, 301, 444, 855, 987, 1151, 1187, 1226, 1367, 1512, 1815, 2112, 2271, 2337]. Utilization [1480]. Utopianism [337].

V [246, 1048, 1079, 1911, 1947]. V-Model [1079]. V2I [432]. V2V [432]. Validating [711]. Validation [711, 733, 855]. Validity [2467]. Value [10, 296, 596, 1899].

Value-Added [10]. Varieties [337]. Vault [2121]. VD [1545]. VD-HEN [1545]. Vector [516]. Vectors [1911]. Vehicle [625, 629, 632, 1375, 1813, 1814]. Vehicle-to-Vehicle [1814]. Vehicles [334, 633, 764, 811, 1063, 1290, 1814, 2046, 2050, 2052]. Vehicular [627]. Velocity [1853]. Vendor [113]. Venture [1799, 1804, 1806, 1812]. Verifiability [1258]. Verification [6, 657, 855, 1454, 1576, 1764, 2017, 2020, 2419]. Versus [13, 1428, 1777, 2519]. Vertical [266]. Very [1198]. via [139, 537, 572, 1778]. Victims [1556]. Video [199, 232, 1583, 1739]. Video-Understanding [232]. View [1110, 1330, 1591]. Views [1062]. VIII [678, 1489, 2444]. Violable [18]. Violence [2135]. Viral [699]. Virtual [41, 105, 256, 727, 789, 896, 1119, 1618, 1620, 1813, 2466, 2468]. Virtualization [1104, 1704]. Virtually [111]. Vision [340, 504, 534, 902, 1372, 1475, 2103, 2294]. Visual [1566]. Visualization [987, 1049]. Visualizing [475, 1480, 2254]. Viva [235]. VoIP [1307]. Voltage [630]. Voltage-Prediction [630]. Volunteer [121, 799, 1111, 1239, 1323, 1412, 1438, 1473, 1515, 1561, 1632, 1682, 1717, 1785, 1866, 1902, 1937, 1978, 2116, 2338, 2383, 2488, 2518]. Voted [388, 1196]. Voting [817]. Vs [2481, 2503]. Vulnerabilities [516, 1905]. Vulnerability [312, 1545].

Waiting [542]. Wake [2339]. Wake-Up [2339]. Walking [1261]. Walter [1015]. Want [1929]. Wanted [838]. Wants [2387]. War [1420, 1436, 2040, 2135]. Warcraft [2424]. Warfare [1979]. Warming [1029]. Warning [2017]. Wars [1226]. Was [1932]. Wasabi [1256]. Washington [1207]. Waste [2177]. Watch [196, 224, 907, 948, 1013]. Water [1299]. Waterfall [1349]. Watts [964, 988, 1024, 1343, 1371, 1404, 1437, 1476]. Wave [707, 1373]. Way [278, 515, 571, 2262, 2489]. Ways [1037]. Weapons [1014, 1067]. Wearable [2520]. Wearables [354]. Web [595, 1886, 2417]. Website [1296]. Week [61, 109, 131, 1053, 1391, 1393, 1425, 1460]. Weight [20]. Well [1823]. Wen [979]. Wen-Mei [979]. Went [2060, 2434]. West [337]. Wheels [37]. Where [271, 413, 621, 803, 976, 1542, 2268]. Which [296]. While [908]. Who [510, 1702, 1733, 2387]. Wholesale [1710]. Whom [335]. Wide [132]. WiF0 [599]. Will [451, 478, 508, 782, 993, 1011, 1206, 1269, 2019]. Winners [889]. Winning [1475]. Winter [2483]. Wireless [2427]. Wise [1453]. Wiser [991]. Without [1674]. Women [102, 141, 1191, 1264]. Wood [2322]. Word [1271]. Words [971]. Work [592, 1151, 1216, 1453, 1666, 1807, 1899]. Workers [1581]. Workflow [872]. Workflows [1680, 1926]. Workforce [305, 1218, 2010, 2221]. Working [819]. Workloads [1888]. Works [713]. World [46, 206, 291, 405, 540, 772, 809, 864, 971, 1206, 1609, 1974, 2055, 2424]. Worlds [592, 1119]. Worth [2191]. Write [50]. Writing [675, 1503, 1668, 1823, 2324]. Wrong [2060, 2434].

Xplainable [2333].

Year [78, 396, 440, 758, 2084, 2085, 2149]. Years [8, 36, 65, 95, 128, 158, 191, 209, 221, 241, 259, 286, 327, 343, 356, 401, 435, 469, 499, 528, 531, 557, 584, 619, 654, 687, 723, 724, 756, 758, 792, 828, 863, 898, 929, 959, 990, 1026, 1058, 1096, 1133, 1169, 1209, 1251, 1283, 1318, 1357, 1397, 1430, 1467, 1500, 1534, 1569, 1606, 1645, 1688, 1724, 1758, 1798, 1839, 1841, 1879, 1920, 1955, 1989, 2038, 2090, 2133, 2145, 2174, 2211, 2248, 2284, 2320, 2360, 2403, 2424, 2456, 2490, 2502]. Yes [675]. Yesterday [673]. Yield [1658].

Young [758, 1806].

Z [2155].  Z80 [1969].  Zero [312,

729, 830, 1046, 1674, 1737, 1777, 2410]. Zero-Knowledge [2410]. Zero-Touch [1737]. Zero-Trust [830].

References

Anonymous:2020:FCa

[1] Anonymous. Front cover. Computer, 53(1):C1, January 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Ca

[2] Anonymous. COMPSAC 2020. Computer, 53(1):C2, January 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCa

[3] Anonymous. Table of contents. Computer, 53(1):1–2, January 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Ma

[4] Anonymous. Masthead. Computer, 53(1):3, January 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSa

[5] Anonymous. Computer highlights Society magazines. Computer, 53(1): 4–6, January 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ricanek:2020:KV

[6] K. Ricanek. Kinship verification. Computer, 53(1):7–8, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITSa

[7] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 53 (1):8, January 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAa

[8] E. Neuhold. 50 & 25 years ago. Computer, 53(1):9–12, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITCa

[9] Anonymous. IEEE Transactions on Computing. Computer, 53(1):12, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:VAG

[10] J. Voas. A value-added gem. Computer, 53(1):13–14, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CSS

[11] Anonymous. Computer Society social networking. Computer, 53(1): 15, January 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:ODT

[12] J. Voas. Outlooks on diverse technologies. Computer, 53(1):16–17, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2020:BBF

[13] N. Kshetri. Blockchain-based financial technologies and cryptocurrencies for low-income people: Technical potential versus practical reality. Computer, 53 (1):18–29, January 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:SPSa

[14] Anonymous. Security and Privacy subscribe. Computer, 53(1):29, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paulsen:2020:FIO

[15] C. Paulsen. The future of IT operational technology supply chains. Computer, 53(1):30–36, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:2020:ISC

[16] R. Kuhn, R. N. Kacker, Y. Lei, and D. Simos. Input space coverage matters. Computer, 53(1):37–44, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peitso:2020:PIS

[17] L. E. Peitso and J. B. Michael. The promise of interactive shared augmented reality. Computer, 53(1):45– 52, January 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singh:2020:CGV

[18] M. P. Singh and A. K. Chopra. Computational governance and violable contracts for blockchain applications. Computer, 53(1):53–62, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:PCC

[19] Anonymous. Pervasive Computing call for articles. Computer, 53(1):62, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hosio:2020:CPW

[20] S. J. Hosio, N. van Berkel, J. Oppenlaender, and J. Goncalves. Crowdsourcing personalized weight loss diets. Computer, 53(1):63–71, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2020:CLC

[21] H. Berghel. A critical look at the 2019 college admissions scandal? Computer, 53(1):72–77, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CSB

[22] Anonymous. Computer Society benefits. Computer, 53(1):77, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:MM

[23] D. A. Grier. The manual Memex. Computer, 53(1):78–80, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lueder:2020:FPS

[24] A. Lueder. Flexibility in production systems by exploiting cyberphysical systems. Computer, 53(1):81– 85, January 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITBa

[25] Anonymous. IEEE Transactions on Big Data. Computer, 53(1):85, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2020:TGD

[26] N. Kshetri and J. Voas. Thoughts on General Data Protection Regulation and online human surveillance. Computer, 53(1):86–90, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCa

[27] Anonymous. CS connection. Computer, 53(1):91–92, January 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSa

[28] Anonymous. IEEE Computer Society. Computer, 53(1):C3, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CSJ

[29] Anonymous. Computer Society Jobs Board. Computer, 53(1):C4, January 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCb

[30] Anonymous. Front cover. Computer, 53(2):C1, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:HSa

[31] Anonymous. HOST Symposium 2020. Computer, 53(2):C2, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCb

[32] Anonymous. Table of contents. Computer, 53(2):1–2, February 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Mb

[33] Anonymous. Masthead. Computer, 53 (2):3, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peng:2020:HDA

[34] X. Peng. Helping developers analyze and debug industrial microservice systems. Computer, 53(2):4–5, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSb

[35] Anonymous. Computer highlights Society magazines. Computer, 53(2): 6–8, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAb

[36] E. Neuhold. 50 & 25 years ago. Computer, 53(2):9–11, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:RW

[37] J. Voas. Reinventing wheels. Computer, 53(2):12–13, February 2020.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2020:CTTa

[38] E. Akleman. Computing through time: Flying. Computer, 53(2):13, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2020:AAI

[39] D. Milojicic. Accelerators for artificial intelligence and high-performance computing. Computer, 53(2):14–22, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Helal:2020:DHC

[40] S. Helal. Digital health: E-coaching and remote monitoring. Computer, 53 (2):23–25, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhao:2020:VAB

[41] W. Zhao, X. Liu, T. Qiu, and X. Luo. Virtual avatar-based life coaching for children with autism spectrum disorder. Computer, 53(2):26–34, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ali:2020:IMP

[42] T. Ali, J. Hussain, M. B. Amin, M. Hussain, U. Akhtar, W. A. Khan, S. Lee, B. H. Kang, M. Hussain, M. Afzal, H. W. Yu, U. U. Rehman, H. Han, J. Y. Choi, and A. Jamshed. The intelligent medical platform: A novel dialogue-based platform for health-care services. Computer, 53(2): 35–45, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Manocha:2020:ITF

[43] A. Manocha, R. Singh, and P. Verma. An Internet of Things fog-assisted sleep-deprivation prediction framework for spinal cord injury patients. Computer, 53(2):46–56, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:SPSb

[44] Anonymous. Security and Privacy subscribe. Computer, 53(2):56, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dhir:2020:UBF

[45] S. Dhir and S. K. A. Devi. The use of biometric fingerprints for onthe-fly digital signing of documents. Computer, 53(2):57–67, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mariani:2020:ICS

[46] R. Mariani. IEEE Computer Society Standards Activities Board: Adapting to a changing world. Computer, 53(2): 68–70, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Theofanos:2020:USO

[47] M. Theofanos. Is usable security an oxymoron? Computer, 53(2):71– 74, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Cb

[48] Anonymous. Compsac 2020. Computer, 53(2):75, February 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barker:2020:RPI

[49] O. Barker. Realizing the promise of the Internet of Things in smart buildings. Computer, 53(2):76–79, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2020:ACQ

[50] M. Campbell. Automated coding: The quest to develop programs that write programs. Computer, 53(2): 80–82, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gustavsson:2020:MOS

[51] T. Gustavsson. Managing the open source dependency. Computer, 53(2): 83–87, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TBD

[52] Anonymous. Transactions on Big Data. Computer, 53(2):87, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:RC

[53] D. A. Grier. Reflecting on CAP. Computer, 53(2):88–90, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TSC

[54] Anonymous. Transactions on Sustainable Computing. Computer, 53 (2):90, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Debenedictis:2020:BQS

[55] E. P. Debenedictis. Beyond quantum supremacy. Computer, 53(2): 91–94, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Cd

[56] Anonymous. Correction. Computer, 53 (2):94, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCb

[57] Anonymous. CS connection. Computer, 53(2):95–96, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSb

[58] Anonymous. IEEE Computer Society. Computer, 53(2):C3, February 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSc

[59] Anonymous. IEEE Computer Society Jobs Board. Computer, 53(2): C4, February 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCc

[60] Anonymous. Front cover. Computer, 53(3):C1, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IQWa

[61] Anonymous. IEEE Quantum Week 2020. Computer, 53(3):C2, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCc

[62] Anonymous. Table of contents. Computer, 53(3):1–2, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Mc

[63] Anonymous. Masthead. Computer, 53(3):3, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSc

[64] Anonymous. Computer highlights Society magazines. Computer, 53(3): 4–6, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAc

[65] E. Neuhold. 50 & 25 years ago. Computer, 53(3):7–10, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CAS

[66] Anonymous. Call for articles Software. Computer, 53(3):10, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xue:2020:DAM

[67] D. Xue, L. Huang, C. Li, and C. Wu. Dapper: An adaptive manager for large-capacity persistent memory. Computer, 53(3):11–13, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IT

[68] Anonymous. IEEE Transactions. Computer, 53(3):13, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolf:2020:DDA

[69] M. Wolf. Distributed data analysis and reliable operation of cyberphysical systems. Computer, 53(3):14–15, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fukuda:2020:ABC

[70] M. Fukuda, C. Gordon, U. Mert, and M. Sell. An agent-based computational framework for distributed data analysis. Computer, 53(3):16–25, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2020:COD

[71] H. Lin, H. Alemzadeh, Z. Kalbarczyk, and R. Iyer. Challenges and opportunities in the detection of safety-critical cyberphysical attacks. Computer, 53 (3):26–37, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISPa

[72] Anonymous. IEEE Security & Privacy. Computer, 53(3):37, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barranco:2020:AHR

[73] M. Barranco, S. Derasevic, and J. Proenza. An architecture for highly reliable fault-tolerant adaptive distributed embedded systems. Computer, 53(3):38–46, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Javed:2020:PDI

[74] Y. Javed, M. Felemban, T. Shawly, J. Kobes, and A. Ghafoor. A partitiondriven integrated security architecture for cyberphysical systems. Computer, 53(3):47–56, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koutsoukos:2020:SSS

[75] X. Koutsoukos. Systems science of secure and resilient cyberphysical systems. Computer, 53(3):57–61, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSd

[76] Anonymous. IEEE Computer Society. Computer, 53(3):61, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2020:SY

[77] N. Kshetri and J. Voas. 5G, security, and you. Computer, 53(3): 62–66, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2020:SHY

[78] H. Berghel. The Sokol hoax: A 25year retrospective. Computer, 53(3):

67–72, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Ce

[79] Anonymous. COMPSAC 2020. Computer, 53(3):73, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:TI

[80] D. A. Grier. Top influence. Computer, 53(3):74–76, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITCb

[81] Anonymous. IEEE Transactions on Computers. Computer, 53(3):76, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2020:NPA

[82] H. Berghel. New perspectives on (anti)Social media. Computer, 53(3): 77–82, March 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCc

[83] Anonymous. CS connection. Computer, 53(3):83–84, March 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolf:2020:MDD

[84] M. Wolf. In memoriam: Dan Dobberpuhl. Computer, 53(3):83, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IP

[85] Anonymous. IT Professional. Computer, 53(3):84, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSe

[86] Anonymous. IEEE Computer Society. Computer, 53(3):C3, March 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICJ

[87] Anonymous. IEEE CS Job Board. Computer, 53(3):C4, March 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCd

[88] Anonymous. Front cover. Computer, 53(4):C1, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CNA

[89] Anonymous. Call for nominations: AAAI Squirrel AI Award for Artificial Intelligence for the Benefit of Humanity. Computer, 53(4):C2, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCd

[90] Anonymous. Table of contents. Computer, 53(4):1–2, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Md

[91] Anonymous. Masthead. Computer, 53(4):3, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSd

[92] Anonymous. Computer highlights Society magazines. Computer, 53(4):4–6, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sorin:2020:CAO

[93] D. J. Sorin. Computer architecture for orbital edge computing. Computer, 53(4):7–8, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IMMa

[94] Anonymous. IEEE Multimedia magazine. Computer, 53(4):8, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAd

[95] E. Neuhold. 50 & 25 years ago. Computer, 53(4):9–12, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2020:CTTb

[96] E. Akleman. Computing through time: Artificial life. Computer, 53 (4):10, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IIC

[97] Anonymous. IEEE 21st International Conference of Information Reuse and Integration for Data Science (IEEE IRI 2020). Computer, 53(4):12, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:TCT

[98] J. Voas. Thoughts on complexity, trust, and truth. Computer, 53(4): 13–14, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICGa

[99] Anonymous. IEEE Computer Graphics & Applications. Computer, 53 (4):14, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:2020:NSS

[100] R. Kuhn and M. S. Raunak. Narrow spectrum software testing addressing complexity and trust. Computer, 53(4):15–18, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:Ia

[101] J. Voas. In this issue. Computer, 53(4):17, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:HWCa

[102] Anonymous. Honoring women in computing. Computer, 53(4):19, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hodges:2020:PCK

[103] S. Hodges, S. Sentance, J. Finney, and T. Ball. Physical computing: A key element of modern computer science education. Computer, 53(4):20–30, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmed:2020:BCI

[104] M. Ahmed and A. K. Pathan. Blockchain: Can it be trusted? Computer, 53(4):31–35, April 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lamberti:2020:IVR

[105] F. Lamberti, A. Cannavo, and P. Montuschi. Is immersive virtual reality the ultimate interface for 3D animators? Computer, 53(4):36–45, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wei:2020:TSM

[106] S. Wei, Y. Ma, R. Li, and L. Hu. Toward smart manufacturing: Key technologies and trends driving standardization. Computer, 53(4):46–50, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPCa

[107] Anonymous. IT Professional call for articles. Computer, 53(4):50, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2020:OTE

[108] D. Drusinsky and J. B. Michael. Obtaining trust in executable derivatives using crowdsourced critiques with blind signatures. Computer, 53(4):51–56, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IQWb

[109] Anonymous. IEEE Quantum Week. Computer, 53(4):57, April 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kubitza:2020:DIS

[110] T. Kubitza, P. Bader, M. Mogerle, and A. Schmidt. Developing IoT systems: It’s all about the software. Computer, 53(4):58–62, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2020:DDS

[111] M. Campbell. DNA data storage: Automated DNA synthesis and sequencing are key to unlocking virtually unlimited data storage. Computer, 53 (4):63–67, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPCb

[112] Anonymous. IEEE Pervasive Computing call for articles. Computer, 53(4):67, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2020:SVO

[113] D. Riehle. Single-vendor open source firms. Computer, 53(4):68–72, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISPb

[114] Anonymous. IEEE Security & Privacy subscribe. Computer, 53(4):72, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:OS

[115] D. A. Grier. The outward spiral. Computer, 53(4):73–76, April 2020.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITBb

[116] Anonymous. IEEE Transactions on Big Data. Computer, 53(4):76, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeBenedictis:2020:QM

[117] E. P. DeBenedictis. The quantum moonshot. Computer, 53(4):77–80, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCd

[118] Anonymous. CS connection. Computer, 53(4):81–84, April 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITCc

[119] Anonymous. IEEE Transactions on Computers. Computer, 53(4):84, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSf

[120] Anonymous. IEEE Computer Society. Computer, 53(4):C3, April 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSg

[121] Anonymous. IEEE Computer Society Volunteer Service awards. Computer, 53(4):C4, April 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCe

[122] Anonymous. Front cover. Computer, 53(5):C1, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSh

[123] Anonymous. IEEE Computer Society member benefits. Computer, 53 (5):C2, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCe

[124] Anonymous. Table of contents. Computer, 53(5):1–2, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Me

[125] Anonymous. Masthead. Computer, 53(5):3, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:EC

[126] Anonymous. Elsewhere in the CS. Computer, 53(5):4–6, May 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IMMb

[127] Anonymous. IEEE Multimedia magazine. Computer, 53(5):6, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAe

[128] E. Neuhold. 50 & 25 years ago. Computer, 53(5):7–9, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSi

[129] Anonymous. IEEE Computer Society call for papers. Computer, 53(5):9, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Konstantinou:2020:CSG

[130] C. Konstantinou and S. P. Mohanty. Cybersecurity for the smart grid. Computer, 53(5):10–12, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IQWc

[131] Anonymous. IEEE Quantum Week. Computer, 53(5):13, May 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chawla:2020:DSR

[132] A. Chawla, P. Agrawal, A. Singh, B. K. Panigrahi, K. Paul, and B. Bhalja. Denial-of-service resilient frameworks for synchrophasor-based wide area monitoring systems. Computer, 53 (5):14–24, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICGb

[133] Anonymous. IEEE Computer Graphics & Applications magazine. Computer, 53(5):24, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:2020:PPO

[134] D. Smith, P. Wang, M. Ding, J. Chan, B. Spak, X. Guan, P. Tyler, T. Rakotoarivelo, Z. Lin, and T. Abbasi. Privacy-preserved optimal energy trading, statistics, and forecasting for a neighborhood area network. Computer, 53(5):25–34, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shrestha:2020:DCE

[135] B. Shrestha and H. Lin. Data-centric edge computing to defend power grids against IoT-based attacks. Computer, 53(5):35–43, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khan:2020:CPS

[136] M. M. S. Khan, A. Palomino, J. Brugman, J. Giraldo, S. K. Kasera, and M. Parvania. The Cyberphysical Power System Resilience Testbed: Architecture and applications. Computer, 53 (5):44–54, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISPc

[137] Anonymous. IEEE Security & Privacy magazine subscribe. Computer, 53(5):54, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barreto:2020:AEM

[138] C. Barreto, H. Neema, and X. Koutsoukos. Attacking electricity markets through IoT devices. Computer, 53 (5):55–62, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xie:2020:SDG

[139] J. Xie and A. P. S. Meliopoulos. Sensitive detection of GPS spoofing attack in phasor measurement units via quasidynamic state estimation. Computer, 53(5):63–72, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Helal:2020:MAE

[140] S. Helal. The monkey, the ant, and the elephant: Addressing safety in smart spaces. Computer, 53(5):73–76, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:HWCb

[141] Anonymous. Honoring women in computing call for nominations. Computer, 53(5):77, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schneier:2020:AML

[142] B. Schneier. Attacking machine learning systems. Computer, 53(5):78–80, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IIS

[143] Anonymous. IEEE Intelligent Systems magazine. Computer, 53(5):80, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2020:OQF

[144] J. B. Michael, G. W. Dinolt, and D. Drusinsky. Open questions in formal methods. Computer, 53(5):81–84, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2020:EHR

[145] H. Berghel. The Equifax hack revisited and repurposed. Computer, 53 (5):85–90, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITBc

[146] Anonymous. IEEE Transactions on Big Data. Computer, 53(5):90, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:MT

[147] D. A. Grier. Mechanical trust. Computer, 53(5):91–92, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCe

[148] Anonymous. CS connection. Computer, 53(5):93–96, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITSb

[149] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 53(5):96, May 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSj

[150] Anonymous. IEEE Computer Society. Computer, 53(5):C3, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:OJC

[151] Anonymous. Open Journal of the Computer Society. Computer, 53(5):C4, May 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCf

[152] Anonymous. Front cover. Computer, 53(6):C1, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCSa

[153] Anonymous. Computer call for special issue proposals. Computer, 53 (6):C2, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCf

[154] Anonymous. Table of contents. Computer, 53(6):1–2, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Mf

[155] Anonymous. Masthead. Computer, 53(6):3, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSe

[156] Anonymous. Computer highlights Society magazines. Computer, 53(6):4–6, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Deng:2020:LPM

[157] X. Deng, T. Xiao, and K. Zhu. Learn to play maximum revenue auction. Computer, 53(6):7–8, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAf

[158] E. Neuhold. 50 & 25 years ago. Computer, 53(6):9–12, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2020:CTTc

[159] E. Akleman. Computing through time. Computer, 53(6):11, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:BRP

[160] Anonymous. BPCNet resources portal. Computer, 53(6):12, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:C

[161] J. Voas and P. Laplante. Cyberpandemics. Computer, 53(6):13–15, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:Ib

[162] J. Voas. In this issue. Computer, 53(6):14, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IAHa

[163] Anonymous. IEEE Annals of the History of Computing. Computer, 53 (6):15, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2020:C

[164] J. B. Michael, R. Kuhn, and J. Voas. Cyberthreats in 2025. Computer, 53 (6):16–27, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parhami:2020:RIC

[165] B. Parhami. Reliability inversion: A cautionary tale. Computer, 53 (6):28–33, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Demir:2020:ETG

[166] M. Demir, O. Turetken, and A. Mashatan. An enterprise transformation guide for the inevitable blockchain disruption. Computer, 53(6):34–43, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bettini:2020:PRS

[167] C. Bettini, S. Kanhere, M. Langheinrich, A. Misra, and D. Reinhardt. Is privacy regulation slowing down research on pervasive computing? Computer, 53(6):44–52, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robson:2020:ISO

[168] R. Robson. IEEE SA Open: Engaging industry, academia, and researchers in open source development. Computer, 53(6):53–56, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICGc

[169] Anonymous. IEEE Computer Graphics And Applications magazine. Computer, 53(6):56, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McGraw:2020:TRM

[170] G. McGraw, R. Bonett, V. Shepardson, and H. Figueroa. The top 10 risks of machine learning security. Computer, 53(6):57–61, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPCc

[171] Anonymous. IT Professional call for articles. Computer, 53(6):61, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:2020:SIT

[172] S. W. Smith. Securing the Internet of Things: An ongoing challenge. Computer, 53(6):62–66, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPCd

[173] Anonymous. IEEE Pervasive Computing call for articles. Computer, 53(6):66, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Caleffi:2020:RQI

[174] M. Caleffi, D. Chandra, D. Cuomo, S. Hassanpour, and A. S. Cacciapuoti. The rise of the quantum Internet. Computer, 53(6):67–72, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISPd

[175] Anonymous. IEEE Security & Privacy magazine subscribe. Computer, 53(6):72, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2020:IAA

[176] M. Campbell. Inference acceleration: Adding brawn to the brains. Computer, 53(6):73–76, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harutyunyan:2020:MYO

[177] N. Harutyunyan. Managing your open source supply chain — why and how? Computer, 53(6):77–81, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITBd

[178] Anonymous. IEEE Transactions on Big Data. Computer, 53(6):81, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:ANN

[179] D. A. Grier. An ambassador for neural networks. Computer, 53(6):82–84, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITCd

[180] Anonymous. IEEE Transactions on Computers call for papers. Computer, 53(6):84, June 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCf

[181] Anonymous. CS connection. Computer, 53(6):C3, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSk

[182] Anonymous. IEEE Computer Society. Computer, 53(6):C3, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICAa

[183] Anonymous. IEEE cybersecurity analysis. Computer, 53(6):C4, June 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCg

[184] Anonymous. Front cover. Computer, 53(7):C1, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSl

[185] Anonymous. IEEE Computer Society member benefits. Computer, 53 (7):C2, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCg

[186] Anonymous. Table of contents. Computer, 53(7):1–2, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Mg

[187] Anonymous. Masthead. Computer, 53(7):3, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSf

[188] Anonymous. Computer highlights Society magazines. Computer, 53(7):4–6, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Andre:2020:AIA

[189] E. Andre. Is artificial intelligence able to help with pain assessment? Computer, 53(7):7–8, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISMa

[190] Anonymous. IEEE Software magazine call for articles. Computer, 53(7):8, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAg

[191] E. Neuhold. 50 & 25 years ago. Computer, 53(7):9–11, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2020:CTTd

[192] E. Akleman. Computing through time: Social distancing. Computer, 53(7):11, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:PM

[193] J. Voas. The “Patching” mentality. Computer, 53(7):12–13, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITCe

[194] Anonymous. IEEE Transactions on Computers call for papers. Computer, 53(7):13, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nandakumar:2020:BTM

[195] Karthik Nandakumar, Nalini Ratha, Sharath Pankanti, Alex Pentland, and Maurice Herlihy. Blockchain: From technology to marketplaces. Computer, 53(7):14–18, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:AWCa

[196] Anonymous. AI’s 10 to watch call for nominations. Computer, 53(7):19, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cao:2020:BAA

[197] Sean Cao, Lin William Cong, Meng Han, Qixuan Hou, and Baozhong Yang. Blockchain architecture for auditing automation and trust building in public markets. Computer, 53(7):20–28, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Saxena:2020:PBC

[198] Neetesh Saxena, Ieuan Thomas, Prosante Gope, Pete Burnap, and Neeraj Kumar. PharmaCrypt: Blockchain for critical pharmaceutical industry to counterfeit drugs. Computer, 53(7):29–44, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barman:2020:BVS

[199] Nabajeet Barman, Deepak G C, and Maria G. Martini. Blockchain for video streaming: Opportunities, challenges, and open issues. Computer, 53(7):45–56, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITBe

[200] Anonymous. IEEE Transactions on Big Data. Computer, 53(7):56, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Biswas:2020:BHC

[201] Sujit Biswas, Kashif Sharif, Fan Li, and Saraju Mohanty. Blockchain for ehealth-care systems: Easier said than done. Computer, 53(7):57–67, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:PPD

[202] D. A. Grier. A prize, a prediction, and a drama. Computer, 53(7):68–70, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISS

[203] Anonymous. IEEE subscribe to Security & Privacy. Computer, 53(7):70, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhattacharyya:2020:RCH

[204] S. S. Bhattacharyya and M. C. Wolf. Research challenges for heterogeneous cyberphysical system design. Computer, 53(7):71–75, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPCe

[205] Anonymous. IEEE Pervasive Computing call for articles. Computer, 53(7):75, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2020:CUH

[206] P. Laplante. Contactless U: Higher education in the postcoronavirus world. Computer, 53(7):76–79, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IIPa

[207] Anonymous. IEEE IT Professional call for articles. Computer, 53(7):79, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2020:CRE

[208] H. Berghel, R. N. Charette, E. G. Happ, and J. L. King. It’s on: COVID-19, risk ecology, and preparedness tips. Computer, 53(7):80–87, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Osterweil:2020:YDDa

[209] E. Osterweil, A. Stavrou, and L. Zhang. 21 years of distributed denial-of service: Current state of affairs. Computer, 53 (7):88–92, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2020:CAP

[210] H. Berghel. A collapsing academy, part 1. Computer, 53(7):93–98, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:HSb

[211] Anonymous. HOST 2020 Symposium. Computer, 53(7):99, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCg

[212] Anonymous. CS connection. Computer, 53(7):100–102, July 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSm

[213] Anonymous. IEEE Computer Society. Computer, 53(7):102, July 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCSb

[214] Anonymous. Computer call for special issue proposals. Computer, 53 (7):C3, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IOJa

[215] Anonymous. IEEE Open Journal of the Computer Society. Computer, 53 (7):C4, July 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCh

[216] Anonymous. Front cover. Computer, 53(8):C1, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCSc

[217] Anonymous. Computer call for special issue proposals. Computer, 53 (8):C2, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCh

[218] Anonymous. Table of contents. Computer, 53(8):1–2, August 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Mh

[219] Anonymous. Masthead. Computer, 53(8):3, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSg

[220] Anonymous. Computer highlights Society magazines. Computer, 53(8): 4–6, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAh

[221] E. Neuhold. 50 & 25 years ago. Computer, 53(8):7–9, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2020:CTTe

[222] E. Akleman. Computing through time. Computer, 53(8):9, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2020:CMB

[223] N. Kshetri. COVID-19 meets big tech. Computer, 53(8):10–13, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:AWCb

[224] Anonymous. AI’s 10 to watch call for nominations. Computer, 53(8):13, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:2020:HHA

[225] L. Chen, H. Ning, C. D. Nugent, and Z. Yu. Hybrid human-artificial intelligence. Computer, 53(8):14–17, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSn

[226] Anonymous. IEEE Computer Society call for papers. Computer, 53 (8):17, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akata:2020:RAH

[227] Z. Akata, D. Balliet, M. de Rijke, F. Dignum, V. Dignum, G. Eiben, A. Fokkens, D. Grossi, K. Hindriks, H. Hoos, H. Hung, C. Jonker, C. Monz, M. Neerincx, F. Oliehoek, H. Prakken, S. Schlobach, L. van der Gaag, F. van Harmelen, H. van Hoof, B. van Riemsdijk, A. van Wynsberghe, R. Verbrugge, B. Verheij, P. Vossen, and M. Welling. A research agenda for hybrid intelligence: Augmenting human intellect with collaborative, adaptive, responsible, and explainable artificial intelligence. Computer, 53(8): 18–28, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wenskovitch:2020:IAI

[228] J. Wenskovitch and C. North. Interactive artificial intelligence: Designing for the “Two black boxes” problem. Computer, 53(8):29–39, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPCf

[229] Anonymous. IEEE Pervasive Computing. Computer, 53(8):39, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feng:2020:MIM

[230] H. Feng, J. Zhou, W. Lin, Y. Zhang, and Z. Qu. Multiple-input, multilayerperception-based classification of traces from side-channel attacks. Computer, 53(8):40–48, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pescetelli:2020:BRT

[231] N. Pescetelli, M. Cebrian, and I. Rahwan. BeeMe: Real-time Internet control of situated human agents. Computer, 53(8):49–58, August 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hu:2020:EHV

[232] B. Hu, K. Guo, J. Zhang, S. Ren, and J. Ma. Exploring a humanoid videounderstanding algorithm guided by behavior. Computer, 53(8):59–67, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:PI

[233] D. A. Grier. Pervasive ideals. Computer, 53(8):68–70, August 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISPe

[234] Anonymous. IEEE Security & Privacy. Computer, 53(8):70, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2020:VR

[235] S. Reisman. Viva la revolución? Computer, 53(8):71–73, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paler:2020:ACQ

[236] A. Paler. Aggregated control of quantum computations: When stacked architectures are too good to be practical soon. Computer, 53(8):74–78, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laporte:2020:DSS

[237] C. Y. Laporte and J. M. Miranda. Delivering software- and systemsengineering standards for small teams. Computer, 53(8):79–83, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITBf

[238] Anonymous. IEEE Transactions on Big Data. Computer, 53(8):83, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hasselbring:2020:OSR

[239] W. Hasselbring, L. Carr, S. Hettrick, H. Packer, and T. Tiropanis. Open source research software. Computer, 53 (8):84–88, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2020:PPP

[240] M. Campbell. Putting the passé into passwords: How passwordless technologies are reshaping digital identity. Computer, 53(8):89–93, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Osterweil:2020:YDDb

[241] E. Osterweil, A. Stavrou, and L. Zhang. 21 years of distributed denial-of-service: A call to action. Computer, 53 (8):94–99, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITSc

[242] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 53 (8):99, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kleidermacher:2020:BIS

[243] D. Kleidermacher. Building an Internet of Secure Things. Computer, 53(8): 100–104, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSo

[244] Anonymous. IEEE Computer Society Jobs Board. Computer, 53(8): 105, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSp

[245] Anonymous. IEEE Computer Society election. Computer, 53(8):106– 124, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IDV

[246] Anonymous. IEEE Division V Delegate/Director Candidates. Computer, 53(8):125–126, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSq

[247] Anonymous. IEEE Computer Society. Computer, 53(8):C3, August 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IOJb

[248] Anonymous. IEEE Open Journal of the Computer Society. Computer, 53 (8):C4, August 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCi

[249] Anonymous. Front cover. Computer, 53 (9):C1, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCSd

[250] Anonymous. Computer call for special issue proposals. Computer, 53(9): C2, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCi

[251] Anonymous. Table of contents. Computer, 53(9):1–2, September 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Mi

[252] Anonymous. Masthead. Computer, 53 (9):3, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFloriani:2020:JDC

[253] L. De Floriani. 2020: A journey of discovery, challenges, and opportunity. Computer, 53(9):4–7, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Iaa

[254] Anonymous. ICCQ 2021. Computer, 53(9):7, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:L

[255] Anonymous. Letters. Computer, 53 (9):8, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowman:2020:EPK

[256] D. A. Bowman. Embracing physical keyboards for virtual reality. Computer, 53(9):9–10, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IAHb

[257] Anonymous. IEEE Annals of the History of Computing. Computer, 53(9): 10, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSh

[258] Anonymous. Computer highlights Society magazines. Computer, 53(9):11– 13, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAi

[259] E. Neuhold. 50 & 25 years ago. Computer, 53(9):14–17, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2020:DL

[260] E. Akleman. Deep learning. Computer, 53(9):17, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:CL

[261] J. Voas. Citations and letters. Computer, 53(9):18–19, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSr

[262] Anonymous. IEEE Computer Society election. Computer, 53(9):19, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2020:SPC

[263] J. B. Michael, R. Kuhn, and J. Voas. Security or privacy: Can you have both? Computer, 53(9):20–30, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Braeken:2020:BCS

[264] A. Braeken, M. Liyanage, S. S. Kanhere, and S. Dixit. Blockchain and cyberphysical systems. Computer, 53(9): 31–35, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bouachir:2020:BFC

[265] O. Bouachir, M. Aloqaily, L. Tseng, and A. Boukerche. Blockchain and fog computing for cyberphysical systems: The case of smart industry. Computer, 53(9):36–45, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garrocho:2020:RTS

[266] C. T. B. Garrocho, M. C. Silva, C. M. S. Ferreira, C. F. M. d. C. Cavalcanti, and R. A. Rabelo Oliveira.

Real-time systems implications in the blockchain-based vertical integration of Industry 4.0. Computer, 53(9):46– 55, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICGd

[267] Anonymous. IEEE Computer Graphics & Applications subscribe. Computer, 53(9):55, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ali:2020:CBE

[268] F. S. Ali, M. Aloqaily, O. Alfandi, and O. Ozkasap. Cyberphysical blockchain-enabled peer-to-peer energy trading. Computer, 53(9):56– 65, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eisele:2020:BTE

[269] S. Eisele, C. Barreto, A. Dubey, X. Koutsoukos, T. Eghtesad, A. Laszka, and A. Mavridou. Blockchains for transactive energy systems: Opportunities, challenges, and approaches. Computer, 53(9):66–76, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSs

[270] Anonymous. IEEE Computer Society Jobs Board. Computer, 53(9): 77, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:IHT

[271] D. A. Grier. If not here, then where? Computer, 53(9):78–79, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tsoutsos:2020:CRM

[272] N. G. Tsoutsos, N. Gupta, and R. Karri. Cybersecurity road map for digital manufacturing. Computer, 53 (9):80–84, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IIPb

[273] Anonymous. IEEE IT Professional call for papers. Computer, 53(9): 84, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2020:ECB

[274] N. Kshetri and J. F. DeFranco. The economics of cyberattacks on Brazil. Computer, 53(9):85–90, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPCg

[275] Anonymous. IEEE Pervasive Computing call for papers. Computer, 53 (9):90, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Linkov:2020:CEA

[276] I. Linkov, S. Galaitsi, B. D. Trump, J. M. Keisler, and A. Kott. Cybertrust: From explainable to actionable and interpretable artificial intelligence. Computer, 53(9):91–96, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISPf

[277] Anonymous. IEEE Security and Privacy magazine subscribe. Computer, 53 (9):96, September 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2020:BHS

[278] P. Laplante. A brief history of software professionalism and the way forward. Computer, 53(9):97–100, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSt

[279] Anonymous. IEEE Computer Society. Computer, 53(9):C3, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSu

[280] Anonymous. IEEE Computer Society election. Computer, 53(9):C4, September 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCj

[281] Anonymous. Front cover. Computer, 53(10):C1, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Cf

[282] Anonymous. CVPR 2020. Computer, 53(10):C2, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCj

[283] Anonymous. Table of contents. Computer, 53(10):1–2, October 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Mj

[284] Anonymous. Masthead. Computer, 53 (10):3, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSi

[285] Anonymous. Computer highlights Society magazines. Computer, 53(10): 4–6, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAj

[286] E. Neuhold. 50 & 25 years ago. Computer, 53(10):7–10, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2020:CTTf

[287] E. Akleman. Computing through time. Computer, 53(10):9, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICGe

[288] Anonymous. IEEE Computer Graphics and Applications subscribe. Computer, 53(10):10, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:RHO

[289] J. Voas and P. Laplante. Rethinking home, office, school. Computer, 53(10): 11–12, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Ib

[290] Anonymous. ICCQ. Computer, 53 (10):12, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anagnostou:2020:COS

[291] A. Anagnostou and S. J. E. Taylor. Can open science change the world? Computer, 53(10):13–22, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jimenez-Gomez:2020:AIG

[292] C. E. Jimenez-Gomez, J. Cano-Carrillo, and F. Falcone Lanas. Artificial intelligence in government. Computer, 53(10):23–27, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITBg

[293] Anonymous. IEEE Transactions on Big Data. Computer, 53(10):27, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campion:2020:MAI

[294] A. Campion, M.-G. Hernandez, S. Mikhaylov Jankin, and M. Esteve. Managing artificial intelligence deployment in the public sector. Computer, 53(10):28–37, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITSd

[295] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 53 (10):37, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Viscusi:2020:PSA

[296] G. Viscusi, A. Rusu, and M.-V. Florin. Public strategies for artificial intelligence: Which value drivers? Computer, 53(10):38–46, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cobbe:2020:CLD

[297] J. Cobbe, M. S. A. Lee, H. Janssen, and J. Singh. Centering the law in the digital state. Computer, 53(10): 47–58, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IAHc

[298] Anonymous. IEEE Annals of the History of Computing. Computer, 53 (10):58, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jordan:2020:TTI

[299] S. B. Jordan, S. L. Fenn, and B. B. Shannon. Transparency as threat at the intersection of artificial intelligence and cyberbiosecurity. Computer, 53 (10):59–68, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSv

[300] Anonymous. IEEE Computer Society member benefits. Computer, 53 (10):68, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

eSilva:2020:RNP

[301] A. M. e Silva, Y. R. Rodrigues, and R. P. Ishii. RIGOR: A new proposal for predicting infant mortality in government health systems using artificial intelligence in Brazil. Computer, 53(10): 69–76, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gonzalez-Briceno:2020:AIB

[302] G. Gonzalez-Briceno, A. Sanchez, S. Ortega-Cisneros, M. S. Garcia Contreras, G. A. Pinedo Diaz, and E. U. Moya-Sanchez. Artificial intelligencebased referral system for patients with diabetic retinopathy. Computer, 53 (10):77–87, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:PP

[303] D. A. Grier. A prize and a prediction. Computer, 53(10):88–90, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSw

[304] Anonymous. IEEE Computer Society call for papers. Computer, 53 (10):90, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haney:2020:SAT

[305] J. Haney and W. Lutters. Security awareness training for the workforce: Moving beyond “Check-the-Box” compliance. Computer, 53(10): 91–95, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IIPc

[306] Anonymous. IEEE IT Professional call for articles. Computer, 53(10): 95, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeBenedictis:2020:QA

[307] E. DeBenedictis. Quantum access. Computer, 53(10):96–100, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPCh

[308] Anonymous. IEEE Pervasive Computing call for articles. Computer, 53 (10):100, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Herrera:2020:PCO

[309] A. Herrera. The promises and challenges of open source hardware. Computer, 53(10):101–104, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISPg

[310] Anonymous. IEEE Security and Privacy subscribe. Computer, 53(10): 104, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ombredanne:2020:FOS

[311] P. Ombredanne. Free and open source software license compliance: Tools for software composition analysis. Computer, 53(10):105–109, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2020:BZT

[312] M. Campbell. Beyond zero trust: Trust is a vulnerability. Computer, 53(10): 110–113, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gaudiot:2020:CEA

[313] J.-L. Gaudiot and H. Kasahara. Computer education in the age of COVID-19. Computer, 53(10):114–118, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITCf

[314] Anonymous. IEEE Transactions on Computers call for papers. Computer, 53(10):118, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2020:SDP

[315] N. Ahmad and P. Chauhan. State of data privacy during COVID-19. Computer, 53(10):119–122, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IOJc

[316] Anonymous. IEEE Open Journal of the Computer Society. Computer, 53 (10):123, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCh

[317] Anonymous. CS connection. Computer, 53(10):124, October 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Cg

[318] Anonymous. Correction. Computer, 53 (10):124, October 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSx

[319] Anonymous. IEEE Computer Society. Computer, 53(10):C3, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IC

[320] Anonymous. IEEE ComputingEdge. Computer, 53(10):C4, October 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCk

[321] Anonymous. Front cover. Computer, 53 (11):C1, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IOA

[322] Anonymous. IEEE Open Access journal of the Computer Society. Computer, 53(11):C2, November 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCk

[323] Anonymous. Table of contents. Computer, 53(11):1–2, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Mk

[324] Anonymous. Masthead. Computer, 53 (11):3, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSj

[325] Anonymous. Computer highlights Society magazines. Computer, 53(11): 4–6, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parashar:2020:PDS

[326] M. Parashar. Parallel and Distributed Systems. Computer, 53(11): 7–8, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2020:YAk

[327] E. Neuhold. 50 & 25 years ago. Computer, 53(11):9–12, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISMb

[328] Anonymous. IEEE Software magazine call for articles. Computer, 53 (11):12, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2020:AAA

[329] J. B. Michael and J. Voas. Algorithms, algorithms, algorithms. Computer, 53(11):13–15, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:Ic

[330] J. Voas. In this issue. Computer, 53 (11):14, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Id

[331] Anonymous. ICCQ. Computer, 53 (11):15, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2020:ECC

[332] D. Milojicic. The edge-to-cloud continuum. Computer, 53(11):16–25, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISPh

[333] Anonymous. IEEE & Security Privacy magazine subscribe. Computer, 53 (11):25, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2020:ALM

[334] B. Li, S. Liu, J. Tang, J.-L. Gaudiot, L. Zhang, and Q. Kong. Autonomous last-mile delivery vehicles in complex traffic environments. Computer, 53(11):26–35, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

vanNuenen:2020:TWA

[335] T. van Nuenen, X. Ferrer, J. M. Such, and M. Cote. Transparency for whom? Assessing discriminatory artificial intelligence. Computer, 53(11): 36–44, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2020:AIC

[336] P. Laplante, D. Milojicic, S. Serebryakov, and D. Bennett. Artificial intelligence and critical systems: From hype to reality. Computer, 53(11): 45–52, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ossewaarde:2020:NVA

[337] M. Ossewaarde and E. Gulenc. National varieties of artificial intelligence discourses: Myth, utopianism, and solutionism in West European policy expectations. Computer, 53(11):53– 61, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:FT

[338] D. A. Grier. Fellow travelers. Computer, 53(11):62–64, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IIPd

[339] Anonymous. IEEE IT Professional call for articles. Computer, 53(11): 64, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2020:EAI

[340] D. Serpanos, G. Ferrari, G. Nikolakopoulos, J. Perez, M. Tauber, and S. Van Baelen. Embedded artificial intelligence: The ARTEMIS vision. Computer, 53(11):65–69, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coughlan:2020:SOS

[341] S. Coughlan. Standardizing open source license compliance with Open-Chain. Computer, 53(11):70–74, November 2020. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSy

[342] Anonymous. IEEE Computer Society call for papers. Computer, 53 (11):74, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Viega:2020:YSS

[343] J. Viega. 20 years of software security. Computer, 53(11):75–78, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Seacord:2020:EC

[344] R. C. Seacord. Effective C. Computer, 53(11):79–82, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:E

[345] Anonymous. Erratum. Computer, 53 (11):82, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CCi

[346] Anonymous. CS connection. Computer, 53(11):83–84, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:OFA

[347] Anonymous. Obituary: Frances E. Allen. Computer, 53(11):84, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSz

[348] Anonymous. IEEE Computer Society. Computer, 53(11):C3, November 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSba

[349] Anonymous. IEEE Computer Society Jobs Board. Computer, 53(11): C4, November 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:FCl

[350] Anonymous. Front cover. Computer, 53 (12):C1, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IOJd

[351] Anonymous. IEEE Open Journal of the Computer Society. Computer, 53 (12):C2, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:TCl

[352] Anonymous. Table of contents. Computer, 53(12):1–2, December 2020. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:Ml

[353] Anonymous. Masthead. Computer, 53 (12):3, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaidya:2020:SCW

[354] J. Vaidya. Securely connecting wearables to displays. Computer, 53(12):

4–5, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:CHSk

[355] Anonymous. Computer highlights Society magazines. Computer, 53(12): 6–8, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:YA

[356] Anonymous. 50 & 25 years ago. Computer, 53(12):9–11, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2020:CTTg

[357] E. Akleman. Computing through time. Computer, 53(12):10, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IAb

[358] Anonymous. IEEE Annals. Computer, 53(12):11, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSbb

[359] Anonymous. IEEE Computer Society awards. Computer, 53(12):12– 17, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2020:AIM

[360] J. Voas and H.-Y. Lin. Is artificial intelligence/machine learning the anticipated silver bullet? Computer, 53(12): 18–19, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICGf

[361] Anonymous. IEEE Computer Graphics and Applications. Computer, 53 (12):19, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mashkoor:2020:IAS

[362] A. Mashkoor, P. Arcaini, and A. Gargantini. Intelligent autonomous systems. Computer, 53(12):20–23, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kabir:2020:CIS

[363] S. Kabir and Y. Papadopoulos. Computational intelligence for safety assurance of cooperative systems of systems. Computer, 53(12):24–34, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Asaadi:2020:DAC

[364] E. Asaadi, E. Denney, J. Menzies, G. J. Pai, and D. Petroff. Dynamic assurance cases: a pathway to trusted autonomy. Computer, 53(12): 35–46, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSbc

[365] Anonymous. IEEE Computer Society call for papers. Computer, 53 (12):46, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garces:2020:ASS

[366] L. Garces, S. Martinez-Fernandez, V. V. Graciano Neto, and E. Y. Nakagawa. Architectural solutions for selfadaptive systems. Computer, 53(12): 47–59, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Christie:2020:PTD

[367] S. H. Christie V, D. Smirnova, A. K. Chopra, and M. P. Singh. Protocols over things: A decentralized programming model for the Internet of Things. Computer, 53(12):60–68, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shi:2020:CRM

[368] H. Shi, W. Dong, R. Li, and W. Liu. Controller resynthesis for multirobot system when changes happen. Computer, 53(12):69–79, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sarkar:2020:RTO

[369] S. Sarkar, M. W. Totaro, A. Kumar, and K. Elgazzar. Real-time object processing and routing for intelligent drones: a novel approach. Computer, 53(12):80–89, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IIPe

[370] Anonymous. IEEE IT Professional call for articles. Computer, 53(12): 89, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2020:SP

[371] D. A. Grier. Stirring the pot. Computer, 53(12):90–91, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zou:2020:ADA

[372] Q. Zou, X. Sun, P. Liu, and A. Singhal. An approach for detection of advanced persistent threat attacks. Computer, 53(12):92–96, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICAb

[373] Anonymous. IEEE Computer Architecture Letters. Computer, 53(12): 97, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jensen:2020:ASE

[374] J. J. Jensen. Applying a “Smart ecosystem” mindset to rethink your products. Computer, 53(12):98–101, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2020:SEL

[375] P. Laplante. Software engineering lineage: How do you know what you know? Computer, 53(12):102– 105, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITSe

[376] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 53 (12):105, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2020:EBF

[377] N. Kshetri and J. DeFranco. The economics behind food supply blockchains. Computer, 53(12):106–110, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ITBh

[378] Anonymous. IEEE Trans on Big Data. Computer, 53(12):110, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mansfield:2020:NMI

[379] S. Mansfield. Network management in IEEE 802 standards. Computer, 53 (12):111–114, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPCi

[380] Anonymous. IEEE Pervasive Computing. Computer, 53(12):114, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Phipps:2020:COS

[381] S. Phipps and S. Zacchiroli. Continuous open source license compliance. Computer, 53(12):115–119, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSbd

[382] Anonymous. IEEE Computer Society generic. Computer, 53(12):120, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2020:BCA

[383] M. Campbell. Beyond conversational artificial intelligence. Computer, 53 (12):121–125, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:IPS

[384] Anonymous. IEEE S & P subscribe. Computer, 53(12):125, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liebowitz:2020:LPA

[385] J. Liebowitz. Life as a professor amid COVID-19. Computer, 53(12):126– 128, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chauhan:2020:PDD

[386] P. Chauhan. Plagiarism detection: a data search challenge. Computer, 53 (12):129–132, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ISN

[387] Anonymous. IEEE social networking. Computer, 53(12):133, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:GVI

[388] Anonymous. Gropp voted 2021 IEEE Computer Society President-Elect. Computer, 53(12):134–136, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSbe

[389] Anonymous. IEEE Computer Society. Computer, 53(12):C3, December 2020. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2020:ICSbf

[390] Anonymous. IEEE Computer Society Jobs Board. Computer, 53(12): C4, December 2020. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCa

[391] Anonymous. Front cover. Computer, 54(1):C1, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IOJa

[392] Anonymous. IEEE Open Journal of the Computer Society. Computer, 54 (1):C2, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCa

[393] Anonymous. Table of contents. Computer, 54(1):1–2, January 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Ma

[394] Anonymous. Masthead. Computer, 54(1):3, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shull:2021:DIR

[395] F. Shull and M. Ben. Driving innovation and removing barriers: Forrest Shull takes the helm as 2021 IEEE Computer Society President. Computer, 54(1):4–9, January 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFloriani:2021:YTC

[396] L. De Floriani. 2020: A year of transformative change and conquering adversity. Computer, 54(1):10– 14, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICAa

[397] Anonymous. IEEE Computer Architecture Letters. Computer, 54(1): 15, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Andre:2021:CAI

[398] E. Andre. Can artificial intelligence detect good (or bad) chemistry between people? Computer, 54(1): 16–17, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHa

[399] Anonymous. IEEE Annals of the History of Computing. Computer, 54 (1):17, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CHSa

[400] Anonymous. Computer highlights Society magazines. Computer, 54(1): 18–20, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2021:YAa

[401] E. Neuhold. 50 & 25 years ago. Computer, 54(1):21–24, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IJBa

[402] Anonymous. IEEE Jobs Board. Computer, 54(1):25, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:S

[403] J. Voas and N. Kshetri. Scarcity. Computer, 54(1):26–28, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICGa

[404] Anonymous. IEEE Computer Graphics Applications subscription. Computer, 54(1):28, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mouha:2021:AFM

[405] N. Mouha and A. Hailane. The application of formal methods to realworld cryptographic algorithms, protocols, and systems. Computer, 54(1): 29–38, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:AIC

[406] Anonymous. ACM/IEEE Computer Society Eckert–Mauchly Award. Computer, 54(1):39, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Williams:2021:O

[407] J. Williams and J. Voas. Outlook 2021. Computer, 54(1):40–41, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Antinyan:2021:HLC

[408] V. Antinyan. Hypnotized by lines of code. Computer, 54(1):42–48, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weiss:2021:UFC

[409] A. Weiss, S. Gautham, A. V. Jayakumar, C. R. Elks, D. R. Kuhn, R. N. Kacker, and T. B. Preusser. Understanding and fixing complex faults in embedded cyberphysical systems. Computer, 54(1):49–60, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSa

[410] Anonymous. IEEE Computer Society call for papers. Computer, 54 (1):60, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Piroumian:2021:DTU

[411] V. Piroumian. Digital twins: Universal interoperability for the digital age. Computer, 54(1):61–69, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:2021:TER

[412] H. M. Kim, M. Laskowski, M. Zargham, H. Turesson, M. Barlin, and D. Kabanov. Token economics in real life: Cryptocurrency and incentives design for Insolar’s blockchain network.

Computer, 54(1):70–80, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:WDW

[413] D. A. Grier. Where do we learn to behave like grownups? Computer, 54 (1):81–83, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIPa

[414] Anonymous. IEEE IT Professional call for articles. Computer, 54(1): 83, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuketa:2021:EAI

[415] H. Fuketa and K. Uchiyama. Edge artificial intelligence chips for the cyberphysical systems era. Computer, 54 (1):84–88, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IPCa

[416] Anonymous. IEEE Pervasive Computing call for articles. Computer, 54 (1):88, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2021:ITS

[417] R. D. Riehle and J. B. Michael. Improving the trustworthiness of software through rigorous data type design. Computer, 54(1):89–95, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISPa

[418] Anonymous. IEEE Security & Privacy subscribe. Computer, 54(1): 95, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Borenstein:2021:AEL

[419] J. Borenstein, F. S. Grodzinsky, A. Howard, K. W. Miller, and M. J. Wolf. AI ethics: a long history and a recent burst of attention. Computer, 54(1):96–102, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITBa

[420] Anonymous. IEEE Transactions on Big Data. Computer, 54(1):102, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Edwards:2021:III

[421] P. N. Edwards and J. L. King. Institutions, infrastructures, and innovation. Computer, 54(1):103–109, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITSa

[422] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 54 (1):109, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:E

[423] Anonymous. Erratum. Computer, 54 (1):110, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CCa

[424] Anonymous. CS connection. Computer, 54(1):111–112, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:RMA

[425] Anonymous. Richard E. Merwin Award. Computer, 54(1):112, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSb

[426] Anonymous. IEEE Computer Society. Computer, 54(1):C3, January 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSc

[427] Anonymous. IEEE Computer Society membership. Computer, 54(1): C4, January 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCb

[428] Anonymous. Front cover. Computer, 54(2):C1, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IOJb

[429] Anonymous. IEEE Open Journal of the Computer Society. Computer, 54 (2):C2, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCb

[430] Anonymous. Table of contents. Computer, 54(2):1–2, February 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Mb

[431] Anonymous. Masthead. Computer, 54 (2):3, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mittal:2021:SVV

[432] S. Mittal. Secure V2V and V2I technologies for the next-generation intelligent transportation systems. Computer, 54(2):4–6, February 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITCa

[433] Anonymous. IEEE Transactions on Computers call for papers. Computer, 54(2):6, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CHSb

[434] Anonymous. Computer highlights Society magazines. Computer, 54(2): 7–9, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2021:YAb

[435] E. Neuhold. 50 & 25 years ago. Computer, 54(2):10–12, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2021:CTTa

[436] E. Akleman. Computing through time. Computer, 54(2):11, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:MY

[437] J. Voas and N. Kshetri. Monetizing you. Computer, 54(2):13–15, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:Ia

[438] J. Voas. In this issue. Computer, 54 (2):14, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIC

[439] Anonymous. IEEE IRI 2021 call for papers. Computer, 54(2):15, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:STC

[440] D. A. Grier, D. Milojicic, S. Reisman, and J. Walz. Special technical communities: a 10-year retrospective. Computer, 54(2):16–25, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2021:CST

[441] N. Kshetri, J. Voas, and R. Sharma. Computing and socioeconomic transformations. Computer, 54(2):26–29, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Victor:2021:TDL

[442] F. Victor, P. Ruppel, and A. Kupper. A taxonomy for distributed ledger analytics. Computer, 54(2):30–38, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISPb

[443] Anonymous. IEEE Security & Privacy magazine subscribe. Computer, 54 (2):38, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mendes:2021:NRS

[444] G. S. Mendes, D. Chen, B. M. C. Silva, C. Serrao, and J. Casal. A novel reputation system for mobile app stores using blockchain. Computer, 54(2): 39–49, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IPCb

[445] Anonymous. IEEE Pervasive Computing call for articles. Computer, 54 (2):49, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:2021:SFF

[446] H.-Y. Kim, L. Xu, W. Shi, and T. Suh. A secure and flexible FPGAbased blockchain system for the IIoT. Computer, 54(2):50–59, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:TSN

[447] D. A. Grier. There is still no silver bullet. Computer, 54(2):60–62, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIPb

[448] Anonymous. IEEE IT Professional call for articles. Computer, 54(2): 62, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Strawn:2021:PPE

[449] G. O. Strawn. Pandemics, public education, and the future. Computer, 54 (2):63–65, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSd

[450] Anonymous. IEEE Computer Society call for papers. Computer, 54 (2):65, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2021:TCC

[451] M. Campbell. Tomorrow’s call center will be neither: Not call focused, nor centralized. Computer, 54(2):66– 69, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICGb

[452] Anonymous. IEEE Computer Graphics & Applications. Computer, 54 (2):69, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2021:DDA

[453] N. Ahmad. Do data almost always eventually leak? Computer, 54(2):

70–74, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHb

[454] Anonymous. IEEE Annals of the History of Computing. Computer, 54 (2):74, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gonzalez-Barahona:2021:BHF

[455] J. M. Gonzalez-Barahona. A brief history of free, open source software and its communities. Computer, 54(2): 75–79, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kott:2021:ICR

[456] A. Kott and I. Linkov. To improve cyber resilience, measure it. Computer, 54 (2):80–85, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITSb

[457] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 54 (2):85, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2021:RMS

[458] J. F. DeFranco, N. Kshetri, and J. Voas. Rearchitecting the malls: The shifting U.S. retail landscape. Computer, 54(2):86–90, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CCb

[459] Anonymous. CS connection. Computer, 54(2):91–92, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSe

[460] Anonymous. IEEE Computer Society. Computer, 54(2):C3, February 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSf

[461] Anonymous. IEEE Computer Society member benefits. Computer, 54 (2):C4, February 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCc

[462] Anonymous. Front cover. Computer, 54(3):C1, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IOJc

[463] Anonymous. IEEE Open Journal of the Computer Society. Computer, 54 (3):C2, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCc

[464] Anonymous. Table of contents. Computer, 54(3):1–2, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Mc

[465] Anonymous. Masthead. Computer, 54(3):3, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sorin:2021:RHP

[466] D. J. Sorin. Reconfigurable hardware in postsilicon microarchitecture. Computer, 54(3):4–5, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CHSc

[467] Anonymous. Computer highlights Society magazines. Computer, 54(3): 6–8, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IEA

[468] Anonymous. IEEE Education Awards nominations. Computer, 54(3):9, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:YAa

[469] Anonymous. 50 & 25 years ago. Computer, 54(3):10–13, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHc

[470] Anonymous. IEEE Annals of the History of Computing. Computer, 54 (3):13, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:OEF

[471] J. Voas and N. Kshetri. Online eats “Food Courts”. Computer, 54(3): 14–17, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:Ib

[472] J. Voas. In this issue. Computer, 54(3):15, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICGc

[473] Anonymous. IEEE Computer Graphics and Applications. Computer, 54 (3):17, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nawrocki:2021:CNC

[474] P. Nawrocki, K. Wrona, M. Marczak, and B. Sniezynski. A comparison of native and cross-platform frameworks for mobile applications. Computer, 54 (3):18–27, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2021:VLC

[475] C. Liu. Visualizing logical correlation in trace data for system debugging. Computer, 54(3):28–36, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sugandhi:2021:ISL

[476] Sugandhi, P. Kumar, and S. Kaur. Indian sign language generation system. Computer, 54(3):37–46, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICAb

[477] Anonymous. IEEE Computer Architecture Letters. Computer, 54(3):47, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2021:WWA

[478] H.-Y. Lin. Will we adopt AI like we adopted electricity? Computer, 54 (3):48–51, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:EM

[479] D. A. Grier. The engineering mindset. Computer, 54(3):52–54, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSg

[480] Anonymous. IEEE Computer Society call for papers. Computer, 54 (3):54, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2021:TFS

[481] J. F. DeFranco, D. F. Ferraiolo, R. Kuhn, and J. Roberts. A trusted federated system to share granular data among disparate database resources. Computer, 54(3):55–62, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIPc

[482] Anonymous. IEEE IT Professional call for articles. Computer, 54(3):62, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kapsalas:2021:RMM

[483] P. Kapsalas, A. S. Lalos, K. Moustakas, and D. Serpanos. The role of modularity in multimodal simultaneous localization and mapping systems. Computer, 54(3):63–67, March 2021.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IPCc

[484] Anonymous. IEEE Pervasive Computing call for articles. Computer, 54 (3):67, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Furnell:2021:PPW

[485] S. Furnell, J. Haney, and M. Theofanos. Pandemic parallels: What can cybersecurity learn from COVID-19? Computer, 54(3):68–72, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISPc

[486] Anonymous. IEEE Security & Privacy subscription. Computer, 54 (3):72, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sangwan:2021:ADA

[487] R. S. Sangwan, R. L. Nord, and I. Ozkaya. Architectural dependency analysis: Addressing the elephant in the room. Computer, 54(3):73–78, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IJBb

[488] Anonymous. IEEE Jobs Board. Computer, 54(3):79, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mariani:2021:IIF

[489] R. Mariani, N. Maor, J. Athavale, and K. Gay. The importance of interoperability in functional safety standards. Computer, 54(3):80–84, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITBb

[490] Anonymous. IEEE Transactions on Big Data. Computer, 54(3):84, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSh

[491] Anonymous. IEEE Computer Society. Computer, 54(3):C3, March 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSi

[492] Anonymous. IEEE Computer Society member benefits. Computer, 54 (3):C4, March 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCd

[493] Anonymous. Front cover. Computer, 54(4):C1, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ASA

[494] Anonymous. 2022 AAAI Squirrel AI Award call for nominations. Computer, 54(4):C2, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCd

[495] Anonymous. Table of contents. Computer, 54(4):1–2, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Md

[496] Anonymous. Masthead. Computer, 54(4):3, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CHSd

[497] Anonymous. Computer highlights Society magazines. Computer, 54(4):4–6, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IJBc

[498] Anonymous. IEEE Jobs Boards. Computer, 54(4):7, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:YAb

[499] Anonymous. 50 & 25 years ago. Computer, 54(4):8–10, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2021:CTTb

[500] E. Akleman. Computing through time. Computer, 54(4):10, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:CYD

[501] J. Voas and K. Miller. Confronting your digital accuser. Computer, 54 (4):11–13, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:Ic

[502] J. Voas. In this issue. Computer, 54(4):12, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISPd

[503] Anonymous. IEEE Security & Privacy subscribe. Computer, 54(4):13, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2021:PRR

[504] L. Liu, J. Tang, S. Liu, B. Yu, Y. Xie, and J.-L. Gaudiot. Π-RT: A runtime framework to enable energyefficient real-time robotic vision applications on heterogeneous architectures. Computer, 54(4):14–25, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2021:CMH

[505] J. Wang, L. Wang, and Y. Song. Crowd-machine hybrid urban sensing and computing. Computer, 54(4): 26–34, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lew:2021:FSS

[506] A. Lew. Flipping the script: A sociotechnical approach to platforms and unanticipated uses. Computer, 54 (4):35–44, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICM

[507] Anonymous. IEEE call for 2021 major award nominations. Computer, 54(4):45, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:LET

[508] D. A. Grier. Linking the elements together: Will NoSQL databases live up to their promise? Computer, 54 (4):46–48, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIPd

[509] Anonymous. IEEE IT Professional call for articles. Computer, 54(4):48, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2021:WAM

[510] D. Drusinsky. Who is authenticating my e-commerce logins? Computer, 54 (4):49–54, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CSEa

[511] Anonymous. Computing in sciences and engineering. Computer, 54(4):54, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2021:CAO

[512] M. Campbell. Computational acceleration: Offense in depth. Computer, 54 (4):55–58, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chauhan:2021:BDP

[513] P. Chauhan and M. Sood. Big data: Present and future. Computer, 54 (4):59–65, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2021:EEA

[514] S. Liu, J.-L. Gaudiot, and H. Kasahara. Engineering education in the age of autonomous machines. Computer, 54(4):66–69, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drost-Fromm:2021:OSC

[515] I. Drost-Fromm and R. Tompkins. Open source community governance the Apache way. Computer, 54(4): 70–75, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lanoue:2021:AVT

[516] M. Lanoue, C. A. Bollmann, J. B. Michael, J. Roth, and D. Wijesekera. An attack vector taxonomy for mobile telephony security vulnerabilities. Computer, 54(4):76–84, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICE

[517] Anonymous. IEEE call for educational awards nominations. Computer, 54(4):85, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2021:EDT

[518] N. Kshetri. The economics of digital twins. Computer, 54(4):86–90, April 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSj

[519] Anonymous. IEEE Computer Society. Computer, 54(4):C3, April 2021.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSk

[520] Anonymous. IEEE Computer Society membership benefits. Computer, 54(4):C4, April 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCe

[521] Anonymous. Front cover. Computer, 54(5):C1, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IOJd

[522] Anonymous. IEEE Open Journal of the Computer Society. Computer, 54 (5):C2, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCe

[523] Anonymous. Table of contents. Computer, 54(5):1–2, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Me

[524] Anonymous. Masthead. Computer, 54(5):3, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dogan:2021:AHA

[525] Gulustan Dogan. Advances in human activity recognition. Computer, 54(5):4–6, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSl

[526] Anonymous. IEEE Computer Society call for papers. Computer, 54(5):6, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CHSe

[527] Anonymous. Computer highlights Society magazines. Computer, 54(5):7–9, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2021:YAc

[528] Erich Neuhold. 50 & 25 years ago. Computer, 54(5):10–12, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cerf:2021:WWD

[529] Vinton G. Cerf. “What we do now, Kemosabe?”. Computer, 54(5):13–14, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITBc

[530] Anonymous. IEEE Transactions on Big Data. Computer, 54(5):14, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2021:MCT

[531] Nir Kshetri and Jeffrey Voas. Major computing technologies of the past 75 years. Computer, 54(5):15–21, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISPe

[532] Anonymous. IEEE Security & Privacy magazine. Computer, 54(5):21, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolf:2021:CAL

[533] Marilyn Wolf and Dimitrios Serpanos. Computing for autonomy: Latency, power, resilience. Computer, 54(5): 22–23, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arulraj:2021:EVE

[534] Joy Arulraj, Abhijit Chatterjee, Alexandros Daglis, Ashutosh Dhekne, and Umakishore Ramachandran. eCloud: A vision for the evolution of the edge– cloud continuum. Computer, 54(5): 24–33, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IPCd

[535] Anonymous. IEEE Pervasive Computing call for articles. Computer, 54(5):33, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rinner:2021:MSM

[536] Bernhard Rinner, Christian Bettstetter, Hermann Hellwagner, and Stephan Weiss. Multidrone systems: More than the sum of the parts. Computer, 54 (5):34–43, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Christie:2021:BIF

[537] Samuel H. Christie V, Amit Khushwant Chopra, and Munindar P. Singh. Bungie: Improving fault tolerance via extensible application-level protocols. Computer, 54(5):44–53, May 2021.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIPe

[538] Anonymous. IEEE IT Professionals call for articles. Computer, 54 (5):53, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:DMM

[539] David Alan Grier. Don’t make me puke. Computer, 54(5):54–55, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2021:AML

[540] Hsiao-Ying Lin and Battista Biggio. Adversarial machine learning: Attacks from laboratories to the real world. Computer, 54(5):56–60, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IJS

[541] Anonymous. IEEE Job Site Board. Computer, 54(5):61, May 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parnas:2021:SEP

[542] David Lorge Parnas. Software engineering: A profession in waiting. Computer, 54(5):62–64, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2021:H

[543] Joanna F. DeFranco, Christopher S. Eagle, James Bret Michael, John Viega, and Jeffrey Voas. Hackathons 101. Computer, 54(5):65–69, May 2021.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSm

[544] Anonymous. IEEE Computing in Science Engineering. Computer, 54(5):69, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eleftheriadou:2021:CCD

[545] Dana Eleftheriadou, Eddy Hartog, and Natalia Gkiaouri. The cities challenge: Driving a green and digital recovery and social resilience. Computer, 54 (5):70–75, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICGd

[546] Anonymous. IEEE Computer Graphics and Applications. Computer, 54 (5):75, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2021:USM

[547] Joanna F. DeFranco and Michael Hutchinson. Understanding smart medical devices. Computer, 54(5): 76–80, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHd

[548] Anonymous. IEEE Annals of the History of Computing. Computer, 54(5):80, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laporte:2021:TSE

[549] Claude Y. Laporte and Mirna Munoz. Not teaching software engineering standards to future software engineers — malpractice? Computer, 54(5):81–88, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSn

[550] Anonymous. IEEE Computer Society. Computer, 54(5):C3, May 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSo

[551] Anonymous. IEEE Computer Society member benefits. Computer, 54 (5):C4, May 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCf

[552] Anonymous. [Front cover]. Computer, 54(6):C1, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IOJe

[553] Anonymous. IEEE Open Journal of the Computer Society. Computer, 54 (6):C2, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCf

[554] Anonymous. Table of contents. Computer, 54(6):1–2, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Mf

[555] Anonymous. Masthead. Computer, 54(6):3, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CHSf

[556] Anonymous. Computer highlights Society magazines. Computer, 54(6):4–6, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:YAc

[557] Anonymous. 50 & 25 years ago. Computer, 54(6):7–9, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2021:CTTc

[558] Ergun Akleman. Computing through time. Computer, 54(6):8, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSp

[559] Anonymous. IEEE Computer Society call for papers. Computer, 54(6):9, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:HEM

[560] Jeffrey Voas and Nir Kshetri. Homework [EIC’S message]. Computer, 54 (6):10–12, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:AI

[561] Jeffrey Voas. About this issue. Computer, 54(6):11, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IJBd

[562] Anonymous. IEEE Job Boards. Computer, 54(6):13, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Canellas:2021:DIS

[563] Marc Canellas. Defending IEEE software standards in federal criminal court. Computer, 54(6):14–23, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Meza:2021:HMC

[564] Jaime Meza, Leticia Vaca-Cardenas, Monica Elva Vaca-Cardenas, Luis Teran, and Edy Portmann. A humanmachine collaboration model for urban planning in smart cities. Computer, 54(6):24–35, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fleming:2021:CST

[565] Cody H. Fleming, Carl Elks, Georgios Bakirtzis, Stephen Adams, Bryan Carter, Peter Beling, and Barry Horowitz. Cyberphysical security through resiliency: a systems-centric approach. Computer, 54(6):36–45, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:FML

[566] David Alan Grier. Forgetting Moore’s Law [body of knowledge]. Computer, 54(6):46–48, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSq

[567] Anonymous. IEEE Computing in Science and Engineering publication. Computer, 54(6):48, June 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jung:2021:QAO

[568] Changmi Jung and Lasse Mertins. Quality and accountability in the online business education environment. Computer, 54(6):49–52, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2021:ECB

[569] Nir Kshetri. The economics of central bank digital currency [computing’s economics]. Computer, 54(6):53–58, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICGe

[570] Anonymous. IEEE Computer Graphics and Applications. Computer, 54 (6):58, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Beaton:2021:OSS

[571] Wayne Beaton. Open source software engineering the Eclipse way. Computer, 54(6):59–63, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kroll:2021:ECA

[572] Joshua A. Kroll, James Bret Michael, and David B. Thaw. Enhancing cybersecurity via artificial intelligence: Risks, rewards, and frameworks. Computer, 54(6):64–71, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CCc

[573] Anonymous. CS connection. Computer, 54(6):72, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHe

[574] Anonymous. IEEE Annals of the History of Computing. Computer, 54 (6):72, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSr

[575] Anonymous. IEEE Computer Society. Computer, 54(6):C3, June 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSs

[576] Anonymous. IEEE Computer Society member benefits. Computer, 54 (6):C4, June 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCg

[577] Anonymous. Front cover. Computer, 54(7):C1, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSt

[578] Anonymous. IEEE Computer Society Open Journal. Computer, 54 (7):C2, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCg

[579] Anonymous. Table of contents. Computer, 54(7):1–2, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Mg

[580] Anonymous. Masthead. Computer, 54(7):3, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parashar:2021:ERR

[581] Manish Parashar. Enabling reproducible research in parallel and distributed systems. Computer, 54(7):4– 5, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISPf

[582] Anonymous. IEEE Security and Privacy subscribe. Computer, 54(7):5, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CHSg

[583] Anonymous. Computer highlights Society magazines. Computer, 54(7):6–8, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:YAd

[584] Anonymous. 50 & 25 years ago. Computer, 54(7):9–11, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:LNF

[585] Jeffrey Voas and Nir Kshetri. Lost and never found. Computer, 54(7): 12–13, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISM

[586] Anonymous. IEEE Software magazine call for articles. Computer, 54(7):13, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2021:TP

[587] Phil Laplante and Dejan Milojicic. Technology predictions. Computer, 54 (7):14–16, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICA

[588] Anonymous. IEEE Computer Architecture Letters. Computer, 54(7):17, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

GarciaMacias:2021:DIT

[589] J. Antonio GarciaMacias, Luis A. Castro, and Arturo Serrano-Santoyo. Diverging images of the technoscapes in developing economies. Computer, 54 (7):18–26, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xu:2021:RLF

[590] Lei Xu, Lin Chen, Zhimin Gao, Hiram Moya, and Weidong Shi. Reshaping the landscape of the future: Softwaredefined manufacturing. Computer, 54 (7):27–36, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Motwani:2021:EHT

[591] Sachin Motwani and Anubha Gupta. Experiencing hyperloops: The transit of the future. Computer, 54 (7):37–45, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Russell:2021:WAT

[592] Aspen Russell and Eitan Frachtenberg. Worlds apart: Technology, remote work, and equity. Computer, 54 (7):46–56, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Broering:2021:TEI

[593] Arne Broering, Christoph Niedermeier, Ioana Olaru, Ulrich Schopp, Kilian Telschig, and Michael Villnow. Toward embodied intelligence: Smart things on the rise. Computer, 54(7):57–68, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHf

[594] Anonymous. IEEE Annals of the History of Computing. Computer, 54(7):68, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mikkonen:2021:IIT

[595] Tommi Mikkonen, Cesare Pautasso, and Antero Taivalsaari. Isomorphic Internet of Things architectures with Web technologies. Computer, 54(7): 69–78, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:VS

[596] David Alan Grier. Value by squares. Computer, 54(7):79–81, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IMM

[597] Anonymous. IEEE Multimedia magazine. Computer, 54(7):81, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSu

[598] Anonymous. IEEE Computer Society call for papers. Computer, 54 (7):81, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chatzoglou:2021:WAY

[599] Efstratios Chatzoglou, Georgios Kambourakis, and Constantinos Kolias. WiF0: All your passphrase are belong to us. Computer, 54(7):82–88, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSv

[600] Anonymous. IEEE Computing in Science and Engineering. Computer, 54(7):88, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brophy:2021:DAS

[601] Dennis Brophy. Design Automation Standards Committee: Setting tomorrow’s standards. Computer, 54 (7):89–93, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gao:2021:NAR

[602] Sicun Gao. Nonlinearity, automation, and reliable cyberphysical systems. Computer, 54(7):94–96, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IJBe

[603] Anonymous. IEEE Jobs Board. Computer, 54(7):97, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Viega:2021:SSC

[604] John Viega and James Bret Michael. Struggling with supply-chain security. Computer, 54(7):98–104, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2021:DD

[605] Norita Ahmad, Nash Milic, and Mohammed Ibahrine. Data and disinformation. Computer, 54(7):105–110, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2021:MPU

[606] Doron Drusinsky and James Bret Michael. Multiagent pathfinding under rigid, optimization, and uncertainty constraints. Computer, 54(7): 111–118, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Muller:2021:TCE

[607] Heimo Muller, Michaela Theresia Mayrhofer, Evert-Ben Van Veen, and Andreas Holzinger. The ten commandments of ethical medical AI. Computer, 54(7):119–123, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIPf

[608] Anonymous. IEEE IT Professional call for articles. Computer, 54(7):123, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CCd

[609] Anonymous. CS connection. Computer, 54(7):124, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IPCe

[610] Anonymous. IEEE Pervasive Computing call for articles. Computer, 54(7):124, July 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSw

[611] Anonymous. IEEE Computer Society. Computer, 54(7):C3, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSx

[612] Anonymous. IEEE Computer Society membership. Computer, 54(7):C4, July 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCh

[613] Anonymous. [Front cover]. Computer, 54(8):C1, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSy

[614] Anonymous. IEEE Computer Society election. Computer, 54(8):C2, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCh

[615] Anonymous. Table of contents. Computer, 54(8):1–2, August 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Mh

[616] Anonymous. Masthead. Computer, 54(8):3, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kelly:2021:L

[617] Terence Kelly. Letters. Computer, 54(8):4, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:EC

[618] Anonymous. Elsewhere in the CS. Computer, 54(8):5–7, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2021:YAd

[619] Erich Neuhold. 50 & 25 years ago. Computer, 54(8):8–10, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2021:CTTd

[620] Ergun Akleman. Compliting through time. Computer, 54(8):9, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2021:WS

[621] Nir Kshetri and Jeffrey Voas. Where’s the silicon? Computer, 54(8):11–12, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Getov:2021:IOI

[622] Vladimir Getov. Insights into the origins of the IEEE Computer Society and the invention of electronic digital computing. Computer, 54(8):13–18, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHg

[623] Anonymous. IEEE Annals of the History of Computing. Computer, 54 (8):18, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:OJI

[624] Anonymous. Open Journal of the IEEE Computer Society. Computer, 54(8):19, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wong:2021:SSR

[625] W. Eric Wong and Zijiang Yang. Safety, security, and reliability of autonomous vehicle software. Computer, 54(8):20–21, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Burton:2021:SCA

[626] Simon Burton, John Alexander Mc-Dermid, Philip Garnett, and Rob Weaver. Safety, complexity, and automated driving: Holistic perspectives on safety assurance. Computer, 54 (8):22–32, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yu:2021:BBC

[627] Haiyang Yu, Shuai Ma, Qi Hu, and Zhen Yang. Blockchain-based continuous auditing for dynamic data sharing in autonomous vehicular networks. Computer, 54(8):33–45, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICGf

[628] Anonymous. IEEE Computer Graphics and Applications. Computer, 54 (8):45, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Motwani:2021:EAV

[629] Sachin Motwani, Tarun Sharma, and Anubha Gupta. Ethics in autonomous vehicle software: The dilemmas. Computer, 54(8):46–55, August 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ni:2021:OMF

[630] Ye Ni, Zhilong Xia, Fangtong Zhao, Chunrong Fang, and Zhenyu Chen. An online multistep-forward voltageprediction approach based on an LSTM-TD model and KF algorithm. Computer, 54(8):56–65, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSz

[631] Anonymous. IEEE Computer Society call for papers. Computer, 54 (8):65, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aslansefat:2021:TIC

[632] Koorosh Aslansefat, Sohag Kabir, Amr Abdullatif, Vinod Vasudevan, and Yiannis Papadopoulos. Toward improving confidence in autonomous vehicle software: a study on traffic sign recognition systems. Computer, 54(8): 66–76, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Toohey:2021:NCS

[633] Jack R. Toohey, M. S. Raunak, and David Binkley. From neuron coverage to steering angle: Testing autonomous vehicles effectively. Computer, 54(8): 77–85, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:CD

[634] David Alan Grier. For the common defense. Computer, 54(8):86–88, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIPg

[635] Anonymous. IEEE IT Professional call for articles. Computer, 54(8):88, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2021:CAI

[636] Mark Campbell and Mladan Jovanovic. Conversational artificial intelligence: Changing tomorrow’s health care today. Computer, 54(8):89–93, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2021:EBB

[637] Dejan Milojicic. Educate back better: a perspective from industry. Computer, 54(8):94–97, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IPCf

[638] Anonymous. IEEE Pervasive Computing call for articles. Computer, 54 (8):97, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2021:EBB

[639] Nir Kshetri. The economics of blockchain-based supply chain traceability in developing countries. Computer, 54(8):98–103, August 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISPg

[640] Anonymous. IEEE Security & Privacy subscribe. Computer, 54(8): 103, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goggins:2021:MOS

[641] Sean P. Goggins, Matt Germonprez, and Kevin Lumbard. Making open source project health transparent. Computer, 54(8):104–111, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITCb

[642] Anonymous. IEEE Transactions on Computers call for papers. Computer, 54(8):111, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kott:2021:CRD

[643] Alexander Kott, Maureen S. Golan, Benjamin D. Trump, and Igor Linkov. Cyber resilience: by design or by intervention? Computer, 54(8):112– 117, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Treleaven:2021:FCD

[644] Philip Treleaven. Future computers: Digital, quantum, biological. Computer, 54(8):118–124, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chauhan:2021:SPD

[645] Preeti S. Chauhan and Nir Kshetri. 2021 state of the practice in data privacy and security. Computer, 54(8): 125–132, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSba

[646] Anonymous. IEEE Computer Society. Computer, 54(8):C3, August 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbb

[647] Anonymous. IEEE Computer Society member benefits. Computer, 54 (8):C4, August 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCi

[648] Anonymous. Front cover. Computer, 54 (9):C1, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CSEb

[649] Anonymous. Computer Society election. Computer, 54(9):C2, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCi

[650] Anonymous. Table of contents. Computer, 54(9):1–2, September 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Mi

[651] Anonymous. Masthead. Computer, 54 (9):3, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaumont:2021:CSF

[652] Patrick Schaumont and Paolo Montuschi. Computer security at the forefront of emerging topics in computing. Computer, 54(9):4–5, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:HSMa

[653] Anonymous. Highlights Society magazines. Computer, 54(9):6–8, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2021:AYA

[654] Erich Neuhold. 50 & 25 years ago. Computer, 54(9):9–12, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Ia

[655] Anonymous. IQQC. Computer, 54 (9):12, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:OB

[656] Jeffrey Voas and Phil Laplante. Opening banking. Computer, 54(9):13– 14, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2021:FVC

[657] James Bret Michael, Doron Drusinsky, and Duminda Wijesekera. Formal verification of cyberphysical systems. Computer, 54(9):15–24, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2021:FMC

[658] James Bret Michael, Doron Drusinsky, and Duminda Wijesekera. Formal methods in cyberphysical systems. Computer, 54(9):25–29, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IPCg

[659] Anonymous. IEEE Pervasive Computing call for articles. Computer, 54 (9):29, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Deng:2021:IFD

[660] Ziquan Deng and Zhaodan Kong. Interpretable fault diagnosis for cyberphysical systems: a learning perspective. Computer, 54(9):30–38, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIPh

[661] Anonymous. IEEE IT Professional call for articles. Computer, 54(9): 38, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ma:2021:TFM

[662] Meiyi Ma, John A. Stankovic, and Lu Feng. Toward formal methods for smart cities. Computer, 54(9):39– 48, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:2021:EMD

[663] Guangyao Chen and Zhihao Jiang. Environment modeling during model checking of cyberphysical systems. Computer, 54(9):49–58, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbc

[664] Anonymous. IEEE Computer Society call for papers. Computer, 54(9): 58, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jablonski:2021:CSF

[665] Matthew Jablonski, Bo Yu, Gabriela Felicia Ciocarlie, and Paulo Costa. A case study in the formal modeling of safe and secure manufacturing automation. Computer, 54(9):59–71, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:TM

[666] David Alan Grier. Time and money. Computer, 54(9):72–74, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IOJf

[667] Anonymous. IEEE Open Journal of the Computer Society. Computer, 54 (9):75, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rokne:2021:S

[668] Jon Rokne. Shoehorning. Computer, 54(9):76–79, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbd

[669] Anonymous. IEEE Computing in Science and Engineering. Computer, 54 (9):79, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghoff:2021:IAB

[670] Christian Berghoff, Matthias Neu, and Arndt von Twickel. The interplay of AI and biometrics: Challenges and opportunities. Computer, 54(9):80– 85, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICGg

[671] Anonymous. IEEE Computer Graphics and Applications. Computer, 54 (9):85, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2021:FCS

[672] Dimitrios Serpanos and Konstantinos Katsigiannis. Fuzzing: Cyberphysical system testing for security and dependability. Computer, 54(9):86– 89, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2021:DAF

[673] James Bret Michael and Thomas C. Wingfield. Defensive AI: The future is yesterday. Computer, 54(9):90– 96, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IJBf

[674] Anonymous. IEEE Job Boards. Computer, 54(9):97, September 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2021:WDS

[675] Michael Zyda. Why did I say, yes to writing this bimonthly column? Computer, 54(9):98–102, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHh

[676] Anonymous. IEEE Annals of the History of Computing. Computer, 54(9): 102, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbe

[677] Anonymous. IEEE Computer Society Election. Computer, 54(9):103– 118, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IDV

[678] Anonymous. IEEE Division VIII Delegate/Director candidates. Computer, 54(9):119–120, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbf

[679] Anonymous. IEEE Computer Society. Computer, 54(9):C3, September 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbg

[680] Anonymous. IEEE Computer Society member benefits. Computer, 54 (9):C4, September 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCj

[681] Anonymous. [Front cover]. Computer, 54(10):C1, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IOJg

[682] Anonymous. IEEE Open Journal of the Computer Society. Computer, 54 (10):C2, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCj

[683] Anonymous. Table of contents. Computer, 54(10):1–2, October 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Mj

[684] Anonymous. Masthead. Computer, 54 (10):3, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhou:2021:CIS

[685] Qihua Zhou, Song Guo, Haodong Lu, Li Li, Minyi Guo, Yanfei Sun, and Kun Wang. A comprehensive inspection of the straggler problem. Computer, 54 (10):4–5, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:HSMb

[686] Anonymous. Highlights Society magazines. Computer, 54(10):6–8, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2021:YAe

[687] Erich Neuhold. 50 & 25 years ago. Computer, 54(10):9–12, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2021:CTTe

[688] Ergun Akleman. Computing through time. Computer, 54(10):11, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:TMS

[689] Jeffrey Voas. Tell me a story. Computer, 54(10):13, October 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2021:AIC

[690] Phil Laplante and Ben Amaba. Artificial intelligence in critical infrastructure systems. Computer, 54(10): 14–24, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISPh

[691] Anonymous. IEEE Security and Privacy subscribe. Computer, 54(10): 24, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raunak:2021:EAI

[692] M. S. Raunak and Rick Kuhn. Explainable artificial intelligence and machine learning [Guest Editors’ introduction]. Computer, 54(10):25–27, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Naiseh:2021:ERC

[693] Mohammad Naiseh, Deniz Cemiloglu, Dena Al Thani, Nan Jiang, and Raian Ali. Explainable recommendations and calibrated trust: Two systematic user errors. Computer, 54(10): 28–37, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Deo:2021:UUC

[694] Sahil Deo and Neha Sameer Sontakke. Usability, user comprehension, and perceptions of explanations for complex decision support systems in finance: a robo-advisory use case. Computer, 54 (10):38–48, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Psychoula:2021:EML

[695] Ismini Psychoula, Andreas Gutmann, Pradip Mainali, Sharon H. Lee, Paul Dunphy, and Fabien Petitcolas. Explainable machine learning for fraud detection. Computer, 54(10):49–59, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Srinivasan:2021:EAC

[696] Seshadhri Srinivasan, Pandarasamy Arjunan, Baihong Jin, Alberto L. Sangiovanni-Vincentelli, Zuraimi Sultan, and Kameshwar Poolla. Explainable AI for chiller fault-detection systems: Gaining human trust. Computer, 54(10):60–68, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DaPoian:2021:SAE

[697] Victoria DaPoian, Eric Lyness, William Brinckerhoff, Ryan Danell, Xiang Li, and Melissa Trainer. Science autonomy and the ExoMars Mission: Machine learning to help find life on Mars. Computer, 54(10):69–77, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holzinger:2021:THA

[698] Andreas T. Holzinger and Heimo Muller. Toward human AI interfaces to support explainability and causability in medical AI. Computer, 54(10): 78–86, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:DWG

[699] David Alan Grier. Do we go viral? Computer, 54(10):87–89, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHi

[700] Anonymous. IEEE Annals of the History of Computing. Computer, 54 (10):89, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mortati:2021:CWP

[701] Joseph Mortati and Erran Carmel. Can we prevent a technological arms race in university student cheating? Computer, 54(10):90–94, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harel:2021:SBA

[702] David Harel, Assaf Marron, and Raz Yerushalmi. Scenario-based algorithmics: Coding algorithms by automatic composition of separate concerns. Computer, 54(10):95–101, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Williams:2021:CAD

[703] Joseph Williams. The computing architectures donnybrook. Computer, 54 (10):102–105, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ligo:2021:ACI

[704] Alexandre K. Ligo, Alexander Kott, and Igor Linkov. Autonomous cyberdefense introduces risk: Can we manage the risk? Computer, 54(10): 106–110, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hannig:2021:OSH

[705] Frank Hannig and Jurgen Teich. Open source hardware. Computer, 54(10): 111–115, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICGh

[706] Anonymous. IEEE Computer Graphics and Applications. Computer, 54 (10):115, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vahdat:2021:NWC

[707] Amin Vahdat and Dejan Milojicic. The next wave in cloud systems architecture. Computer, 54(10):116– 120, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbh

[708] Anonymous. IEEE Computer Society membership. Computer, 54(10): 121, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2021:OBD

[709] Phil Laplante and Nir Kshetri. Open banking: Definition and description. Computer, 54(10):122–128, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbi

[710] Anonymous. IEEE Computing in Science & Engineering. Computer, 54 (10):128, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2021:VDV

[711] Norita Ahmad and Kevin Dias. Validating data validation. Computer, 54 (10):129–132, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2021:FI

[712] Joanna DeFranco. 12 flavors of IoT. Computer, 54(10):133–137, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2021:CAP

[713] Hal Berghel. A collapsing academy, Part II: How cancel culture works on the academy. Computer, 54(10):138– 144, October 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbj

[714] Anonymous. IEEE Computer Society. Computer, 54(10):C3, October 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IJBg

[715] Anonymous. IEEE Jobs Board. Computer, 54(10):C4, October 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCk

[716] Anonymous. Front cover. Computer, 54 (11):C1, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IOJh

[717] Anonymous. IEEE Open Journal of the Computer Society. Computer, 54 (11):C2, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCk

[718] Anonymous. Table of contents. Computer, 54(11):1–2, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Mk

[719] Anonymous. Masthead. Computer, 54 (11):3, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ju:2021:DS

[720] Tao Ju. Deforming with sparsity. Computer, 54(11):4–5, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CHSh

[721] Anonymous. Computer highlights Society magazines. Computer, 54(11): 6–8, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbk

[722] Anonymous. IEEE Computer Society member benefits. Computer, 54 (11):9, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:YAe

[723] Anonymous. 50 & 25 years ago. Computer, 54(11):10–12, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:CAY

[724] David Alan Grier. Counting the accomplishments, not the years. Computer, 54(11):13–14, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2021:AL

[725] Jeffrey Voas. Is anybody listening? Computer, 54(11):15–16, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Ib

[726] Anonymous. In this issue. Computer, 54(11):16, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ernst:2021:ASD

[727] Rolf Ernst and Xiaobo Sharon Hu. Autonomous systems design a virtual roundtable. Computer, 54(11): 17–25, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Id

[728] Anonymous. IQQC. Computer, 54 (11):25, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shore:2021:ZTW

[729] Malcolm Shore, Sherali Zeadally, and Astha Keshariya. Zero trust: The what, how, why, and when. Computer, 54(11):26–35, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gope:2021:CSD

[730] Prosanta Gope and Biplab Sikdar. A comparative study of design paradigms for PUF-based security protocols for IoT devices: Current progress, challenges, and future expectation. Computer, 54(11):36–46, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chernyshev:2021:CNA

[731] Maxim Chernyshev, Zubair Baig, and Sherali Zeadally. Cloud-native application security: Risks, opportunities, and challenges in securing the evolving attack surface. Computer, 54(11): 47–57, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Linares:2021:HCA

[732] Michael Linares, Nishant Aswani, Gary Mac, Chenglu Jin, Fei Chen, Nikhil Gupta, and Ramesh Karri. Hack3D: Crowdsourcing the assessment of cybersecurity in digital manufacturing. Computer, 54(11):58–67, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2021:TFT

[733] Christof Ebert and Ruschil Ray. Toward a formal traceability model for efficient security validation. Computer, 54(11):68–78, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHj

[734] Anonymous. IEEE Annals of the History of Computing. Computer, 54(11): 78, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:OMC

[735] David Alan Grier. Objects in the mirror are closer than they appear. Computer, 54(11):79–81, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICGi

[736] Anonymous. IEEE Computer Graphics and Applications. Computer, 54 (11):81, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2021:SL

[737] Phil Laplante. Software labels. Computer, 54(11):82–86, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbl

[738] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. Computer, 54(11):87, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cron:2021:ISM

[739] Adam Cron and Erik Jan Marinissen. IEEE Standard 1838 is on the move. Computer, 54(11):88– 94, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbm

[740] Anonymous. IEEE Computer Society call for papers. Computer, 54 (11):94, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2021:CAI

[741] Hsiao-Ying Lin. Colors of artificial intelligence. Computer, 54(11):95– 99, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITSc

[742] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 54 (11):99, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nikolakopoulos:2021:PLA

[743] George Nikolakopoulos and Ali Agha. Pushing the limits of autonomy for enabling the next generation of space robotics exploration missions. Computer, 54(11):100–103, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2021:THM

[744] James Bret Michael. Trusting human machine teaming. Computer, 54(11): 104–107, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ISPi

[745] Anonymous. IEEE Security and Privacy magazine subscription. Computer, 54(11):107, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhowmik:2021:HNF

[746] Achintya K. Bhowmik, David A. Fabry, Phillip Armour, Hal Berghel, Robert N. Charette, and John L. King. Hear, now, and in the future: Transforming hearing aids into multipurpose devices. Computer, 54(11):108– 120, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbn

[747] Anonymous. IEEE Computer Society. Computer, 54(11):C3, November 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbo

[748] Anonymous. IEEE Computer Society Jobs Board. Computer, 54(11): C4, November 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:FCl

[749] Anonymous. [Front cover]. Computer, 54(12):C1, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IOJi

[750] Anonymous. IEEE Open Journal of the Computer Society. Computer, 54 (12):C2, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:TCl

[751] Anonymous. Table of contents. Computer, 54(12):1–2, December 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Ml

[752] Anonymous. Masthead. Computer, 54 (12):3, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2021:DIL

[753] Qian Zhang. DROO: Integrated learning and optimization for edge computing offloading. Computer, 54(12): 4–6, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IAHk

[754] Anonymous. IEEE Annals of the History of Computing. Computer, 54 (12):6, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:CHSi

[755] Anonymous. Computer highlights Society magazines [elsewhere in the CS]. Computer, 54(12):7–9, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:YAf

[756] Anonymous. 50 & 25 years ago. Computer, 54(12):10–12, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2021:CTTf

[757] Ergun Akleman. Computing through time. Computer, 54(12):11, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shull:2021:YYR

[758] Forrest Shull and Megha Ben. 75 years young: Reflecting on a milestone year for the Society. Computer, 54(12): 13–17, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbp

[759] Anonymous. IEEE Computing in Science & Engineering. Computer, 54 (12):17, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbq

[760] Anonymous. IEEE Computer Society awards. Computer, 54(12):18– 23, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2021:RFF

[761] Joanna F. DeFranco and Jeffrey Voas. Reproducibility, fabrication, and falsification. Computer, 54(12):24–26, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:Ie

[762] Anonymous. In this issue. Computer, 54(12):25, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICb

[763] Anonymous. IEEE COMPSAC 2022. Computer, 54(12):27, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koopman:2021:ESA

[764] Philip Koopman, Benjamin Kuipers, William H. Widen, and Marilyn Wolf. Ethics, safety, and autonomous vehicles. Computer, 54(12):28–37, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbr

[765] Anonymous. IEEE Computer Society call for papers. Computer, 54 (12):37, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Henriquez:2021:ACA

[766] Valeria Henriquez, Ana M. Moreno, Jose A. Calvo-Manzano, and Tomas San Feliu. Agile–CMMI alignment: Contributions and to-dos for organizations. Computer, 54(12):38–49, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IIPi

[767] Anonymous. IEEE IT Professional call for articles. Computer, 54(12): 49, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bakirtzis:2021:CTC

[768] Georgios Bakirtzis, Eswaran Subrahmanian, and Cody H. Fleming. Compositional thinking in cyberphysical systems theory. Computer, 54(12):50– 59, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IPCh

[769] Anonymous. IEEE Pervasive Computing call for articles. Computer, 54 (12):59, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Leroy:2021:WSS

[770] Dorian Leroy, June Sallou, Johann Bourcier, and Benoit Combemale. When scientific software meets software engineering. Computer, 54(12): 60–71, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhu:2021:BBS

[771] Yan Zhu, Qian Guo, Hongjian Yin, Kaitai Liang, and Stephen S. Yau. Blockchain-based software architecture development for service requirements with smart contracts. Computer, 54 (12):72–80, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grier:2021:KWS

[772] David Alan Grier. “The known world” and its successors. Computer, 54(12): 81–83, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schmeelk:2021:PDG

[773] Suzanna Schmeelk. Perspectives on directing a graduate cybersecurity program. Computer, 54(12):84–87, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2021:PAW

[774] Mark Campbell. Protecting AI: We built the brains, but what about helmets? Computer, 54(12):88–91, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hanafi:2021:EAI

[775] Mamduh M. Hanafi, Nir Kshetri, and Ravi Sharma. Economics of artificial intelligence in the Gulf Cooperation Council countries. Computer, 54(12): 92–98, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITBd

[776] Anonymous. IEEE Transactions on Big Data. Computer, 54(12):98, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2021:OSD

[777] Dirk Riehle. The open source distributor business model. Computer, 54(12): 99–103, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2021:FCS

[778] Joanna F. DeFranco and Dimitrios Serpanos. The 12 flavors of cyberphysical systems. Computer, 54(12):104– 108, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITCc

[779] Anonymous. IEEE Transactions on Computers call for papers. Computer, 54(12):108, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2021:ODO

[780] Hal Berghel. The online disinformation opera. Computer, 54(12):109– 115, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ITSd

[781] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 54 (12):115, December 2021. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2021:WSS

[782] Michael Zyda. Will I see sensor-based games happen before I die? Computer, 54(12):116–120, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:ICSbs

[783] Anonymous. IEEE Computer Society. Computer, 54(12):C3, December 2021. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2021:IJBh

[784] Anonymous. IEEE Jobs Board. Computer, 54(12):C4, December 2021. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCa

[785] Anonymous. [Front cover]. Computer, 55(1):C1, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IOJa

[786] Anonymous. IEEE Open Journal of the Computer Society. Computer, 55 (1):C2, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCa

[787] Anonymous. Table of contents. Computer, 55(1):1–2, January 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Ma

[788] Anonymous. Masthead. Computer, 55(1):3, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebrahimi:2022:CSB

[789] Elham “Ellie” Ebrahimi. Can shooter bias be measured in immersive virtual reality? Computer, 55(1):4–5, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSa

[790] Anonymous. IEEE Computing in Sciences and Engineering. Computer, 55 (1):5, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSa

[791] Anonymous. Computer highlights Society magazines. Computer, 55(1):

6–8, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2022:YAa

[792] Erich Neuhold. 50 & 25 years ago. Computer, 55(1):9–11, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gropp:2022:ST

[793] William Bill Gropp and Forrest Shull. Succeeding together. Computer, 55(1): 12–17, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:C

[794] Nir Kshetri and Jeffrey Voas. Cryptojacking. Computer, 55(1):18–19, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coughlin:2022:DSM

[795] Tom Coughlin. Digital storage and memory. Computer, 55(1):20–29, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSb

[796] Anonymous. IEEE Computer Society call for papers. Computer, 55 (1):29, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Williams:2022:OFT

[797] Joseph Williams and Jeffrey Voas. Outlook: Forward thinking for 2022. Computer, 55(1):30–32, January 2022.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IIPa

[798] Anonymous. IEEE IT Professionals call for articles. Computer, 55 (1):32, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSc

[799] Anonymous. IEEE Computer Society Volunteer Service Awards. Computer, 55(1):33, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

VerWey:2022:OAI

[800] John VerWey. The other artificial intelligence hardware problem. Computer, 55(1):34–42, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Crawford:2022:EFS

[801] Rod F. Crawford and Andrew N. Sloss. Is expressiveness the future of software? Computer, 55(1):43–52, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IPCa

[802] Anonymous. IEEE Pervasive Computing call for articles. Computer, 55 (1):52, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kassab:2022:OBW

[803] Mohamad Kassab and Phillip A. Laplante. Open banking: What it is, where it’s at, and where it’s going.

Computer, 55(1):53–63, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2022:RAM

[804] Shaoshan Liu and Jean-Luc Gaudiot. Rise of the autonomous machines. Computer, 55(1):64–73, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:2022:CNA

[805] Keith W. Miller, Deborah G. Johnson, and Perry D. Drake. The Connection-Room: A new analogy for understanding the ethical dimensions of social media. Computer, 55(1):74–81, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2022:CCL

[806] Joanna F. DeFranco. From calculations to computations: A look back at the first computer. Computer, 55 (1):82–87, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2022:HEC

[807] Doron Drusinsky. On the high energy consumption of Bitcoin mining. Computer, 55(1):88–93, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McCalman:2022:AAF

[808] Lachlan McCalman, Daniel Steinberg, Grace Abuhamad, Marc-Etienne Brunet, Robert C. Williamson, and Richard Zemel. Assessing AI fairness in finance. Computer, 55(1): 94–97, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horton:2022:BAD

[809] Sarah Horton. Building an accessible digital world. Computer, 55(1): 98–102, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSd

[810] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. Computer, 55(1):103, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alexakos:2022:RIT

[811] Christos Alexakos, Konstantinos Votis, Dimitrios Tzovaras, and Dimitrios Serpanos. Reshaping the intelligent transportation scene: Challenges of an operational and safe Internet of Vehicles. Computer, 55(1): 104–107, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISPa

[812] Anonymous. IEEE Security and Privacy magazine subscribe. Computer, 55 (1):107, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Keckler:2022:A

[813] Steve Keckler and Dejan Milojicic. Accelerators. Computer, 55(1):108– 112, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITBa

[814] Anonymous. IEEE Transactions on Big Data. Computer, 55(1):112, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2022:CWR

[815] Michael Zyda. Can we replace all, or mostly all, in-class teaching with educational games? Computer, 55(1): 113–117, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITSa

[816] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 55 (1):117, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2022:SAV

[817] Hal Berghel. The state of the art in voting machine technology: Just how reliable are they? Computer, 55(1): 118–126, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSe

[818] Anonymous. IEEE Computer Society member benefits. Computer, 55 (1):127, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weiss:2022:CSC

[819] Joseph Weiss, Rob Stephens, and Nadine Miller. Control system cyber incidents are real — and current prevention and mitigation strategies are not working. Computer, 55(1):128–137, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCHa

[820] Anonymous. TC Chair highlight: Erich Neuhold of the Technical Community on Data Engineering (TCDE). Computer, 55(1):138, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSf

[821] Anonymous. IEEE Computer Society. Computer, 55(1):C3, January 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBa

[822] Anonymous. IEEE Jobs Board. Computer, 55(1):C4, January 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCb

[823] Anonymous. Front cover. Computer, 55(2):C1, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IOJb

[824] Anonymous. IEEE Open Journal of the Computer Society. Computer, 55 (2):C2, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCb

[825] Anonymous. Table of contents. Computer, 55(2):1–2, February 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Mb

[826] Anonymous. Masthead. Computer, 55 (2):3, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSb

[827] Anonymous. Computer highlights Society magazines. Computer, 55(2): 4–6, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2022:YAb

[828] Erich Neuhold. 50 & 25 years ago. Computer, 55(2):7–9, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CTTa

[829] Anonymous. Computing through time. Computer, 55(2):8, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2022:ZTA

[830] Phil Laplante and Jeffrey Voas. Zerotrust artificial intelligence? Computer, 55(2):10–12, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Ia

[831] Anonymous. In this issue. Computer, 55(2):11, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IAHa

[832] Anonymous. IEEE Annals of the History of Computing. Computer, 55 (2):12, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICCa

[833] Anonymous. IEEE Computer call for special issues. Computer, 55(2): 13, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Park:2022:MDS

[834] Jungheum Park, Hyunji Chung, and Joanna F. DeFranco. Multilayered diagnostics for smart cities. Computer, 55 (2):14–22, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICGa

[835] Anonymous. IEEE Computer Graphics and Applications. Computer, 55 (2):22, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yang:2022:DOU

[836] Zhou Yang, Spencer Bradshaw, Rattikorn Hewett, and Fang Jin. Discovering opioid use patterns from social media for relapse prevention. Computer, 55(2):23–33, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSg

[837] Anonymous. IEEE Computing in Science and Engineering. Computer, 55 (2):33, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Preusser:2022:EYA

[838] Thomas B. Preußer, Smitha Gautham, Abhi D. Rajagopala, Carl R. Elks, and Alexander Weiss. Everything you always wanted to know about embedded trace. Computer, 55(2):34– 43, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSh

[839] Anonymous. IEEE Computer Society call for papers. Computer, 55 (2):43, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Salama:2022:MCF

[840] Usama Salama, Lina Yao, and Hye-Young Paik. A multilevel collective framework for Internet of Things digital forensic investigation. Computer, 55 (2):44–53, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IIPb

[841] Anonymous. IEEE IT Professional call for articles. Computer, 55(2): 53, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmed:2022:PGB

[842] Salman Ahmed, Ya Xiao, Taejoong Tijay Chung, Carol Fung, Moti Yung, and Danfeng Daphne Yao. Privacy guarantees of Bluetooth low energy contact tracing: a case study on COVIDWISE. Computer, 55(2):54– 62, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bonnell:2022:CID

[843] Jerry Bonnell, Mitsunori Ogihara, and Yelena Yesha. Challenges and issues in data science education. Computer, 55 (2):63–66, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bernstein:2022:HAI

[844] Phillip Bernstein, Mark Greaves, Steve McConnell, and Clifford Pearson. Harnessing artificial intelligence to design healthy, sustainable, and equitable places. Computer, 55(2):67– 71, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IPCb

[845] Anonymous. IEEE Pervasive Computing call for articles. Computer, 55 (2):71, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:GAC

[846] Nir Kshetri and Maya Chhetri. Gender asymmetry in cybersecurity: Socioeconomic causes and consequences. Computer, 55(2):72–77, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISPb

[847] Anonymous. IEEE Security & Privacy subscribe. Computer, 55(2): 77, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2022:CPO

[848] Yuxia Zhang, Hao He, and Minghui Zhou. Commercial participation in OpenStack: Two sides of a coin. Computer, 55(2):78–84, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITBb

[849] Anonymous. IEEE Transactions on Big Data. Computer, 55(2):84, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mee:2022:AMD

[850] Sarah Jane Mee and Salam Daher. Alexa, M.D. Computer, 55(2):85– 89, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITCa

[851] Anonymous. IEEE Transactions on Computing. Computer, 55(2):89, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jovanovi:2022:EUR

[852] Mla an Jovanovi and Mia Schmitz. Explainability as a user requirement for artificial intelligence systems. Computer, 55(2):90–94, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2022:DSH

[853] Norita Ahmad, Areeba Hamid, and Vian Ahmed. Data science: Hype and reality. Computer, 55(2):95– 101, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITSb

[854] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 55 (2):101, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2022:LVV

[855] Doron Drusinsky, Matthew Litton, and James Bret Michael. Lightweight verification and validation of cyberphysical systems using machine-learned correctness properties. Computer, 55(2): 102–108, February 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSi

[856] Anonymous. IEEE Computer Society. Computer, 55(2):C3, February 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJ

[857] Anonymous. IEEE JobBoards. Computer, 55(2):C4, February 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCc

[858] Anonymous. [Front cover]. Computer, 55(3):C1, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IOJc

[859] Anonymous. IEEE Open Journal of the Computer Society. Computer, 55 (3):C2, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCc

[860] Anonymous. Table of contents. Computer, 55(3):1–2, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Mc

[861] Anonymous. Masthead. Computer, 55(3):3, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSc

[862] Anonymous. Computer highlights Society magazines. Computer, 55(3): 4–6, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:YAa

[863] Anonymous. 50 & 25 years ago. Computer, 55(3):7–10, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CPI

[864] Anonymous. Call for papers: 2022 IEEE World Congress on Services. Computer, 55(3):10, March 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:RPP

[865] Nir Kshetri and Jeffrey Voas. Ransomware: Pay to play? Computer, 55 (3):11–13, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2022:SCS

[866] Joanna F. DeFranco. Should cyberphysical systems and the Internet of Things get married? Computer, 55 (3):14–23, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mashkoor:2022:AIS

[867] Atif Mashkoor, Tim Menzies, Alexander Egyed, and Rudolf Ramler. Artificial intelligence and software engineering: Are we ready? Computer, 55 (3):24–28, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSj

[868] Anonymous. IEEE Computing in Science and Engineering. Computer, 55 (3):28, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSk

[869] Anonymous. IEEE Computer Society Diversity and Inclusive Fund. Computer, 55(3):29, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Washizaki:2022:SED

[870] Hironori Washizaki, Foutse Khomh, Yann-Gaël Guéhéneuc, Hironori Takeuchi, Naotake Natori, Takuo Doi, and Satoshi Okuda. Software-engineering design patterns for machine learning applications. Computer, 55(3):30–39, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSl

[871] Anonymous. IEEE Computer Society call for paper. Computer, 55 (3):39, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Warnett:2022:ADD

[872] Stephen John Warnett and Uwe Zdun. Architectural design decisions for the machine learning workflow. Computer, 55(3):40–51, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adigun:2022:CAI

[873] Jubril Gbolahan Adigun, Matteo Camilli, Michael Felderer, Andrea Giusti, Dominik T. Matt, Anna Perini, Barbara Russo, and Angelo Susi. Collaborative artificial intelligence needs stronger assurances driven by risks. Computer, 55(3):52–63, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IIPc

[874] Anonymous. IEEE IT Professionals call for papers. Computer, 55 (3):63, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khaliq:2022:TGT

[875] Zubair Khaliq, Sheikh Umar Farooq, and Dawood Ashraf Khan. Transformers for GUI testing: a plausible solution to automated test case generation and flaky tests. Computer, 55 (3):64–73, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IPCc

[876] Anonymous. IEEE Pervasive Computing call for articles. Computer, 55 (3):73, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siegmund:2022:GCC

[877] Norbert Siegmund, Johannes Dorn, Max Weber, Christian Kaltenecker, and Sven Apel. Green configuration: Can artificial intelligence help reduce energy consumption of configurable software systems? Computer, 55(3):74–81, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Freeman:2022:PCA

[878] Laura Freeman, Feras A. Batarseh, D. Richard Kuhn, M. S. Raunak, and Raghu N. Kacker. The path to a consensus on artificial intelligence assurance. Computer, 55(3):82–86, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Curtis:2022:MSQ

[879] Bill Curtis, Robert A. Martin, and Philippe-Emmanuel Douziech. Measuring the structural quality of software systems. Computer, 55(3):87–90, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ella:2022:HBS

[880] Lior Ella. How to build a scalable quantum controller. Computer, 55(3): 91–94, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:DPC

[881] Nir Kshetri and Elena Loukoianova. Data privacy considerations for central bank digital currencies in Asia Pacific countries. Computer, 55(3): 95–100, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISPc

[882] Anonymous. IEEE Security and Privacy magazine subscribe. Computer, 55(3):100, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sangwan:2022:TAI

[883] Raghvinder S. Sangwan, Youakim Badr, Satish Srinivasan, and Partha Mukherjee. On the testability of artificial intelligence and machine learning systems. Computer, 55(3):101– 105, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weiss:2022:CPC

[884] Joseph Weiss, Rob Stephens, and Nadine Miller. Changing the paradigm of control system cybersecurity. Computer, 55(3):106–116, March 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITBc

[885] Anonymous. IEEE Transactions on Big Data. Computer, 55(3):116, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2022:CAP

[886] Hal Berghel. A collapsing academy, Part III: Scientometrics and metric mania. Computer, 55(3):117–123, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITSc

[887] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 55 (3):123, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2022:LRE

[888] Michael Zyda. Let’s rename everything the metaverse! Computer, 55 (3):124–129, March 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:GSC

[889] Anonymous. 2021 Global Student Challenge Winners announced. Computer, 55(3):130–132, March 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSm

[890] Anonymous. IEEE Computer Society. Computer, 55(3):C3, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBb

[891] Anonymous. IEEE Jobs Boards. Computer, 55(3):C4, March 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCd

[892] Anonymous. Front cover. Computer, 55(4):C1, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IOJd

[893] Anonymous. IEEE Open Journal of the CS. Computer, 55(4):C2, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCd

[894] Anonymous. Table of contents. Computer, 55(4):1–2, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Md

[895] Anonymous. Masthead. Computer, 55(4):3, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Andre:2022:HVA

[896] Elisabeth André. How virtual agents learn to speak expressively. Computer, 55(4):4–5, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSd

[897] Anonymous. Computer highlights Society magazines. Computer, 55(4):6–8, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:YAb

[898] Anonymous. 50 & 25 years ago. Computer, 55(4):9–11, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2022:CTTa

[899] Ergun Akleman. Computing through time. Computer, 55(4):10, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2022:ASI

[900] Joanna F. DeFranco, Jeffrey Voas, and Nir Kshetri. Algorithms: Society’s invisible puppeteers. Computer, 55 (4):12–14, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Ib

[901] Anonymous. In this issue. Computer, 55(4):13, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thiruvathukal:2022:ECV

[902] George K. Thiruvathukal and Yung-Hsiang Lu. Efficient computer vision for embedded systems. Computer, 55 (4):15–19, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Treleaven:2022:FLP

[903] Philip Treleaven, Malgorzata Smietanka, and Hirsh Pithadia. Federated learning: The pioneering distributed machine learning and privacypreserving data technology. Computer, 55(4):20–29, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSn

[904] Anonymous. IEEE Computer Society call for papers. Computer, 55 (4):29, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tran:2022:DCR

[905] Dai Hoang Tran, Quan Z. Sheng, Wei Emma Zhang, Salma Abdalla Hamad, Nguyen Lu Dang Khoa, and Nguyen H. Tran. Deep conversational recommender systems: Challenges and opportunities. Computer, 55(4):30–39, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koshiyama:2022:AAM

[906] Adriano Koshiyama, Emre Kazim, and Philip Treleaven. Algorithm auditing: Managing the legal, ethical, and technological risks of artificial intelligence, machine learning, and associated algorithms. Computer, 55(4):40–50, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IAWa

[907] Anonymous. IEEE AI 10 to watch. Computer, 55(4):51, April 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaffer:2022:CAC

[908] Henry E. Schaffer. Computerized analysis can improve education while taking advantage of economies of scale. Computer, 55(4):52–55, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2022:IT

[909] Joanna F. DeFranco and Michael J. Metro. Internet of Telemedicine. Computer, 55(4):56–59, April 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:SFC

[910] Nir Kshetri. Scams, frauds, and crimes in the nonfungible token market. Computer, 55(4):60–64, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IDIa

[911] Anonymous. IEEE diversity & inclusion. Computer, 55(4):65, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gonzalez-Barahona:2022:SDM

[912] Jesus M. Gonzalez-Barahona, Daniel Izquierdo-Cortázar, and Gregorio Robles. Software development metrics with a purpose. Computer, 55(4): 66–73, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fitzgerald:2022:ECP

[913] Joshua Brian Fitzgerald. Elliptic curve pairings. Computer, 55(4):74–77, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bergman:2022:SP

[914] Keren Bergman, Ray Beausoleil, and Dejan Milojicic. Silicon photonics. Computer, 55(4):78–81, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2022:DSN

[915] Mark Campbell and Mla an Jovanovi . Digital self: The next evolution of the digital human. Computer, 55 (4):82–86, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITSd

[916] Anonymous. IEEE Transaction on Sustainable Computing. Computer, 55(4):86, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICP

[917] Anonymous. IEEE call for papers. Computer, 55(4):87, April 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chauhan:2022:RDA

[918] Preeti S. Chauhan and Nir Kshetri. The role of data and artificial intelligence in driving diversity, equity, and inclusion. Computer, 55(4):88–93, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISPd

[919] Anonymous. IEEE Security & Privacy. Computer, 55(4):93, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vucinic:2022:LAK

[920] Mališa Vučinić, Göran Selander, John Preuss Mattsson, and Thomas Watteyne. Lightweight authenticated key exchange with EDHOC. Computer, 55(4):94–100, April 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSo

[921] Anonymous. IEEE Computer Society. Computer, 55(4):C3, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBc

[922] Anonymous. IEEE Jobs Board. Computer, 55(4):C4, April 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCe

[923] Anonymous. Front cover. Computer, 55(5):C1, May 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IOJe

[924] Anonymous. IEEE Open Journal CS. Computer, 55(5):C2, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCe

[925] Anonymous. Table of contents. Computer, 55(5):1–2, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Me

[926] Anonymous. Masthead. Computer, 55(5):3, May 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parashar:2022:ARP

[927] Manish Parashar. Advancing reproducibility in parallel and distributed systems research. Computer, 55(5):4– 5, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSe

[928] Anonymous. Computer highlights Society magazines. Computer, 55(5):6–8, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2022:YAc

[929] Erich Neuhold. 50 & 25 years ago. Computer, 55(5):9–12, May 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICCb

[930] Anonymous. IEEE CS CFP. Computer, 55(5):12, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2022:D

[931] Jeffrey Voas. Dumbing down. Computer, 55(5):13, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Ic

[932] Anonymous. In this issue. Computer, 55(5):14, May 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IDIb

[933] Anonymous. IEEE diversity inclusion. Computer, 55(5):15, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2022:WNS

[934] Michael Zyda. We need something like Bell Labs for the games industry. Computer, 55(5):16–21, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jenkins:2022:TNT

[935] Barry L. Jenkins, John Scott, Francois Malassenet, and Kshitij Patil. Toward a new type of stream for interactive content. Computer, 55(5):22–30, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IPCd

[936] Anonymous. IEEE Pervasive Computing filler. Computer, 55(5):30, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bayliss:2022:DOD

[937] Jessica D. Bayliss. The data-oriented design process for game development. Computer, 55(5):31–38, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Radeta:2022:DLO

[938] Marko Radeta, Agustin Zuniga, Naser Hossein Motlagh, Mohan Liyanage, Ruben Freitas, Moustafa Youssef, Sasu Tarkoma, Huber Flores, and Petteri Nurmi. Deep learning and the oceans. Computer, 55(5):39–50, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Deng:2022:DRS

[939] Shuiguang Deng, Pan Lv, Ouwen Jin, Schahram Dustdar, Ying Li, De Ma, Zhaohui Wu, and Gang Pan. Darwin-S: a reference software architecture for brain-inspired computers. Computer, 55(5):51–63, May 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sobel:2022:CBP

[940] Ann Sobel and Ron Vetter. Cybersecurity best practices for CISE programs. Computer, 55(5):64–72, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grottke:2022:AFS

[941] Michael Grottke and Kishor S. Trivedi. Aging, fast and slow. Computer, 55 (5):73–75, May 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IRF

[942] Anonymous. IEEE over the rainbow filler. Computer, 55(5):75, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2022:LSA

[943] Hsiao-Ying Lin. Large-scale artificial intelligence models. Computer, 55 (5):76–80, May 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IGFa

[944] Anonymous. IT general filler. Computer, 55(5):80, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Komninos:2022:CSK

[945] Theodoros Komninos, Michael Paraskevas, Zacharoula Smyrnaiou, and Dimitrios Serpanos. Cyberphysical systems in K 12 education. Computer, 55(5):81–84, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shieh:2022:STA

[946] Shiuhpyng Shieh, Shanhsin Lee, and Mengru Tsai. Strategy and tactics against ransomware. Computer, 55 (5):85–92, May 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2022:QPS

[947] Hal Berghel. The QAnon phenomenon: The storm has always been among us. Computer, 55(5):93–100, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IAWb

[948] Anonymous. IEEE AI to watch filler. Computer, 55(5):101, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2022:HDG

[949] Michael Zyda. How do I get a position in the games industry? The FAQ. Computer, 55(5):102–108, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSp

[950] Anonymous. IEEE Computing Science Engineering filler. Computer, 55(5):108, May 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSq

[951] Anonymous. IEEE Computer Society. Computer, 55(5):C3, May 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBd

[952] Anonymous. IEEE Jobs Board filler. Computer, 55(5):C4, May 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCf

[953] Anonymous. Front cover. Computer, 55(6):C1, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:OCF

[954] Anonymous. OJ CS filler. Computer, 55(6):C2, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCf

[955] Anonymous. Table of contents. Computer, 55(6):1–2, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Mf

[956] Anonymous. Masthead. Computer, 55(6):3, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yang:2022:FSI

[957] Chia-Lin Yang. A forward speculative interference attack. Computer, 55(6):4–5, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSf

[958] Anonymous. Computer highlights Society magazines. Computer, 55(6):6–8, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:YAc

[959] Anonymous. 50 & 25 years ago. Computer, 55(6):9–11, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2022:CTTb

[960] Ergun Akleman. Computing through time. Computer, 55(6):10, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2022:RSM

[961] Joanna F. DeFranco and Jeffrey Voas. Revisiting software metrology. Computer, 55(6):12–14, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Id

[962] Anonymous. In this issue. Computer, 55(6):13, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2022:DDA

[963] James Bret Michael and Marko Orescanin. Developing and deploying artificial intelligence systems. Computer, 55(6):15–17, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:WHAa

[964] Anonymous. Watts S. Humphrey Award. Computer, 55(6):17, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Muiruri:2022:PIM

[965] Dennis Muiruri, Lucy Ellen Lwakatare, Jukka K. Nurminen, and Tommi Mikkonen. Practices and infrastructures for machine learning systems: an interview study in Finnish organizations. Computer, 55(6):18–29, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISR

[966] Anonymous. IEEE SP over the rainbow. Computer, 55(6):29, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2022:LLR

[967] Kunai Zhang, Yahui Liu, Weiqiang Zhang, and Shaoshan Liu. π-learner: a lifelong roadside learning framework for infrastructure augmented autonomous driving. Computer, 55(6):30–39, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICCc

[968] Anonymous. IEEEE CS CFP. Computer, 55(6):39, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Solaiman:2022:AML

[969] K. M. A. Solaiman, Tao Sun, Alina Nesen, Bharat Bhargava, and Michael Stonebraker. Applying machine learning and data fusion to the “Missing person” problem. Computer, 55 (6):40–55, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahvanooey:2022:IHD

[970] Milad Taleby Ahvanooey, Mark Xuefang Zhu, Wojciech Mazurczyk, and Malika Bendechache. Information hiding in digital textual contents: Techniques and current challenges. Computer, 55(6):56–65, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reviriego:2022:LGC

[971] Pedro Reviriego, Elena Merino-Gómez, and Fabrizio Lombardi. Latin and Greek in computing: Ancient words in a new world. Computer, 55(6): 66–72, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pastor:2022:TMT

[972] Oscar Pastor, Alfonso Pierantonio, and Gustavo Rossi. Teaching modeling in the time of agile development. Computer, 55(6):73–76, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zulkifli:2022:TRD

[973] Arief Zulkifli. TikTok in 2022: Revisiting data and privacy. Computer, 55(6):77–80, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vancsa:2022:FOO

[974] Ildikó Vancsa. The four opens: Open source beyond the code. Computer, 55 (6):81–84, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:SBG

[975] Nir Kshetri. El Salvador’s bitcoin gamble. Computer, 55(6):85–89, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Iaboni:2022:WTG

[976] Craig Iaboni, Himanshu Patel, Deepan Lobo, Ji won Choi, and Pramod Abichandani. Where are they going? Clustering event camera data to detect and track moving objects. Computer, 55(6):90–94, June 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IDa

[977] Anonymous. IEEE Dataport. Computer, 55(6):95, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2022:CBS

[978] Doron Drusinsky. Cryptographic– biometric self-sovereign personal identities. Computer, 55(6):96–102, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCHb

[979] Anonymous. TC Chair highlight: Wen-Mei Hwu and Hyesoon Kim of the Technical Community on Microprogramming and Microarchitecture (TCuARCH). Computer, 55(6):103– 104, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISS

[980] Anonymous. IEEE SP subscribe. Computer, 55(6):104, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSr

[981] Anonymous. IEEE Computer Society. Computer, 55(6):C3, June 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBe

[982] Anonymous. IEEE Jobs Board filler. Computer, 55(6):C4, June 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCg

[983] Anonymous. Front cover. Computer, 55(7):C1, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:D

[984] Anonymous. Dataport. Computer, 55(7):C2, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCg

[985] Anonymous. Table of contents. Computer, 55(7):1–2, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Mg

[986] Anonymous. Masthead. Computer, 55(7):3, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shen:2022:SVU

[987] Han-Wei Shen. Scientific visualization using neural networks. Computer, 55(7):4–6, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:WH

[988] Anonymous. Watts Humphrey. Computer, 55(7):6, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSg

[989] Anonymous. Computer highlights Society magazines. Computer, 55(7):7–9, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2022:YAd

[990] Erich Neuhold. 50 & 25 years ago. Computer, 55(7):10–12, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2022:SSW

[991] Joanna F. DeFranco and Jeffrey Voas. Smarter seeking wiser. Computer, 55 (7):13–14, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2022:PTA

[992] Phil Laplante and Dejan Milojicic. Predicting technologies that advance humanity. Computer, 55(7):15–17, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patterson:2022:CFM

[993] David Patterson, Joseph Gonzalez, Urs Hölzle, Quoc Le, Chen Liang, Lluis-Miquel Munguia, Daniel Rothchild, David R. So, Maud Texier, and Jeff Dean. The carbon footprint of machine learning training will plateau, then shrink. Computer, 55(7):18–28, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoefler:2022:CHD

[994] Torsten Hoefler, Ariel Hendel, and Duncan Roweth. The convergence of hyperscale data center and highperformance computing networks. Computer, 55(7):29–37, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huo:2022:SBA

[995] Yiming Huo. Space broadband access: The race has just begun. Computer, 55(7):38–45, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICCd

[996] Anonymous. IEEE-CS — call for papers filler. Computer, 55(7):45, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cheruvu:2022:UCH

[997] Ria Cheruvu. Unconventional concerns for human-centered artificial intelligence. Computer, 55(7):46–55, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marimuthu:2022:CDD

[998] Ramalatha Marimuthu, Shally Gupta, Larry Stapleton, Dominique Duncan, and Bozenna Pasik-Duncan. Challenging the digital divide: Factors affecting the availability, adoption, and acceptance of future technology in elderly user communities. Computer, 55 (7):56–66, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Badia:2022:MRT

[999] Rosa M. Badia, Ian Foster, and Dejan Milojicic. More real than real: The race to simulate everything. Computer, 55(7):67–72, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:RFa

[1000] Anonymous. Over the Rainbow filler. Computer, 55(7):72, July 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2022:TDT

[1001] Phil Laplante. Trusting digital twins. Computer, 55(7):73–77, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSs

[1002] Anonymous. IEEE Computer Society election. Computer, 55(7):77, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2022:PTA

[1003] James Bret Michael. Placing trust in automated software development processes. Computer, 55(7):78–81, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coughlin:2022:CND

[1004] Thomas M. Coughlin and William R.

Tonti. Computing nearer to data. Computer, 55(7):82–87, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICCe

[1005] Anonymous. IEEE CS CISE filler.

Computer, 55(7):87, July 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2022:CU

[1006] Dimitrios Serpanos and Theodoros Komninos. The cyberwarfare in Ukraine. Computer, 55(7):88–91, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICCf

[1007] Anonymous. IEEE CS CGA filler.

Computer, 55(7):91, July 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sipper:2022:AGI

[1008] Moshe Sipper. Artificial general intelligence: Pressure cooker or crucible? Computer, 55(7):92–96, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rowe:2022:AAI

[1009] Neil C. Rowe. Algorithms for artificial intelligence. Computer, 55(7): 97–102, July 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICD

[1010] Anonymous. IEEE CS diversity inclusion. Computer, 55(7):103, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2022:PTG

[1011] Hal Berghel. Prosperity theology goes online: Will this be a fifth great awakening? Computer, 55(7):104–110, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IAF

[1012] Anonymous. IEEE Annals filler. Computer, 55(7):110, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICA

[1013] Anonymous. IEEE CS AI — 10 to watch. Computer, 55(7):111, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2022:WMD

[1014] Michael Zyda. Weapons of mass distraction the America’s army game at 20. Computer, 55(7):112–122, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCHc

[1015] Anonymous. TC Chair highlight: Bryan Morse and Walter Scheirer of the Technical Community on Pattern Analysis and Machine Intelligence. Computer, 55(7):123–124, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSt

[1016] Anonymous. IEEE Computer Society. Computer, 55(7):C3, July 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBf

[1017] Anonymous. IEEE Jobs Board filler.

Computer, 55(7):C4, July 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCh

[1018] Anonymous. Front cover. Computer, 55(8):C1, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSu

[1019] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. Computer, 55(8):C2, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCi

[1020] Anonymous. Table of contents. Computer, 55(8):1–2, August 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Mh

[1021] Anonymous. Masthead. Computer, 55(8):3, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IRD

[1022] Anonymous. The international roadmap for devices and systems: a beacon for the electronics industry. Computer, 55(8):4, August 2022.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferner:2022:CQF

[1023] Clayton Ferner. Contributions of quantum factoring on quantum research. Computer, 55(8):5–7, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:WHAb

[1024] Anonymous. Watts Humphrey Award nomination. Computer, 55(8):7, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSh

[1025] Anonymous. Computer highlights Society magazines. Computer, 55(8): 8–10, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2022:YAe

[1026] Erich Neuhold. 50 & 25 years ago. Computer, 55(8):11–13, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2022:CTTc

[1027] Ergun Akleman. Computing through time. Computer, 55(8):12, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSv

[1028] Anonymous. IEEE Computer Society call for papers. Computer, 55 (8):13, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2022:WYC

[1029] Jeffrey Voas and Elena Loukoianova.

Warming your chicken. Computer, 55 (8):14–15, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2022:I

[1030] Jeffrey Voas. In this issue. Computer, 55(8):15, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parashar:2022:RR

[1031] Manish Parashar, Michael A. Heroux, and Victoria Stodden. Research reproducibility. Computer, 55(8):16–18, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSw

[1032] Anonymous. IEEE Computer Society election. Computer, 55(8):19, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foster:2022:CLC

[1033] Ian Foster and Carl Kesselman. CUF-Links: Continuous and Ubiquitous FAIRness Linkages for reproducible research. Computer, 55(8):20–30, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISPe

[1034] Anonymous. IEEE Security & Privacy subscribe. Computer, 55(8):30, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Halbert:2022:ARN

[1035] Martin D. Halbert. Advancing reproducibility at the NSF. Computer, 55 (8):31–39, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barba:2022:DRO

[1036] Lorena A. Barba. Defining the role of open source software in research reproducibility. Computer, 55(8):40–48, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoefler:2022:BDS

[1037] Torsten Hoefler. Benchmarking data science: 12 ways to lie with statistics and performance on parallel computers. Computer, 55(8):49–56, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Majeed:2022:PAD

[1038] Abdul Majeed and Seong Oun Hwang. A privacy-assured data lifecycle for epidemic-handling systems. Computer, 55(8):57–69, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flores:2022:CSQ

[1039] Huber Flores, Agustin Zuniga, Sasu Tarkoma, Leonardo Tonetto, Tristan Braud, Pan Hui, Yong Li, Mostafa Ammar, and Petteri Nurmi. Collaboration stability: Quantifying the success and failure of opportunistic collaboration. Computer, 55(8):70–81, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stefanov:2022:SSF

[1040] Krassen Stefanov. Software systems and frameworks for competency-based learning. Computer, 55(8):82–88, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:BCE

[1041] Nir Kshetri and Jeffrey Voas. Blockchain’s carbon and environmental footprints. Computer, 55(8):89–94, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSx

[1042] Anonymous. IEEE Computing in Science & Engineering. Computer, 55 (8):94, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2022:PPC

[1043] Mark Campbell. Privacy-preserving computation: Doomed to succeed. Computer, 55(8):95–99, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IIPd

[1044] Anonymous. IEEE IT Professional call for articles. Computer, 55(8):99, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2022:CSA

[1045] Joanna F. DeFranco and Bob Maley.

Closing the security agility gap. Computer, 55(8):100–102, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2022:CYT

[1046] James Bret Michael, George C. Dinolt, Frederick B. Cohen, and Duminda Wijesekera. Can you trust zero trust? Computer, 55(8):103– 105, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSy

[1047] Anonymous. IEEE Computer Society election. Computer, 55(8):106– 120, August 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IDV

[1048] Anonymous. IEEE Division V delegate/director candidates. Computer, 55(8):121–122, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCCa

[1049] Anonymous. Technical Committee Chair highlight: Jim Ahrens and Bongshin Lee of the visualization and graphics technical community. Computer, 55(8):123–124, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSz

[1050] Anonymous. IEEE Computer Society. Computer, 55(8):C3, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBg

[1051] Anonymous. IEEE Jobs Board. Computer, 55(8):C4, August 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCi

[1052] Anonymous. Front cover. Computer, 55 (9):C1, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IQW

[1053] Anonymous. IEEE Quantum Week.

Computer, 55(9):C2, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCj

[1054] Anonymous. Table of contents. Computer, 55(9):1–2, September 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Mi

[1055] Anonymous. Masthead. Computer, 55 (9):3, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schaumont:2022:ECC

[1056] Patrick Schaumont. Emerging computing challenges in the interaction of hardware and software. Computer, 55 (9):4–5, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSi

[1057] Anonymous. Computer highlights Society magazines. Computer, 55(9): 6–8, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2022:YAf

[1058] Erich Neuhold. 50 & 25 years ago. Computer, 55(9):9–12, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICCg

[1059] Anonymous. IEEE CS CFP filler. Computer, 55(9):12, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:COG

[1060] Nir Kshetri and Jeffrey Voas. Cyber Olympics: Game on! Computer, 55(9): 13–15, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Ie

[1061] Anonymous. In this issue. Computer, 55(9):14, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DiMarzoSerugendo:2022:DTC

[1062] Giovanna Di Marzo Serugendo, Anne-Françoise Cutting-Decelle, Laurent Guise, Thierry Cormenier, Imran Khan, and Luc Hossenlopp. Digital twins: From conceptual views to industrial applications in the electrical domain. Computer, 55(9):16–25, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yu:2022:AVD

[1063] Bo Yu, Chongyu Chen, Jie Tang, Shaoshan Liu, and Jean-Luc Gaudiot. Autonomous vehicles digital twin: A practical paradigm for autonomous driving system development. Computer, 55(9):26–34, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:RFb

[1064] Anonymous. Over the Rainbow filler.

Computer, 55(9):34, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kriens:2022:WML

[1065] Peter Kriens and Tim Verbelen. What machine learning can learn from software modularity. Computer, 55(9):35– 42, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TFa

[1066] Anonymous. TBD filler. Computer, 55 (9):42, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kertesz:2022:BCS

[1067] Attila Kertesz. Block the chain:

Software weapons of fighting against COVID-19. Computer, 55(9):43–53, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:SSF

[1068] Anonymous. SP subscribe filler. Computer, 55(9):53, September 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Puliafito:2022:SFS

[1069] Carlo Puliafito, Claudio Cicconetti, Marco Conti, Enzo Mingozzi, and Andrea Passarella. Stateful function as a service at the edge. Computer, 55(9): 54–64, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CFa

[1070] Anonymous. CISE filler. Computer, 55 (9):64, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cicconetti:2022:NCF

[1071] Claudio Cicconetti, Marco Conti, and Andrea Passarella. In-network computing with function as a service at the edge. Computer, 55(9):65– 73, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2022:SST

[1072] Hsiao-Ying Lin. Staying stronger together: Synergy from multiple sensors. Computer, 55(9):74–78, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parashar:2022:DST

[1073] Manish Parashar. Democratizing science through advanced cyberinfrastructure. Computer, 55(9):79–84, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CFb

[1074] Anonymous. CGA filler. Computer, 55 (9):84, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maglaras:2022:RSP

[1075] Leandros A. Maglaras, Mohamed Amine Ferrag, Helge Janicke, Nick Ayres, and Leandros Tassiulas. Reliability, security, and privacy in power grids. Computer, 55(9):85–88, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carleton:2022:AFS

[1076] Anita Carleton, Forrest Shull, and Erin Harper. Architecting the future of software engineering. Computer, 55(9): 89–93, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mehra:2022:TMD

[1077] Pankaj Mehra and Tom Coughlin.

Taming memory with disaggregation. Computer, 55(9):94–98, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:GF

[1078] Anonymous. AN general filler. Computer, 55(9):98, September 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bauer:2022:DLE

[1079] Bernhard Bauer, Mouadh Ayache, Saleh Mulhem, Meirav Nitzan, Jyotika Athavale, Rainer Buchty, and Mladen Berekovic. On the dependability lifecycle of electrical/electronic product development: The dual-cone V-model. Computer, 55(9):99–106, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TFb

[1080] Anonymous. TSUSC filler. Computer, 55(9):106, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2022:CAI

[1081] Hal Berghel. A collapsing academy IV: How did memory and gag laws gain admission? Computer, 55(9):107– 114, September 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TGF

[1082] Anonymous. TC generic filler. Computer, 55(9):114, September 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2022:UCA

[1083] James Bret Michael. Understanding conversational artificial intelligence. Computer, 55(9):115–119, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:PGF

[1084] Anonymous. PVC general filler. Computer, 55(9):119, September 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2022:BHI

[1085] Michael Zyda. Building a humanintelligent metaverse. Computer, 55 (9):120–128, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IGFb

[1086] Anonymous. IT general filler. Computer, 55(9):128, September 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSba

[1087] Anonymous. IEEE Computer Society.

Computer, 55(9):C3, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBh

[1088] Anonymous. IEEE Jobs Board filler.

Computer, 55(9):C4, September 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCj

[1089] Anonymous. Front cover. Computer, 55(10):C1, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbb

[1090] Anonymous. IEEE Computer Society benefits. Computer, 55(10):C2, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCk

[1091] Anonymous. Table of contents. Computer, 55(10):1–2, October 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Mj

[1092] Anonymous. Masthead. Computer, 55 (10):3, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Montuschi:2022:CAC

[1093] Paolo Montuschi, Jean-Michel Muller, and Florent de Dinechin. Computer arithmetic: Continuing a long and steady emergence. Computer, 55(10): 4–6, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbc

[1094] Anonymous. IEEE Computer Society call for papers. Computer, 55 (10):6, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSj

[1095] Anonymous. Computer highlights Society magazines. Computer, 55(10): 7–9, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2022:YAg

[1096] Erich Neuhold. 50 & 25 years ago. Computer, 55(10):10–12, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CTTb

[1097] Anonymous. Computing through time. Computer, 55(10):11, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IOJf

[1098] Anonymous. IEEE Open Journal of the Computer Society. Computer, 55 (10):13, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2022:CC

[1099] Jeffrey Voas. Crying cybersecurity!

Computer, 55(10):14–15, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:If

[1100] Anonymous. In this issue. Computer, 55(10):15, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2022:SSC

[1101] Joanna F. DeFranco and Nir Kshetri.

Software supply chains. Computer, 55 (10):16–17, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buchicchio:2022:ISC

[1102] Emanuele Buchicchio, Luca Grilli, Eros Capobianco, Salvatore Cipriano, and Diego Antonini. Invisible supply chain attacks based on Trojan source. Computer, 55(10):18–25, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Soto-Valero:2022:MDS

[1103] César Soto-Valero, Martin Monperrus, and Benoit Baudry. The multibillion dollar software supply chain of Ethereum. Computer, 55(10):26– 34, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cilardo:2022:VMS

[1104] Alessandro Cilardo, Marcello Cinque, Luigi De Simone, and Nicola Mazzocca. Virtualization over multiprocessor systems-on-chip: an enabling paradigm for the industrial Internet of Things. Computer, 55(10):35–47, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fokkink:2022:SSB

[1105] Wan Fokkink, Martijn Goorden, Joanna van de Mortel-Fronczak, Ferdie Reijnen, and Jacobus Rooda. Supervisor synthesis: Bridging theory and practice. Computer, 55(10):48–54, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISPf

[1106] Anonymous. IEEE Security and Privacy subscribe. Computer, 55(10): 54, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kimovski:2022:DML

[1107] Dragi Kimovski, Sasko Ristov, and Radu Prodan. Decentralized machine learning for intelligent health-care systems on the computing continuum. Computer, 55(10):55–65, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liebowitz:2022:DTU

[1108] Jay Liebowitz. Digital transformation for the university of the future: a perspective. Computer, 55(10):66– 69, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IPCe

[1109] Anonymous. IEEE Pervasive Computing call for articles. Computer, 55 (10):69, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kassab:2022:PIC

[1110] Mohamad Kassab and Joanna F. De-Franco. IoTility: a contemporary view. Computer, 55(10):70–74, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbd

[1111] Anonymous. IEEE Computer Society Volunteer Awards. Computer, 55 (10):75, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sharma:2022:TAA

[1112] Ravi Sharma, Norita Ahmad, Saqib Ali, Adil Bilal, Samar Fatima, Nir Kshetri, Fatima Abu Salem, and Prateek Sibal. Technomoral affordances of artificial intelligence in data-driven systems. Computer, 55(10):76–81, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITBd

[1113] Anonymous. IEEE Transaction on Big Data. Computer, 55(10):81, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Talia:2022:MIH

[1114] Domenico Talia. Machine intelligence for human health: a few noteworthy cases. Computer, 55(10):82–86, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IIPe

[1115] Anonymous. IEEE IT Professional call for articles. Computer, 55(10): 86, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbe

[1116] Anonymous. IEEE Computer Society social networking. Computer, 55 (10):87, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mitra:2022:SBA

[1117] Sayan Mitra and Hussein Sibai. Symmetry for boosting algorithmic proofs of cyberphysical systems. Computer, 55(10):88–93, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:ENT

[1118] Nir Kshetri. Economics of nonfungible tokens. Computer, 55(10):94– 99, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Faraboschi:2022:VWM

[1119] Paolo Faraboschi, Eitan Frachtenberg, Phil Laplante, Dejan Milojicic, and Roberto Saracco. Virtual worlds (Metaverse): From skepticism, to fear, to immersive opportunities. Computer, 55(10):100–106, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jovanovi:2022:GAI

[1120] Mla an Jovanovi and Mark Campbell. Generative artificial intelligence:

Trends and prospects. Computer, 55 (10):107–112, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICGb

[1121] Anonymous. IEEE Computer Graphics and Application. Computer, 55 (10):112, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Williams:2022:PE

[1122] Joseph Williams and Lisa Nee. Privacy engineering. Computer, 55(10): 113–118, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCHd

[1123] Anonymous. TC Chair highlight: Anne Benoit of the Technical Community on Parallel Processing (TCPP). Computer, 55(10):119–120, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IAHb

[1124] Anonymous. IEEE Annals of the History of Computing. Computer, 55 (10):120, October 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbf

[1125] Anonymous. IEEE Computer Society.

Computer, 55(10):C3, October 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBi

[1126] Anonymous. IEEE Job Board. Computer, 55(10):C4, October 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCk

[1127] Anonymous. Front cover. Computer, 55 (11):C1, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbg

[1128] Anonymous. IEEE Computer Society benefits. Computer, 55(11): C2, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCl

[1129] Anonymous. Table of contents. Computer, 55(11):1–2, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Mk

[1130] Anonymous. Masthead. Computer, 55 (11):3, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSk

[1131] Anonymous. Computer highlights Society magazines. Computer, 55(11): 4–7, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbh

[1132] Anonymous. IEEE Computer Society call for papers. Computer, 55 (11):7, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2022:YAh

[1133] Erich Neuhold. 50 & 25 years ago. Computer, 55(11):8–11, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITSe

[1134] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 55 (11):11, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2022:CFI

[1135] Charles Li and Jeffrey Voas. Carbon footprinting IT. Computer, 55(11): 12–14, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Ig

[1136] Anonymous. In this issue. Computer, 55(11):13, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICF

[1137] Anonymous. IEEE CS filler. Computer, 55(11):14, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IOJg

[1138] Anonymous. IEEE Open Journal of the Computer Society. Computer, 55 (11):15, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:2022:CLT

[1139] Anubha Gupta, Sachin Motwani, Abhishek Agarwal, Vishaal Udandarao, and Tanmoy Chakraborty. Changing landscape of technical education pedagogy from traditional to practical e-learning. Computer, 55(11):16– 28, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fritz:2022:PMW

[1140] Connor Fritz, Dylan Bray, Grace Lee, Christine Julien, Sheri Burson, Darla Castelli, Carol Ramsey, and Jamie Payton. Project moveSMART: When physical education meets computational thinking in elementary classrooms. Computer, 55(11):29–39, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITCb

[1141] Anonymous. IEEE Transactions on Computers call for papers. Computer, 55(11):39, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nassi:2022:OSR

[1142] Ben Nassi, Yaron Pirutin, Jacob Shams, Raz Swissa, Yuval Elovici, and Boris Zadov. Optical speech recovery from desktop speakers. Computer, 55 (11):40–51, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

LopesdosSantos:2022:TUI

[1143] Uélison Jean Lopes dos Santos, Cristiano André da Costa, André Henrique Mayer, Eduardo Souza dos Reis, Juan Eduardo Cruz Maldonado, Jorge Luis Victória Barbosa, Rodolfo Stoffel Antunes, Rodrigo da Rosa Righi, and Nelson Eduardo Flores. Trends in user identity and continuous authentication. Computer, 55(11):52– 61, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Perez-Sanchez:2022:DCI

[1144] Juanjo Pérez-Sánchez, Joaquín Nicolás Ros, Juan Manuel Carrillo de Gea, and José Luis Fernández-Alemán. DevOps certifications for IT professionals. Computer, 55(11):62–73, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Roman:2022:BDP

[1145] Dumitru Roman, Radu Prodan, Nikolay Nikolov, Ahmet Soylu, Mihhail Matskin, Andrea Marrella, Dragi Kimovski, Brian Elvesæter, Anthony Simonet-Boulogne, Giannis Ledakis, Hui Song, Francesco Leotta, and Evgeny Kharlamov. Big Data pipelines on the computing continuum: Tapping the dark data. Computer, 55(11): 74–84, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Armour:2022:UA

[1146] Phillip Glen Armour, Hal Berghel, Robert N. Charette, and John Leslie King. Ukraine aftershocks. Computer, 55(11):85–93, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oprea:2022:PAA

[1147] Alina Oprea, Anoop Singhal, and Apostol Vassilev. Poisoning attacks against machine learning: Can machine learning be trustworthy? Computer, 55(11):94–99, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITBe

[1148] Anonymous. IEEE Transactions on Big Data. Computer, 55(11):99, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2022:SMB

[1149] Hal Berghel. Social media and the banality of (online) crowds. Computer, 55 (11):100–105, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISPg

[1150] Anonymous. IEEE Security and Privacy subscription. Computer, 55(11): 105, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weiss:2022:UML

[1151] Joseph Weiss, Michael Hood, Nadine Miller, Chris Potorieko, and James Bret Michael. Using machine learning to work around the operational and cybersecurity limitations of legacy process sensors. Computer, 55 (11):106–111, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2022:SEA

[1152] Phil Laplante. Software engineering’s adolescent growing pains. Computer, 55(11):112–114, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IPCf

[1153] Anonymous. IEEE Pervasive Computing call for articles. Computer, 55(11): 114, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mamduhi:2022:IOT

[1154] Mohammad H. Mamduhi, Efe C. Balta, Alisa Rupenyan, and John Lygeros. Information-operation technology integration in industrial cyberphysical systems. Computer, 55(11):115–118, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbi

[1155] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. Computer, 55(11):119, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2022:PLG

[1156] Michael Zyda. Patent litigation in the games space. Computer, 55(11):120– 129, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:RP

[1157] Anonymous. Over The Rainbow podcast. Computer, 55(11):129, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCHe

[1158] Anonymous. TC Chair highlight:

Maiga Chang of the Technical Community on Learning Technology. Computer, 55(11):130–132, November 2022.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IIPf

[1159] Anonymous. IEEE IT Professional call for articles. Computer, 55(11): 132, November 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbj

[1160] Anonymous. IEEE Computer Society. Computer, 55(11):C3, November 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBj

[1161] Anonymous. IEEE Jobs Board. Computer, 55(11):C4, November 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:FCl

[1162] Anonymous. Front cover. Computer, 55 (12):C1, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbk

[1163] Anonymous. IEEE Computer Society benefits. Computer, 55(12): C2, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCm

[1164] Anonymous. Table of contents. Computer, 55(12):1–2, December 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:Ml

[1165] Anonymous. Masthead. Computer, 55 (12):3, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2022:EC

[1166] Jeffrey Voas, Patrick McConnell, and Nir Kshetri. Energy considerations. Computer, 55(12):4–5, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IAHc

[1167] Anonymous. IEEE Annals of the History of Computing. Computer, 55 (12):5, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:CHSl

[1168] Anonymous. Computer highlights Society magazines. Computer, 55(12): 6–8, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2022:YAi

[1169] Erich Neuhold. 50 & 25 years ago. Computer, 55(12):9–11, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2022:CTTd

[1170] Ergun Akleman. Computing through time. Computer, 55(12):10, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbl

[1171] Anonymous. IEEE Computer Society call for papers. Computer, 55 (12):17, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbm

[1172] Anonymous. IEEE Computer Society awards. Computer, 55(12):18– 23, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IRP

[1173] Anonymous. IEEE Over the Rainbow podcast. Computer, 55(12): 23, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:RR

[1174] Nir Kshetri and Jeffrey Voas. RegTech’s rise. Computer, 55(12):24–26, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITSf

[1175] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 55 (12):26, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolf:2022:CEE

[1176] Marilyn Wolf. Computer engineering education. Computer, 55(12):27– 37, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITBf

[1177] Anonymous. IEEE Transactions on Big Data. Computer, 55(12):37, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2022:SCC

[1178] Dimitrios Serpanos, Luis Muñoz, and Ioannis Chatzigiannakis. Smart and circular cities. Computer, 55(12):38– 41, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ISPh

[1179] Anonymous. IEEE Security and Privacy subscription. Computer, 55(12): 41, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eleftheriadou:2022:RCC

[1180] Dana Eleftheriadou, William Neale, Natalia Gkiaouri, Andrea Halmos, and Vasilis Liogkas. Redefining circular cities: Regulation, governance, infrastructure, and technology. Computer, 55 (12):42–53, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Motlagh:2022:ECC

[1181] Naser Hossein Motlagh, Lauri Lovén, Jacky Cao, Xiaoli Liu, Petteri Nurmi, Schahram Dustdar, Sasu Tarkoma, and Xiang Su. Edge computing: The computing infrastructure for the smart megacities of the future. Computer, 55 (12):54–64, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITCc

[1182] Anonymous. IEEE Transactions on Computers call for papers. Computer, 55(12):64, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matheu:2022:FCD

[1183] Sara N. Matheu, Enrique Mármol, José L. Hernández-Ramos, Antonio Skarmeta, and Gianmarco Baldini. Federated cyberattack detection for Internet of Things-enabled smart cities. Computer, 55(12):65–73, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kosti:2022:SCI

[1184] Zoran Kosti , Alex Angus, Zhengye Yang, Zhuoxu Duan, Ivan Seskar, Gil Zussman, and Dipankar Raychaudhuri. Smart city intersections: Intelligence nodes for future metropolises. Computer, 55(12):74–85, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kazmi:2022:IIT

[1185] Aqeel Kazmi, Saqib Rasool Chaudhry, Hadi Tabatabaee Malazi, Sangita Dhara, and Siobhán Clarke. Interoperable internet of things for smart transportation systems in circular cities. Computer, 55(12):86–97, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

De:2022:UML

[1186] Suparna De, Wei Wang, Usamah Jassat, and Klaus Moessner. Usage mining of the London Santander bikesharing system. Computer, 55(12):98– 108, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conde:2022:OBU

[1187] Javier Conde, Andres Munoz-Arcentales, Johnny Choque, Gabriel Huecas, and Álvaro Alonso. Overcoming the barriers of using linked open data in smart city applications. Computer, 55(12): 109–118, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anthopoulos:2022:BMC

[1188] Leonidas G. Anthopoulos and Marijn Janssen. Business model canvas for big and open linked data in smart and circular cities: Findings from Europe. Computer, 55(12):119–133, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:DAI

[1189] Nir Kshetri. Data and artificial intelligence as tools to fight poverty: Some notable applications in agriculture and health care. Computer, 55(12):134– 139, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Graciano-Neto:2022:SIA

[1190] Valdemar V. Graciano-Neto, Mohamad Kassab, Vinicius Lopes, Roberto Oliveira, and Renato Bulcão-Neto. The state of IoT for agribusiness in Brazil. Computer, 55(12):140–144, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trinkenreich:2022:WOS

[1191] Bianca Trinkenreich, Marco Aurelio Gerosa, and Igor Steinmacher. Women in open source: We need to talk about it. Computer, 55(12):145–149, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2022:TM

[1192] Nir Kshetri. A typology of metaverses. Computer, 55(12):150–155, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IPCg

[1193] Anonymous. IEEE Pervasive Computing call for articles. Computer, 55(12): 155, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ang:2022:NHH

[1194] James A. Ang and David J. Mountain. New horizons for high-performance computing. Computer, 55(12):156– 162, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IOJh

[1195] Anonymous. IEEE Open Journal Computer Society. Computer, 55(12): 163, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:JAV

[1196] Anonymous. Jyotika Athavale voted 2023 Computer Society President-Elect. Computer, 55(12):164–166, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IIPg

[1197] Anonymous. IEEE IT Professional call for articles. Computer, 55(12): 166, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:TCHf

[1198] Anonymous. TC Chair highlight:

Mondira (Mandy) Deb Pant of the technical community on very large scale integration. Computer, 55(12):167– 168, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ITCd

[1199] Anonymous. IEEE Transactions on Computing in Science. Computer, 55 (12):168, December 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:ICSbn

[1200] Anonymous. IEEE Computer Society.

Computer, 55(12):C3, December 2022. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2022:IJBk

[1201] Anonymous. IEEE Jobs Board. Computer, 55(12):C4, December 2022. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCa

[1202] Anonymous. Front cover. Computer, 56(1):C1, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSa

[1203] Anonymous. IEEE Computer Society benefits. Computer, 56(1):C2, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCa

[1204] Anonymous. Table of contents. Computer, 56(1):1–2, January 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Ma

[1205] Anonymous. Masthead. Computer, 56(1):3, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patel:2023:YCW

[1206] Nita Patel. It’s your commitment that will change the world. Computer, 56 (1):4–7, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:GWU

[1207] Anonymous. The George Washington University. Computer, 56(1):7, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSa

[1208] Anonymous. Computer highlights Society magazines. Computer, 56(1): 8–10, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAa

[1209] Erich Neuhold. 50 & 25 years ago. Computer, 56(1):11–13, January 2023.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:MSS

[1210] Joanna F. DeFranco and Jeffrey Voas.

Machining a sixth sense: Intuition. Computer, 56(1):14–15, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:O

[1211] Jeffrey Voas and Ron Vetter. Outlook 2023. Computer, 56(1):16–18, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IOJa

[1212] Anonymous. IEEE Open Journal of the Computer Society. Computer, 56 (1):19, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Byrd:2023:QCP

[1213] Gregory T. Byrd and Yongshan Ding.

Quantum computing: Progress and innovation. Computer, 56(1):20–29, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITSa

[1214] Anonymous. IEEE transactions on sustainable computing. Computer, 56 (1):29, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tennenhouse:2023:SIN

[1215] David Tennenhouse. Surprise-inspired networking. Computer, 56(1):30–41, January 2023. CODEN CPTRB4.

ISSN 0018-9162 (print), 1558-0814 (electronic).

Piroumian:2023:MDT

[1216] Vartan Piroumian. Making digital twins work. Computer, 56(1):42– 51, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITCa

[1217] Anonymous. IEEE transactions of computers. Computer, 56(1):51, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arlitt:2023:FW

[1218] Martin Arlitt, Thomas Coughlin, Paolo Faraboschi, Eitan Frachtenberg, Phil Laplante, Dejan Milojicic, Nita Patel, and Roberto Saracco. Future of the workforce. Computer, 56(1): 52–63, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITBa

[1219] Anonymous. IEEE Transactions of Big Data. Computer, 56(1):63, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:RTS

[1220] Nir Kshetri. Regulatory technology and supervisory technology: Current status, facilitators, and barriers. Computer, 56(1):64–75, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ISPa

[1221] Anonymous. IEEE security and privacy subscription. Computer, 56(1): 75, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:2023:FRT

[1222] Keith W. Miller. Facial recognition technology: Navigating the ethical challenges. Computer, 56(1):76– 81, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCa

[1223] Anonymous. IEEE pervasive computing call for papers. Computer, 56 (1):81, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robie:2023:RSI

[1224] John Robie, Alireza Famili, and Angelos Stavrou. Revisiting the spaceborne illuminators of opportunity for airborne object tracking. Computer, 56 (1):82–92, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:RPa

[1225] Anonymous. Over the Rainbow podcast. Computer, 56(1):92, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wendzel:2023:ART

[1226] Steffen Wendzel, Luca Caviglione, and Wojciech Mazurczyk. Avoiding research tribal wars using taxonomies. Computer, 56(1):93–96, January 2023.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCb

[1227] Anonymous. IT professional call for articles. Computer, 56(1):96, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2023:SSA

[1228] Hsiao-Ying Lin. Standing on the shoulders of AI giants. Computer, 56(1): 97–101, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICGa

[1229] Anonymous. IEEE computer graphics applications. Computer, 56(1): 101, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2023:ETL

[1230] Doron Drusinsky. Encrypting for time- and/or location-based decryption. Computer, 56(1):102–110, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IAHa

[1231] Anonymous. IEEE annals of the history of computing. Computer, 56 (1):110, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hands:2023:NIM

[1232] Jonmichael Hands and Tom Coughlin. New IEEE media sanitization specification enables circular economy for storage. Computer, 56(1):111–116, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oppliger:2023:HMC

[1233] Rolf Oppliger and Andreas Grünert.

How to manage cyber risks: Lessons learned from medical science. Computer, 56(1):117–119, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tauber:2023:PPF

[1234] Markus Tauber, Benedikt Gollan, Christoph Schmittner, and Philipp Knopf. Passive precision farming reshapes the agricultural sector. Computer, 56(1):120–124, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bash:2023:SFB

[1235] Cullen Bash, Ninad Hogade, Dejan Milojicic, Gourav Rattihalli, and Chandrakant D. Patel. Sustainability: Fundamentals-based approach to paying it forward. Computer, 56(1): 125–132, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Valerdi:2023:LFS

[1236] Ricardo Valerdi. Lessons from the father of software engineering. Computer, 56(1):133–136, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IIC

[1237] Anonymous. IEEE internet computing. Computer, 56(1):137, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Werthner:2023:DHT

[1238] Hannes Werthner, Allison Stanger, Viola Schiaffonati, Peter Knees, Lynda Hardman, and Carlo Ghezzi. Digital humanism: The time is now. Computer, 56(1):138–142, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSb

[1239] Anonymous. IEEE Computer Society volunteer service awards. Computer, 56 (1):143, January 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2023:DDS

[1240] Hal Berghel. The dictator’s dilemma and the sovereignty of cyberspace delusion. Computer, 56(1):144–149, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2023:MU

[1241] Michael Zyda. The metaverse university. Computer, 56(1):150–156, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSc

[1242] Anonymous. IEEE Computer Society. Computer, 56(1):C3, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICAa

[1243] Anonymous. 2023 IEEE Conference on Artificial Intelligence (IEEE CAI). Computer, 56(1):C4, January 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCb

[1244] Anonymous. Front cover. Computer, 56(2):C1, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSd

[1245] Anonymous. IEEE Computer Society member benefits. Computer, 56 (2):C2, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCb

[1246] Anonymous. Table of contents. Computer, 56(2):1–2, February 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Mb

[1247] Anonymous. Masthead. Computer, 56 (2):3, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vetter:2023:SII

[1248] Ron Vetter. Security issues for the internet of drones. Computer, 56(2): 4–5, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSe

[1249] Anonymous. IEEE Computer Society fall for papers. Computer, 56 (2):5, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSb

[1250] Anonymous. Computer highlights Society magazines. Computer, 56(2):

6–8, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAb

[1251] Erich Neuhold. 50 & 25 years ago. Computer, 56(2):9–11, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2023:CTTa

[1252] Ergun Akleman. Computing through time. Computer, 56(2):10, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:CNM

[1253] Joanna F. DeFranco and Jeffrey Voas. Computing new medicines. Computer, 56(2):12–13, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mariani:2023:TAPa

[1254] Riccardo Mariani, Francesca Rossi, Rita Cucchiara, Marco Pavone, Barnaby Simkin, Ansgar Koene, and Jochen Papenbrock. Trustworthy AI part 1. Computer, 56(2):14–18, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IOJb

[1255] Anonymous. IEEE Open Journal of the Computer Society. Computer, 56 (2):19, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Singh:2023:WCM

[1256] Amika M. Singh and Munindar P. Singh. Wasabi: a conceptual model for trustworthy artificial intelligence. Computer, 56(2):20–28, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fernandez-Llorca:2023:TAI

[1257] David Fernández-Llorca and Emilia Gómez. Trustworthy artificial intelligence requirements in the autonomous driving domain. Computer, 56(2): 29–39, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mousavi:2023:TAS

[1258] Mohammad Reza Mousavi, Ana Cavalcanti, Michael Fisher, Louise Dennis, Rob Hierons, Bilal Kaddouh, Effie Lai-Chong Law, Rob Richardson, Jan Oliver Ringer, Ivan Tyukin, and Jim Woodcock. Trustworthy autonomous systems through verifiability. Computer, 56(2):40–47, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chen:2023:AMR

[1259] Pin-Yu Chen and Payel Das. AI maintenance: a robustness perspective. Computer, 56(2):48–56, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ruospo:2023:SDL

[1260] Annachiara Ruospo, Ernesto Sanchez, Lucas Matana Luza, Luigi Dilillo, Marcello Traiola, and Alberto Bosio. A survey on deep learning resilience assessment methodologies. Computer, 56 (2):57–66, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Martinez-Vargas:2023:EWT

[1261] Anabel Martínez-Vargas, Jesús A. Gómez-Avilés, MA Cosío-León, and Ángel G. Andrade. Explaining the walking through of a team of algorithms. Computer, 56(2):67–81, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Speakman:2023:DSD

[1262] Skyler Speakman, Girmaw Abebe Tadesse, Celia Cintas, William Ogallo, Tanya Akumu, and Adebayo Oshingbesan. Detecting systematic deviations in data and models. Computer, 56(2): 82–92, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Haresamudram:2023:TLA

[1263] Kashyap Haresamudram, Stefan Larsson, and Fredrik Heintz. Three levels of AI transparency. Computer, 56(2): 93–100, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Richter:2023:WTF

[1264] Alexis Richter, Josh Yamamoto, and Eitan Frachtenberg. Why are there so few women in computer systems research? Computer, 56(2):101– 105, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rajanen:2023:OSU

[1265] Mikko Rajanen. Open source usability and user experience. Computer, 56(2): 106–110, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2023:MBI

[1266] Sorel Reisman. Metaverse-based instructional settings: Matters that matter. Computer, 56(2):111–115, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IRPa

[1267] Anonymous. IEEE Over the Rainbow podcast. Computer, 56(2):115, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Andruetto:2023:ACS

[1268] Claudia Andruetto, Rafia Inam, and Martin Törngren. Adding cyberphysical systems to the engineering education pi. Computer, 56(2):116– 120, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2023:WDS

[1269] Norita Ahmad and Areeba Hamid.

Will data science outrun the data scientist? Computer, 56(2):121– 128, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IAHb

[1270] Anonymous. IEEE Annals of the History of Computing. Computer, 56(2): 128, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2023:TAR

[1271] Mark Campbell. Tomorrow’s applications require IT operations that are autonomous, ubiquitous, and smarter in a word, invisible. Computer, 56(2): 129–133, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ISPb

[1272] Anonymous. IEEE Security and Privacy subscribe. Computer, 56(2): 133, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ford:2023:QCT

[1273] Pete Ford. The quantum cybersecurity threat may arrive sooner than you think. Computer, 56(2):134– 136, February 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:NMS

[1274] Nir Kshetri. National metaverse strategies. Computer, 56(2):137–142, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mariani:2023:FSS

[1275] Riccardo Mariani, Jyotika Athavale, Andrea Matteucci, Darren Galpin, and Rob Schaaf. Functional safety standards committee: Results and perspectives. Computer, 56(2):143–148, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSf

[1276] Anonymous. IEEE Computer Society. Computer, 56(2):C3, February 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IJBa

[1277] Anonymous. IEEE Job Board. Computer, 56(2):C4, February 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCc

[1278] Anonymous. Front cover. Computer, 56(3):C1, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSg

[1279] Anonymous. IEEE Computer Society member benefits. Computer, 56 (3):C2, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCc

[1280] Anonymous. Table of contents. Computer, 56(3):1–2, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Mc

[1281] Anonymous. Masthead. Computer, 56(3):3, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSc

[1282] Anonymous. Computer highlights Society magazines. Computer, 56(3): 4–6, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAc

[1283] Erich Neuhold. 50 & 25 years ago.

Computer, 56(3):7–9, March 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

deFranco:2023:PS

[1284] Joanna F. deFranco and Jeffrey Voas. Perfect strangers. Computer, 56(3): 10–12, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:Ia

[1285] Jeffrey Voas. In this issue. Computer, 56(3):11–12, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:TAI

[1286] Joanna F. DeFranco and Jeffrey Voas. Thoughts on automotive IoT. Computer, 56(3):13–18, March 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CCA

[1287] Anonymous. CAI 2023 — conference on artificial intelligence. Computer, 56(3):19, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhattacharjee:2023:BUD

[1288] Shameek Bhattacharjee and Sajal K. Das. Building a unified data falsification threat landscape for Internet of Things/cyberphysical systems applications. Computer, 56(3):20–31, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:2023:FSR

[1289] Varun Gupta, Jose Maria Fernandez-Crehuet, and Chetna Gupta. A freelancer-supported requirement engineering framework for software startups. Computer, 56(3):32–46, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dick:2023:RCE

[1290] Robert P. Dick, Rob Aitken, Jace Mogill, John Paul Strachan, Kirk Bresniker, Wei Lu, Yorie Nakahira, Zhiyong Li, Matthew J. Marinella, William Severa, A. Alec Talin, Craig M. Vineyard, Suhas Kumar, Christian Mailhiot, and Lennie Klebanoff. Research challenges for energy-efficient computing in automated vehicles. Computer, 56(3):47–58, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Allcroft:2023:OHM

[1291] Shane Allcroft, Mohammed Metwaly, Zachery Berg, Isha Ghodgaonkar, Fischer Bordwell, Xinxin Zhao, Xinglei Liu, Jiahao Xu, Subhankar Chakraborty, Vishnu Banna, Akhil Chinnakotla, Abhinav Goel, Caleb Tung, Gore Kao, Wei Zakharov, David A. Shoham, George K. Thiruvathukal, and Yung-Hsiang Lu. Observing human mobility internationally during COVID-19. Computer, 56 (3):59–69, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSh

[1292] Anonymous. IEEE Computer Society call for papers. Computer, 56 (3):69, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Venieris:2023:MDN

[1293] Stylianos I. Venieris, Christos-Savvas Bouganis, and Nicholas D. Lane. Multiple-deep neural network accelerators for next-generation artificial intelligence systems. Computer, 56(3): 70–79, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hewa:2023:BGT

[1294] Tharaka Hewa, Pawani Porambage, Anshuman Kalla, Diana Pamela Moya Osorio, Madhusanka Liyanage, and Mika Ylianttila. Blockchain and game theory convergence for network slice brokering. Computer, 56(3):80–91, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCc

[1295] Anonymous. IEEE Pervasive Computing call for articles. Computer, 56 (3):91, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krieger:2023:RAC

[1296] Don N. Krieger. Random access concatenated libraries and dd enable a short-latency, content-rich website on an inexpensive shared server. Computer, 56(3):92–100, March 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rosenfeld:2023:DGC

[1297] Ariel Rosenfeld. Is DBLP a good computer science journals database? Computer, 56(3):101–108, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:SPS

[1298] Anonymous. Security and privacy subscribe. Computer, 56(3):108, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Batarseh:2023:AW

[1299] Feras A. Batarseh and Ajay Kulkarni. AI for water. Computer, 56(3): 109–113, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CGAa

[1300] Anonymous. Computer Graphics & Applications. Computer, 56(3):113, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Perez-Cerrolaza:2023:UMD

[1301] Jon Perez-Cerrolaza, Francisco J. Cazorla, and Jaume Abella. Uncertainty management in dependable and intelligent embedded software. Computer, 56 (3):114–117, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reviriego:2023:BEG

[1302] Pedro Reviriego, Elena Merino-Gómez, and Fabrizio Lombardi. The Byzantine Empire and its generals: an ancient empire back to life in computer security. Computer, 56(3):118–124, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:RPb

[1303] Anonymous. Over the Rainbow podcast. Computer, 56(3):124, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Handy:2023:ODN

[1304] Jim Handy and Tom Coughlin. Optane’s dead: Now what? Computer, 56 (3):125–130, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSi

[1305] Anonymous. IEEE Computing in Science & Engineering. Computer, 56 (3):130, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2023:IEC

[1306] George Hurlburt. The importance of educating computational educators. Computer, 56(3):131–134, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2023:SVD

[1307] Hal Berghel. 911 swatting, VoIP, and doxxing. Computer, 56(3):135– 139, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2023:AP

[1308] Michael Zyda. American professor. Computer, 56(3):140–149, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IJBb

[1309] Anonymous. IEEE Job Boards. Computer, 56(3):150, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CC

[1310] Anonymous. CS connection. Computer, 56(3):151–152, March 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSj

[1311] Anonymous. IEEE Computer Society. Computer, 56(3):C3, March 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IOJc

[1312] Anonymous. IEEE Open Journal of the Computer Society. Computer, 56 (3):C4, March 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCd

[1313] Anonymous. Front cover. Computer, 56(4):C1, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSk

[1314] Anonymous. IEEE Computer Society benefits. Computer, 56(4):C2, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCd

[1315] Anonymous. Table of contents. Computer, 56(4):1–2, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Md

[1316] Anonymous. Masthead. Computer, 56(4):3, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSd

[1317] Anonymous. Computer highlights Society magazines. Computer, 56(4):4–6, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAd

[1318] Erich Neuhold. 50 & 25 years ago. Computer, 56(4):7–10, April 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2023:CTTb

[1319] Ergun Akleman. Computing through time. Computer, 56(4):8, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:GE

[1320] Jeffrey Voas. Good enough. Computer, 56(4):11–12, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lu:2023:SER

[1321] Qinghua Lu, Liming Zhu, Jon Whittle, and James Bret Michael. Software engineering for responsible AI. Computer, 56(4):13–16, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IAHc

[1322] Anonymous. IEEE Annals of the History of Computing. Computer, 56 (4):16, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSl

[1323] Anonymous. IEEE Computer Society Volunteer Service Awards. Computer, 56(4):17, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maalej:2023:TRE

[1324] Walid Maalej, Yen Dieu Pham, and Larissa Chazette. Tailoring requirements engineering for responsible AI. Computer, 56(4):18–27, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSm

[1325] Anonymous. IEEE Computer Society call for papers. Computer, 56 (4):27, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2023:QAT

[1326] Gongyuan Li, Bohan Liu, and He Zhang. Quality attributes of trustworthy artificial intelligence in normative documents and secondary studies: a preliminary review. Computer, 56 (4):28–37, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITSb

[1327] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 56(4):37, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bao:2023:SAR

[1328] Zhicheng Bao, Weishan Zhang, Xingjie Zeng, Hongwei Zhao, Cihao Dong, Yuming Nie, Yuru Liu, Yuange Liu, and Junzhong Wu. Software architecture for responsible artificial intelligence systems: Practice in the digitization of industrial drawings. Computer, 56(4):38–49, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITCb

[1329] Anonymous. IEEE Transactions on Computers call for papers. Computer, 56(4):49, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2023:TVE

[1330] Ding Li, Yan Liu, Jun Huang, and Zerui Wang. A trustworthy view on explainable artificial intelligence method evaluation. Computer, 56(4):50–60, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITBb

[1331] Anonymous. IEEE Transactions on Big Data. Computer, 56(4):60, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raja:2023:AAA

[1332] Ashwin Kumar Raja and Jianlong Zhou. AI accountability: Approaches, affecting factors, and challenges. Computer, 56(4):61–70, April 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Badran:2023:CEP

[1333] Khaled Badran, Pierre-Olivier Côté, Amanda Kolopanis, Rached Bouchoucha, Antonio Collante, Diego Elias Costa, Emad Shihab, and Foutse Khomh. Can ensembling preprocessing algorithms lead to better machine learning fairness? Computer, 56 (4):71–79, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2023:DDS

[1334] Shunli Zhang, Laurence T. Yang, Yue Zhang, Xiaokang Zhou, and Zongmin Cui. Data-driven system-level design framework for responsible cyberphysical-social systems. Computer, 56 (4):80–91, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IRPb

[1335] Anonymous. IEEE Over the Rainbow podcast. Computer, 56(4):91, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xia:2023:LLA

[1336] Xiaoya Xia, Wei Wang, Shengyu Zhao, Sikang Bian, and Rong Wang. Lessons learned from the Ant Group Open Source Program Office. Computer, 56 (4):92–97, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:LT

[1337] Joanna F. DeFranco and Phil Laplante.

Labeling things. Computer, 56(4): 98–101, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCd

[1338] Anonymous. IEEE Pervasive Computing call for articles. Computer, 56 (4):101, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shraga:2023:OAR

[1339] Roee Shraga and Avigdor Gal. One algorithm to rule them all: On the changing roles of humans in data integration. Computer, 56(4):102–109, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:NPT

[1340] Nir Kshetri, Jeffrey Voas, and Ralf Bebenroth. Nations play their chips. Computer, 56(4):110–114, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICAb

[1341] Anonymous. 2023 IEEE Conference on Artificial Intelligence. Computer, 56(4):115, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2023:AIS

[1342] Kin Long Kelvin Lee and Nalini Kumar. Artificial intelligence for scientific discovery at high-performance computing scales. Computer, 56(4): 116–122, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSn

[1343] Anonymous. IEEE Computer Society Watts S. Humphrey Software Quality Award. Computer, 56(4):122, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Faraboschi:2023:DTL

[1344] Paolo Faraboschi, Eitan Frachtenberg, Phil Laplante, Dejan Milojicic, and Roberto Saracco. Digital transformation: Lights and shadows. Computer, 56(4):123–130, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ISPc

[1345] Anonymous. IEEE Security and Privacy subscription. Computer, 56 (4):130, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Talia:2023:UCP

[1346] Domenico Talia and Paolo Trunfio. Urgent computing for protecting people from natural disasters. Computer, 56 (4):131–134, April 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Farahmand:2023:QCC

[1347] Fariborz Farahmand. Quantum cognition: a cognitive architecture for human-AI and in-memory computing. Computer, 56(4):135–138, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chan:2023:MHU

[1348] Henry Chan. 3E model: How to use OERs to enhance teaching/learning. Computer, 56(4):139–142, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wingfield:2023:WCE

[1349] Thomas C. Wingfield and James Bret Michael. Waterfall: Cascading effects of a strategic cyber campaign. Computer, 56(4):143–148, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSo

[1350] Anonymous. IEEE Computer Society. Computer, 56(4):C3, April 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICCa

[1351] Anonymous. IEEE Career Center. Computer, 56(4):C4, April 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCe

[1352] Anonymous. Front cover. Computer, 56(5):C1, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICAc

[1353] Anonymous. IEEE Conference on Artificial Intelligence. Computer, 56 (5):C2, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCe

[1354] Anonymous. Table of contents. Computer, 56(5):1–2, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Me

[1355] Anonymous. Masthead. Computer, 56(5):3, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSe

[1356] Anonymous. Computer highlights Society magazines. Computer, 56(5):4–6, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAe

[1357] Erich Neuhold. 50 & 25 years ago.

Computer, 56(5):7–10, May 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSp

[1358] Anonymous. IEEE Computer Society call for papers. Computer, 56 (5):10, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:SB

[1359] Joanna F. DeFranco, Nir Kshetri, and Jeffrey Voas. Sustainable blockchain? Computer, 56(5):11–12, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mariani:2023:TAPb

[1360] Riccardo Mariani, Francesca Rossi, Rita Cucchiara, Marco Pavone, Barnaby Simkin, Ansgar Koene, and Jochen Papenbrock. Trustworthy AI part II. Computer, 56(5):13–16, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSq

[1361] Anonymous. IEEE Computer Society benefits. Computer, 56(5):16, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSr

[1362] Anonymous. IEEE Computer Society Career Center. Computer, 56 (5):17, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hupont:2023:DHR

[1363] Isabelle Hupont, Marina Micheli, Blagoj Delipetrev, Emilia Gómez, and Josep Soler Garrido. Documenting high-risk AI: a European regulatory perspective. Computer, 56(5):18–27, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mazumder:2023:FTA

[1364] Sourav Mazumder, Subhankar Dhar, and Avinash Asthana. A framework for trustworthy AI in credit risk management: Perspectives and practices. Computer, 56(5):28–40, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brando:2023:NNR

[1365] Axel Brando, Isabel Serra, Enrico Mezzetti, Francisco J. Cazorla, Jon Perez-Cerrolaza, and Jaume Abella. On neural networks redundancy and diversity for their use in safety-critical systems. Computer, 56(5):41–50, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IRPc

[1366] Anonymous. IEEE Over the Rainbow podcast. Computer, 56(5):50, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tariq:2023:ETR

[1367] Shahroz Tariq, Sowon Jeon, and Simon S. Woo. Evaluating trustworthiness and racial bias in face recognition APIs using deepfakes. Computer, 56(5):51–61, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tong:2023:TAE

[1368] Jiarui Richard Tong and Timothy Xueqian Lee. Trustworthy AI that engages humans as partners in teaching and learning. Computer, 56(5):62–73, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CGAb

[1369] Anonymous. Computer Graphics & Applications. Computer, 56(5):73, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Degefe:2023:FAI

[1370] Elizabeth Demissie Degefe, Yulius Denny Prabowo, Krishna Savani, and Abhishek Sheetal. Functional analogies increase trust in black-box AI systems among lay consumers: The case of GeNose C-19. Computer, 56(5):74–83, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:WHAa

[1371] Anonymous. Watts Humphrey Award call for nominations. Computer, 56 (5):83, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Migliorelli:2023:ADL

[1372] Lucia Migliorelli, Simona Tiribelli, Alessandro Cacciatore, Benedetta Giovanola, Emanuele Frontoni, and Sara Moccia. Accountable deep-learning-based vision systems for preterm infant monitoring. Computer, 56(5): 84–93, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grudin:2023:CCH

[1373] Jonathan Grudin. ChatGPT and chat history: Challenges for the new wave. Computer, 56(5):94–100, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITSc

[1374] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 56 (5):100, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koopman:2023:UWI

[1375] Philip Koopman. UL 4600: What to include in an autonomous vehicle safety case. Computer, 56(5):101–104, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2023:FCA

[1376] Phil Laplante and Jeffrey Voas. Frameworking carbon-aware computing research. Computer, 56(5):105–108, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITBc

[1377] Anonymous. IEEE Transactions on Big Data. Computer, 56(5):108, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Majeed:2023:DCA

[1378] Abdul Majeed and Seong Oun Hwang.

Data-centric artificial intelligence, preprocessing, and the quest for transformative artificial intelligence systems development. Computer, 56(5): 109–115, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ISPd

[1379] Anonymous. IEEE Security and Privacy subscription. Computer, 56 (5):115, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Murugesan:2023:RGA

[1380] San Murugesan and Aswani Kumar Cherukuri. The rise of generative artificial intelligence and its impact on education: The promises and perils. Computer, 56(5):116–121, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IAHd

[1381] Anonymous. IEEE Annals of the History of Computing. Computer, 56 (5):121, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barua:2023:HSC

[1382] Anomadarshi Barua and Mohammad Abdullah Al Faruque. Hardware/ software co-design for sensor security. Computer, 56(5):122–125, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITCc

[1383] Anonymous. IEEE Transactions on Computers. Computer, 56(5):125, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Handy:2023:SAE

[1384] Jim Handy and Tom Coughlin. Semiconductor architectures enable compute in memory. Computer, 56(5): 126–129, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2023:CAE

[1385] Hal Berghel. ChatGPT and AIChat epistemology. Computer, 56(5):130– 137, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCe

[1386] Anonymous. IEEE Pervasive Computing call for articles. Computer, 56(5):137, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laporte:2023:SPP

[1387] Claude Y. Laporte, Guillaume Verret, and Mirna Muñoz. A software project that partially failed: a small organization that ignored the management and technical practices of software standards. Computer, 56(5):138–144, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2023:DWN

[1388] Michael Zyda. Do we need universities anymore? Computer, 56(5):

145–150, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSs

[1389] Anonymous. IEEE Computing in Science & Engineering. Computer, 56 (5):150, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSt

[1390] Anonymous. IEEE Computer Society. Computer, 56(5):C3, May 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IQWa

[1391] Anonymous. IEEE Quantum Week call for participation. Computer, 56 (5):C4, May 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCf

[1392] Anonymous. Front cover. Computer, 56(6):C1, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IQWb

[1393] Anonymous. IEEE quantum week.

Computer, 56(6):C2, June 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCf

[1394] Anonymous. Table of contents. Computer, 56(6):1–2, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Mf

[1395] Anonymous. Masthead. Computer, 56(6):3, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSf

[1396] Anonymous. Computer highlights Society magazines. Computer, 56(6):4–6, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAf

[1397] Erich Neuhold. 50 & 25 years ago. Computer, 56(6):7–10, June 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2023:CTTc

[1398] Ergun Akleman. Computing through time. Computer, 56(6):9, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:D

[1399] Jeffrey Voas. Disappearing. Computer, 56(6):11–12, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:Ib

[1400] Jeffrey Voas. In this issue. Computer, 56(6):12, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2023:GSC

[1401] Norita Ahmad and Joseph Williams. Green and sustainable computing. Computer, 56(6):13–15, June 2023.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSu

[1402] Anonymous. IEEE Computer Society has you covered! Computer, 56(6):15, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shin:2023:HCA

[1403] Donghee Shin and Emily Y. Shin.

Human-centered AI: A framework for green and sustainable AI. Computer, 56(6):16–25, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:WHSa

[1404] Anonymous. Watts S. Humphrey Software Quality Award nominations. Computer, 56(6):25, June 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Galantino:2023:APE

[1405] Stefano Galantino, Fulvio Risso, Vlad C. Coroamă, and Antonio Manzalini. Assessing the potential energy savings of a fluidified infrastructure. Computer, 56(6):26–34, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSv

[1406] Anonymous. IEEE Computer Society over the rainbow podcast. Computer, 56(6):34, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gutierrez:2023:GML

[1407] María Gutiérrez, Ma Ángeles Moraga, Félix García, and Coral Calero. Green-IN machine learning at a glance. Computer, 56(6):35–43, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Elagovan:2023:FDS

[1408] Ramanujam Elagovan, Thinagaran Perumal, and Shankar Krishnan. Fall detection systems at night. Computer, 56(6):44–51, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jensen:2023:BDT

[1409] Maria Hoffmann Jensen, Peter Axel Nielsen, and John Stouby Persson. From big data technologies to big data benefits. Computer, 56(6):52–61, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kirkham:2023:EPI

[1410] Reuben Kirkham. The ethical problems with IT experts in the legal system. Computer, 56(6):62–71, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2023:GAI

[1411] Norita Ahmad, San Murugesan, and Nir Kshetri. Generative artificial intelligence and the education sector. Computer, 56(6):72–76, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSw

[1412] Anonymous. IEEE Computer Society volunteer service awards. Computer, 56(6):77, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Levy:2023:CRU

[1413] Matthew L. Levy. Cybersecurity risks unique to open source and what communities are doing to reduce them. Computer, 56(6):78–83, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Speicher:2023:IDT

[1414] Terrance L. Speicher and Joanna F.

DeFranco. Industry 4.0 and digital twins. Computer, 56(6):84–88, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSx

[1415] Anonymous. IEEE Computer Society call for 2023 major awards nominations. Computer, 56(6):89, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shin:2023:DIA

[1416] Donghee Shin and Emily Y. Shin.

Data’s impact on algorithmic bias. Computer, 56(6):90–94, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICGb

[1417] Anonymous. IEEE Computer Graphics and Applications. Computer, 56 (6):94, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CIE

[1418] Anonymous. Celebrating the invention of electronic digital computing. Computer, 56(6):95, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2023:RDI

[1419] Mark Campbell. The road to decentralized identity: The techniques, promises, and challenges of tomorrow s digital identity. Computer, 56(6): 96–100, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:ECW

[1420] Nir Kshetri. The economics of chip war: China’s struggle to develop the semiconductor industry. Computer, 56 (6):101–106, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Whitmore:2023:IDS

[1421] Jim Whitmore. Information-driven security analysis: Tools and techniques for the study and practice of security engineering. Computer, 56(6): 107–120, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSy

[1422] Anonymous. IEEE Computer Society. Computer, 56(6):C3, June 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSz

[1423] Anonymous. IEEE Computer Society Career Center. Computer, 56 (6):C4, June 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCg

[1424] Anonymous. Front cover. Computer, 56(7):C1, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IQWc

[1425] Anonymous. IEEE Quantum Week 2023. Computer, 56(7):C2, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCg

[1426] Anonymous. Table of contents. Computer, 56(7):1–2, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Mg

[1427] Anonymous. Masthead. Computer, 56(7):3, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bodmann:2023:SEE

[1428] Pablo R. Bodmann, George Papadimitriou, Rubens L. Rech Junior, Dimitris Gizopoulos, and Paolo Rech. Soft error effects on Arm microprocessors: Early estimations versus chip measurements. Computer, 56(7):4–6, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSg

[1429] Anonymous. Computer highlights Society magazines. Computer, 56(7):7–9, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAg

[1430] Erich Neuhold. 50 & 25 years ago. Computer, 56(7):10–13, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2023:CTTd

[1431] Ergun Akleman. Computing through time. Computer, 56(7):11, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSba

[1432] Anonymous. IEEE Computer Society call for papers. Computer, 56 (7):13, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:ILI

[1433] Nir Kshetri, Joanna F. DeFranco, and Jeffrey Voas. Is it live, or is it deepfake? Computer, 56(7):14–16, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbb

[1434] Anonymous. IEEE Computer Society member benefits. Computer, 56(7):16, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2023:PTI

[1435] Dejan Milojicic and Phil Laplante. Predicting technology and its impact on humanity. Computer, 56(7):17–20, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michael:2023:BSF

[1436] Katina Michael. Biometric surveillance and the future of war. Computer, 56 (7):21–30, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:WHSb

[1437] Anonymous. Watts S. Humphrey Software Quality Award nominations. Computer, 56(7):30, July 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbc

[1438] Anonymous. IEEE Computer Society volunteer service awards. Computer, 56(7):31, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pasricha:2023:EDM

[1439] Sudeep Pasricha. Ethics for digital medicine: a path for ethical emerging medical IoT design. Computer, 56 (7):32–40, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Watkins:2023:RAA

[1440] Lanier Watkins, Denzel Hamilton, Tyler A. Young, Sebastian Zanlongo, Louis L. Whitcomb, Andrew R. Spielvogel, and Barbara Kobzik-Juul. The roles of autonomy and assurance in the future of uncrewed aircraft systems in low-altitude airspace operations. Computer, 56(7):41–53, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hammi:2023:SSC

[1441] Badis Hammi and Sherali Zeadally. Software supply-chain security: Issues and countermeasures. Computer, 56 (7):54–66, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoefler:2023:DCE

[1442] Torsten Hoefler, Duncan Roweth, Keith Underwood, Robert Alverson, Mark Griswold, Vahid Tabatabaee, Mohan Kalkunte, Surendra Anubolu, Siyuan Shen, Moray McLaren, Abdul Kabbani, and Steve Scott. Data center Ethernet and remote direct memory access: Issues at hyperscale. Computer, 56(7):67–77, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Landsman:2023:DDS

[1443] Dave Landsman and Karin Strauss. The DNA data storage model. Computer, 56(7):78–85, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:RPc

[1444] Anonymous. Over the Rainbow podcast. Computer, 56(7):85, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2023:FOS

[1445] Hal Berghel. Free online services and Gonzo capitalism. Computer, 56 (7):86–92, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bash:2023:M

[1446] Cullen Bash, Paolo Faraboschi, Eitan Frachtenberg, Phil Laplante, Dejan Milojicic, and Roberto Saracco. Megatrends. Computer, 56(7):93–100, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:2023:CEC

[1447] Keith W. Miller. Computer-enhanced communication among deaf, hard of hearing, and hearing people. Computer, 56(7):101–105, July 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Belli:2023:SCU

[1448] Laura Belli, Luca Davoli, and Gianluigi Ferrari. Smart city as an urban intelligent digital system: The case of Parma. Computer, 56(7):106–109, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ISPe

[1449] Anonymous. IEEE Security and Privacy subscription. Computer, 56 (7):109, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2023:TFB

[1450] Phil Laplante. Tribute to Fred Brooks.

Computer, 56(7):110–111, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Strawn:2023:SCS

[1451] George Strawn. Some computer science education reminiscences by an old educator. Computer, 56(7):112–115, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IAHe

[1452] Anonymous. IEEE Annals of the History of Computing. Computer, 56 (7):115, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2023:TSW

[1453] Hsiao-Ying Lin. Team with the strong, work with the wise. Computer, 56 (7):116–120, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2023:MLV

[1454] Doron Drusinsky, Matthew Litton, and James Bret Michael. Machine-learned verification and advance notice oracles for autonomous systems. Computer, 56 (7):121–130, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbd

[1455] Anonymous. IEEE Computing in Science and Engineering. Computer, 56 (7):130, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2023:HMC

[1456] Michael Zyda. Hubris from the metaverse to ChatGPT to TikTok’s time of tribulation. Computer, 56(7):131–136, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbe

[1457] Anonymous. IEEE Computer Society. Computer, 56(7):C3, July 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbf

[1458] Anonymous. IEEE Computer Society Career Center. Computer, 56 (7):C4, July 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCh

[1459] Anonymous. Front cover. Computer, 56(8):C1, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IQWd

[1460] Anonymous. IEEE quantum week.

Computer, 56(8):C2, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCh

[1461] Anonymous. Table of contents. Computer, 56(8):1–2, August 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Mh

[1462] Anonymous. Masthead. Computer, 56(8):3, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yang:2023:IAS

[1463] Hailong Yang, Yi Liu, Zhongzhi Luan, Lin Gan, Guangwen Yang, and Depei Qian. Input-aware sparse tensor storage format selection for optimizing MTTKRP. Computer, 56 (8):4–7, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbg

[1464] Anonymous. IEEE Computer Society benefits. Computer, 56(8):7, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSh

[1465] Anonymous. Computer highlights Society magazines. Computer, 56(8): 8–10, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbh

[1466] Anonymous. IEEE Computer Society call for papers. Computer, 56 (8):10, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAh

[1467] Erich Neuhold. 50 & 25 years ago. Computer, 56(8):11–14, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2023:CTTe

[1468] Ergun Akleman. Computing through time. Computer, 56(8):12, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICGc

[1469] Anonymous. IEEE Computer Graphics and Applications. Computer, 56 (8):14, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:DYR

[1470] Nir Kshetri, Joanna F. DeFranco, and Jeffrey Voas. Dusting off your resume. Computer, 56(8):15–18, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:Ic

[1471] Jeffrey Voas. In this issue. Computer, 56(8):16–17, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IRPd

[1472] Anonymous. IEEE Over the Rainbow podast. Computer, 56(8):18, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbi

[1473] Anonymous. IEEE Computer Society volunteer service awards. Computer, 56(8):19, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McGregor:2023:SMR

[1474] Sean McGregor. A scaled multiyear responsible artificial intelligence impact assessment. Computer, 56(8): 20–27, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hu:2023:EWS

[1475] Xiao Hu, Ziteng Jiao, Ayden Kocher, Zhenyu Wu, Junjie Liu, James C. Davis, George K. Thiruvathukal, and Yung-Hsiang Lu. Evolution of winning solutions in the 2021 low-power computer vision challenge. Computer, 56 (8):28–37, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:WHAb

[1476] Anonymous. Watts Humphrey Award nominations. Computer, 56(8):37, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ry:2023:TTL

[1477] Maciej Ry and Peter A. Gloor. There is truth in legends: Lessons for team performance from hackathons. Computer, 56(8):38–47, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tan:2023:DLM

[1478] Tianxiang Tan and Guohong Cao. Deep learning on mobile devices with neural processing units. Computer, 56 (8):48–57, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tan:2023:PQA

[1479] Teik Guan Tan, Jianying Zhou, Vishal Sharma, and Saraju P. Mohanty. Postquantum adversarial modeling: a user’s perspective. Computer, 56(8):58–67, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kathiravelu:2023:VSU

[1480] Pradeeban Kathiravelu, Nan Li, Nishchal Singi, Ananth Reddy Bhimireddy, Ryan Birmingham, Judy Wawira Gichoya, Hari Trivedi, Nabile Safdar, Ashish Sharma, and Puneet Sharma. Visualizing scanner utilization from MRI metadata and clinical data. Computer, 56(8):68–76, August 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:EGA

[1481] Nir Kshetri. The economics of generative artificial intelligence in the academic industry. Computer, 56(8): 77–83, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ISPf

[1482] Anonymous. IEEE Security and Privacy subscription. Computer, 56 (8):83, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Daniels:2023:BFF

[1483] Jeff Daniels. Building the factory of the future with the Industrial Internet of Things. Computer, 56(8): 84–88, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCf

[1484] Anonymous. IEEE Pervasive Computing call for articles. Computer, 56 (8):88, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:ED

[1485] Nir Kshetri. The economics of deepfakes. Computer, 56(8):89–94, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shin:2023:ANA

[1486] Donghee Shin and Norita Ahmad. Algorithmic nudge: an approach to designing human-centered generative artificial intelligence. Computer, 56(8): 95–99, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2023:DAI

[1487] Mark Campbell and Mla an Jovanovi . Detecting artificial intelligence: a new cyberarms race begins. Computer, 56(8):100–105, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbj

[1488] Anonymous. IEEE Computer Society election. Computer, 56(8):106– 121, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IDV

[1489] Anonymous. IEEE Division VIII delegate/director candidates. Computer, 56(8):122–123, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbk

[1490] Anonymous. IEEE Computer Society call for 2023 major awards nominations. Computer, 56(8):124, August 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbl

[1491] Anonymous. IEEE Computer Society. Computer, 56(8):C3, August 2023.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbm

[1492] Anonymous. IEEE Computer Society Career Center. Computer, 56 (8):C4, August 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCi

[1493] Anonymous. Front cover. Computer, 56 (9):C1, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbn

[1494] Anonymous. IEEE Computer Society has you covered! Computer, 56(9): C2, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCi

[1495] Anonymous. Table of contents. Computer, 56(9):1–2, September 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Mi

[1496] Anonymous. Masthead. Computer, 56 (9):3, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yang:2023:UCA

[1497] Chia-Lin Yang. Understanding computer architecture sustainability. Computer, 56(9):4–5, September 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbo

[1498] Anonymous. IEEE Computer Society — call for papers. Computer, 56 (9):5, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSi

[1499] Anonymous. Computer highlights Society magazines. Computer, 56(9): 6–8, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAi

[1500] Erich Neuhold. 50 & 25 years ago. Computer, 56(9):9–12, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2023:CTTf

[1501] Ergun Akleman. Computing through time. Computer, 56(9):10, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IIS

[1502] Anonymous. IEEE Intelligent Systems. Computer, 56(9):12, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:WWB

[1503] Joanna F. DeFranco, Nir Kshetri, and Jeffrey Voas. Are we writing for bots or humans? Computer, 56(9):13– 14, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:CT

[1504] Joanna F. DeFranco. Computing in telemedicine. Computer, 56(9):15– 17, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:RCSa

[1505] Anonymous. Over the Rainbow: 21st Century security & privacy podcast. Computer, 56(9):17, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ristov:2023:SES

[1506] Sashko Ristov, Marjan Gusev, Armin Hohenegger, Radu Prodan, Dimitar Mileski, Pano Gushev, and Goran Temelkov. Serverless electrocardiogram stream processing in federated clouds with lambda architecture. Computer, 56(9):18–27, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balzer:2023:ERT

[1507] Jeffrey R. Balzer, Julia Caviness, and Don Krieger. The evolution of realtime remote intraoperative neurophysiological monitoring. Computer, 56(9): 28–38, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kathiravelu:2023:TDH

[1508] Pradeeban Kathiravelu, Dalibor Fonovi , Tihana Galinac Grbac, Zachary Zaiman, Luís Veiga, Judy Wawira Gichoya, Saptarshi Purkayastha, and Babak Mahmoudi. The telehealth dilemma health-care deserts meet the Internet’s remote regions. Computer, 56(9):39–49, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCg

[1509] Anonymous. IEEE pervasive computing - call for articles. Computer, 56 (9):49, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grayson:2023:MPC

[1510] Nakia Grayson, Ronald Pulivarti, Bronwyn Hodges, Kevin Littlefield, Jeremy Miller, Chris Peloquin, Julie Snyder, Sue Wang, and Ryan Williams. Mitigating privacy and cybersecurity risks affecting telehealth remote patient monitoring ecosystems. Computer, 56 (9):50–61, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ISPg

[1511] Anonymous. IEEE Security and Privacy subscription. Computer, 56(9): 61, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Watanabe:2023:SCT

[1512] Takuro Watanabe, Chengshuo Xia, Koji Fujita, and Yuta Sugiura. Screening for carpal tunnel syndrome using daily behavior on mobile devices. Computer, 56(9):62–70, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:2023:NFL

[1513] Varun Gupta and Chetna Gupta. Navigating foreign language-taught degrees: Embracing artificial intelligence-driven language translators to overcome linguistic challenges. Computer, 56(9): 71–76, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IIPa

[1514] Anonymous. IEEE IT Professional — call for articles. Computer, 56(9): 76, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbp

[1515] Anonymous. IEEE Computer Society volunteer service awards. Computer, 56 (9):77, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2023:FFC

[1516] Hal Berghel. Fatal flaws in ChatAI as a content generator. Computer, 56(9): 78–82, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goodloe:2023:ASC

[1517] Alwyn E. Goodloe. Assuring safetycritical machine learning-enabled systems: Challenges and promise. Computer, 56(9):83–88, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbq

[1518] Anonymous. IEEE Computer Society diversity and inclusion. Computer, 56 (9):89, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2023:EAI

[1519] Hsiao-Ying Lin. Embedded artificial intelligence: Intelligence on devices. Computer, 56(9):90–93, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flynn:2023:OPB

[1520] David Flynn and Thomas Coughlin. Overcoming performance bottlenecks with a network file system in solid state drives. Computer, 56(9):94– 98, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IOJd

[1521] Anonymous. IEEE Open Journal of the Computer Society. Computer, 56 (9):99, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mylonas:2023:DTP

[1522] Georgios Mylonas, Athanasios Kalogeras, George Pavlidis, Aris Lalos, and Ana García-López. Digital twins for protecting cultural heritage against climate change. Computer, 56(9):100– 104, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbr

[1523] Anonymous. IEEE Computing in Science and Engineering. Computer, 56 (9):104, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2023:MJS

[1524] Michael Zyda. The metaverse is just a software layer. Computer, 56(9):105– 108, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbs

[1525] Anonymous. IEEE Computer Society.

Computer, 56(9):C3, September 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbt

[1526] Anonymous. IEEE Computer Society Career Center. Computer, 56(9): C4, September 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCj

[1527] Anonymous. Front cover. Computer, 56(10):C1, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbu

[1528] Anonymous. IEEE Computer Society D&I fund. Computer, 56(10): C2, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCj

[1529] Anonymous. Table of contents. Computer, 56(10):1–2, October 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Mj

[1530] Anonymous. Masthead. Computer, 56 (10):3, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Montuschi:2023:MCE

[1531] Paolo Montuschi, Yuan-Hao Chang, and Vincenzo Piuri. In-memory computing: The emerging computing topic in the post-von Neumann era. Computer, 56(10):4–6, October 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbv

[1532] Anonymous. IEEE Computer Society — call for papers. Computer, 56 (10):6, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSj

[1533] Anonymous. Computer highlights Society magazines. Computer, 56(10): 7–9, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAj

[1534] Erich Neuhold. 50 & 25 years ago. Computer, 56(10):10–13, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2023:CTTg

[1535] Ergun Akleman. Computing through time. Computer, 56(10):11, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbw

[1536] Anonymous. IEEE Computer Society has you covered! Computer, 56 (10):13, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2023:LEL

[1537] Hsiao-Ying Lin and Jeffrey Voas. Lower energy large language models (LLMs). Computer, 56(10):14–16, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:Id

[1538] Jeffrey Voas. In this issue. Computer, 56(10):15, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IS

[1539] Anonymous. IEEE Software. Computer, 56(10):16, October 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:CTP

[1540] Joanna F. DeFranco. Computing in telemedicine. Part II. Computer, 56 (10):17–19, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:RCSb

[1541] Anonymous. Over the Rainbow: 21st Century security & privacy podcast. Computer, 56(10):19, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ko:2023:RTT

[1542] Htet Yamin Ko Ko, Chitrini Mozumder, Nitin Kumar Tripathi, and Indrajit Pal. Recent trends in telecardiology and telemedicine: Where are they leading now? Computer, 56(10): 20–30, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Minhas:2023:DLT

[1543] Noman Nasir Minhas, Muhammad Wahaj Mubeen, and Hasnain Khawaja. Distributed ledger technologies for electronic health care: IOTA-based remote patient monitoring and telemedicine system. Computer, 56(10):31–39, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nikolov:2023:CBD

[1544] Nikolay Nikolov, Arnor Solberg, Radu Prodan, Ahmet Soylu, Mihhail Matskin, and Dumitru Roman. Container-based data pipelines on the computing continuum for remote patient monitoring. Computer, 56(10): 40–48, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hao:2023:VCS

[1545] Jingwei Hao, Senlin Luo, Limin Pan, and Chuantao Chen. VD-HEN: Capturing semantic dependencies for source code vulnerability detection with a hierarchical embedding network. Computer, 56(10):49–61, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

kov:2023:ECB

[1546] Novak Bo kov, Ari Trachtenberg, and David Starobinski. Enabling costbenefit analysis of data sync protocols. Computer, 56(10):62–71, October 2023.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IIPb

[1547] Anonymous. IEEE IT Professional — call for articles. Computer, 56 (10):71, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Andriole:2023:WCS

[1548] Stephen J. Andriole. What C-suites should expect from business schools. Computer, 56(10):72–76, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCh

[1549] Anonymous. IEEE Pervasive Computing — call for articles. Computer, 56 (10):76, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:MRD

[1550] Nir Kshetri, Jeffrey Voas, Yogesh K.

Dwivedi, Diana Rojas Torres, and Gayle O Keefe. Metaverse’s rise and decline. Computer, 56(10):77–82, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:SMM

[1551] Joanna F. DeFranco and Jeffrey Voas.

Social media’s mental health quagmire. Computer, 56(10):83–85, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maitra:2023:OCR

[1552] Subhamoy Maitra and Pantelimon Stănică. An overview and concerns related to quantum cryptanalysis: Past, present, and future. Computer, 56(10): 86–92, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ISPh

[1553] Anonymous. IEEE Security & Privacy. Computer, 56(10):92, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Faraboschi:2023:AGI

[1554] Paolo Faraboschi, Eitan Frachtenberg, Phil Laplante, Dejan Milojicic, and Roberto Saracco. Artificial general intelligence: Humanity s downturn or unlimited prosperity. Computer, 56(10): 93–101, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CSE

[1555] Anonymous. Computing in Science & Engineering. Computer, 56(10): 101, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chauhan:2023:RSK

[1556] Preeti S. Chauhan and Nir Kshetri. Ransomware as a service kit: a novel cybercrime strategy to monetize victims data. Computer, 56(10):102– 106, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IAHf

[1557] Anonymous. IEEE Annals of the History of Computing. Computer, 56 (10):106, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbx

[1558] Anonymous. IEEE Computer Society Career Center. Computer, 56(10): 107, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2023:SF

[1559] Joanna F. DeFranco, Nir Kshetri, and Jeffrey Voas. Sensoring the farm. Computer, 56(10):108–112, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSby

[1560] Anonymous. IEEE Computer Society.

Computer, 56(10):C3, October 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSbz

[1561] Anonymous. IEEE Computer Society volunteer service awards. Computer, 56 (10):C4, October 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCk

[1562] Anonymous. Front cover. Computer, 56 (11):C1, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICM

[1563] Anonymous. IEEE Computer magazine seeks Editor in Chief. Computer, 56 (11):C2, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCk

[1564] Anonymous. Table of contents. Computer, 56(11):1–2, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Mk

[1565] Anonymous. Masthead. Computer, 56 (11):3, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2023:ETD

[1566] Shixia Liu. Enhancing training data quality with visual analytics. Computer, 56(11):4–6, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSca

[1567] Anonymous. IEEE Computer Society call for papers. Computer, 56 (11):6, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSk

[1568] Anonymous. Computer highlights Society magazines. Computer, 56(11): 7–9, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAk

[1569] Erich Neuhold. 50 & 25 years ago. Computer, 56(11):10–13, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2023:CTTh

[1570] Ergun Akleman. Computing through time. Computer, 56(11):11, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:HSD

[1571] Jeffrey Voas. Hiding in the shadows:

Do shadows still exist? Computer, 56 (11):14–16, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:Ie

[1572] Jeffrey Voas. In this issue. Computer, 56(11):15, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICScb

[1573] Anonymous. IEEE Computer Society member benefits. Computer, 56 (11):16, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wolf:2023:CEE

[1574] Marilyn Wolf and Jan Madsen. Computer engineering education. Computer, 56(11):17–19, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICGd

[1575] Anonymous. IEEE Computer Graphics and Applications. Computer, 56 (11):19, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schoeberl:2023:CDV

[1576] Martin Schoeberl and Luca Pezzarossa.

Chip design and verification in a computer engineering education. Computer, 56(11):20–29, November 2023.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Farcas:2023:TEA

[1577] Allen-Jasmin Farcas and Radu Marculescu. Teaching edge AI at the undergraduate level: a hardware software co-design approach. Computer, 56(11): 30–38, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2023:AEM

[1578] Mingxuan Liu, Jianhua Gu, Zhenhao Guo, and Tianhai Zhao. AempOS: an educational multiprocessor operating system. Computer, 56(11):39– 47, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wu:2023:MNB

[1579] Guangfu Wu, Enning Li, and Taigui Wei. Multimaster node Byzantine fault-tolerant consensus algorithm based on consistent hash algorithm. Computer, 56(11):48–63, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICScc

[1580] Anonymous. IEEE Computing in Science and Engineering. Computer, 56 (11):63, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dreesbach:2023:ILA

[1581] Tobias Dreesbach, Jan Heinrich Beinke, Arne Dethloff, Fabian van Dülmen, and Oliver Thomas. Impact and limitations of AR-based guidance for assembly workers. Computer, 56 (11):64–72, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IIPc

[1582] Anonymous. IEEE IT Professional call for articles. Computer, 56(11): 72, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wu:2023:CVS

[1583] Junrui Wu, Wenyong Wang, Sai Zou, Wei Ni, Lisheng Huang, and Yong Tang. Can video as a service paradigm lead to the future Internet of Video Things? Computer, 56(11):73– 84, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IRPe

[1584] Anonymous. IEEE Over the Rainbow podcast. Computer, 56(11): 84, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parashar:2023:SDA

[1585] Manish Parashar, Tess DeBlanc-Knowles, Erwin Gianchandani, and Lynne E. Parker. Strengthening and democratizing artificial intelligence research and development. Computer, 56 (11):85–90, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITSd

[1586] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 56 (11):90, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kassab:2023:SCP

[1587] Mohamad Kassab, Ben Amaba, Edward S. Pound, and Brian Fox. Is the solution for cybercrime also a path to greater productivity? Computer, 56 (11):91–94, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2023:TTP

[1588] Hal Berghel. The thousand talents program prosecutions in context. Computer, 56(11):95–102, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jovanovi:2023:CAM

[1589] Mla an Jovanovi and Mark Campbell. Connecting AI: Merging large language models and knowledge graph. Computer, 56(11):103–108, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2023:WYS

[1590] Sorel Reisman. What you should expect from B-schools. Computer, 56(11): 109–112, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Azzoni:2023:CSC

[1591] Paolo Azzoni, Jean-Luc di Paola Galloni, and Dimitrios Serpanos. Chips shortage and chips acts: The inside view. Computer, 56(11):113– 117, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IAHg

[1592] Anonymous. IEEE Annals of the History of Computing. Computer, 56(11): 117, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Farahmand:2023:SEA

[1593] Fariborz Farahmand. A system engineering approach to AI security and safety. Computer, 56(11):118– 122, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITBd

[1594] Anonymous. IEEE Transactions on Big Data. Computer, 56(11):122, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IOJe

[1595] Anonymous. IEEE Open Journal of the Computer Society. Computer, 56 (11):123, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2023:GTT

[1596] Michael Zyda. Game technology takes over the film industry. Computer, 56 (11):124–128, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCi

[1597] Anonymous. IEEE Pervasive Computing call for articles. Computer, 56(11): 128, November 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICScd

[1598] Anonymous. IEEE Computer Society. Computer, 56(11):C3, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICCb

[1599] Anonymous. IEEE career center. Computer, 56(11):C4, November 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:FCl

[1600] Anonymous. Front cover. Computer, 56 (12):C1, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSce

[1601] Anonymous. IEEE Computer Society search for a new Editor In Chief. Computer, 56(12):C2, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:TCl

[1602] Anonymous. Table of contents. Computer, 56(12):1–2, December 2023. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:Ml

[1603] Anonymous. Masthead. Computer, 56 (12):3, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yang:2023:DYH

[1604] Simin Yang, Mathieu Barthet, Courtney Reed, and Elaine Chew. Do you hear what I hear? Computer, 56(12):

4–6, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:CHSl

[1605] Anonymous. Computer highlights Society magazines. Computer, 56(12): 7–9, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2023:YAl

[1606] Erich Neuhold. 50 & 25 years ago.

Computer, 56(12):10–13, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2023:CTTi

[1607] Ergun Akleman. Computing through time. Computer, 56(12):11, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICScf

[1608] Anonymous. IEEE Computer Society call for papers. Computer, 56 (12):13, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patel:2023:ECL

[1609] Nita Patel. Engaging, changing, leading: Opportunities that shape our world. Computer, 56(12):14–17, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ISPi

[1610] Anonymous. IEEE Security and Privacy subscription. Computer, 56(12):

17, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICScg

[1611] Anonymous. IEEE Computer Society awards. Computer, 56(12):18– 24, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITSe

[1612] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 56 (12):24, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2023:PD

[1613] Jeffrey Voas. Propagating disinformation. Computer, 56(12):25–26, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bojanova:2023:EDT

[1614] Irena Bojanova, Rick Kuhn, and Jeffrey Voas. Emerging disruptive technologies. Computer, 56(12):27–30, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ITBe

[1615] Anonymous. IEEE Transactions on Big Data. Computer, 56(12):30, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSch

[1616] Anonymous. IEEE Computer Society member benefits. Computer, 56 (12):31, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Duarte:2023:APR

[1617] Carlos Henrique C. Duarte. Authorship and peer review in the era of artificial intelligence. Computer, 56(12): 32–41, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xia:2023:VSD

[1618] Chengshuo Xia and Yuta Sugiura. Virtual sensors with 3D digital human motion for interactive simulation. Computer, 56(12):42–54, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IPCj

[1619] Anonymous. IEEE Pervasive Computing call for articles. Computer, 56 (12):54, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fleury:2023:AVR

[1620] Sylvain Fleury, Fabien Bernard, Raphaël Paquin, Philippe Blanchard, and Simon Richir. Augmented and virtual reality simulation in industry. Computer, 56(12):55–64, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lin:2023:AII

[1621] Ying-Dar Lin, Yuan-Cheng Lai, Didik Sudyana, and Ren-Hung Hwang. Artificial intelligence for internet of things as a service: Small or big data, private or public model, centralized or federated learning? Computer, 56(12):

65–79, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSci

[1622] Anonymous. IEEE Computers In Science & Engineering. Computer, 56 (12):79, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mulligan:2023:RAN

[1623] Catherine Mulligan, Tiago Silva, and Joaquim Jorge. A research agenda for NFTs. Computer, 56(12):80– 90, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kassab:2023:UPI

[1624] Mohamad Kassab and Joanna F. De-Franco. Unleashing the potential: Integrating ChatGPT and the Internet of Things for enhanced user experiences and automation. Computer, 56(12): 91–94, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2023:FOS

[1625] Dirk Riehle. The future of the open source definition. Computer, 56(12): 95–99, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IRPf

[1626] Anonymous. IEEE Over the Rainbow podcast. Computer, 56(12): 99, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sanchez-Gordon:2023:EAP

[1627] Mary Sánchez-Gordón, Edmundo Tovar, Ricardo Colomo-Palacios, Nelson Piedra, and Manuel Castro. Educating augmented programmers. Computer, 56(12):100–104, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:IAC

[1628] Anonymous. IEEE Annals in Computing. Computer, 56(12):104, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICCc

[1629] Anonymous. IEEE career center.

Computer, 56(12):105, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2023:SDA

[1630] Norita Ahmad and Salwa Beheiry. Sustainability in the data age: Harnessing the power of data for a greener future. Computer, 56(12):106–110, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICGe

[1631] Anonymous. IEEE Computer Graphics and Applications. Computer, 56 (12):110, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICScj

[1632] Anonymous. IEEE Computer Society volunteer services awards. Computer, 56(12):111, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2023:GAI

[1633] Nir Kshetri. Generative artificial intelligence and the economics of effective prompting. Computer, 56(12):112– 118, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICSck

[1634] Anonymous. IEEE Computer Society diversity and inclusion. Computer, 56 (12):119, December 2023. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weiss:2023:TNC

[1635] Joseph Weiss, James Bret Michael, and Michael T. Swearingen. There is no chilling when your control system cybersecurity is unfulfilling. Computer, 56(12):120–124, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2023:ICScl

[1636] Anonymous. IEEE Computer Society. Computer, 56(12):C3, December 2023. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCa

[1637] Anonymous. Front cover. Computer, 57(1):C1, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSa

[1638] Anonymous. IEEE Computer Society diversity and inclusion fund. Computer, 57(1):C2, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCa

[1639] Anonymous. Table of contents. Computer, 57(1):1–2, January 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Ma

[1640] Anonymous. Masthead. Computer, 57(1):3, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athavale:2024:MYM

[1641] Jyotika Athavale. Make your mark.

Computer, 57(1):4–7, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:I

[1642] Anonymous. ICCQ 2024. Computer, 57(1):7, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSa

[1643] Anonymous. Computer highlights Society magazines. Computer, 57(1): 8–11, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSb

[1644] Anonymous. IEEE Computer Society benefits. Computer, 57(1):11, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAa

[1645] Erich Neuhold. 50 & 25 years ago.

Computer, 57(1):12–15, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTa

[1646] Ergun Akleman. Computing through time. Computer, 57(1):13, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:NNA

[1647] Jeffrey Voas. Niceties, necessities, antiquities, and ubiquities. Computer, 57 (1):16–17, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPa

[1648] Anonymous. IEEE Security & Privacy subscribe. Computer, 57(1): 17, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:O

[1649] Jeffrey Voas and Ron Vetter. Outlook 2024. Computer, 57(1):18–20, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICGa

[1650] Anonymous. IEEE Computer Graphics and Applications. Computer, 57 (1):20, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2024:ESAa

[1651] George Hurlburt and Sorel Reisman. Experts speak out about artificial intelligence in education: First of two roundtables. Computer, 57(1):21– 26, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSc

[1652] Anonymous. IEEE Computers In Science & Engineering. Computer, 57 (1):26, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Layman:2024:GAI

[1653] Lucas Layman and Ron Vetter. Generative artificial intelligence and the future of software testing. Computer, 57 (1):27–32, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IPCa

[1654] Anonymous. IEEE Pervasive Computing call for articles. Computer, 57 (1):32, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSd

[1655] Anonymous. IEEE Computer Society. Computer, 57(1):33, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:2024:FT

[1656] Brian David Johnson. The future of threatcasting. Computer, 57(1): 34–39, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cascone:2024:SBC

[1657] Lucia Cascone, Vincenzo Loia, Michele Nappi, and Fabio Narducci. Soft biometrics for cybersecurity: Ongoing revolution for Industry 4.0. Computer, 57 (1):40–50, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neisser:2024:HLM

[1658] Mark Neisser, Ndubuisi G. Orji, Harry J. Levinson, Umberto Celano, James Moyne, Supika Mashiro, Dan Wilcox, and Slava Libman. How lithography and metrology are enabling yield in the next generation of semiconductor patterning. Computer, 57(1): 51–58, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weir:2024:ICS

[1659] Charles Weir, Anna Dyson, Olamide Jogunola, Louise Dennis, and Katie Paxton-Fear. Interlinked computing in 2040: Safety, truth, ownership, and accountability. Computer, 57(1): 59–68, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Majeed:2024:FSK

[1660] Abdul Majeed and Seong Oun Hwang. Feature stores: a key enabler for feature reusability and availability across machine learning pipelines. Computer, 57(1):69–74, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sharma:2024:UCI

[1661] Debendra Das Sharma and Thomas Coughlin. Universal chiplet interconnect express: an open industry standard for memory and storage applications. Computer, 57(1):75–81, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:DCY

[1662] Nir Kshetri and Jeffrey Voas. Data centers in your backyard? Computer, 57 (1):82–87, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITBa

[1663] Anonymous. IEEE Transactions on Big Data. Computer, 57(1):87, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akers:2024:FAE

[1664] Matthew Akers and Armon Barton.

Forming adversarial example attacks against deep neural networks with reinforcement learning. Computer, 57(1): 88–99, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IRPa

[1665] Anonymous. IEEE Over the Rainbow podcast. Computer, 57(1):99, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arora:2024:CFF

[1666] Payal Arora and Usha Raman. The case for femwork: Feminist design principles in tech and tech-enabled work. Computer, 57(1):100–104, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IACa

[1667] Anonymous. IEEE Annals in Computing. Computer, 57(1):104, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conde:2024:UIA

[1668] Javier Conde, Pedro Reviriego, Joaquín Salvachúa, Gonzalo Martínez, José Alberto Hernández, and Fabrizio Lombardi. Understanding the impact of artificial intelligence in academic writing: Metadata to the rescue. Computer, 57 (1):105–109, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSe

[1669] Anonymous. IEEE Computer Society call for papers. Computer, 57 (1):109, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2024:DIS

[1670] Phil Laplante. Deception and intuition in software engineering. Computer, 57 (1):110–113, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSa

[1671] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57 (1):113, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2024:HCG

[1672] Michael Zyda. How to create a games industry in your country. Computer, 57(1):114–119, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITCa

[1673] Anonymous. IEEE Transactions on Computers. Computer, 57(1):119, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Swearingen:2024:RZT

[1674] Michael T. Swearingen, James Bret Michael, Joseph Weiss, and Robert Radvanovsky. Resilient without zero trust. Computer, 57(1):120–122, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:SNE

[1675] Anonymous. Search for new Editor In Chief. Computer, 57(1):123, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2024:IIT

[1676] Dimitrios Serpanos. Industrial internet of things: Trends and challenges. Computer, 57(1):124–128, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IIPa

[1677] Anonymous. IEEE IT Professional call for articles. Computer, 57(1): 128, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISNa

[1678] Anonymous. IEEE search for new Editor In Chief (generic). Computer, 57 (1):129, January 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2024:GAI

[1679] Hal Berghel. Generative artificial intelligence, semantic entropy, and the big sort. Computer, 57(1):130–135, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bruel:2024:PHS

[1680] Pedro Bruel, Sai Rahul Chalamalasetti, Aditya Dhakal, Eitan Frachtenberg, Ninad Hogade, Rolando Pablo Hong Enriquez, Alok Mishra, Dejan Milojicic, Pavana Prakash, and Gourav Rattihalli. Predicting heterogeneity and serverless principles of converged highperformance computing, artificial intelligence, and workflows. Computer, 57 (1):136–144, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICCa

[1681] Anonymous. IEEE career center. Computer, 57(1):C3, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IVSa

[1682] Anonymous. IEEE volunteer service awards. Computer, 57(1):C4, January 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCb

[1683] Anonymous. Front cover. Computer, 57(2):C1, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISNb

[1684] Anonymous. IEEE search for new Editor In Chief. Computer, 57(2): C2, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCb

[1685] Anonymous. Table of contents. Computer, 57(2):1–2, February 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Mb

[1686] Anonymous. Masthead. Computer, 57 (2):3, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSb

[1687] Anonymous. Computer highlights Society magazines. Computer, 57(2): 4–6, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAb

[1688] Erich Neuhold. 50 & 25 years ago.

Computer, 57(2):7–10, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTb

[1689] Ergun Akleman. Computing through time: Text and drive. Computer, 57 (2):8, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSf

[1690] Anonymous. IEEE Computer Society benefits. Computer, 57(2):10, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2024:TC

[1691] Joanna F. DeFranco and Jeffrey Voas.

Thirsty computing. Computer, 57(2):

11–13, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:Ia

[1692] Jeffrey Voas. In this issue. Computer, 57(2):12, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2024:ESAb

[1693] George Hurlburt and Sorel Reisman. Experts speak out about artificial intelligence in education: Second of two panel roundtables. Computer, 57(2): 14–20, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSg

[1694] Anonymous. IEEE Computer Society call for papers. Computer, 57 (2):20, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hinchey:2024:EC

[1695] Mike Hinchey. Ethics and computing. Computer, 57(2):21–23, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IRPb

[1696] Anonymous. IEEE Over the Rainbow podcast. Computer, 57(2):23, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kreps:2024:IFI

[1697] David Kreps and Moira de Roche. The international federation for information processing code of ethics in context.

Computer, 57(2):24–32, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramos:2024:MAE

[1698] Gabriela Ramos, Mariagrazia Squicciarini, and Eleonora Lamm. Making AI ethical by design: The UN-ESCO perspective. Computer, 57(2): 33–43, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grodzinsky:2024:EIE

[1699] Frances S. Grodzinsky, Marty J. Wolf, and Keith W. Miller. Ethical issues from emerging AI applications: Harms are happening. Computer, 57 (2):44–52, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPb

[1700] Anonymous. IEEE security & privacy subscribe. Computer, 57(2): 52, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Greene:2024:APM

[1701] Kristen K. Greene, Mary F. Theofanos, Craig Watson, Anne Andrews, and Eyeisha Barron. Avoiding past mistakes in unethical human subjects research: Moving from artificial intelligence principles to practice. Computer, 57(2):53–63, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gotterbarn:2024:CBS

[1702] Donald Gotterbarn and Keith W.

Miller. Cui bono? Software professionals should always ask who benefits? Computer, 57(2):64–69, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSb

[1703] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57 (2):69, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cerveira:2024:RV

[1704] Frederico Cerveira, António Howcroft Ferreira, and Raul Barbosa. Resilient virtualization. Computer, 57(2):70– 78, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kim:2024:SED

[1705] Jisoo Kim, Mungyu Bae, Yeonsung Chu, Moon Chan Park, Dongwon Park, Dawoon Jung, and Sangyeun Cho. Selfencrypting drive evolving toward multitenant cloud computing. Computer, 57 (2):79–90, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nettstrater:2024:HFR

[1706] Andreas Nettsträter and Marius Brehler. How to find the right scope for open source developments? Computer, 57(2):91–94, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tommarello:2024:CIT

[1707] Nina DeFranco Tommarello and Rebecca A. Deek. The convergence of the Internet of Things and artificial intelligence in medicine: Assessing the benefits, challenges, and risks. Computer, 57 (2):95–99, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSh

[1708] Anonymous. IEEE Computer Society.

Computer, 57(2):100, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eeckhout:2024:TSC

[1709] Lieven Eeckhout. Toward sustainable computer systems. Computer, 57(2): 101–104, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:WCB

[1710] Nir Kshetri. Wholesale central bank digital currencies in Asia: Current status and motivation. Computer, 57(2): 105–110, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oppliger:2024:HMC

[1711] Rolf Oppliger and Andreas Grünert.

How to measure cybersecurity and why heuristics matter. Computer, 57(2): 111–115, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2024:DAC

[1712] Mark Campbell and Mla an Jovanovi . Directing AI: Charting a roadmap of AI opportunities and risks. Computer, 57(2):116–120, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nordquist:2024:EHI

[1713] Christopher D. Nordquist and Stanley S. Chou. Exploring heterogeneous integration: Its essence and future path. Computer, 57(2):121– 124, February 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IACb

[1714] Anonymous. IEEE Annals in Computing. Computer, 57(2):124, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:GAIa

[1715] Nir Kshetri. Generative artificial intelligence and e-commerce. Computer, 57 (2):125–128, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICCb

[1716] Anonymous. IEEE career center. Computer, 57(2):C3, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IVSb

[1717] Anonymous. IEEE volunteer service awards. Computer, 57(2):C4, February 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCc

[1718] Anonymous. Front cover. Computer, 57(3):C1, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSi

[1719] Anonymous. IEEE Computer Society diversity & integrity fund. Computer, 57(3):C2, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCc

[1720] Anonymous. Table of contents. Computer, 57(3):1–2, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Mc

[1721] Anonymous. Masthead. Computer, 57(3):3, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSc

[1722] Anonymous. Computer highlights Society magazines. Computer, 57(3): 4–6, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSc

[1723] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57(3):6, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAc

[1724] Erich Neuhold. 50 & 25 years ago.

Computer, 57(3):7–10, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTc

[1725] Ergun Akleman. Computing through time: Trustworthy AI. Computer, 57(3):8, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:Ib

[1726] Jeffrey Voas. The influencer. Computer, 57(3):11–13, March 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:Ic

[1727] Jeffrey Voas. Also in this issue. Computer, 57(3):12, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSj

[1728] Anonymous. IEEE Computer Society benefits. Computer, 57(3):13, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2024:AAA

[1729] Phil Laplante, Joanna F. DeFranco, and Rick Kuhn. Assured autonomy, artificial intelligence, and machine learning: a roundtable discussion. Computer, 57(3):14–21, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mariani:2024:TAP

[1730] Riccardo Mariani, Francesca Rossi, Rita Cucchiara, Marco Pavone, Barnaby Simkin, Ansgar Koene, and Jochen Papenbrock. Trustworthy AI part III. Computer, 57(3):22–24, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSk

[1731] Anonymous. IEEE Computer Society. Computer, 57(3):25, March 2024.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cina:2024:MLS

[1732] Antonio Emanuele Cinà, Kathrin Grosse, Ambra Demontis, Battista Biggio, Fabio Roli, and Marcello Pelillo. Machine learning security against data poisoning: Are we there yet? Computer, 57(3):26–34, March 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

T:2024:TAB

[1733] Padmapritha T, Korkut Bekiroglu, Subathra Seshadhri, and Seshadhri Srinivasan. Trustworthy AI-based personalized insulin recommender for elderly people who have type-2 diabetes. Computer, 57(3):35–45, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSl

[1734] Anonymous. IEEE Computer Society call for papers. Computer, 57 (3):45, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sousa:2024:HCT

[1735] Sonia Sousa, David Lamas, José Cravino, and Paulo Martins. Humancentered trustworthy framework: a human computer interaction perspective. Computer, 57(3):46–58, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peckham:2024:AHG

[1736] Jeremy B. Peckham. An AI harms and governance framework for trustworthy AI. Computer, 57(3):59–68, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hazra:2024:DAZ

[1737] Abhishek Hazra, Andrea Morichetta, Ilir Murturi, Lauri Lovén, Chinmaya Kumar Dehury, Víctor Casamayor Pujol, Praveen Kumar Donta, and Schahram Dustdar. Distributed AI in zero-touch provisioning for edge networks: Challenges and research directions. Computer, 57(3):69–78, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Qu:2024:LUI

[1738] Youyang Qu, Xin Yuan, Ming Ding, Wei Ni, Thierry Rakotoarivelo, and David Smith. Learn to unlearn: Insights into machine unlearning. Computer, 57(3):79–90, March 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cucchiara:2024:VSP

[1739] Rita Cucchiara, Lorenzo Baraldi, Marcella Cornia, and Sara Sarto. Video surveillance and privacy: a solvable paradox? Computer, 57(3):91–100, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Majeed:2024:MLL

[1740] Abdul Majeed and Seong Oun Hwang. Making large language models more reliable and beneficial: Taking Chat-GPT as a case study. Computer, 57 (3):101–106, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IRPc

[1741] Anonymous. IEEE Over the Rainbow podcast. Computer, 57(3):106, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2024:SAC

[1742] Yuntao Liu, Daniel Xing, Isaac Mc-Daniel, Olsan Ozbay, Abir Akib, Mumtahina Islam Sukanya, Sanjay Rekhi, and Ankur Srivastava. Security advantages and challenges of 3D heterogeneous integration. Computer, 57(3): 107–112, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IACc

[1743] Anonymous. IEEE Annals in Computing. Computer, 57(3):112, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Handy:2024:HEM

[1744] Jim Handy and Tom Coughlin. How emerging memories extend battery life. Computer, 57(3):113–116, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barana:2024:ESP

[1745] Alice Barana and Marina Marchisio Conte. Eight sustainable practices for digital activity development: Drivers and barriers in international higher education collaboration. Computer, 57 (3):117–122, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSm

[1746] Anonymous. IEEE Computing in Science & Engineering. Computer, 57 (3):122, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2024:SAI

[1747] Dimitrios Serpanos and Marilyn Wolf. Safety and artificial intelligence in cyberphysical systems. Computer, 57(3): 123–126, March 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IIPb

[1748] Anonymous. IEEE IT professionals. Computer, 57(3):126, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2024:LLM

[1749] Michael Zyda. Large language models and generative AI, oh my! Computer, 57(3):127–132, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICCc

[1750] Anonymous. IEEE career center. Computer, 57(3):C3, March 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICEa

[1751] Anonymous. IEEE Computing Edge. Computer, 57(3):C4, March 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCd

[1752] Anonymous. Front cover. Computer, 57(4):C1, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSn

[1753] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. Computer, 57(4):C2, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCd

[1754] Anonymous. Table of contents. Computer, 57(4):1–2, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Md

[1755] Anonymous. Masthead. Computer, 57(4):3, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSd

[1756] Anonymous. Computer highlights Society magazines. Computer, 57(4):4–6, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSd

[1757] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57(4):6, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAd

[1758] Erich Neuhold. 50 & 25 years ago.

Computer, 57(4):7–10, April 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTd

[1759] Ergun Akleman. Computing through time: Dependability. Computer, 57 (4):8, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:TD

[1760] Jeffrey Voas. Thoughts on dependability. Computer, 57(4):11–12, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2024:DC

[1761] Joanna F. DeFranco, Phil Laplante, Rick Kuhn, and Steven Li. Dependable computing. Computer, 57 (4):13–15, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McDermid:2024:SAS

[1762] John A. McDermid, Radu Calinescu, Ibrahim Habli, Richard Hawkins, Yan Jia, John Molloy, Matt Osborne, Colin Paterson, Zoe Porter, and Philippa Ryan Conmy. The safety of autonomy: a systematic approach. Computer, 57(4):16–25, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yano:2024:PCS

[1763] Hiromu Yano, Tadashi Dohi, and Hiroyuki Okamura. Performance comparison of software reliability estimation algorithms. Computer, 57(4): 26–36, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goodloe:2024:HIR

[1764] Alwyn E. Goodloe and Klaus Havelund.

High-integrity runtime verification. Computer, 57(4):37–45, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huang:2024:ETD

[1765] Ren-Yi Huang, Dumindu Samaraweera, and J. Morris Chang. Exploring threats, defenses, and privacypreserving techniques in federated learning: a survey. Computer, 57 (4):46–56, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSo

[1766] Anonymous. IEEE Computer Society — call for papers. Computer, 57(4):56, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hatton:2024:DID

[1767] Les Hatton. Dependability improvement depends on dependable measurement. Computer, 57(4):57–67, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Binder:2024:SDM

[1768] Robert V. Binder. Software dependability measurement at the age of 36. Computer, 57(4):68–81, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSp

[1769] Anonymous. IEEE Computer Society has you covered! Computer, 57(4):81, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trivedi:2024:RAA

[1770] Kishor Trivedi and Andrea Bobbio. Reliability and availability analysis in practice: Toward multilevel models for complex systems. Computer, 57 (4):82–92, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2024:ITI

[1771] Christof Ebert. Industry Trends 2024: Innovation needs competence. Computer, 57(4):93–98, April 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPc

[1772] Anonymous. IEEE Security and Privacy. Computer, 57(4):98, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McGregor:2024:ODS

[1773] Sean McGregor. Open digital safety. Computer, 57(4):99–103, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cesario:2024:MLE

[1774] Eugenio Cesario. Machine learning for enhancing public safety in modern cities. Computer, 57(4):104–107, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patt:2024:CFC

[1775] Yale N. Patt. The CORRECT first course in computing for serious students. Computer, 57(4):108–112, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:EAI

[1776] Nir Kshetri. Economics of artificial intelligence governance. Computer, 57 (4):113–118, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2024:PVN

[1777] Doron Drusinsky. From P versus NP to probabilistic and zero knowledge proof systems. Computer, 57(4): 119–130, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tang:2024:DQC

[1778] Wei Tang and Margaret Martonosi.

Distributed quantum computing via integrating quantum and classical computing. Computer, 57(4):131–136, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Agarwal:2024:ABD

[1779] Sugandha Agarwal, Norita Ahmad, and Dima Jamali. AI and big data in contemporary marketing. Computer, 57(4):137–142, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITPa

[1780] Anonymous. IEEE Transactions on Privacy. Computer, 57(4):142, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Frachtenberg:2024:DP

[1781] Eitan Frachtenberg, Viyom Mittal, Pedro Bruel, Michalis Faloutsos, and Dejan Milojicic. The distribution is the performance. Computer, 57(4): 143–149, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CSEa

[1782] Anonymous. Computing in Science & Engineering. Computer, 57(4):149, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yu:2024:SMT

[1783] Shimeng Yu and Tae-Hyeon Kim. Semiconductor memory technologies: State-of-the-art and future trends. Computer, 57(4):150–154, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IAHa

[1784] Anonymous. IEEE Annals of the History of Computing. Computer, 57 (4):154, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSq

[1785] Anonymous. IEEE Computer Society Volunteer Service Awards. Computer, 57(4):155, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2024:ITF

[1786] Joanna F. DeFranco and Jeffrey Voas. Internet of Things-flavored chips. Computer, 57(4):156–158, April 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSr

[1787] Anonymous. IEEE Computer Society. Computer, 57(4):159, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McGraw:2024:SRB

[1788] Gary McGraw, Richie Bonett, Harold Figueroa, and Katie McMahon. 23 security risks in black-box large language model foundation models. Computer, 57(4):160–164, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITBb

[1789] Anonymous. IEEE Transactions on Big Data. Computer, 57(4):164, April 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSs

[1790] Anonymous. IEEE Computer Society Career Center. Computer, 57 (4):C3, April 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICEb

[1791] Anonymous. IEEE Computing Edge.

Computer, 57(4):C4, April 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCe

[1792] Anonymous. Front cover. Computer, 57(5):C1, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSt

[1793] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. Computer, 57(5):C2, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCe

[1794] Anonymous. Table of contents. Computer, 57(5):1–2, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Me

[1795] Anonymous. Masthead. Computer, 57(5):3, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSe

[1796] Anonymous. Computer highlights Society magazines. Computer, 57(5):4–6, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSu

[1797] Anonymous. IEEE Computer Society call for papers. Computer, 57(5):6, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAe

[1798] Erich Neuhold. 50 & 25 years ago. Computer, 57(5):7–11, May 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTe

[1799] Ergun Akleman. Computing through time: Venture scientists. Computer, 57(5):8, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITPb

[1800] Anonymous. IEEE Transaction on Privacy. Computer, 57(5):11, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:DI

[1801] Jeffrey Voas. Dissertation, Inc. Computer, 57(5):12–14, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:Id

[1802] Jeffrey Voas. Also in this issue. Computer, 57(5):13, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IAHb

[1803] Anonymous. IEEE Annals of the History of Computing. Computer, 57(5):14, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Helal:2024:VST

[1804] Sumi Helal and Philip Treleaven. Venture scientists and technology entrepreneurship. Computer, 57(5):15– 19, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICGb

[1805] Anonymous. IEEE Computer Graphics and Applications. Computer, 57 (5):19, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Treleaven:2024:EVS

[1806] Philip Treleaven. Educating venture scientists: Inspiring students and young professionals to commercialize their deep tech research supported by their institutions. Computer, 57 (5):20–28, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Aiello:2024:ARF

[1807] Marco Aiello. The AI revolution and the future of work: Threats and opportunities. Computer, 57(5):29–34, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balassone:2024:FIE

[1808] Sergio Balassone. Fostering innovation and entrepreneurship in Europe: The EIT digital journey. Computer, 57(5):35–38, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wong:2024:HPR

[1809] Chan-Yuan Wong and Jui-Jan Chan. How public research institutions propagate industrial niches and configure new industrial structure: The case of ITRI (Taiwan) for semiconductor technologies. Computer, 57(5):39–43, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Solvak:2024:CSP

[1810] Mihkel Solvak and Ave Lauringson. A case study of the public sector digital ecosystem in Estonia. Computer, 57(5):44–49, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CSEb

[1811] Anonymous. Computing in Science & Engineering. Computer, 57(5):49, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alexiev:2024:VBD

[1812] Victor Alexiev. Venture building in the digital era: Unraveling the past, present, and future of corporate innovation. Computer, 57(5):50–57, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rassolkin:2024:SRS

[1813] Anton Rassõlkin, Pāvels Maksimkins, Andrejs Stupāns, Viktor Rjabtšikov, Armands Šenfelds, and Vladimir Kuts. The spatial representation of a selfdriving vehicle for the virtual entity of a digital twin. Computer, 57 (5):58–66, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khan:2024:CSC

[1814] Ajmal Khan, Amer Saeed, Farman Ullah, Muhammad Bilal, Yasir Muhammad, and Hesham El-Sayed. Charging support communication system based on vehicle-to-vehicle communication for electric vehicles. Computer, 57(5):67–77, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zou:2024:UEA

[1815] Qingtian Zou, Lan Zhang, Xiaoyan Sun, Anoop Singhal, and Peng Liu. Using explainable AI for neural networkbased network attack detection. Computer, 57(5):78–85, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IIPc

[1816] Anonymous. IEEE IT Professional. Computer, 57(5):85, May 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:2024:AAT

[1817] D. Richard Kuhn, M. S. Raunak, Raghu N. Kacker, Jaganmohan Chandrasekaran, Erin Lanus, Tyler Cody, and Laura Freeman. Assured autonomy through combinatorial methods. Computer, 57(5):86–90, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IPCb

[1818] Anonymous. IEEE Pervasive Computing call for articles. Computer, 57(5):90, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2024:APS

[1819] George Hurlburt. Addressing the perceived skills gap. Computer, 57 (5):91–96, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPd

[1820] Anonymous. IEEE Security & Privacy subscribe. Computer, 57(5):96, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSv

[1821] Anonymous. IEEE Computer Society information. Computer, 57(5):97, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2024:TGA

[1822] Michael Zyda. Today generative AI is just a parlor trick. Computer, 57 (5):98–101, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2024:UPW

[1823] Phil Laplante. Use a pencil: On writing software documentation well and the role of autodocumentation. Computer, 57(5):102–105, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Calabrese:2024:PIS

[1824] Matteo Calabrese and Gregory Falco.

Physics-informed satellite cybersecurity. Computer, 57(5):106–109, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSe

[1825] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57 (5):109, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IOJa

[1826] Anonymous. IEEE Open Journal of the Computer Society. Computer, 57 (5):110, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campbell:2024:DAM

[1827] Mark Campbell and Mla an Jovanovi . Disinfecting AI: Mitigating generative AI’s top risks. Computer, 57 (5):111–116, May 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSw

[1828] Anonymous. IEEE Computer Society benefits. Computer, 57(5):117, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2024:SMC

[1829] Hal Berghel. Social media, cognitive dysfunction, and social disruption. Computer, 57(5):118–124, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITBc

[1830] Anonymous. IEEE Transactions on Big Data. Computer, 57(5):124, May 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICCd

[1831] Anonymous. IEEE Career Center. Computer, 57(5):C3, May 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICEc

[1832] Anonymous. IEEE Computing Edge. Computer, 57(5):C4, May 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCf

[1833] Anonymous. Front cover. Computer, 57(6):C1, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITPc

[1834] Anonymous. IEEE Transactions on Privacy. Computer, 57(6):C2, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCf

[1835] Anonymous. Table of contents. Computer, 57(6):1–2, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Mf

[1836] Anonymous. Masthead. Computer, 57(6):3, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSf

[1837] Anonymous. Computer highlights Society magazines. Computer, 57(6):4–6, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSx

[1838] Anonymous. IEEE Computer Society call for papers. Computer, 57(6):6, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAf

[1839] Erich Neuhold. 50 & 25 years ago.

Computer, 57(6):7–10, June 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTf

[1840] Ergun Akleman. Computing through time: Biometrics. Computer, 57(6):8, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2024:GYU

[1841] Joanna F. DeFranco and Jeffrey Voas. “Golden Years” users. Computer, 57 (6):11–14, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:Ie

[1842] Jeffrey Voas. In this issue. Computer, 57(6):12–13, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSy

[1843] Anonymous. IEEE Computer Society benefits. Computer, 57(6):14, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSz

[1844] Anonymous. IEEE Computer Society information. Computer, 57(6):15, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macia-Perez:2024:ASE

[1845] Francisco Mácia-Pérez, Iren Lorenzo-Fonseca, José-Vicente Berná-Martinez, and Alex Maciá-Fiteni. Accompaniment services for the elderly: a comparison of decision-making support architectures in distributed IT. Computer, 57(6):16–28, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ran:2024:PTE

[1846] Dezhi Ran, Yingjie Fu, Yibo He, Tianyu Chen, Xing Tang, and Tao Xie. Path toward elderly friendly mobile apps. Computer, 57(6):29–39, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alonso-Fernandez:2024:PBM

[1847] Fernando Alonso-Fernandez, Josef Bigun, Julian Fierrez, Naser Damer, Hugo Proença, and Arun Ross. Periocular biometrics: a modality for unconstrained scenarios. Computer, 57 (6):40–49, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jamshidi:2024:GRB

[1848] Kiavash Jamshidi, Ramin Toosi, and Mohammad Ali Akhaee. Gender recognition based on hand images employing local and global shape information. Computer, 57(6):50–61, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2024:TPH

[1849] Jiangtao Wang, Long Chen, Deborah Lycett, Duncan Vernon, and Dingchang Zheng. Toward population health intelligence: When artificial intelligence meets population health research. Computer, 57(6):62–72, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Naqvi:2024:IAC

[1850] Bilal Naqvi, Joakim Kävrestad, and A. K. M. Najmul Islam. Inclusive and accessible cybersecurity: Challenges and future directions. Computer, 57(6):73–81, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSf

[1851] Anonymous. IEEE Transactions on Susatainable Computing. Computer, 57(6):81, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Verma:2024:DER

[1852] Akriti Verma, Shama Islam, Valeh Moghaddam, and Adnan Anwar. Digital emotion regulation on social media. Computer, 57(6):82–89, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zaks:2024:IVC

[1853] Irina Zaks, Timothy Lehnen, and Aaron Zaks. Improving velocity of code contributions in open source. Computer, 57(6):90–96, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2024:CAI

[1854] Christof Ebert and Xuya Wang. China artificial intelligence industry insights. Computer, 57(6):97–101, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPe

[1855] Anonymous. IEEE Security & Privacy subscribe. Computer, 57(6):101, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:GAIb

[1856] Nir Kshetri. Generative artificial intelligence in the financial services industry. Computer, 57(6):102–108, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITCb

[1857] Anonymous. IEEE Transactions on Computers. Computer, 57(6):108, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSba

[1858] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. Computer, 57(6):109, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2024:PCU

[1859] Sorel Reisman. Practical classroom use of generative artificial intelligence: a case study. Computer, 57(6):110–114, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITBd

[1860] Anonymous. IEEE Transactions on Big Data. Computer, 57(6):114, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IOJb

[1861] Anonymous. IEEE Open Journal of the Computer Society. Computer, 57 (6):115, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jenkins:2024:SIM

[1862] Barry Jenkins, Warren Mayoss, and Michael Zyda. Streamed interactive media: How changing the paradigm on streaming game content enables a new media category. Computer, 57 (6):116–121, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IPCc

[1863] Anonymous. IEEE Pervasive Computing call for articles. Computer, 57(6):121, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Panda:2024:GLB

[1864] Biswabandan Panda. The game of latency, bandwidth, and hardware prefetching. Computer, 57(6):122–126, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IIPd

[1865] Anonymous. IEEE IT Professional. Computer, 57(6):126, June 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IVSc

[1866] Anonymous. IEEE Volunteer Service Awards. Computer, 57(6):127, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Taufer:2024:HCC

[1867] Michela Taufer, Daniel Milroy, Todd Gamblin, Andrew Jones, Bill Magro, Heidi Poxon, and Seetharami Seelam. HPC and cloud convergence beyond technical boundaries: Strategies for economic sustainability, standardization, and data accessibility. Computer, 57(6):128–136, June 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kassab:2024:UIT

[1868] Mohamad Kassab and Joanna F. De-Franco. Unlocking the Internet of Things engineering job market. Computer, 57(6):137–140, June 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbb

[1869] Anonymous. IEEE Computers in Science & Engineering. Computer, 57 (6):140, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chauhan:2024:DRO

[1870] Preeti S. Chauhan. Deepfake: Risks and opportunities. Computer, 57(6): 141–144, June 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICCe

[1871] Anonymous. IEEE Career Center.

Computer, 57(6):C3, June 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICEd

[1872] Anonymous. IEEE Computing Edge.

Computer, 57(6):C4, June 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCg

[1873] Anonymous. Front cover. Computer, 57(7):C1, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITPd

[1874] Anonymous. IEEE Transactions on Privacy. Computer, 57(7):C2, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCg

[1875] Anonymous. Table of contents. Computer, 57(7):1–2, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Mg

[1876] Anonymous. Masthead. Computer, 57(7):3, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSg

[1877] Anonymous. Computer highlights Society magazines. Computer, 57(7):4–6, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbc

[1878] Anonymous. IEEE Computer Society call for papers. Computer, 57(7):6, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAg

[1879] Erich Neuhold. 50 & 25 years ago. Computer, 57(7):7–10, July 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTg

[1880] Ergun Akleman. Computing through time: Security. Computer, 57(7):8, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:AD

[1881] Jeffrey Voas and Keith Miller. Algorithms as defendants? Computer, 57 (7):11–14, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:If

[1882] Jeffrey Voas. In this issue. Computer, 57(7):12–13, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbd

[1883] Anonymous. IEEE Computer Society information. Computer, 57(7):15, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chandrasekaran:2024:LCC

[1884] Jaganmohan Chandrasekaran, Erin Lanus, Tyler Cody, Laura J. Freeman, Raghu N. Kacker, M S Raunak, and D. Richard Kuhn. Leveraging combinatorial coverage in the machine learning product lifecycle. Computer, 57 (7):16–26, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yao:2024:CDM

[1885] Wei Yao, Nicholas Gorlewski, Fadi P.

Deek, and Guiling Wang. Considerations for decision makers and developers toward the adoption of decentralized key management systems technology in emerging applications. Computer, 57(7):27–38, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abdo:2024:CDU

[1886] Jacques Bou Abdo, Sherali Zeadally, and Jacques Demerjian. Compartmentalizationby-design: Usability-aware web privacy. Computer, 57(7):39–52, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karati:2024:EDS

[1887] Arijit Karati and Sajal K. Das. Effective data sharing in an edge cloud model: Security challenges and solutions. Computer, 57(7):53–65, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

deCastro:2024:RFP

[1888] Manuel de Castro, David L. Vilariño, Yuri Torres, and Diego R. Llanos. The role of field-programmable gate arrays in the acceleration of modern high-performance computing workloads. Computer, 57(7):66–76, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zheng:2024:MFT

[1889] Zheng Zheng, Daixu Ren, Huai Liu, and Tsong Yueh Chen. Metamorphic fault tolerance: Addressing the oracle problem of reliability assurance for contemporary software systems. Computer, 57(7):77–86, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbe

[1890] Anonymous. IEEE Computer Society benefits. Computer, 57(7):86, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hilowle:2024:LHC

[1891] Malyun Hilowle, William Yeoh, Marthie Grobler, Graeme Pye, and Frank Jiang. Leveraging human-centric cybersecurity to improve usage of national digital identity systems in Australia. Computer, 57(7):87–98, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPf

[1892] Anonymous. IEEE Security & Privacy subscribe. Computer, 57(7):98, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sipper:2024:WAM

[1893] Moshe Sipper and Raz Lapid. What’s in an AI’s mind’s eye? We must know. Computer, 57(7):99–103, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSg

[1894] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57 (7):103, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baudry:2024:PAD

[1895] Benoit Baudry and Martin Monperrus.

Programming art with drawing machines. Computer, 57(7):104–108, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITCc

[1896] Anonymous. IEEE Transactions on Computers. Computer, 57(7):108, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Talia:2024:ACD

[1897] Domenico Talia. Algorithms that can deny care, and a call for AI explainability. Computer, 57(7):109–112, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bordis:2024:CRL

[1898] Tabea Bordis, Maximilian Kodetzki, and Ina Schaefer. From concept to reality: Leveraging correctness-by-construction for better algorithm design. Computer, 57(7):113–119, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramesh:2024:WIS

[1899] S. K. Ramesh. Why instructors should care: ABET and CSAB, how they work, IEEE’s role, and their value. Computer, 57(7):120–125, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IIPe

[1900] Anonymous. IEEE IT Professionals call for articles. Computer, 57 (7):125, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bash:2024:ESF

[1901] Cullen Bash, Jessica Bian, Dejan Milojicic, Chandrakant D. Patel, Luka Strezoski, and Vladimir Terzija. Energy supplies for future data centers. Computer, 57(7):126–134, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IVSd

[1902] Anonymous. IEEE Volunteer Service Awards. Computer, 57(7):135, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2024:COS

[1903] Michael Zyda. Can OpenAI’s Sora generate Pixar’s Toy Story? Computer, 57 (7):136–140, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbf

[1904] Anonymous. IEEE Computers In Science & Engineering. Computer, 57 (7):140, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lee:2024:ECI

[1905] Shanhsin Lee, Shiuhpyng Winston Shieh, and Mengru Tsai. Enterprisecentric intelligence: a prioritization scheme for cyberthreat intelligence and common vulnerabilities and exposures. Computer, 57(7):141–152, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICGc

[1906] Anonymous. IEEE Computer Graphics and Applications. Computer, 57 (7):152, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbg

[1907] Anonymous. IEEE Computer Society Diversity and Inclusion Fund. Computer, 57(7):153, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bagchi:2024:RRC

[1908] Saurabh Bagchi, Mahsa Ghasemi, Kang G. Shin, Nalini Venkatasubramanian, Dongyan Xu, and Saman Zonouz. Role of resilience in cyber-physical systems: a roundtable panel. Computer, 57(7):154–159, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:FAC

[1909] Nir Kshetri, Elena Loukoianova, and Jeffrey Voas. Fintech applications for climate finance and sustainable development. Computer, 57(7):160–165, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IACd

[1910] Anonymous. IEEE Annals in Computing. Computer, 57(7):165, July 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghe:2024:CAV

[1911] Hal Berghe. A collapsing academy V: Internal and external threat vectors facing higher education. Computer, 57 (7):166–172, July 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICCf

[1912] Anonymous. IEEE Career Center. Computer, 57(7):C3, July 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICEe

[1913] Anonymous. IEEE Computing Edge.

Computer, 57(7):C4, July 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCh

[1914] Anonymous. Front cover. Computer, 57(8):C1, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITPe

[1915] Anonymous. IEEE Transaction on Privacy filler. Computer, 57(8):C2, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCh

[1916] Anonymous. Table of contents. Computer, 57(8):1–2, August 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Mh

[1917] Anonymous. Masthead. Computer, 57(8):3, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:LE

[1918] Jeffrey Voas. Letters to the editor. Computer, 57(8):4, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSh

[1919] Anonymous. Computer highlights Society magazines. Computer, 57(8): 5–7, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAh

[1920] Erich Neuhold. 50 & 25 years ago. Computer, 57(8):8–11, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTh

[1921] Ergun Akleman. Computing through time: Technology predictions. Computer, 57(8):9, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2024:CGD

[1922] Phil Laplante, Dejan Milojicic, and Jeffrey Voas. “Cyberpanics” and global disinformation campaigns. Computer, 57(8):12–14, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2024:CRD

[1923] Joanna F. DeFranco, Nir Kshetri, George K. Thiruvathukal, and George Hurlburt. Crowdsourcing: a roundtable discussion. Computer, 57(8):15–21, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2024:TPT

[1924] Dejan Milojicic and Phil Laplante. Technology predictions: Technologies that advance humanity and broad AI adoption. Computer, 57(8):22–25, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shan:2024:LAI

[1925] Richard Shan. Language artificial intelligence at a crossroads: Deciphering the future of small and large language models. Computer, 57(8): 26–35, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

daSilva:2024:FSW

[1926] Rafael Ferreira da Silva, Rosa M.

Badia, Deborah Bard, Ian T. Foster, Shantenu Jha, and Frédéric Suter. Frontiers in scientific workflows: Pervasive integration with high-performance computing. Computer, 57(8):36–44, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cianca:2024:ATS

[1927] Ernestina Cianca, Mauro De Sanctis, Tommaso Rossi, and Marina Ruggieri. Advanced technologies for a sustainable future of space. Computer, 57 (8):45–54, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flores:2024:ASD

[1928] Huber Flores. AI sensors and dashboards. Computer, 57(8):55–64, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shirali-Shahreza:2024:SMS

[1929] S. Shirali-Shahreza. Set up my smart home as I want. Computer, 57(8): 65–73, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conde:2024:SCL

[1930] Javier Conde, Miguel González, Pedro Reviriego, Zhen Gao, Shanshan Liu, and Fabrizio Lombardi. Speed and conversational large language models: Not all is about tokens per second. Computer, 57(8):74–80, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbh

[1931] Anonymous. IEEE Computer Society information. Computer, 57(8):81, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kott:2024:RCO

[1932] Alexander Kott, George Yegor Dubynskyi, Andrii Paziuk, Stephanie E. Galaitsi, Benjamin D. Trump, and Igor Linkov. Russian cyber onslaught was blunted by Ukrainian cyber resilience, not merely security. Computer, 57(8): 82–89, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbi

[1933] Anonymous. IEEE Computer Society call for papers filler. Computer, 57(8):89, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chan:2024:CMS

[1934] Henry Chan. CARES model for supporting computing education and 4C framework for teaching in hybrid classrooms. Computer, 57(8):90–95, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbj

[1935] Anonymous. IEEE Computer Society benefits filler. Computer, 57 (8):95, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jovanovi:2024:CAS

[1936] Mla an Jovanovi and Mark Campbell. Compacting AI: In search of the small language model. Computer, 57 (8):96–100, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IVSe

[1937] Anonymous. IEEE Volunteer Service Awards filler. Computer, 57 (8):101, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2024:BIE

[1938] Norita Ahmad and Sabir Haque. Balancing innovation and ethics in AI- and big data-driven marketing. Computer, 57(8):102–107, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2024:CIO

[1939] Christof Ebert. Convergence of IT, OT, and the IoT. Computer, 57(8): 108–112, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbk

[1940] Anonymous. IEEE Computer Society Diversity & Inclusion Fund filler. Computer, 57(8):113, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2024:FOS

[1941] Dirk Riehle. Free and open source software. Computer, 57(8):114–118, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IOJc

[1942] Anonymous. IEEE Open Journal of the Computer Society filler. Computer, 57(8):119, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:ESM

[1943] Nir Kshetri. The economics of semiconductor manufacturing in the United States. Computer, 57(8):120–126, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Das:2024:SCO

[1944] Reetuparna Das and Satish Narayanasamy. Systems challenges and opportunities for genomics. Computer, 57(8):127– 135, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CIC

[1945] Anonymous. Candidates for 2025 IEEE Computer Society Officers and Board of Governors positions. Computer, 57 (8):136–150, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICGd

[1946] Anonymous. IEEE Computer Graphics and Applications filler. Computer, 57(8):150, August 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IDV

[1947] Anonymous. IEEE Division V Delegate/Director candidates. Computer, 57(8):151–152, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICCg

[1948] Anonymous. IEEE Career Center filler. Computer, 57(8):C3, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICEf

[1949] Anonymous. IEEE Computing Edge filler. Computer, 57(8):C4, August 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCi

[1950] Anonymous. Front cover. Computer, 57 (9):C1, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITPf

[1951] Anonymous. IEEE Transaction on Privacy. Computer, 57(9):C2, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCi

[1952] Anonymous. Table of contents. Computer, 57(9):1–2, September 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Mi

[1953] Anonymous. Masthead. Computer, 57 (9):3, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSi

[1954] Anonymous. Computer highlights Society magazines. Computer, 57(9): 4–6, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAi

[1955] Erich Neuhold. 50 & 25 years ago. Computer, 57(9):7–10, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTi

[1956] Ergun Akleman. Computing through time: Playing chess. Computer, 57 (9):8, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Loukoianova:2024:DS

[1957] Elena Loukoianova and Jeffrey Voas. Data submarines. Computer, 57(9): 11–12, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbl

[1958] Anonymous. IEEE Computer Society social networking. Computer, 57 (9):13, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jorge:2024:ERA

[1959] Joaquim Jorge and Wolfgang Hürst.

Extended reality and artificial intelligence’s ethical crossroads: From sensory manipulation to creative disruption. Computer, 57(9):14–20, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wong:2024:CRA

[1960] W. Eric Wong. Certifying and regulating AI/ML-centric applications. Computer, 57(9):21–23, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Granlund:2024:TRC

[1961] Tuomas Granlund, Vlad Stirbu, and Tommi Mikkonen. Toward a regulatory-compliant lifecycle for artificial-intelligence-based medical devices in the European Union: Industry perspectives. Computer, 57(9):24– 34, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shan:2024:CGA

[1962] Richard Shan. Certifying generative AI: Retrieval-augmented generation chatbots in high-stakes environments. Computer, 57(9):35–44, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vasudevan:2024:CAM

[1963] Vinod Vasudevan, Amr Abdullatif, Sohag Kabir, and Felician Campean. Certifiability analysis of machine learning systems for low-risk automotive applications. Computer, 57(9):45– 56, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbm

[1964] Anonymous. IEEE Computer Society call for papers. Computer, 57(9): 56, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tenn:2024:TTA

[1965] Kíng-Pîng Tēnn, Yu-Wei Chang, Hong-Yen Chen, Teng-Kai Fan, and Tsungnan Lin. Toward trustworthy artificial intelligence: an integrated framework approach mitigating threats. Computer, 57(9):57–67, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hessami:2024:AIB

[1966] Ali G. Hessami, Alexander Kriebitz, Gerlinde Weger, Eleanor Nell Watson, and Patricia Shaw. Artificial intelligence for the benefit of everyone. Computer, 57(9):68–79, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chesterman:2024:EAG

[1967] Simon Chesterman, Yuting Gao, Jungpil Hahn, and Valerie Sticher. The evolution of AI governance. Computer, 57 (9):80–92, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbn

[1968] Anonymous. IEEE Computer Society information. Computer, 57(9):

93, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schuurman:2024:TZ

[1969] Derek C. Schuurman. A tribute to the Z80. Computer, 57(9):94– 95, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hsu:2024:CFA

[1970] D. Frank Hsu, Bruce S. Kristal, and Christina Schweikert. Combinatorial fusion analysis. Computer, 57(9):96– 100, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Perez:2024:ENS

[1971] Elsa López Pérez, Göran Selander, John Preuß Mattsson, Thomas Watteyne, and Mališa Vučinić. EDHOC is a new security handshake standard: an overview of security analysis. Computer, 57(9):101–110, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coughlin:2024:IRO

[1972] Tom Coughlin and Roger Hoyt. IEEE roadmap outlines development of mass digital storage technology. Computer, 57(9):111–116, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbo

[1973] Anonymous. IEEE Computer Society benefits. Computer, 57(9):117, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2024:GAC

[1974] Michael Zyda. Generative AI changes the world, maybe. Computer, 57(9): 118–123, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2024:GAI

[1975] Sorel Reisman. Generative artificial intelligence and problematic student e-mails. Computer, 57(9):124– 127, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPg

[1976] Anonymous. IEEE Security and Privacy subscription. Computer, 57(9): 127, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:AGA

[1977] Nir Kshetri and Jeffrey Voas. Adapting to generative artificial intelligence: Approaches in higher education institutions. Computer, 57(9):128–133, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IVSf

[1978] Anonymous. IEEE Volunteer Service Awards. Computer, 57(9):134, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Komninos:2024:CNF

[1979] Theodoros Komninos and Dimitrios Serpanos. Cyberconflict: The new frontier of global warfare and social influence. Computer, 57(9):135– 138, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbp

[1980] Anonymous. IEEE Computer Society call for special issue proposal. Computer, 57(9):139, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bannon:2024:IAI

[1981] Tracy Trac Bannon. Infusing artificial intelligence into software engineering and the DevSecOps continuum. Computer, 57(9):140–148, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICGe

[1982] Anonymous. IEEE Computer Graphics and Applications. Computer, 57(9): 148, September 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICCh

[1983] Anonymous. IEEE Career Center.

Computer, 57(9):C3, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICEg

[1984] Anonymous. IEEE Computing Edge.

Computer, 57(9):C4, September 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCj

[1985] Anonymous. Table of contents. Computer, 57(10):1–2, October 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Mj

[1986] Anonymous. Masthead. Computer, 57 (10):3, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSj

[1987] Anonymous. Computer highlights Society magazines. Computer, 57(10): 4–6, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbq

[1988] Anonymous. IEEE Computer Society: Call for papers. Computer, 57 (10):6, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAj

[1989] Erich Neuhold. 50 & 25 years ago. Computer, 57(10):7–10, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTj

[1990] Ergun Akleman. Computing through time: Security. Computer, 57(10): 8, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:MFR

[1991] Jeffrey Voas. Mockery, fake reviews, and fake reviewers. Computer, 57(10): 11–13, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:Ig

[1992] Jeffrey Voas. In this issue. Computer, 57(10):12, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbr

[1993] Anonymous. IEEE Computer Society has you covered! Computer, 57 (10):13, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wong:2024:RSR

[1994] W. Eric Wong and Jeffrey Voas. Revisiting software reliability modeling and testing. Computer, 57(10):14– 25, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koball:2024:MLS

[1995] Carson Koball, Yong Wang, Bhaskar P.

Rimal, and Varghese Vaidyan. Machine learning security: Threat model, attacks, and challenges. Computer, 57 (10):26–35, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hristozov:2024:ECD

[1996] Anton Dimov Hristozov. Engineering challenges for developing reliable and secure heterogeneous computing systems. Computer, 57(10):36–44, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSh

[1997] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57 (10):44, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boudouaia:2024:DSS

[1998] Mohammed Amine Boudouaia and Samir Ouchani. Designing secure and smart supply chains: a roadmap. Computer, 57(10):45–55, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITCd

[1999] Anonymous. IEEE Transactions on Computers. Computer, 57(10):55, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ry:2024:UPM

[2000] Maciej Ry . Unraveling participation motivations in pandemic-related hackathons: a COVID-19 study. Computer, 57(10):56–65, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Naqvi:2024:PPP

[2001] Bilal Naqvi and Kari Smolander. Practitioners’ perspectives on and prospects for usable security. Computer, 57(10): 66–74, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbs

[2002] Anonymous. IEEE Computer Society D&I Fund. Computer, 57(10): 75, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fister:2024:AIE

[2003] Theresa Fister and George K. Thiruvathukal. Artificial intelligence employment interviews: Examining limitations, biases, and perceptions. Computer, 57(10):76–81, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITBe

[2004] Anonymous. IEEE Transactions on Big Data. Computer, 57(10):81, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koch:2024:PDO

[2005] Christian Koch. The prisoner’s dilemma of open-source software security. Computer, 57(10):82–85, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IPCd

[2006] Anonymous. IEEE Pervasive Computing: Call for articles. Computer, 57 (10):85, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grunert:2024:WPC

[2007] Andreas Grünert, James Bret Michael, Rolf Oppliger, and Ruedi Rytz. Why probabilities cannot be used in cyber risk management. Computer, 57(10): 86–89, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2024:HCC

[2008] Joanna F. DeFranco. Health Care 4.0 and Clinical Internet of Things.

Computer, 57(10):90–92, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CSI

[2009] Anonymous. Call for special issue proposals. Computer, 57(10):93, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2024:AIW

[2010] George F. Hurlburt. Artificial intelligence, the workforce and educational realities. Computer, 57(10): 94–98, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IPCe

[2011] Anonymous. IT Professional: Call for articles. Computer, 57(10):98, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhatt:2024:WSA

[2012] Umang Bhatt and Holli Sargeant. When should algorithms resign? A proposal for AI governance. Computer, 57 (10):99–103, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPh

[2013] Anonymous. IEEE Security and Privacy. Computer, 57(10):103, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Macwan:2024:IBT

[2014] Kevin Macwan. Integrating blockchain technology in online gaming ecosystems. Computer, 57(10):104–111, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:NER

[2015] Nir Kshetri. Navigating EU regulations: Challenges for U.S. technology firms and the rise of Europe’s generative AI ecosystem. Computer, 57(10): 112–117, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CSEc

[2016] Anonymous. Computing in Science & Engineering. Computer, 57(10): 117, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Litton:2024:MLC

[2017] Matthew Litton, Doron Drusinsky, London Bridget, and James Bret Michael. Machine-learned correctness properties, runtime verification, and advance-warning oracles for autonomous systems. Computer, 57(10): 118–130, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICGf

[2018] Anonymous. IEEE Computer Graphics and Applications. Computer, 57 (10):130, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2024:EIU

[2019] Hal Berghel. Election integrity in the United States: How will 2024 compare to 2020? Computer, 57(10): 131–137, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yan:2024:FVS

[2020] Mengjia Yan, Thomas Bourgeat, and Sharad Malik. Formal verification for secure processors: a guide for computer architects. Computer, 57(10): 138–143, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IAHc

[2021] Anonymous. IEEE Annals of the History of Computing. Computer, 57 (10):143, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2024:GYA

[2022] Christof Ebert and Ulrich Hemel. Grow your artificial intelligence competence. Computer, 57(10):144–150, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IIS

[2023] Anonymous. IEEE Intelligent Systems. Computer, 57(10):150, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IM

[2024] Anonymous. IEEE Multimedia. Computer, 57(10):150, October 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athavale:2024:DTD

[2025] Jyotika Athavale, Cullen Bash, Wesley Brewer, Matthias Maiterth, Dejan Milojicic, Harry Petty, and Soumyendu Sarkar. Digital twins for data centers. Computer, 57(10):151–158, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbt

[2026] Anonymous. IEEE Computer Society information. Computer, 57(10): 159, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Soures:2024:LCS

[2027] Nicholas Soures, Jayanta Dey, and Dhireesha Kudithipudi. Learning continually in silicon. Computer, 57(10): 160–164, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCj

[2028] Anonymous. Front cover. Computer, 57(10):C1, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITPg

[2029] Anonymous. IEEE Transactions on Privacy. Computer, 57(10):C2, October 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbu

[2030] Anonymous. IEEE Computer Society Career Center. Computer, 57 (10):C3, October 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CE

[2031] Anonymous. Computing Edge. Computer, 57(10):C4, October 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCk

[2032] Anonymous. Table of contents. Computer, 57(11):1–2, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCl

[2033] Anonymous. Table of contents. Computer, 57(11):1–2, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Mk

[2034] Anonymous. Masthead. Computer, 57 (11):3, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Ml

[2035] Anonymous. Masthead. Computer, 57 (11):3, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSk

[2036] Anonymous. Computer highlights Society magazines. Computer, 57(11): 4–6, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSl

[2037] Anonymous. Computer highlights Society magazines. Computer, 57(11): 4–6, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YAk

[2038] Erich Neuhold. 50 & 25 years ago. Computer, 57(11):7–10, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTk

[2039] Ergun Akleman. Computing through time: Intelligence failures. Computer, 57(11):8, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:PTW

[2040] Nir Kshetri and Jeffrey Voas. 3D printing the “Tools” of war. Computer, 57 (11):11–13, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbv

[2041] Anonymous. IEEE Computer Society has you covered! Computer, 57 (11):13, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbz

[2042] Anonymous. IEEE Computer Society has you covered! Computer, 57 (11):13, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raunak:2024:AF

[2043] M S Raunak and D. Richard Kuhn.

AI failures. Computer, 57(11):14– 16, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbw

[2044] Anonymous. IEEE Computer Society information. Computer, 57(11): 17, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSca

[2045] Anonymous. IEEE Computer Society information. Computer, 57(11): 17, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Joshi:2024:AIF

[2046] Hem Chandra Joshi and Sandeep Kumar. Artificial intelligence failures in autonomous vehicles: Causes, implications, and prevention. Computer, 57 (11):18–30, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cummings:2024:RCA

[2047] Mary Missy Cummings, Benjamin Wheeler, and John Kliem. A root cause analysis of a self-driving car dragging a pedestrian. Computer, 57(11): 31–40, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSbx

[2048] Anonymous. IEEE Computer Society: Call for papers. Computer, 57 (11):40, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICScb

[2049] Anonymous. IEEE Computer Society: Call for papers. Computer, 57 (11):40, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ngo:2024:SAT

[2050] Anh Tu Ngo, Alka Luqman, Jonathan Chan Chew Meng, and Anupam Chattopadhyay. A systematic approach toward robust perception in (semi- )Autonomous vehicles. Computer, 57 (11):41–48, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tsikerdekis:2024:DAA

[2051] Michail Tsikerdekis, Sherali Zeadally, and Iyad Katib. Defenses against artificial intelligence attacks. Computer, 57 (11):49–59, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Molloy:2024:HSA

[2052] John Molloy, Sepeedeh Shahbeigi, and John A. McDermid. Hazard and safety analysis of machine-learning-based perception capabilities in autonomous vehicles. Computer, 57(11): 60–70, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSi

[2053] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57 (11):70, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSj

[2054] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57 (11):70, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeMiguelVelazquez:2024:DRW

[2055] Julia De Miguel Velázquez, Sanja Šćepanović, Andrés Gvirtz, and Daniele Quercia. Decoding real-world artificial intelligence incidents. Computer, 57(11):71–81, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITBf

[2056] Anonymous. IEEE Transactions on Big Data. Computer, 57(11):81, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITBg

[2057] Anonymous. IEEE Transactions on Big Data. Computer, 57(11):81, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maung:2024:GAD

[2058] Barani Maung Maung, Keegan McBride, Jason S. Lucas, Maryam Tabar, and Dongwon Lee. Generative AI disproportionately harms long tail users. Computer, 57(11):82–85, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

King:2024:HS

[2059] John Leslie King, Hal Berghel, Phillip Glen Armour, and Robert N. Charette. Healthy skepticism. Computer, 57(11):86–91, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chillarege:2024:SRW

[2060] Ram Chillarege. Software reliability: What went wrong? How to fix it? Computer, 57(11):92–94, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Widen:2024:RCG

[2061] William Widen and Marilyn Wolf.

Redux: Corporate governance and management of AI-driven product development. Computer, 57(11):95– 98, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IIPf

[2062] Anonymous. IEEE IT Professional.

Computer, 57(11):98, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IIPg

[2063] Anonymous. IEEE IT Professional.

Computer, 57(11):98, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghe:2024:BSA

[2064] Hal Berghe. Bruce Schneier on AI security (interview). Computer, 57(11): 99–101, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2024:BSA

[2065] Hal Berghel. Bruce Schneier on AI security (interview). Computer, 57(11): 99–101, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2024:TCL

[2066] Michael Zyda. Twenty columns later . . .. Computer, 57(11):102– 117, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:2024:NLA

[2067] Varun Gupta and Chetna Gupta. Navigating the landscape of AI-generated text detection: Issues and solutions for upholding academic integrity. Computer, 57(11):118–123, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jovanovic:2024:FAM

[2068] Mla– dan Jovanović and Mark Campbell. Fusing AI: Multimodal language models inference across diverse inputs. Computer, 57(11):124–130, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IPCf

[2069] Anonymous. IEEE Pervasive Computing: Call for articles. Computer, 57 (11):130, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IPCg

[2070] Anonymous. IEEE Pervasive Computing: Call for articles. Computer, 57 (11):130, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sharma:2024:GIS

[2071] Ravi S. Sharma, Samia Loucif, Nir Kshetri, and Jeffrey Voas. Global initiatives on “Safer” and more “Responsible” artificial intelligence. Computer, 57(11):131–137, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPi

[2072] Anonymous. IEEE Security and Privacy. Computer, 57(11):137, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPj

[2073] Anonymous. IEEE Security and Privacy. Computer, 57(11):137, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:OTD

[2074] Anonymous. Obituary: Tribute to Doris L. Carver. Computer, 57(11): 138–139, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CC

[2075] Anonymous. CS connection. Computer, 57(11):138–140, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCk

[2076] Anonymous. Front cover. Computer, 57 (11):C1, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:GPN

[2077] Anonymous. Get published in the new IEEE Transactions on Privacy. Computer, 57(11):C2, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSby

[2078] Anonymous. IEEE Computer Society Career Center. Computer, 57(11): C3, November 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICEh

[2079] Anonymous. IEEE Computing Edge. Computer, 57(11):C4, November 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:FCl

[2080] Anonymous. Front cover. Computer, 57 (12):C1, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITPh

[2081] Anonymous. IEEE Transactions on Privacy. Computer, 57(12):C2, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:TCm

[2082] Anonymous. Table of contents. Computer, 57(12):1–2, December 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:Mm

[2083] Anonymous. Masthead. Computer, 57 (12):3, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athavale:2024:EDI

[2084] Jyotika Athavale and Shivam. Empowering diversity and innovation: Jyotika Athavale led the charge this year as 2024 IEEE Computer Society President. Computer, 57(12):4–9, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CYI

[2085] Anonymous. Congratulations to this year’s IEEE Computer Society award recipients. Computer, 57(12):10–15, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IAHd

[2086] Anonymous. IEEE Annals of the History of Computing. Computer, 57(12): 15, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:L

[2087] Jeffrey Voas. Letters. Computer, 57 (12):16, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:CHSm

[2088] Anonymous. Computer highlights Society magazines. Computer, 57(12): 17–19, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICScc

[2089] Anonymous. IEEE Computer Society benefits. Computer, 57(12): 19, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2024:YA

[2090] Erich Neuhold. 50 & 25 years ago.

Computer, 57(12):20–24, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2024:CTTl

[2091] Ergun Akleman. Computing through time: Agriculture. Computer, 57(12): 23, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISM

[2092] Anonymous. IEEE Software magazine call for articles. Computer, 57 (12):24, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IMM

[2093] Anonymous. IEEE Multimedia magazine. Computer, 57(12):24, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:S

[2094] Jeffrey Voas and D. Richard Kuhn. “Over spooked”. Computer, 57(12): 25–26, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2024:I

[2095] Jeffrey Voas. Also in this issue. Computer, 57(12):26, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bannon:2024:GAS

[2096] Tracy Trac Bannon and Phil Laplante. Generative AI in the software development lifecycle. Computer, 57(12): 27–34, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2024:CAU

[2097] Joanna F. DeFranco and Jeffrey Voas.

Computing for agricultural and urban environments. Computer, 57(12):35– 37, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICScd

[2098] Anonymous. IEEE Computer Society call for papers. Computer, 57 (12):37, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:2024:ART

[2099] Ashish Gupta, Vishesh Kumar Tanwar, Amit Nath Jha, and Sajal K. Das. Analyzing real-time insect detection in smart connected farms. Computer, 57(12):38–46, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITSk

[2100] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 57 (12):46, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ottun:2024:ARA

[2101] Abdul-Rasheed Ottun, Adeyinka Akintola, Mohan Liyanage, Michell Boerger, Pan Hui, Sasu Tarkoma, Nikolay Tcholtchev, Petteri Nurmi, and Huber Flores. AI robustness against attacks in city-scale autonomous drone deployments. Computer, 57(12):47– 57, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ITBh

[2102] Anonymous. IEEE Transactions on Big Data. Computer, 57(12):57, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

He:2024:VTI

[2103] Yuankai He and Weisong Shi. A vision for transformative intersections. Computer, 57(12):58–68, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Silvis-Cividjian:2024:TAW

[2104] Natalia Silvis-Cividjian. Therac-25 accidents: We keep on learning from them. Computer, 57(12):69–78, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rivadeneira:2024:PDC

[2105] Jorge Eduardo Rivadeneira, José Marcelo Fernandes, André Rodrigues, Fernando Boavida, and Jorge Sá Silva. Passive data collection in a pandemic situation: a case study. Computer, 57(12): 79–89, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISPk

[2106] Anonymous. IEEE Security and Privacy subscription. Computer, 57(12): 89, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2024:OSP

[2107] Dirk Riehle. Open source program offices. Computer, 57(12):90–94, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Majeed:2024:DIP

[2108] Abdul Majeed and Seong Oun Hwang.

The data island problem and its mitigation: Are we there yet? Computer, 57(12):95–103, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2024:SMF

[2109] Christof Ebert and David A. Gustafson.

Software measurement: Four steps to unlock success. Computer, 57(12):104– 109, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2024:DMA

[2110] Nir Kshetri. Disinformation and misinformation in the age of artificial intelligence and the metaverse. Computer, 57(12):110–116, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICSce

[2111] Anonymous. IEEE Computer Society search for an EIC. Computer, 57(12): 117, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laufer:2024:EMS

[2112] Konstantin Läufer, Gunda Mertin, and George K. Thiruvathukal. Engaging more students in formal methods education: a practical approach using temporal logic of actions. Computer, 57 (12):118–123, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IPCh

[2113] Anonymous. IEEE Pervasive Computing. Computer, 57(12):123, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Young:2024:BQA

[2114] Cliff Young. Beyond the quantitative approach. Computer, 57(12):124– 127, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IIPh

[2115] Anonymous. IEEE IT Professionals. Computer, 57(12):127, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:IVSg

[2116] Anonymous. IEEE Volunteer Service Awards. Computer, 57(12):128, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Au:2024:IIC

[2117] Edward Au. Introduction to IEEE Computer Society standards activities. Computer, 57(12):129–132, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICScf

[2118] Anonymous. IEEE Computers In Science and Engineering. Computer, 57 (12):132, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICScg

[2119] Anonymous. IEEE Computer Society information. Computer, 57(12): 133, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shankar:2024:CTE

[2120] Sadasivan Shankar. Challenging trends in energy of computing for data centers. Computer, 57(12):134–142, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ISV

[2121] Anonymous. IEEE Swiss vault. Computer, 57(12):143, December 2024. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Megas:2024:NCI

[2122] Katerina Megas, Barbara Cuthill, Michael Fagan, and Paul Watrobski. NIST cybersecurity for IoT program. Computer, 57(12):144–148, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICGg

[2123] Anonymous. IEEE Computer Graphics and Applications. Computer, 57 (12):148, December 2024. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICCi

[2124] Anonymous. IEEE Career Center.

Computer, 57(12):C3, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2024:ICEi

[2125] Anonymous. IEEE Computing Edge. Computer, 57(12):C4, December 2024. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCa

[2126] Anonymous. Front cover. Computer, 58(1):C1, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IACa

[2127] Anonymous. IEEE author’s choice. Computer, 58(1):C2, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCa

[2128] Anonymous. Table of contents. Computer, 58(1):1–2, January 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Ma

[2129] Anonymous. Masthead. Computer, 58(1):3, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Washizaki:2025:LEI

[2130] Hironori Washizaki. Leadership and engagement: IEEE Computer Society 2025 key strategies. Computer, 58 (1):4–8, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CHSa

[2131] Anonymous. Computer highlights Society magazines. Computer, 58(1):

9–11, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSa

[2132] Anonymous. IEEE Computer Society benefits. Computer, 58(1):11, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YAa

[2133] Erich Neuhold. 50 & 25 years ago.

Computer, 58(1):12–15, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTa

[2134] Ergun Akleman. Computing through time: Augmented AI. Computer, 58 (1):14, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:SUP

[2135] Jeffrey Voas and Keith Miller. Social unrest, political violence, and war. Computer, 58(1):16–19, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSb

[2136] Anonymous. IEEE Computer Society call for papers. Computer, 58 (1):19, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:O

[2137] Jeffrey Voas. Outlook 2025. Computer, 58(1):20–22, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSc

[2138] Anonymous. IEEE Computer Society call for special issue proposal. Computer, 58(1):23, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sinha:2025:AAI

[2139] Saurabh Sinha, J. Wynand Lambrechts, Aba Bimana, and Aili Ashipala. Augmented artificial intelligence in 5G, 6G, and beyond: a quantum leap. Computer, 58(1):24–32, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davis:2025:PDC

[2140] Keith M. Davis III and Tuukka Ruotsalo. Physiological data: Challenges for privacy and ethics. Computer, 58 (1):33–44, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAHa

[2141] Anonymous. IEEE Annals of the History of Computing. Computer, 58 (1):44, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Qu:2025:FDE

[2142] Youyang Qu, Ming Ding, Nan Sun, Kanchana Thilakarathna, Tianqing Zhu, and Dusit Niyato. The frontier of data erasure: a survey on machine unlearning for large language models. Computer, 58(1):45–57, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wang:2025:MLL

[2143] Lu Wang and K. L. Eddie Law. Model to learn like humans. Computer, 58 (1):58–65, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lopez-Pernas:2025:MCE

[2144] Sonsoles López-Pernas, Mohammed Saqr, Miguel Á. Conde, Mikko Apiola, and Matti Tedre. Mapping computer engineering education research: a topic analysis. Computer, 58(1): 66–75, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adhinarayanan:2025:LBL

[2145] Vignesh Adhinarayanan and Wu chun Feng. Looking back to look forward: 15 years of the Green500. Computer, 58 (1):76–86, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSd

[2146] Anonymous. IEEE Computer Society information. Computer, 58(1): 87, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Milojicic:2025:TP

[2147] Dejan Milojicic. Technology predictions 2025. Computer, 58(1):88– 90, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSe

[2148] Anonymous. IEEE Computing In Science & Engineering. Computer, 58 (1):90, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Duran:2025:PPS

[2149] H. B. Duran. Plug (in) and play: The state of BCI gaming in 2025: Why this 25-year-old technology is finally getting the attention it deserves. Computer, 58 (1):91–96, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blanco-Justicia:2025:ULL

[2150] Alberto Blanco-Justicia, Josep Domingo-Ferrer, Najeeb Moharram Jebreel, Benet Manzanares-Salor, and David Sánchez. Unlearning in large language models: We are not there yet. Computer, 58(1):97–100, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISPa

[2151] Anonymous. IEEE Security and Privacy subscription. Computer, 58(1): 100, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Porter:2025:RSA

[2152] Dylan Porter, Joanna F. DeFranco, and Phil Laplante. Requirements specification automated quality analysis: Past, present, and future. Computer, 58(1): 101–104, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Devadas:2025:TUC

[2153] Srinivas Devadas and Daniel Sanchez. Toward a universal cryptographic accelerator. Computer, 58(1):105–108, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITSa

[2154] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 58 (1):108, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Khaderi:2025:HHG

[2155] Khizer Khaderi, Yusuf Ahmed, and Michael Zyda. How to hire a Gen Z through gaming. Computer, 58(1): 109–116, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Talia:2025:RIS

[2156] Domenico Talia. Are robots and intelligent systems changing our mind? Computer, 58(1):117–120, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weiss:2025:TEA

[2157] Brian A. Weiss, Donald H. Costello III, and Matt Scassero. Test and evaluation of autonomous systems: Developing the educational landscape. Computer, 58(1):121–125, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITCa

[2158] Anonymous. IEEE Transactions on Computers call for papers. Computer, 58(1):125, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clark:2025:MFA

[2159] Abram H. Clark and Ishan Srivastava. Modeling flow–arrest transitions in granular media. Computer, 58(1): 126–133, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:EOS

[2160] Nir Kshetri. Economics of open source software and AI models. Computer, 58 (1):134–141, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siponen:2025:EBC

[2161] Mikko Siponen. Evidence-based cybersecurity management: How to use sanctions effectively and avoid pitfalls. Computer, 58(1):142–145, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2025:SATa

[2162] Hal Berghel. A story arc from TRON to Sora: an interview with Judson Rosebush — Part 1. Computer, 58(1): 146–150, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bash:2025:CEI

[2163] Cullen Bash, John McDonald, Dejan Milojicic, Juliane Mueller, Tom Wilson, and Ying Zhang. Convergence of energy and IT sectors. Computer, 58 (1):151–158, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IOJa

[2164] Anonymous. IEEE Open Journal of the Computer Society. Computer, 58 (1):159, January 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITPa

[2165] Anonymous. IEEE Transactions on Privacy. Computer, 58(1):160, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCa

[2166] Anonymous. IEEE Career Center.

Computer, 58(1):C3, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEa

[2167] Anonymous. IEEE Computing Edge.

Computer, 58(1):C4, January 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCb

[2168] Anonymous. Front cover. Computer, 58(2):C1, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IACb

[2169] Anonymous. IEEE author’s choice.

Computer, 58(2):C2, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCb

[2170] Anonymous. Table of contents. Computer, 58(2):1–2, February 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Mb

[2171] Anonymous. Masthead. Computer, 58 (2):3, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CHSb

[2172] Anonymous. Computer highlights Society magazines. Computer, 58(2): 4–6, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:UYP

[2173] Anonymous. Unlock your potential. Computer, 58(2):6, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YAb

[2174] Erich Neuhold. 50 & 25 years ago. Computer, 58(2):7–11, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTb

[2175] Ergun Akleman. Computing through time: Intelligence assistance. Computer, 58(2):8, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSf

[2176] Anonymous. IEEE Computer Society: Call for papers. Computer, 58 (2):11, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2025:GWE

[2177] Joanna F. DeFranco and Jeffrey Voas. Global e-waste exposed. Computer, 58 (2):12–15, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:Ia

[2178] Jeffrey Voas. In this issue. Computer, 58(2):13, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITSb

[2179] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 58 (2):15, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hole:2025:IAS

[2180] Kjell Jørgen Hole. Intelligent assistants strengthening personhood. Computer, 58(2):16–24, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Al-Madani:2025:IDD

[2181] Basem Al-Madani, Shihab Hasan, Amer Abualhassan, and Farouq Aliyu. Integrating data distribution service (DDS) in smart traffic systems: a comprehensive review. Computer, 58(2): 25–34, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITCb

[2182] Anonymous. IEEE Transactions on Computers: Call for papers. Computer, 58(2):34, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paul:2025:PCS

[2183] JoAnn M. Paul and Isaac Bettendorf. Paradox, conflict, and structural intelligence. Computer, 58(2): 35–41, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rys:2025:BBI

[2184] Maciej Ryś and Anna Maria Górska. Breaking barriers: Increasing female participation in hackathons. Computer, 58(2):42–51, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Albayyali:2025:SAD

[2185] Waad Albayyali and Aymen Belghith. Strengthening active directory security: Detecting and mitigating Kerberoasting attacks. Computer, 58(2): 52–62, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITBa

[2186] Anonymous. IEEE Transactions on Big Data. Computer, 58(2):62, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:BTA

[2187] Nir Kshetri. Building trust in AI: How blockchain enhances data integrity, security, and privacy. Computer, 58(2): 63–70, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IPCa

[2188] Anonymous. IEEE Pervasive Computing. Computer, 58(2):70, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jovanovic:2025:SDA

[2189] Mla– dan Jovanović and Mark Campbell.

Self-directing AI: The road to fully autonomous AI agents. Computer, 58(2): 71–77, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICGa

[2190] Anonymous. IEEE Computer Graphics and Applications. Computer, 58 (2):77, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mead:2025:OSS

[2191] Nancy R. Mead, Carol Woody, and Scott Hissam. Open source software: The ultimate in reuse or a risk not worth taking? Computer, 58(2): 78–83, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAHb

[2192] Anonymous. IEEE Annals of the History of Computing. Computer, 58 (2):83, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Freeman:2025:UAS

[2193] Michael Freeman and Amy Pritchett. Uncrewed aerial systems integration in coast guard search and rescue operations. Computer, 58(2):84– 89, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISPb

[2194] Anonymous. IEEE Security and Privacy. Computer, 58(2):89, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athavale:2025:SRK

[2195] Jyotika Athavale. Standards in review: Key developments with IEEE Computer Society Standards in 2024. Computer, 58(2):90–92, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wu:2025:UAOa

[2196] Yanyi Wu. The university as an AI Ombudsman — Part 1. Computer, 58(2): 93–97, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Price:2025:AAE

[2197] Christopher Price, Ichiro Takeuchi, James Rondinelli, Wei Chen, and Christopher Hinkle. AI-accelerated electronic materials discovery and development. Computer, 58(2):98–104, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IPT

[2198] Anonymous. IT Professional technology solutions for the enterprise. Computer, 58(2):104, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSg

[2199] Anonymous. IEEE Computer Society information. Computer, 58(2): 105, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Majeed:2025:SDN

[2200] Abdul Majeed and Seong Oun Hwang.

Synthetic data: a new frontier for democratizing artificial intelligence and data access. Computer, 58(2):106– 114, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSh

[2201] Anonymous. IEEE Computer Society call for special issue proposals. Computer, 58(2):115, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reddi:2025:AFA

[2202] Vijay Janapa Reddi and Amir Yazdanbakhsh. Architecture 2.0: Foundations of artificial intelligence agents for modern computer system design. Computer, 58(2):116–124, February 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSi

[2203] Anonymous. IEEE Computer Society Career Center. Computer, 58(2): C3, February 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEb

[2204] Anonymous. IEEE Computing Edge.

Computer, 58(2):C4, February 2025.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCc

[2205] Anonymous. Front cover. Computer, 58(3):C1, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IACc

[2206] Anonymous. IEEE author’s choice. Computer, 58(3):C2, March 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCc

[2207] Anonymous. Table of contents. Computer, 58(3):1–2, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Mc

[2208] Anonymous. Masthead. Computer, 58(3):3, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CHSc

[2209] Anonymous. Computer highlights Society magazines. Computer, 58(3): 4–6, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSj

[2210] Anonymous. IEEE Computer Society benefits. Computer, 58(3):6, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YAc

[2211] Erich Neuhold. 50 & 25 years ago. Computer, 58(3):7–11, March 2025.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTc

[2212] Ergun Akleman. Computing through time: Metaverse. Computer, 58(3):8, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTd

[2213] Ergun Akleman. Computing through time: Low code/no code. Computer, 58(3):10, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSk

[2214] Anonymous. IEEE Computer Society call for papers. Computer, 58 (3):11, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:2025:DFP

[2215] Keith W. Miller and Jeffrey Voas. Data feudalism and e-personhood — are we becoming digital serfs? Computer, 58 (3):12–15, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:PC

[2216] Jeffrey Voas. Pros and cons. Computer, 58(3):16–21, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2025:LCN

[2217] George F. Hurlburt, George K. Thiruvathukal, Nir Kshetri, and Norita Ahmad. Low code/no code meets the metaverse. Computer, 58(3):22–28, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSl

[2218] Anonymous. IEEE Computer Society information. Computer, 58(3):29, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Naqvi:2025:QLC

[2219] Bilal Naqvi, Damian Kedziora, Lanqi Zhang, and Shola Oyedeji. Quality of low-code/no-code development platforms through the lens of ISO 25010:2023. Computer, 58(3):30–40, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Viera-Gonzalez:2025:LCA

[2220] Perla Marlene Viera-González, Osvaldo Habib González-González, and Guillermo Ezequiel Sánchez-Guerrero. Low-code automation for resolving incidents during class schedule registration. Computer, 58(3):41–48, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhong:2025:LPL

[2221] Chen Zhong, John Oruongo, and J. B. Joo Baek Kim. LLM-powered lowcode/no-code data analytics in education and workforce development. Computer, 58(3):49–59, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mekki:2025:MEM

[2222] Yosra Magdi Mekki, Leslie V. Simon, William D. Freeman, and Junaid Qadir. Medical education metaverses (MedEd metaverses): Opportunities, use case, and guidelines. Computer, 58 (3):60–70, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wei:2025:MSS

[2223] Zhe Wei, Sen Wang, Fang Wang, and Lingqiao Shui. Metaverse for sustainable smart civil aviation: Technologies, applications, and challenges. Computer, 58(3):71–81, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2025:SATb

[2224] Hal Berghel. A story arc from Tron to Sora: an interview with Judson Rosebush — Part II. Computer, 58 (3):82–86, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2025:RAD

[2225] Phil Laplante. Reasoning about disinformation. Computer, 58(3):87–91, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISPc

[2226] Anonymous. IEEE Security and Privacy subscription. Computer, 58 (3):91, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2025:DYT

[2227] Michael Zyda. Do you think you own stock in the power company??? Computer, 58(3):92–95, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICGb

[2228] Anonymous. IEEE Computer Graphics and Applications. Computer, 58 (3):95, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conde:2025:CCL

[2229] Javier Conde, Gonzalo Martínez, Pedro Reviriego, Zhen Gao, Shanshan Liu, and Fabrizio Lombardi. Can ChatGPT learn to count letters? Computer, 58 (3):96–99, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kassab:2025:FPW

[2230] Mohamad Kassab and João Paulo Fernandes. The forgotten potential: Why aspect-oriented programming faded away. Computer, 58(3): 100–104, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IIPa

[2231] Anonymous. IEEE IT Professional. Computer, 58(3):104, March 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tortonesi:2025:CCT

[2232] Mauro Tortonesi. The compute continuum: Trends and challenges. Computer, 58(3):105–108, March 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAHc

[2233] Anonymous. IEEE Annals of the History of Computing. Computer, 58 (3):108, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wu:2025:UAOb

[2234] Yanyi Wu. The university as an AI Ombudsman — Part II. Computer, 58 (3):109–114, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITBb

[2235] Anonymous. IEEE Transactions on Big Data. Computer, 58(3):114, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hands:2025:ADS

[2236] Jonmichael Hands and Tom Coughlin.

Advancing data security and sustainability: Establishing a circular economy for storage. Computer, 58(3): 115–118, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSm

[2237] Anonymous. IEEE Computer Society call for special issue proposal. Computer, 58(3):119, March 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grunert:2025:MTI

[2238] Andreas Grünert, James Bret Michael, Rolf Oppliger, and Ruedi Rytz. On the measurability and testability of IT security. Computer, 58(3):120– 126, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITSc

[2239] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 58 (3):126, March 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCb

[2240] Anonymous. IEEE Career Center. Computer, 58(3):C3, March 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEc

[2241] Anonymous. IEEE Computing Edge. Computer, 58(3):C4, March 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCd

[2242] Anonymous. Front cover. Computer, 58(4):C1, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IACd

[2243] Anonymous. IEEE author’s choice. Computer, 58(4):C2, April 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCd

[2244] Anonymous. Table of contents. Computer, 58(4):1–2, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Md

[2245] Anonymous. Masthead. Computer, 58(4):3, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CHSd

[2246] Anonymous. Computer highlights Society magazines. Computer, 58(4):4–6, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSn

[2247] Anonymous. IEEE Computer Society call for papers. Computer, 58 (4):6, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YAd

[2248] Erich Neuhold. 50 & 25 years ago.

Computer, 58(4):7–10, April 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTe

[2249] Ergun Akleman. Computing through time: Intelligent security. Computer, 58(4):8, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:BTB

[2250] Nir Kshetri and Jeffrey Voas. Big tech, big energy, and AI. Computer, 58 (4):11–13, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hsu:2025:SIS

[2251] D. Frank Hsu and Christina Schweikert. Secure and intelligent systems. Computer, 58(4):14–16, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSo

[2252] Anonymous. IEEE Computer Society benefits. Computer, 58(4):16, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSp

[2253] Anonymous. IEEE Computer Society information. Computer, 58(4):17, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuhn:2025:MVD

[2254] D. Richard Kuhn, M S Raunak, and Raghu N. Kacker. Measuring and visualizing dataset coverage for machine learning. Computer, 58(4):18–26, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gavrilova:2025:IFD

[2255] Marina L. Gavrilova. Information fusion: a decade of innovations in biometric multimodal research. Computer, 58(4):27–36, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IPCb

[2256] Anonymous. IEEE Pervasive Computing. Computer, 58(4):36, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nakao:2025:UIT

[2257] Koji Nakao, Daisuke Inoue, and Katsunari Yoshioka. Unveiling IoT threats: a case study on darknet and honeypot analysis. Computer, 58(4):37–45, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Owusu:2025:RID

[2258] Evans Owusu, Mariyam Mapkar, Mohamed Rahouti, and Dinesh C. Verma. Robust intrusion detection with combinatorial fusion and generative artificial intelligence. Computer, 58(4): 46–57, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nisiotis:2025:EMI

[2259] Louis Nisiotis, Cecilia Metra, Jyotika Athavale, Dejan Milojcic, Stephen Dukes, Jeewika Ranaweera, and Sin-Kuen Hawkins. The emerging metaverse: IEEE ISEMV 2024 retrospective and future directions. Computer, 58(4):58–64, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tootaghaj:2025:DSS

[2260] Diman Zad Tootaghaj, Yiqian Wu, Ting He, Nilanjan Ray Chaudhuri, Dejan Milojicic, Luka Strezoski, James Anderson, and Puneet Sharma. Distributed, sustainable, and secure energy future through IT modernization. Computer, 58(4):65–73, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IIPb

[2261] Anonymous. IEEE IT Professionals. Computer, 58(4):73, April 2025.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2025:ENW

[2262] Sorel Reisman. Engineering? No way. Computer, 58(4):74–78, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Palmer:2025:IAY

[2263] Russell S. A. Palmer and Andrew M. A. Palmer. Intention is all you need to create your own Hollywood blockbuster movies: Revolutionizing filmmaking with generative AI and Saga. Computer, 58(4):79–94, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hajek:2025:HFR

[2264] Kim M. Hajek, Paul Trauttmansdorff, Sabina Leonelli, Stephan Guttinger, and Silvia Milano. How to foster responsible and resilient data: The ethical data initiative. Computer, 58 (4):95–99, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2025:ACS

[2265] Christof Ebert, Sriram Panyam, and Praveen Gujar. AI for cloud and SaaS: Technologies and business models. Computer, 58(4):100–105, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICGc

[2266] Anonymous. IEEE Computer Graphics and Applications. Computer, 58 (4):105, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oliveira:2025:SFP

[2267] Roberto Oliveira, Carla Nébele Ferreira Hanau, Valdemar Vicente Graciano-Neto, Eliomar Araújo de Lima, Vinícius Carvalho Lopes, José Maria David, Regina Braga, Wagner Arbex, and Mohamad Kassab. Smart farming for poultry: Enhancing growth and efficiency with low-cost Internet of Things solutions. Computer, 58(4): 106–110, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McHugh:2025:OSE

[2268] John McHugh. The origins of software errors: Where do they come from? Where do they go? Computer, 58 (4):111–114, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2025:SBM

[2269] Dirk Riehle. The software bill of materials. Computer, 58(4):115–120, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:PGA

[2270] Nir Kshetri. From predictive and generative to agentic AI: Shaping the future of marketing operations and strategies. Computer, 58(4):121–129, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Daniel:2025:MAF

[2271] Christopher Bud Daniel, Doron Drusinsky, and Loren E. Peitso. Multiagent aircraft flight route planning in a 3D threat-contested environment using cross-entropy search. Computer, 58 (4):130–146, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Julaihi:2025:GGR

[2272] Aqilah Julaihi and Norita Ahmad. Genomic gold rush or ethical minefield? Rethinking data practices in health tech giants. Computer, 58(4): 147–151, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Takamaeda-Yamazaki:2025:UML

[2273] Shinya Takamaeda-Yamazaki. Uncertainty in machine learning and future computers. Computer, 58(4): 152–156, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSq

[2274] Anonymous. IEEE Computer Society call for special issue proposal. Computer, 58(4):157, April 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athavale:2025:FSS

[2275] Jyotika Athavale and Darren Galpin. Functional safety standards: IEEE P2851 road map. Computer, 58(4): 158–160, April 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCc

[2276] Anonymous. IEEE Career Center. Computer, 58(4):C3, April 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEd

[2277] Anonymous. IEEE Computing Edge.

Computer, 58(4):C4, April 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCe

[2278] Anonymous. Front cover. Computer, 58(5):C1, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IACe

[2279] Anonymous. IEEE author’s choice.

Computer, 58(5):C2, May 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCe

[2280] Anonymous. Table of contents. Computer, 58(5):1–2, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Me

[2281] Anonymous. Masthead. Computer, 58(5):3, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CHSe

[2282] Anonymous. Computer highlights Society magazines. Computer, 58(5):4–6, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSr

[2283] Anonymous. IEEE Computer Society benefits. Computer, 58(5):6, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YAe

[2284] Erich Neuhold. 50 & 25 years ago. Computer, 58(5):7–10, May 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTf

[2285] Ergun Akleman. Computing through time: Architecture. Computer, 58(5):8, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeFranco:2025:YCF

[2286] Joanna F. DeFranco and Jeffrey Voas. Your coworker’s first name is artificial and last name is intelligence. Computer, 58(5):11–14, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:Ib

[2287] Jeffrey Voas. In this issue. Computer, 58(5):12, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IC

[2288] Anonymous. 2025 IEEE COMP-SAC. Computer, 58(5):15, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2025:KFK

[2289] Ying Li, Praveen Kumar Donta, Xingwei Wang, Ilir Murturi, Min Huang, and Schahram Dustdar. KDN-FLB: Knowledge-defined networking through federated learning and blockchain. Computer, 58(5):16–26, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gheisari:2025:FSD

[2290] Mehdi Gheisari, Hamid Tahaei, Christian Fernández-Campusano, Mazhar Malik, Ernest Mnkandla, Zenghui Wang, Malusi Sibiya, Julian L. Webber, Muhammad Rizwan Mughal, Cheng-Chi Lee, Panjun Sun, and Abolfazl Mehbodniya. A flexible software-defined networking-based privacy-preserving method for Internet of Things-based smart city environment based on the neighbors situation. Computer, 58(5):27–36, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSs

[2291] Anonymous. IEEE Computer Society call for papers. Computer, 58 (5):36, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patterson:2025:EID

[2292] David Patterson, Jeremy Bennett, Mary Bennett, Hélène Chelin, David Harris, Jennifer Hellar, William Jones, Konrad Moron, Paolo Savini, Roger Shepherd, Ray Simar, Zachary Susskind, and Stefan Wallentowitz. Embench IOT 2.0 and DSP 1.0: Modern embedded computing benchmarks. Computer, 58(5):37–47, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICGd

[2293] Anonymous. IEEE Computer Graphics and Applications. Computer, 58 (5):47, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gajjar:2025:DDV

[2294] Dulari B. Gajjar, Prisha Faldu, Deep Rameshbhai Kothadiya, Aayushi Pushpakant Chaudhari, and Nikita M. Bhatt. DeViTC: Deep-vision transformer to recognize originality of currency. Computer, 58(5):48–56, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Campos:2025:PFC

[2295] João R. Campos, Ernesto Costa, and Marco Vieira. Predicting failures in complex systems. Computer, 58(5): 57–64, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hurlburt:2025:WBS

[2296] George F. Hurlburt. Why bring science to the forefront? Computer, 58 (5):65–70, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IIPc

[2297] Anonymous. IEEE IT Professionals. Computer, 58(5):70, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Panyam:2025:HAA

[2298] Sriram Panyam and Praveen Gujar. How AI agents are transforming software engineering and the future of product development. Computer, 58 (5):71–77, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IPCc

[2299] Anonymous. IEEE Pervasive Computing. Computer, 58(5):77, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2025:MAA

[2300] Michael Zyda. Much ado about DeepSeek . . .. Computer, 58(5):78–81, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISPd

[2301] Anonymous. IEEE Security and Privacy subscription. Computer, 58(5):81, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Oesch:2025:AAC

[2302] Sean Oesch, Jack Hutchins, Phillipe Austria, and Amul Chaulagain. Agentic AI and the cyber arms race. Computer, 58(5):82–85, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:AAI

[2303] Nir Kshetri and Jeffrey Voas. Agentic artificial intelligence for cyber threat management. Computer, 58(5):86–90, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jovanovic:2025:EAW

[2304] Mla– dan Jovanović and Mark Campbell. Evolving AI: What lies beyond today’s language models. Computer, 58(5):91–96, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Renieri:2025:IKF

[2305] Nikolina Renieri and Dimitrios Serpanos. Interoperability: a key to the future. Computer, 58(5):97–100, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sodano:2025:CDL

[2306] J. T. Sodano and Joanna F. De-Franco. Citizen development, lowcode/no-code platforms, and the evolution of generative AI in software development. Computer, 58(5):101–104, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSt

[2307] Anonymous. IEEE Computer Society information. Computer, 58(5):105, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2025:CNE

[2308] Hal Berghel. Cyberdidacticism: The new epistemic paradigm for cognitive minimalism and generative artificial intelligence. Computer, 58(5): 106–112, May 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSu

[2309] Anonymous. IEEE Computer Society call for special issue proposal. Computer, 58(5):113, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CCa

[2310] Anonymous. CS connection. Computer, 58(5):114, May 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCd

[2311] Anonymous. IEEE Career Center.

Computer, 58(5):C3, May 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEe

[2312] Anonymous. IEEE Computing Edge.

Computer, 58(5):C4, May 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCf

[2313] Anonymous. Front cover. Computer, 58(6):C1, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITPb

[2314] Anonymous. IEEE Transactions on Privacy. Computer, 58(6):C2, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCf

[2315] Anonymous. Table of contents. Computer, 58(6):1–2, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Mf

[2316] Anonymous. Masthead. Computer, 58(6):3, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mastropaolo:2025:HAI

[2317] Antonio Mastropaolo. How artificial intelligence is reshaping our lives: a framework to unravel its complexities. Computer, 58(6):4–5, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CHSf

[2318] Anonymous. Computer highlights Society magazines. Computer, 58(6):6–8, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSv

[2319] Anonymous. IEEE Computer Society benefits. Computer, 58(6):8, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YAf

[2320] Erich Neuhold. 50 & 25 years ago. Computer, 58(6):9–12, June 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTg

[2321] Ergun Akleman. Computing through time: Data storage. Computer, 58 (6):10, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chapiro:2025:AIC

[2322] Michael Chapiro and Jeffrey Voas. Artificial intelligence, concrete, and wood. Computer, 58(6):13–16, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:Ic

[2323] Jeffrey Voas. In this issue. Computer, 58(6):14, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:WC

[2324] Jeffrey Voas. Writing for Computer.

Computer, 58(6):15, June 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSw

[2325] Anonymous. IEEE Computer Society call for papers. Computer, 58 (6):16, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2025:AAA

[2326] Phil Laplante. Assured autonomy, artificial intelligence, and machine learning. Computer, 58(6):17–28, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITBc

[2327] Anonymous. IEEE Transactions on Big Data. Computer, 58(6):28, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSx

[2328] Anonymous. IEEE Computer Society information. Computer, 58(6):29, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abramson:2025:CMM

[2329] David Abramson and Jake Carroll. A capability maturity model for research data storage systems. Computer, 58 (6):30–39, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2025:RKG

[2330] Xiang Li, Qing Liu, Quan Bai, and Xiwei Xu. The role of knowledge graphs on responsible artificial intelligence realization: Research opportunities and challenges. Computer, 58 (6):40–48, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Saxena:2025:RST

[2331] Anurag Saxena. Rethinking software testing for modern development. Computer, 58(6):49–58, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nath:2025:SEA

[2332] Suman Nath, Ryen W. White, Fazle E. Faisal, Morris E. Sharp, Robert W. Gruen, and Lenin Ravindranath Sivalingam. From search engines to action engines. Computer, 58 (6):59–68, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stavrou:2025:GAA

[2333] Angelos Stavrou and Jeffrey Voas. General and agentic AI, and the challenges of xplainable reliability. Computer, 58 (6):69–72, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSy

[2334] Anonymous. IEEE Computer Society call for special issue proposal. Computer, 58(6):73, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foster:2025:DAB

[2335] Dawn Foster. From data to action: Building healthy and sustainable open source projects. Computer, 58 (6):74–78, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2025:ITS

[2336] Christof Ebert. Innovation turns smart and green. Computer, 58(6):79–82, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kulkarni:2025:UCU

[2337] Jaydeep P. Kulkarni, Siddhartha Raman Sundara Raman, Shanshan Xie, and Chieh-Pu Lo. Unconventional computing using Ising accelerators. Computer, 58(6):83–86, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSz

[2338] Anonymous. IEEE Computer Society volunteer service awards. Computer, 58(6):87, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maley:2025:BPA

[2339] Bob Maley and Joanna F. DeFranco.

The blue pill attack as a wake-up call. Computer, 58(6):88–90, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berry:2025:WLL

[2340] Daniel M. Berry. Why large language models appear to be intelligent and creative: Because they generate bullsh*t! Computer, 58(6):91–96, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISPe

[2341] Anonymous. IEEE Security and Privacy subscription. Computer, 58(6):96, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rokne:2025:TCC

[2342] Jon Rokne. Teaching a compiler course. Computer, 58(6):97–100, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gizopoulos:2025:DSC

[2343] Dimitris Gizopoulos. The dark side of computing: Silent data corruptions. Computer, 58(6):101–106, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:BSD

[2344] Nir Kshetri. Blockchain and stablecoins: Driving the future of digital finance. Computer, 58(6):107–114, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICGe

[2345] Anonymous. IEEE Computer Graphics and Applications. Computer, 58 (6):114, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IOJb

[2346] Anonymous. IEEE Open Journal of the Computer Society. Computer, 58 (6):115, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athavale:2025:SFT

[2347] Jyotika Athavale and Richard Tong.

Shaping the future through artificial intelligence standardization efforts. Computer, 58(6):116–118, June 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IIPd

[2348] Anonymous. IEEE IT Professional.

Computer, 58(6):118, June 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:RHR

[2349] Nir Kshetri. Redefining human resource practices with AI agents and agentic AI: Automated compliance and enhanced productivity. Computer, 58 (6):119–124, June 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCe

[2350] Anonymous. IEEE Career Center.

Computer, 58(6):C3, June 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEf

[2351] Anonymous. IEEE Computing Edge.

Computer, 58(6):C4, June 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCg

[2352] Anonymous. Front cover. Computer, 58(7):C1, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IETa

[2353] Anonymous. IEEE emerging technologies. Computer, 58(7):C2, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCg

[2354] Anonymous. Table of contents. Computer, 58(7):1–2, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Mg

[2355] Anonymous. Masthead. Computer, 58(7):3, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mastropaolo:2025:PDH

[2356] Antonio Mastropaolo. Pixels of deception: How evolutionary algorithms break AI reliability. Computer, 58 (7):4–6, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSba

[2357] Anonymous. IEEE Computer Society call for papers. Computer, 58(7):6, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CHSg

[2358] Anonymous. Computer highlights Society magazines. Computer, 58(7):7–9, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITCc

[2359] Anonymous. IEEE Transactions on Computers. Computer, 58(7):9, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YAg

[2360] Erich Neuhold. 50 & 25 years ago.

Computer, 58(7):10–13, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTh

[2361] Ergun Akleman. Computing through time: Software development. Computer, 58(7):11, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rupe:2025:CTC

[2362] Jason Rupe and Jeffrey Voas. Is critical thinking critical? Computer, 58 (7):14–16, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbb

[2363] Anonymous. IEEE Computer Society benefits. Computer, 58(7):16, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bannon:2025:GAS

[2364] Tracy Bannon and Phil Laplante. Generative AI in the software development lifecycle. Computer, 58(7):17–19, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITSd

[2365] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 58(7):19, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wu:2025:KMR

[2366] Yanyi Wu, Xinyu Lu, and Chenghua Lin. Knowledge monopoly risks in generative AI-assisted software development lifecycle. Computer, 58(7): 20–30, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITBd

[2367] Anonymous. IEEE Transactions on Big Data. Computer, 58(7):30, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Muratovic:2025:GAS

[2368] F. Muratovic, D. Kearns-Manolatos, and A. Alibage. Generative AI in software development: Challenges, opportunities, and new paradigms for quality assurance. Computer, 58(7): 31–39, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IPCd

[2369] Anonymous. IEEE Pervasive Computing. Computer, 58(7):39, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jain:2025:APT

[2370] Amit Jain and Rohit Tiwari. AI for productivity: Transforming enterprise software development. Computer, 58 (7):40–52, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IIPe

[2371] Anonymous. IEEE IT Professional.

Computer, 58(7):52, July 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yalciner:2025:TSE

[2372] Aybüke Yalçıner, Ebru Gökalp, and Ahmet Dikici. Transforming software engineering processes through generative AI: a framework for integration and implementation. Computer, 58 (7):53–65, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Swaraj:2025:CCD

[2373] Aman Swaraj, Sandeep Kumar, and Lalit Mohan Sharma. ChatGPT choreography: Discovering developer dialogues and potential software development lifecycle applications. Computer, 58(7):66–78, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISPf

[2374] Anonymous. IEEE Security and Privacy subscription. Computer, 58(7):78, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vieira:2025:LLT

[2375] Marco Vieira. Leveraging LLMs for trustworthy software engineering: Insights and challenges. Computer, 58 (7):79–90, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICGf

[2376] Anonymous. IEEE Computer Graphics and Applications. Computer, 58 (7):90, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2025:RMT

[2377] Christof Ebert. The return of the magic triangle. Computer, 58(7):91–94, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbc

[2378] Anonymous. IEEE Computer Society information. Computer, 58(7):95, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sadler:2025:CEM

[2379] Brian M. Sadler. Computer education in the machine learning era: Intelligent systems. Computer, 58(7):96–99, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:BSP

[2380] Nir Kshetri. Blockchain standardization in practice: Contrasting European Union and U.S. approaches. Computer, 58(7):100–108, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbd

[2381] Anonymous. IEEE Computer Society call for special issue proposal. Computer, 58(7):109, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atchley:2025:PFS

[2382] Scott Atchley, Rosa M. Badia, Bronis R. de Supinski, Joshua Fryman, Dieter Kranzlmüller, Srilatha Manne, Pekka Manninen, Satoshi Matsuoka, Dejan Milojicic, Galen Shipman, Eric Van Hensbergen, and Robert W. Wisniewski. Predicting the future of supercomputing. Computer, 58(7): 110–120, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IVSa

[2383] Anonymous. IEEE Volunteer Service Awards. Computer, 58(7):121, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2025:CAF

[2384] Lan Zhang, Anoop Singhal, Qingtian Zou, Xiaoyan Sun, Peng Liu, and Hsiao-Ying Lin. Can AI fix buggy code? Exploring the use of large language models in automated program repair. Computer, 58(7):122–128, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IOJc

[2385] Anonymous. IEEE Open Journal of the Computer Society. Computer, 58 (7):129, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Naqvi:2025:ENC

[2386] Bilal Naqvi and Joakim Kävrestad.

Embracing neurodiversity in cybersecurity: What lies ahead? Computer, 58(7):130–133, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2025:HMM

[2387] Michael Zyda. Hype me to the moon! Let me dream among the AI stars! Let’s give up on that guy who wants to go to Mars . . .. Computer, 58 (7):134–137, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clark:2025:GAF

[2388] Christopher Clark and Armon Barton. GTS attack: Finding adversarial examples with greedy tree search. Computer, 58(7):138–146, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schmidt:2025:STG

[2389] Douglas C. Schmidt. Software testing in the generative AI era: a practitioner’s playbook. Computer, 58(7): 147–152, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAHd

[2390] Anonymous. IEEE Annals of the History of Computing. Computer, 58 (7):152, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rada:2025:GAH

[2391] Roy Rada. Generative AI and health care: Brief survey. Computer, 58(7): 153–156, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weiss:2025:BGB

[2392] Joseph Weiss and James Bret Michael. Bridging the gap between engineering and cybersecurity to better protect critical infrastructures. Computer, 58 (7):157–160, July 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2025:CSC

[2393] Dimitrios Serpanos and Panayotis Antoniadis. Cyberphysical systems in control: Risks of digital transformation. Computer, 58(7):161–164, July 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCf

[2394] Anonymous. IEEE Career Center.

Computer, 58(7):C3, July 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEg

[2395] Anonymous. IEEE Computing Edge.

Computer, 58(7):C4, July 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCh

[2396] Anonymous. Front cover. Computer, 58(8):C1, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IETb

[2397] Anonymous. IEEE emerging technologies. Computer, 58(8):C2, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCh

[2398] Anonymous. Table of contents. Computer, 58(8):1–2, August 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Mh

[2399] Anonymous. Masthead. Computer, 58(8):3, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mastropaolo:2025:SHE

[2400] Antonio Mastropaolo. Smarter, not harder: Efficient artificial intelligence training with selective data. Computer, 58(8):4–6, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CHSh

[2401] Anonymous. Computer highlights Society magazines. Computer, 58(8): 7–9, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbe

[2402] Anonymous. IEEE Computer Society call for papers. Computer, 58 (8):9, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YAh

[2403] Erich Neuhold. 50 & 25 years ago. Computer, 58(8):10–13, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTi

[2404] Ergun Akleman. Computing through time: Playing in Metaverse. Computer, 58(8):11, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:CJR

[2405] Nir Kshetri and Jeffrey Voas. Is citizen journalism reshaping “the news”? Computer, 58(8):14–18, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:Id

[2406] Jeffrey Voas. In this issue. Computer, 58(8):17, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbf

[2407] Anonymous. IEEE Computer Society information. Computer, 58(8):19, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atalay:2025:ENG

[2408] Tolga O. Atalay, Angelos Stavrou, and Lixia Zhang. Envisioning the nextgeneration cellular architecture with named data networking. Computer, 58 (8):20–29, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbg

[2409] Anonymous. IEEE Computer Society benefits. Computer, 58(8):29, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Karaarslan:2025:TRD

[2410] Enis Karaarslan and Beste Akdik. Toward reliable disaster data sharing with blockchain and zero-knowledge proofs. Computer, 58(8):30–40, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IPCe

[2411] Anonymous. IEEE Pervasive Computing. Computer, 58(8):40, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gupta:2025:MSE

[2412] Vishwas Gupta and Shikha Sharma. Meta’s strategic evolution: How Metaverse and AI are shaping the future of technology. Computer, 58(8): 41–52, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IIPf

[2413] Anonymous. IEEE IT Professionals. Computer, 58(8):52, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Villegas:2025:AAA

[2414] William Eduardo Villegas, Rommel Gutierrez, Alexandra Maldonado Navarro, and Aracely Mera-Navarrete. Adaptive authentication and access control system in dynamic educational environments based on AI. Computer, 58(8):53–63, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITSe

[2415] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 58 (8):63, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Salam:2025:CSM

[2416] Abdus Salam, Rasel Mahmud, Tohedul Islam, Saddam Mukta, and Swakkhar Shatabda. A comprehensive survey on Mamba: Architectures, challenges, and opportunities. Computer, 58(8): 64–76, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kollander:2025:INH

[2417] Mary Kollander, Robert Hollowood, Maria Williams, and Pradeeban Kathiravelu. An interactive native history web map for epidemic diseases in Alaska. Computer, 58(8):77–85, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISPg

[2418] Anonymous. IEEE Security and Privacy subscription. Computer, 58 (8):85, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trippel:2025:SFV

[2419] Caroline Trippel. Specification and formal verification of hardware–software contracts for high-assurance computer architectures. Computer, 58(8):86–91, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dey:2025:ROS

[2420] Somdip Dey. The rise of open source models and implications of democratizing AI. Computer, 58(8):92–96, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITCd

[2421] Anonymous. IEEE Transactions on Computers. Computer, 58(8):96, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2025:SDI

[2422] Christof Ebert. Software-defined innovation. Computer, 58(8):97–101, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rizvi:2025:LRU

[2423] Syed Rizvi and Anthony Demeri. Life at risk: Uncovering the urgent security gaps in Internet of Things–integrated cloud infrastructures. Computer, 58(8): 102–106, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:2025:GAT

[2424] Brooke Smith and Elizabeth Stringer. The greatest of all time: World of Warcraft’s 20 years of player commitment: Part 1: The mastery of efficacy. Computer, 58(8):107–113, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Xie:2025:IA

[2425] Xihao Xie, Jia Zhang, and Jeffrey Voas. Immunology for AI. Computer, 58(8): 114–116, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITBe

[2426] Anonymous. IEEE Transactions on Big Data. Computer, 58(8):116, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sengupta:2025:AED

[2427] Kaushik Sengupta, Jonathan Zhou, Zheng Liu, and Emir Ali Karahan. AI enabling discovery and design of radio and high-frequency wireless chips beyond human intuition. Computer, 58 (8):117–122, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICGg

[2428] Anonymous. IEEE Computer Graphics and Applications. Computer, 58 (8):122, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:GAA

[2429] Nir Kshetri. Governing agentic AI:

Security, identity, and oversight in the age of autonomous intelligent systems. Computer, 58(8):123–129, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbh

[2430] Anonymous. IEEE Computing in Science and Engineering. Computer, 58 (8):129, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ahmad:2025:DDR

[2431] Norita Ahmad, Aqilah Julaihi, and Preman Rajalingam. From data to diagnosis: Rethinking clinical decision support with explainable AI. Computer, 58(8):130–135, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISS

[2432] Anonymous. IEEE Software subscription. Computer, 58(8):135, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IM

[2433] Anonymous. IEEE Multimedia. Computer, 58(8):135, August 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barba:2025:EEG

[2434] Lorena A. Barba. Experience embracing GenAI in an engineering computations course: What went wrong and what’s next. Computer, 58(8): 136–141, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akib:2025:MMH

[2435] Abir Akib, Sanjay Rekhi, Kostas Amberiadis, and Ankur Srivastava. Metrics and methodology for hardware security constructs. Computer, 58(8): 142–148, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAI

[2436] Anonymous. IEEE artificial intelligence. Computer, 58(8):148, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISC

[2437] Anonymous. IEEE Software call for articles. Computer, 58(8):148, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Riehle:2025:OSL

[2438] Dirk Riehle. Open source license obligations: Attribution and copyleft. Computer, 58(8):149–152, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAHe

[2439] Anonymous. IEEE Annals of the History of Computing. Computer, 58 (8):152, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jovanovic:2025:BAT

[2440] Mla– dan Jovanović and Mark Campbell. Benchmarking AI: Toward inclusive evaluation of language models. Computer, 58(8):153–159, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athavale:2025:BFS

[2441] Jyotika Athavale and Louis J. Gullo.

Bridging functional safety and reliability: IEEE’s predictive maintenance standardization efforts. Computer, 58 (8):160–162, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CIC

[2442] Anonymous. Candidates for 2026 IEEE Computer Society officers and Board of Governors positions. Computer, 58(8): 163–176, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAa

[2443] Anonymous. IEEE app. Computer, 58(8):176, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IDV

[2444] Anonymous. IEEE Division VIII Delegate/Director candidates. Computer, 58(8):177–178, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITSf

[2445] Anonymous. IEEE Transactions on Security. Computer, 58(8):179, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCg

[2446] Anonymous. IEEE Computer call for special issues. Computer, 58(8): 180, August 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCh

[2447] Anonymous. IEEE Career Center. Computer, 58(8):C3, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEh

[2448] Anonymous. IEEE Computing Edge. Computer, 58(8):C4, August 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCi

[2449] Anonymous. Front cover. Computer, 58 (9):C1, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCi

[2450] Anonymous. IEEE 2026 COMPSAC call for papers. Computer, 58(9): C2, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCi

[2451] Anonymous. Table of contents. Computer, 58(9):1–2, September 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Mi

[2452] Anonymous. Masthead. Computer, 58 (9):3, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mastropaolo:2025:WDA

[2453] Antonio Mastropaolo. When databases age: How SQL Server and MySQL handle the test of time. Computer, 58(9): 4–6, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CHSi

[2454] Anonymous. Computer highlights Society magazines. Computer, 58(9): 7–9, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbi

[2455] Anonymous. IEEE Computer Society call for papers. Computer, 58 (9):9, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YAi

[2456] Erich Neuhold. 50 & 25 years ago. Computer, 58(9):10–12, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbj

[2457] Anonymous. IEEE Computer Society benefits. Computer, 58(9):12, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTTj

[2458] Ergun Akleman. Computing through time: Movie making. Computer, 58 (9):13, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rupe:2025:TRI

[2459] Jason Rupe and Jeffrey Voas. Thoughts on the reliability of information. Computer, 58(9):14–17, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:Ie

[2460] Jeffrey Voas. Also in this issue. Computer, 58(9):15, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISPh

[2461] Anonymous. IEEE Security and Privacy subscription. Computer, 58(9): 17, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zyda:2025:GFI

[2462] Michael Zyda. Game, film, and interactive experience technologies. Computer, 58(9):18–20, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IPCf

[2463] Anonymous. IEEE Pervasive Computing. Computer, 58(9):20, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbk

[2464] Anonymous. IEEE Computer Society information. Computer, 58(9): 21, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zhang:2025:SGD

[2465] Xinyue Zhang, Yi Wang, Donita Slawson, Maria Williams, Robert Boeckmann, and Pradeeban Kathiravelu. A serious game for Dena’ina language and heritage. Computer, 58(9):22– 30, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fotiadis:2025:AVR

[2466] Thomas Fotiadis, Ioanna Theophilou, George Angelos Papadopoulos, Anna-Maria Arapostathi Galani, Sarantos Psycharis, Marilena Andrikopoulou, Katerina Papagiannaki, Vassiliki Kotroni, Ioannis Koletsis, Kyriakoula Makri, Evangelos Spyrou, Giannis Saitis, Niki Evelpidou, Michalis Fitros, Konstantinos Laskaridis, and Vasilis Skliros. Augmented and virtual reality in STEAM: Virtual field trips. Computer, 58(9):31–45, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stringer:2025:MEV

[2467] Elizabeth H. Stringer. Mediated ecological validity in anthropomorphic avatar dyad parasocial interactions. Computer, 58(9):46–56, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Forrester:2025:IAI

[2468] Paul Forrester and Tae Wan Kim. Immersive artificial intelligence and the ethics of virtual experience. Computer, 58(9):57–69, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dewri:2025:IGL

[2469] Rinku Dewri. Interpretation gaps in large language model-assisted comprehension of privacy documents. Computer, 58(9):70–79, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAHf

[2470] Anonymous. IEEE Annals of the History of Computing. Computer, 58(9): 79, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krsmanovic:2025:DMT

[2471] Gala Krsmanovic, Rebecca A. Deek, and Fadi P. Deek. The development and maintenance of thinking processes and learning abilities in the era of artificial intelligence. Computer, 58(9):

80–90, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICGh

[2472] Anonymous. IEEE Computer Graphics and Applications. Computer, 58 (9):90, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IETc

[2473] Anonymous. IEEE emerging technologies. Computer, 58(9):91, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Piroumian:2025:TTY

[2474] Vartan Piroumian. The top 10 things you must do to build a (meaningful) digital twin. Computer, 58(9): 92–96, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITSg

[2475] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 58 (9):96, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCj

[2476] Anonymous. IEEE computer call for special issues. Computer, 58(9): 97, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shin:2025:PDM

[2477] Donghee Shin and Norita Ahmad. Privacy decision making in artificial intelligence. Computer, 58(9):98–103, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITCe

[2478] Anonymous. IEEE Transactions on Computers. Computer, 58(9):103, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laplante:2025:MFC

[2479] Phil Laplante. Making and formalizing claims: This is not an article about disinformation! Computer, 58(9):104– 107, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITBf

[2480] Anonymous. IEEE Transactions on Big Data. Computer, 58(9):107, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Faibish:2025:MAA

[2481] Sorin Faibish. Much ado about Chat-GPT vs. DeepSeek. Computer, 58(9): 108–111, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IIPg

[2482] Anonymous. IEEE IT Professionals. Computer, 58(9):111, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:CCE

[2483] Nir Kshetri. From crisis to comeback: The evolution and resurgence of blockchain post-2022 crypto winter. Computer, 58(9):112–117, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Coughlin:2025:CF

[2484] Tom Coughlin, Todd Kelsey, and Vinton G. Cerf. Communicating with the future. Computer, 58(9):118– 122, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbl

[2485] Anonymous. IEEE Computing in Science and Engineering. Computer, 58 (9):122, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Beale:2025:MAJ

[2486] Russell Beale. In memoriam: The academic journal. Computer, 58(9):123– 126, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berghel:2025:OGA

[2487] Hal Berghel. An overview of generative AI acceptable use policies by universities with top 25 computer science programs (2025). Computer, 58(9):127– 134, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbm

[2488] Anonymous. IEEE Computer Society volunteer service awards. Computer, 58 (9):135, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Serpanos:2025:SDP

[2489] Dimitrios Serpanos. Security by design: Paving the way to cybersecurity and safety certification. Computer, 58 (9):136–139, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Majeed:2025:CPL

[2490] Abdul Majeed and Seong Oun Hwang. Changes in the privacy landscape in recent years: Analysis from AI and non-AI contexts. Computer, 58(9):140– 146, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ID

[2491] Anonymous. IEEE DataPort. Computer, 58(9):146, September 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:CCb

[2492] Anonymous. CS connection. Computer, 58(9):147–148, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAb

[2493] Anonymous. IEEE app. Computer, 58 (9):148, September 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCk

[2494] Anonymous. IEEE Career Center. Computer, 58(9):C3, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEi

[2495] Anonymous. IEEE Computing Edge.

Computer, 58(9):C4, September 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:FCj

[2496] Anonymous. Front cover. Computer, 58(10):C1, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCl

[2497] Anonymous. IEEE 2026 COMPSAC call for papers. Computer, 58(10): C2, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:TCj

[2498] Anonymous. Table of contents. Computer, 58(10):1–2, October 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:Mj

[2499] Anonymous. Masthead. Computer, 58 (10):3, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:EC

[2500] Anonymous. Elsewhere in the CS.

Computer, 58(10):4–6, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbn

[2501] Anonymous. IEEE Computer Society call for papers. Computer, 58 (10):6, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Neuhold:2025:YA

[2502] Erich Neuhold. 50 & 25 years ago. Computer, 58(10):7–9, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akleman:2025:CTT

[2503] Ergun Akleman. Computing through time: Invention vs. application. Computer, 58(10):10, October 2025. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2025:WDY

[2504] Christof Ebert and Jeffrey Voas. What did you learn today? Computer, 58 (10):11–13, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Athavale:2025:MBH

[2505] Jyotika Athavale, Dejan Milojicic, and Nita Patel. Megatrends for the benefit of humanity: AI, health, robotics, and space. Computer, 58(10):14– 18, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCm

[2506] Anonymous. IEEE Computer call for special issues. Computer, 58(10): 19, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pauloski:2025:ADC

[2507] J. Gregory Pauloski, Kyle Chard, and Ian Foster. Agentic discovery: Closing the loop with cooperative agents. Computer, 58(10):20–27, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shan:2025:ALT

[2508] Richard Shan. AI that learns, thinks, and acts: The next frontier of generative AI. Computer, 58(10):28– 39, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Qin:2025:EIM

[2509] Yinshuang Qin, Yufei Zhao, and Lingxue Zhao. Exploring the ideas of meta-learning and knowledge distillation: Core concepts and future development trends of large models. Computer, 58(10):40–48, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Velikic:2025:FAM

[2510] Gordana Velikić, Ivan Velikić, Milan Vidaković, Ivan Kaštelan, Jovana Vidaković, and Dusan Marić. Forecasting AI-Modular decision support system evolution in regenerative health care. Computer, 58(10):49–60, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbo

[2511] Anonymous. IEEE Computer Society benefits. Computer, 58(10):60, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Amjad:2025:RCP

[2512] Maheera Amjad, Mahnoor Gulbin, and Muhammad Usman Shahid Khan.

Robotic capsules: Precision in ingestible robots. Computer, 58(10): 61–68, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ISPi

[2513] Anonymous. IEEE Security and Privacy subscription. Computer, 58(10): 68, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liu:2025:CSA

[2514] Fengyuan Liu, Lizy Kurian John, and Ravinder Dahiya. Cognizant and socially aware robotics. Computer, 58 (10):69–77, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhardwaj:2025:ULE

[2515] Ketan Bhardwaj, Vaibhav Bhosale, and Ada Gavrilovska. Ushering in the low earth orbit compute cloud era. Computer, 58(10):78–86, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shah:2025:AE

[2516] Chirag Shah and Ryen W. White. Agents are not enough. Computer, 58 (10):87–92, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Koch:2025:ROS

[2517] Christian Koch. The role of open source in an AI arms race. Computer, 58(10):93–96, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IVSb

[2518] Anonymous. IEEE Volunteer Service Awards. Computer, 58(10):97, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Faibish:2025:GVC

[2519] Sorin Faibish. Gemini versus ChatGPT and DeepSeek: Much ado about crawling. Computer, 58(10):98–101, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mascolo:2025:AWS

[2520] Cecilia Mascolo. AI on wearable sensing data for human health and fitness. Computer, 58(10):102–105, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITBg

[2521] Anonymous. IEEE Transactions on Big Data. Computer, 58(10):105, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voas:2025:DT

[2522] Jeffrey Voas and Vartan Piroumian.

Digital “Twinlets”. Computer, 58(10): 106–108, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITPc

[2523] Anonymous. IEEE Transactions on Privacy. Computer, 58(10):109, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reisman:2025:AWL

[2524] Sorel Reisman. Acronyms we love: STEM, AI. what else? Computer, 58 (10):110–115, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITSh

[2525] Anonymous. IEEE Transactions on Sustainable Computing. Computer, 58 (10):115, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moradi:2025:SAI

[2526] Morteza Moradi, Mohammad Moradi, and Ali Borji. Scaling alone isn’t enough: Toward building bridges between AI models. Computer, 58(10): 116–121, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drusinsky:2025:GAR

[2527] Doron Drusinsky and James Bret Michael. Generative AI regulation is an undecidable problem. Computer, 58 (10):122–130, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ITCf

[2528] Anonymous. IEEE Transactions on Computers. Computer, 58(10):130, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rizvi:2025:CCU

[2529] Syed Rizvi, Anthony Demeri, and Mohammad R. Rizvi. Convenience at a cost: The urgent need for data privacy standards. Computer, 58(10): 131–135, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kshetri:2025:NMG

[2530] Nir Kshetri. From niche to mainstream:

The global acceleration of cryptocurrency payments. Computer, 58(10): 136–145, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICGi

[2531] Anonymous. IEEE Computer Graphics and Applications. Computer, 58 (10):145, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mastropaolo:2025:CCC

[2532] Antonio Mastropaolo. Code, chaos, and clarity: Neurosymbolic approaches to trustworthy software automation. Computer, 58(10):146–151, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IAHg

[2533] Anonymous. IEEE Annals of the History of Computing. Computer, 58 (10):151, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ebert:2025:GAD

[2534] Christof Ebert. The generative AI dividend: Cost reductions in software and IT. Computer, 58(10):152– 156, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weiss:2025:PBC

[2535] Joseph Weiss, James Bret Michael, and Michael T. Swearingen. Physics-based cyberattacks against electric power grids and alternating current equipment. Computer, 58(10):157–161, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:IPCg

[2536] Anonymous. IEEE Pervasive Computing. Computer, 58(10):161, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Li:2025:HRS

[2537] Sixu Li, Chaojian Li, and Yingyan Lin. Hardware race for scalable real-time 3D neural reconstruction and rendering. Computer, 58(10):162–167, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICSbp

[2538] Anonymous. IEEE Computer Society information. Computer, 58(10): 168, October 2025. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICCn

[2539] Anonymous. IEEE Career Center. Computer, 58(10):C3, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:2025:ICEj

[2540] Anonymous. IEEE Computing Edge. Computer, 58(10):C4, October 2025. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).